A Bibliography of Publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer

(1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 June 2016 Version 3.14

Title word cross-reference

3 [Jan99b]. < [Vil96]. > [Vil96]. C^{3} [Ste98b].

* [Mey99].

-/C [Mir92].

.HLP [Sch93k, Sch93d, Dav93a, Dav93b]. .INI [Cou94a, Cou94b, Gra94]. .RES [Sch93l, FR93].

/C [Mir92].

1 [Eri99b, Sch94v]. 1.0 [Tha97]. 1.1 [HL99]. 1.2 [DDJ98f, HL99]. 128-bit [DKR97].

1394 [Ale99a, Ale99b]. 16-bit [Abr90b, Huf93c, Sip96b, Sip96a]. 160 [BDP97]. 1620 [DDJ99g].

2 [And90a, And90b, AJ90, Bar93a, BDE93, Cla97, Fol94, Gig93b, Gig93a, Hal93, Jan99a, Kan94, Kin91a, Mak90, NRG^{+}99, Old94, Pro94, Ric91, Rod94b, Rod94c, Rot90, Ruc96, Sch90g, SRA^{+}94, Ste98d, Wil94e, Wri94]. 2-D [Cla97, Hal93]. 2.0 [Fai96]. 2.x [BDE93]. 2000 [MSE^{+}99, Bha98, MF98, Ste98f]. 232 [Cro94a]. 2B [Eri99a, DDJ99a]. 2D [Loe99]. 2PANE [Wil94d].

3 [Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Abr96a, Fre95, Hal93, Har92, Har94a, Koc97, Lew94, Mar94a, Naj95, SL96b, SL96a, SC97, Spi94]. 3-D [Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Abr96a, Hal93, Koc97, Lew94, Naj95, SL96b, SL96a, SC97, Spi94]. 32 [Bur94]. 32-Bit [Nic93a, Huf93c, Nel92, Sch92b, Tew97, Wil93a, Bur94]. 386 [Bro94b]. 386BSD [JJ92m]. 3D [DDJ99f, SWF^{+}99].

4 [Col96e, Low98a]. 4-MB [Col96e]. 4M [CJ94].

5 [Wil97d]. 5.0 [Ste97h].

6 [Lom98]. 64-Bit [Nic93a, Hir92c]. 68HC05 [Dic95a, Dic95b, FN98, Van91]. 68HC05-based [Dic95a, Dic95b, FN98]. 68HC11 [ABCE99].

7.0/386 [Sch90f].

80*86 [Abr91a]. 80*87 [Wil92c].

9 [Dib96, PPTT91, Yam96]. ’94 [Sch94i, Vli94, Koo94a]. ’95 [Eri95n, Red95a, Sch95n, Sip96a, Sip97, Smy97, Wil98b]. 95/NT [Sip97, Wil98b]. 96 [Ste96n, Ste96o, Ste96p, Ste96q, Ste96r]. 9660 [JJ92f]. ’97 [SW95a]. 99 [Ste98f, Swa99n].

= [Wil90a].

A. [McC96b, Val94c]. AARD [DDJ98g, Sch93c]. Above [GGL^{+}95]. Above-real-time [GGL^{+}95]. Abrash [Gul95]. Abstract [Hem96b, Sha97]. Access [Ber97b, ML93, Nee99, Rob95b, Sip96d, Ham93, Hig97, RC97a, RC97b]. Accessing [Bal99, Bau92a, Fri90a, Fri90b, Rei93b]. accuracy [Woe97c]. Achieving [NK98]. ACM [Bar99d, DDJ99b]. Acquisition [Cop98, FV93b, GE98, Lin94a, Nol90].

Acronyms [Wil98c]. across [Len97]. Act [Mül94, Sch94m]. Active [Bet98, BN97, dSW98]. ActiveX [HT98, SW97f, SW97d, SW98b, Wil97c, Wil97d, Wil97i]. activity [Kno93a, RC97a]. Actor [Fra90, Hat91]. Acyclic [PH95b, PH95a]. Ada [Smy97, Gal91, Moo95, SvA^{+}98, Smy97]. Adaptec [QG97]. Adapter [Gad96, GS97]. Adaptive [Dee94, Kie99a, Kno98, PB98b, Che92]. Adding [Bit91, Boy98c, Fra91, Hoe98, Ken94, Ros95a]. Addison [Cha95d]. Addison-Wesley [Cha95d]. Address [Fer95]. Adjusting [Lia90]. ADPCM [Kie97a]. Adrian [McC97]. Advanced [McB97, Saw94]. Advances [DDJ98e]. Adventure [Val95k]. Adventures [Swa98b]. After [Fos92]. Again [Eri95n, Tre94]. Against [Bro95c]. Age [BCP^{+}99, Cro98b, Wal95b, War91]. Agent [DL99, LRT99, Gra97]. Agent-Based [DL99]. Agents [SNC97]. Ah [QTC^{+}99]. AHA [QG97]. AHA-1540 [QG97]. aid [vdW97b]. Aided [Bak94]. Air [SvA^{+}98]. Airport [Swa97y]. Airtime [Die93]. Al [Ste97g, Ste97e, Ste97f]. Alan [Wal97]. Algebra [Jaf93, VP96]. Algebraic [HC90, Liu90]. Algorithm [All90, Ben96, BS98a, Ben98, Ben99a, Ben99b, Ben99c, Boy97, Boy98a, Boy98b, Byl99, Cla97, Col96a, Col99, Col96c, Cor96, CC96, Cou96, DKR97, DeV96, GH99, GS94, Gun98, Gut99, Hei96, Hof95b, Jan99b, Jen97, KR96, Kau97, Kie94a, Kie97a, Kie97b, Kie99a, Kie99b, Kis97, Kno98, Mon97, Nel96a, Neu98, Pes96, PB98b, Pig99, PS97a, PRB98, Riv95, Sav98, Sch93b, Sch94b, Sch94c, Sch94e, Sch94d, Sch95k, Sch95l, Sch95c, Sch95d, Sch95e, Sch95f, Sch95g, Sch95h, Sch95i, Sch95j, Sch96d, Sch96e, Sch96f, Sch98b, SC97, SS99, Sit96, Sta98a, Sta99, Ste97a, Swa93a, Swa93y, Swa93z, Swa93-27, Swa93-28, Swa93-29, Swa93-31, Swa93-30, Swa93-32, Swa94z, Swa94-27, Swa94-28, Swa94-31, Swa94-29, Swa94-30, Swa96y, Swa97x, Wil97a, Wit97a, Cla95, GGL^{+}95, Nel97, Sch94f, Sta94c]. Algorithms [BAC^{+}99, Ben99a, BR96, Cel93, Cro94b, DR93, Fla96, Her98, Jar90, KP99, Kno98, Man99a, Man94a, Mat96, ME96, McR96c, Och94b, Pan99, PRB98, Sch94g, Spi93, Sta93a, Syl93, TKG^{+}99, Val93f, WEM^{+}99, Fer94, Kis97, Mor91e, SC97, Zig96a, Och94c]. Allaire [Hig97]. Alley [BS98a, Ben98, Ben99a, Ben99b, Ben99c, Boy97, Boy98a, Boy98b, Byl99, Col96a, Col99, Cor96, CC96, GH99, GS94, Gun98, Gut99, Hei96, Jan99b, Jen97, KR96, Kau97, Kie94a, Kie97a, Kie97b, Kie99a, Kie99b, Kis97, Kno98, Mon97, Nel96a, Neu98, Pes96, PB98b, Pig99, PS97a, PRB98, Sav98, Sch94b, Sch94c, Sch94e, Sch94d, Sch95k, Sch95l, Sch95c, Sch95d, Sch95e, Sch95f, Sch95g, Sch95h, Sch95i, Sch95j, Sch96d, Sch96e, Sch96f, SC97, Sta98a, Ste97a, Swa93a, Swa93y, Swa93z, Swa93-27, Swa93-28, Swa93-29, Swa93-31, Swa93-30, Swa93-32, Swa94z, Swa94-27, Swa94-28, Swa94-31, Swa94-29, Swa94-30, Swa96y, Swa97x, Wil97a, Ben96, Sit96]. Alliance [Flo95g, Swa99d]. Allocation [Bol97, GH99, Ric98, Sch90c, Spe90]. Alone [DDJ98e]. Alpha [SW95c, Wit95a]. also [RC97a, Sto97]. Alternative [Smi93b]. Although [Zha97]. Always [BCP^{+}99, Bak97b]. Am29000 [Low92]. America [Val94d]. Amiga [DDJ98c]. analogies [Swa95n]. Analysis [Ben99a, Che96, Hof95b, McG94b, MK97, Mot91]. Analytic [Col99]. analyze [MK97]. Analyzer [Cod93, Pat90b]. Analyzers [MK97]. Analyzing [BAC^{+}99, TKG^{+}99, WEM^{+}99]. Andrew [Abr92a]. Animal [Swa98q]. animating [Wes94]. Animation [Dav95a, Jam90, Shi94, Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Ken94].

Annealing [Pri94]. Announced [Bar99c, DDJ98h]. antilock [Dic95b]. AntiPatterns [MM98a, Ste98d]. Antitrust [Sch94n]. Antlr [Sch99b]. any [Hig97]. Anyway [Swa99u]. Apache [BEK99]. API [MSE^{+}99, CPRB93, Con94a, DAM96, Han93, Jon98, Kis96, Lis98, Loe99, Low98a, Low98b, OK98, Rad98, Sax92a, Sch96b, Sch98c, Tuc98, Wil97f]. APIs [CP99, Nor96, She93]. App [AW99, DDJ99f, Kru94b, Val95a, Ger93]. Apple [Swa95a, Swa97i]. AppleScript [Sim95]. Applet [BC98, Ber96d, Ber97b, Ber97j, Ber97e, GS98, Kie99d, Mec97, Moo98]. Applets [BC99, Bri99, CM97, Rod95c]. Appliances [DY99a, DY99b]. Application [Bar99e, Ber97l, Boo94, Dow93, Man99b, Rob95b, Row97, SM93, Tha97, Val92e, Fir93b, Kno90, Kry92b, RC97a, Tro91b]. Application-Access [Rob95b]. Applications [Alm96, Bis96, Bit91, Bun94, Dol95, Dru95, DR97, Gut99, Hla94, Mor98, Rod94a, Sel99, Sue95, Tet94, Tew97, TSR98, Val93b, Van93, Var94b, Wil96c, Bon93, Boy98c, Esp97, Gla91, Hof97, Kle91c, Mir92, Pro92, Sha93a, Ste97f, Tha97, Wal97, Zha97, van98b]. Applicators [Cha93]. Applying [NA96, SW97a]. Approach [FV93b, JJ91d, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ92g, JJ92h, JJ92i, JJ92j, JJ92k, JJ92l, Ste93w, JJ91o]. Approximation [Ple95, SC97]. approximations [Nic91b]. Apps [CM97, Hla95c, Ros95a, Qui97]. Aptitude [SDB^{+}99]. Arbitrary [Cal95b, Mot93]. Architecture [Ber98e, Con99, Hou99, Pat97, Sch94x, SW98b, Ste94y, Tre99a, Var95a, Var95b, RK92, Con98]. Architectures [Day92, Val93f]. Area [Cho99]. areas [Nic91b, SW96b]. aren’t [Qui97]. Arguments [Ste99b, TKG^{+}99, Whi96]. Arithmetic [Ham95a, Mot93, Nel91a, Kal92, Kal93]. Arms [Ste96l]. Array [Pea91, Rei93d, Bau92a]. Arrays [Bau92a, Web95, But99, Ros93]. Art [Eri99j, Kep94, McG93]. Artificial [Swa94a]. Ask [Flo94i]. Asleep [BAC^{+}99]. ASM [Hyd92, WS90]. ASP [Bet98]. Assembler [Ive90]. Assembly [Abr90a, Ber98f, Hyd90, Pat90a, Ski91, Wil92c, Hof97, Maz92, McS92]. Assembly-Language [Ber98f]. Assertions [PSS98a]. Asserts [Ros97]. Assistive [Laf99]. Association [Pap95]. Associations [For94, Hel98, Lin96]. Associative [Blu90, Web95]. Astronomy [DDJ99e]. Asynchronous [Fag98a, Ver96]. AT89C2051 [Gad97]. ATA [CS95]. ATASPI [Cal95c]. ATL [AN98, SW97c, SW97b, SW98c, SW98b, SW98a]. ATM [Che96]. Atmel [Gad97]. Attached [Gru95]. Attack [Sch99c]. Attribute [Cha94, Cha95c, Cho95]. Auctions [GW97, OK98]. Audio [Cod94a, Cro94b, Kno98, Moo98, Rat92, Rat93, Rat94, Rei94c, Row94a, Zud96, TWH93]. auditory [Ram97a]. Authentication [Wal95b, Sta99]. author [Car97c]. Authoring [TMP^{+}99]. authors [MK97, QG97, Ste94p]. autoconfiguration [JJ91a]. Automata [Nay93]. automate [McC97]. Automated [BD97, Bea98, Sel99, Swa97y, vdW97b]. Automatic [FHW97, Kie98]. Automatically [Ait96a]. Automating [Ban98b, Fly96, MR99, Obl90, Pap95, Ohl90]. automation [Ban98a, Hig96]. Automotive [Cop98]. autonomous [Car97c]. Auxiliary [Ros95a]. Available [DDJ98c, vdW97b]. Averaging [Gun98]. AVK [Ful92]. avoid [Kis96]. Avoiding [McR93, New94a]. Award [DDJ99b]. Awarded [DDJ98b]. Awards [Eri96m, Eri98a, Eri95a, Eri97a]. Aware [Dah93]. Away [DDJ99d]. Awk [Bal90]. AWT [Gat97].

B [Cha95d, DDJ99b, JS92]. B-trees [JS92]. B. [Och94b]. Baby [Eri91a]. Back [Eri94m, Eri95n, DDJ99a, Ste92a, ML97]. backdoor [Sch90g, Swi90]. background [Bau90a]. Baggage [Swa97y]. Bamba [Swa98k]. Bandwagon [Bar99c]. Banker [Eri98b]. Bankers [PSS^{+}98b]. Bar [Rei94a, Swa98r, Swa98s]. Based [DL99, Flo91b, MC98, Sch90a, Uhl99, Bet91a, Dic95b, FN98, JJ95b, Lea93, Pap95, Day92, Dic95a]. Bashing [Swa97f]. Basic [App97, Ham95a, Lom98, RB97, Wil97d, JJ91s, JJ91b, JJ91c, JJ91a, Esp96, Flo95f, LaO98, Lic92, Nor94, Sar95, Sch92a, Sch94x, Ste91w, Xia95]. Basketful [Wil99d]. batch [Law97]. batch-oriented [Law97]. Be [Bro95b, Eri99i, Nel97, Tha97]. Bean [Hou98]. Beans [Swa97f, Gig97]. Beautifier [Mah92]. Beautiful [Eri99i, Sha98c, Wil99e]. bed [CT90]. Bedfellows [Eri94l]. before [Ste97e]. Begin [Hol96, Swa99f]. Behavior [Ana94, Rob98a, FV93a, McC97]. Behind [Swa98v]. bellatlantic.net [PJR^{+}98]. Bemer [Swa98a]. Bench [Oro98]. Benchmark [Kar90, Oro98]. BenchmarkBook [MR96]. Benchmarking [McR96a, McR96b, Swa98b]. Benchmarks [Col98b, McR96c]. benefit [Hof97]. Beowulf [DDJ98a]. BeRays [Rus99]. Berkeley [DDJ99c, Sha93b]. Best [Bro95d, Eri97c, Och94c, Val95c]. Beta [Man93]. Better [BEK99, BCP^{+}99, CM97, Hes93, TMP^{+}99]. Between [Tes96a, Hav98a, Kis97, Mir92, Pan99, Ste97e, Tha97, Yam96]. bevy [MK97]. Beyond [Mal96, SHR^{+}99, Wil96a]. Bézier [Kin90a]. Bi [Gil95a, Gil95b]. Bi-Endian [Gil95a, Gil95b]. Bicubic [LP90]. Bidirectional [Blu90]. Big [Ste99j, Swa96x, Kal93, Val91b]. big-number [Kal93]. Bill [SWF^{+}99]. Binary [Cop95, Hof96, Luu94, Boy97, Lin96, Rod91, You98]. Binary-Data [Luu94]. Binding [Cre98]. Biochemical [Pea95].

Biocomputing [DDJ98b, Val91d]. BIOS [BH95, Cod94a, Eri90m, RT97]. bis [CH97b]. Bit [Blu95, Nic93a, Rog95a, Abr90b, DKR97, Hir92c, Huf93c, Nel92, Sch92b, Sip96b, Sip96a, Tew97, Wil93a, Bur94, Gil95a, Gil95b]. BitBlt [Nic91b]. Bitmaps [Dix95]. Bits [Bak93]. Bjarne [Ste92q, Ste92-32]. Black [Kep94, Sha98a, Sch92c]. Blarney [Swa99c]. Blending [Sta98c, Wit95a]. Blind [Eri97b]. bloating [Mye97]. BLOBs [Fel92]. Block [DKR97, Dee94, Kru92, Sch94w, Sch94v, Zyt94a, Zyt94b]. blocking [FY96]. Blood [Swa97g]. Bloom [Sta94b]. Blow [Ste99j]. Blowfish [Sch94f, Sch95k]. Blur [Wit97b]. BMP [Cha95a, Cha95b, Lus94]. Bob [Bet95a, Dut95, Swa98a]. Book [DDJ99b, MBK^{+}98, Nic96a, NFP96, Nic96b, Swa96l, Swa99k, Wil98d]. Books [GG95a, Gur95, Swa97f, Wil99d, Bar99c]. Bookshelf [Ams99, BF97, Ber98e, BDT97, Cha95d, Cro98b, Cro98a, Dun92a, Dun93a, Dun93b, Dun93c, Dun95, Eri93m, Eri93n, Eri93o, Eri94n, Eri94o, Eri95o, Fer98, FV97, Flo94h, Gab96, Gra98, GG95b, Gri96a, GF99, Gul95, JL98, Jol92a, Jol93c, Jol95b, Kim97l, Kim99i, LS99, McG98b, McR96c, ML97, Och94a, One94c, Pfe96, Rei94b, Rot98, Rot99, Sch92e, Sch92f, Sta98b, Ste93x, Ste93y, Ste94q, Ste94r, Ste95m, Swa95b, Tay99, Val93d, Val93e, Val95j, Var94a, Var94c, Var94d, Var95d, vdWS97, Wil97i, Wil97j, Wil97h, Wil97g, Wil97k, Wil98d, Wil98c, Wil99d, WC99, Wil99c, WS99, Wil99e, You96, You97, Cla96b, Gri96b, McC96b, Mit96, Sal96, Sta96]. Boom [Val94d]. Boost [The98b]. Boot [RT97]. Bootstrap [McG94b]. Borland [Bet95c, Bor96, Bro95c, Fai95, Fai96, Flo94b, Flo95a, Ste94a]. Boston [Bar99c]. bottleneck [Fly97]. Bounding [Sea90]. bounds [Pea91]. Bowl [Swa99n]. Box [Pro94, Sea90, New94b]. Boxes [Sar92, New93c]. Bozoki [Woe97a]. brake [Dic95b]. branch [Joh97b]. Breach [JJ95a]. Break [DDJ98d]. Breakthrough [DDJ98g]. Breathing [Ste98f]. Bridge [Lu94]. Bridges [Swa97k]. Bridging [Gri97a, Mir92]. BRIEF [Rod91]. Browse [Ber97c]. Browser [Dav95a, GW96, Tic97, Kle91b]. Browser-side [Tic97]. Browsing [Eri97c]. Bryan [One94d]. Bryce [RC97a, RC97b]. BSD [Fag98b]. BSP [Abr95a, Abr95d, Bet93b, Dwy95]. Btrieve [Rei93c]. Buffer [BH95]. buffering [Ton91]. buffers [Tho91b]. Bug [AB97, Bak97a, Coe95, Col97a, Col98c, DDJ98c, The98b, TKG^{+}99, McC97]. bugs [Nel93a, PSS98a]. Build [Cro94a, Hav98b, McC97, Sto97]. Builders [JLC98]. Building [Ano96a, Bro96a, CM97, Col96a, Dol95, Fal98, Flo93b, Flo94c, Ges90, GW97, Gor93, Hla95a, Kry92b, LaP95, Li94a, MC98, Mir96, Mor98, NV90, OH95, Reg94, Row97, Var94b, Wei96a, Wei96b, Yor96, Car97c, Kis96]. builds [Nac97]. Bumper [Eri99b]. Burns [DDJ98b]. Burrows [Nel96b]. bus [Mor91d]. Bushel [Wil98c]. Business [Eri98j, van98b]. Buy [Eri97c]. Buzz [Val95e]. Byte [Hof97]. Byte-oriented [Hof97].

C [Cul94, Eck90, Fai95, Gil95a, Gil95b, Hel93, LaP95, Lad90a, Moo93, Rob97, Rod95a, Sak92, Ste95l, Ste95n, Ste97h, Ste98i, vdW97a, vdWS97, Wen99, Yam94, Ait96b, Ste94j, Ano96b, Arn96, Arn98, Aus97, Bak97b, Bal90, Bar93b, Bau92b, Bea98, BR93, Bet91b, Bev98, BR96, Bla96, Blu90, Blu95, But93, Car92, Car96a, CR92, Cav95, Cha96, Cha93, Cha94, Chu92a, Chu91, Con95, Dav91, Dud90, Ell90, EJ95, Esp97, Fin95a, Fla96, Fle93, For94, For95, Fra99, Gib94, Gra98, Gro97b, Hav98a, Hew92, Hil95, HC99, Jae91, Jae93, Jae95, Jer92, JM92, Kal92, KD97, Kay96, Kef92, Kim95, Kof95, Lad90b, Lad91, Lak97, Lan90, Lea94, Lew94, Lin96, Lip99a, Lyp97a, MGL^{+}98, Mah92, Man94b, Maz92, McB97, ME96, Med94, Mer94, Mey92, MK97, Mil91, Mir92, MBK^{+}98]. C [Mud96, Mye97, Mye98, Nac97, Nel97, Nes94, Ozi91, Pap95, Par95a, Par95b, Pea91, Pen94, PR91, Pla92, Pom98, Ran94, Rei93d, Rei94a, Ric91, Rog95a, Ros93, Rot99, Row94b, Sae95a, Sae95b, Sar92, Sch90b, Sch95a, Sch90c, Sch90e, Sch90f, SW95b, Sem95, Shi92, dSW98, Sin92, Spe97, Sta92, Ste93a, Ste90a, Ste90l, Ste90b, Ste90c, Ste90d, Ste90e, Ste90f, Ste90g, Ste90h, Ste90i, Ste90j, Ste90k, Ste91a, Ste91b, Ste91c, Ste91d, Ste91e, Ste91f, Ste91g, Ste91h, Ste91i, Ste91j, Ste91k, Ste91l, Ste92p, Ste92c, Ste92d, Ste92e, Ste92f, Ste92g, Ste92h, Ste92i, Ste92j, Ste92k, Ste92l, Ste92m, Ste92n, Ste92o, Ste92r, Ste92s, Ste92t, Ste92u, Ste92-34, Ste93c, Ste93d, Ste93e, Ste93f, Ste93g, Ste93h, Ste93i, Ste93j, Ste93k, Ste93m, Ste93l, Ste93n, Ste93o, Ste93p]. C [Ste93q, Ste93r, Ste93s, Ste93t, Ste93u, Ste94b, Ste94c, Ste94d, Ste94e, Ste94f, Ste94g, Ste94h, Ste94i, Ste94k, Ste94l, Ste94m, Ste94n, Ste94o, Ste94y, Ste94s, Ste94t, Ste94u, Ste94v, Ste94w, Ste94x, Ste95l, Ste95a, Ste95b, Ste95c, Ste95d, Ste95e, Ste95f, Ste95g, Ste95h, Ste95i, Ste95j, Ste95k, Ste96m, Ste96l, Ste96a, Ste96b, Ste96c, Ste96d, Ste96e, Ste96f, Ste96g, Ste96h, Ste96i, Ste96j, Ste96k, Ste97g, Ste97j, Ste97l, Ste97h, Ste97i, Ste97k, Ste97e, Ste97f, Ste97b, Ste97c, Ste97d, Ste98d, Ste98e, Ste98b, Ste98c, Ste98f, Ste98g, Ste98i, Ste98j, Ste98h, Ste98k, Ste98l, Ste99j, Ste99a, Ste99b, Ste99c, Ste99k, Ste99d, Ste99e, Ste99f, Ste99g, Ste99i, Ste99h, Sto92a, Sto92b, Sto92c, Sto92d, Sto93, Sue95, Sul94a, Sul94b, Sum91a, Sum91b, Sum91c, Sum91d, Sum92a]. C [Sum92b, Sum92c, Sum92d, Sum92e, Sum92f, Sum92g, Sum92h, Tar91, Tea91, TMP^{+}99, TFS94, Van91, VP96, Vel97, Vig93, Vil95, Vio93, Vog91, Vos94, Web95, Wee90, Wes94, Wil96a, Wil98a, Wil92d, WSM95b, WSM95a, Wri96, Yam95, Yam96, Yue96a, Yue97, Zig95]. C-language [Bak97b]. C-tree [Mud96]. C/C [Bea98, HC99, Jae93, Par95a, Par95b, Esp97, McB97]. C4.5 [Mon97]. C talk [Ber90]. C talk/Views [Ber90]. CA [dS95, da 95]. CA-Visual [dS95, da 95]. Cabinet [Sch97a]. cables [Mus99]. Cache [Dan96, Hel93, Cro93a]. Cached [McC96a, Ano96i]. caching [Cha92]. Café [Ano96c, Hem96a]. Calabaza [Sha99c]. Calculus [Jaf93]. Calendars [Lam99, Mey93]. Call [Sip96d, Ste97l, WSM95b, WSM95a, RC97b]. Call-Level [Sip96d]. Callable [Bet95a]. Callbacks [Ses98, Sta92]. called [RC97a]. Calling [Ros93, Sip96b, Sip96a]. Calls [Fag98b]. Cambridge [Sch94g]. Can [Ber96f, Ber97a, Bro95b, Eri98b, Eri99i, Pit99, PSS^{+}98b, Ran99b, Hig97, McC97, Mye97, Nel97, Nic91a, Zha97]. Can’t [Bar99a, DDJ99b]. Capabilities [Gil95a, Gil95b, Esp97]. Capacities [The98b]. Capture [Sym94]. Capture/Playback [Sym94]. Capturing [Gre92, Con91]. Car [The98b, Eri91a]. Card [Bar99e]. Cards [Con98, Con99, DDJ98d, DDJ98b, DDJ99d]. Care [Eri98j]. Cargo [Wil97i]. Cargo-Cult [Wil97i]. Carpal [DDJ99b]. Cars [Eri99b]. cart [BW96, Ste97e]. Case [Wit95b, RK92, Wes94, Arm95, Hag91, Smi97]. Castanet [Ber98a]. Casting [Sem95]. Catastrophes [Har96b]. categories [Soo98b]. caught [Vel97]. Caveats [Col97e]. CCD [FNL91a]. CD [Ano96i, Jol93a, Har93a, Jol93b, Li94c, ML93, McC96a, Ste96l, TMP^{+}99]. CD-ROM [Jol93a, Har93a, Jol93b, Li94c, Ste96l]. CD-ROMS [ML93, Ano96i, McC96a]. CE [Rad98, Sch98a, Tuc98, Bar99d, Die99, Hav98b, Hav99b, LHB^{+}98, SDB^{+}99, Ste98l, Wil98e, Wil98f]. Celestial [FNL91b, FNL91a]. Cellular [Nay93].

CeRapiInvoke [MSE^{+}99]. Certificates [Tre99c]. CGI [Dut96, Joh97a, Kim95, LaO98, Sim95, Ste97i]. Chad [Swa99n]. Chain [Rot90]. Challenge [DDJ98f]. Challenges [MF98, PSS^{+}98b, Car97c]. Changed [Abr95b]. Changes [DDJ98e, McC97]. Changing [Eri91a, Flo91b, Val91a, Sch91b]. Channel [Ber97a, Ber98a]. Channels [BC98, Ler97]. Chaos [KS97, Swa99o, All91]. Character [Kar93, Rez94a, Rez94b, Val94c, Kin90a, Pla92]. Chart [Rus93, Wil96h]. Chat [Ber96e, Swa98a, Ser93]. Cheaper [DDJ99e]. Cheaps [Eri98d]. check [Nel92]. Checking [Hol97, Oua93, Pea91]. checksum [Kod92]. Chicago [Val94e]. Child [Wei93]. Chip [Wu97]. Chips [DDJ99e]. Choice [Val94g, Wil97g]. Chuck [Woe92a]. Cilwa [Flo94h]. Cipher [DKR97]. Circles [Kie94a, Pat90b, Zig90]. Circuit [Col97c, Bar92]. circuiting [DAM96]. Civic [CS94]. Clara [Swa97k]. Clarifying [GF99]. Clashing [Eri99d]. Class [Bet94d, Con95, Gar96, Gib94, Hil95, Jol96, JLC98, Mus93, Nor93, Rei93d, Rei94a, Ste95n, Tea91, Val92e, Wri96, You98, Cra90, Mey92, Yam96]. Classes [Dav91, Gri99, HHW^{+}99, How98, Ran99b, Sch98a, Bev98, Lam95a, Sae95a, SW96b, Sin92, Ste95l]. classic [GHJV96]. Classics [Och94c]. Classifying [Mon97]. Clear [Eri99c]. Clickable [Dav95b]. Client [Bis96, Bla96, Car96b, Day92, Duf95, Gam96b, Tou99, Law95, Wit95b, Sto97]. Client/Server [Bis96, Car96b, Day92, Gam96b, Law95, Wit95b, Bla96]. clients [Len97]. Climbing [Swa98c]. Clinical [Gam96b]. Clipping [Abr96a, DRG90, DRG96]. Clock [Swa97t, Roa91]. CLOS [Swa90a]. Closing [Swi90]. Clustering [DDJ98a]. Clusters [DDJ98h]. CNS [Ber90]. Cobol [HHW^{+}99, SW95a, SDB^{+}99]. Cocktail [Bat96]. Code [Bal90, Ben99b, Boo92, Fal98, Fom99, Fra99, Gri97a, Gul95, Hen97a, Mar94b, McG94a, Nes94, Sch94a, Swa98n, van98a, Dri94, Loy92, MK97, Sch93c, Sto97, Vog91, Wil96b, Wil98b]. CODER [Sto90]. Coderunner [And91]. Codes [Can94b, HC90, Lyp97a, Mak90, McM90, Rei94a, Can94a, Liu90]. CodeStore [Nac97]. CodeWarrior [Ste97k]. CodeWizard [Ste98b]. Coding [Dav95c, Dee94, PB98b, Wit97a, Nel91a]. COFF [Mor91a]. Coffee [Eri99d, DDJ99a]. cognitive [Joh93]. Cold [Col98e, Hig97, Swa97v, Hig97]. Collection [CR92, PB98a, Spe97, Jac90, vdW97b]. Collections [Eck90]. Collision [Rob95a]. Color [Cla96a, SS93, Kru94a]. ColoradOS [Ruc96]. ColoradOS/2 [Ruc96]. colored [Var91]. Colors [Smi93a]. COM [Ban98a]. Combinations [Swa96y]. Combinatorial [Pea95]. comes [QG97]. Coming [Col98e]. Comm [Flo95f]. Command [TKG^{+}99, Whi96, Kyl91]. Command-Line [TKG^{+}99, Whi96]. Commands [KD97]. comment [Ait96a]. Commerce [Asl98, DDJ98d]. commercial [JJ91r]. commercially [vdW97b]. Committees [Ste98b]. Common [Ana95, Ple95, Ber99]. Common-Fraction [Ple95]. Commons [Fel93]. Communication [BC98, Bol97, Gri97a, Mec97, Mus99, Pan99, Rod94c, Sta95b, TKG^{+}99, Bla96, Fre92, Lam95a, McC91]. Communications [Die99, Fag98a, GS98, Sch95m, Bra94c, Gas90, Sax92a, Sch92d]. communicator [Jam92]. Compact [DeV96]. Comparing [Arm95, Flo93a, Flo93c, Joh99]. Comparison [BL95]. Compatibility [Tho92]. Competing [Bro95a]. Competition [DDJ98g]. Compilation [Hof95a, Eps96]. compile [Arn98]. compile-time [Arn98]. Compiler [App97, Dri94, Law97, Sch99b, Sch90d, Ste92p, Bet91b, Lan90, Sch90f, Smy97]. Compiler-specific [Ste92p]. Compilers [Dah93, HQ98, LHB^{+}98, vdW97a, Par95b]. Compiling [Abr95a]. complete [Kis97]. completeness [Sch94j]. Completing [JJ92m]. completion [vdW97b]. complex [Bon93]. Compliant [Pud98, Hig97]. complicated [SW96b]. Complying [Bak95]. Component [Flo91b, GC98, Rus96, Val94a]. Component-Based [Flo91b]. Components [Hoe98, LHB^{+}98, PH99, Shi99, Wil95c, Swa97y]. Compound [Wil93d]. compress [NN97, Nel97]. Compressing [Row94a]. Compression [Ano91, Deg94, Dio93b, Dio93a, Kno98, Lin95a, Lin95b, Mat96, MFCM96, MF98, Nel91b, Nel97, Pig99, Rei94c, Sea90, Syl93, Yor95, Yu96, Bri91, Kru92, Nel96b, Rat92, Reg90, Tho91a]. Computer [BAC^{+}99, Bak94, Ber98e, DDJ99g, Dor95, HT98, Kae94, Tre93, Swa97y, Woe97b]. Computer-Aided [Bak94]. Computers [DDJ98h, ML97, Sch95m, Woe92b]. Computing [DDJ98d, DDJ98b, DDJ99d, DL99, Lar95, Rus96, SM93, Sch93a, Vel97, Blo95, Joh93, PPTT91]. Con [DDJ99f]. concealing [Joh97b]. Concept [Eri99c]. Concepts [Bro95b, Swa95n]. concurrency [Ver96]. Concurrent [Ano96b, KD97, Hol97, Yue96a, Yue97]. Conditional [Eps96]. Condor [TL95a, TL95b]. Conference [DDJ98a, Dor95, Flo95a, Ruc96, Sch94k, Flo94b]. Conferences [Sch95o, Ste98b]. Conferencing [Bet94a]. Configuration [Ale99a, Ale99b, Har91]. confused [Tha97]. confusing [SW96b]. Congestion [Sta95a]. Congratulations [DDJ99e]. Connection [SW98a, Ful92]. Connections [SW97d]. connectivity [Tro91a, Tro91b]. Consequences [Bar99c]. constant [Gra91]. Constrain [Med94]. Constrained [Sto95, Pat97]. Constructing [CM95, Swa96y, Wil95c]. Construction [Bet94a, Sch99b, Won92]. constructors [But99]. Container [Dav91, Pro94, Sha97, Bat94]. Content [RR98]. Contest [Ano91, DDJ98a, Nel91b, Val92b, Val93a]. Contests [DDJ98h]. Context [Ben99b]. Continued [Ben96]. Continues [DDJ98d]. Contouring [Gil92]. Contours [SL96b, SL96a]. Contract [MBK^{+}98, PJR^{+}98, Ran99a]. Contracts [Bro95e]. Control [Ber98f, Flo95f, HHW^{+}99, Hav99a, HT98, Kry92a, Mel98, MC98, NRG^{+}99, Sar95, Sch94w, Sch94v, SW98b, Sta95a, Ste98a, Tro97, WEM^{+}99, Wil96h, WS96d, Zyt94a, Zyt94b, Car97c, Kry92b, Lyp97a]. Control-Flow [Ber98f]. Controller [Coh92, Moe90, Low92]. controllers [Gol90]. Controlling [Bau90a, Har99]. Controls [Ana95, AN98, Bri93, Wil96g, Wil97d, Flo94h]. Conventional [DDJ99a]. Conversation [Flo94d, Kim98a, Ste92q, Woe96a, Woe97a, Woe97b, Woe97c, Woe99]. Conversion [Fly96, But93, Hig96, Mor91b]. Conversions [Kie98]. Converting [Del97, Ste92a]. Convolutional [Lyp97a]. Cookies [Tic97]. Cooperative [Tar91]. Coping [Sch99a]. coprocessor [DY91]. coprocessors [Let91]. copy [Bur95b]. Copying [Bro95b]. Copyright [Bro95c]. CORBA [Bar99a, Ber97i, Bet95b, FHW97, Ham99, Hou98, Mor98, Pud98]. CORBA-Compliant [Pud98]. Cordic [Jar90]. Core [CH97a, Sch96h, Sch97b, Szc93, Zha97]. Corner [Col96c, Col97e, Col97c, Col97d, Col97b, Col97f, Col98a, Col98d, Col98e, Sch93k, SNH93, Sch93e, Sch93f, Sch93g, Sch93h, Sch93l, Sch93i, Sch93j, Sch94w, Sch94v, SRA^{+}94, Sch94o, Sch94p, Sch94q, Sch94r, Sch94s, Sch94t, Sch94u, Sch95p, Sha98i, Sha98g, Sha98c, Sha98f, Sha98d, Sha98e, Sha98b, Sha98h, Sha99e, Sha99d, Sha99l, Sha99b, Sha99k, Sha99f, Sha99h, Sha99i, Sha99g, Sha99c, Sha99j, SW96a, SW96b, SW97a, SW97f, SW97c, SW97b, SW97d, SW97e, SW98c, SW98b, SW98a, WS96c, WS96d, Col96b, Col96d, Col96e, WS96b]. Correction [Dor99, HC90, Lyp97b, Kod92, Liu90]. Cost [Ben98, Kod92, Smy97]. Could [Swa99g]. counts [Gri97b, McC97]. County [Swa97k]. coupled [McC97]. Courier [GH96]. Cowlishaw [Woe96a]. CP [Ste91m, Ste91n, Ste91o, Ste91p, Ste91q, Ste91r, Ste91s, Ste91t, Ste91u, Ste92r, Ste92s, Ste92t, Ste92u, Ste92v, Ste92w, Ste92x, Ste92y, Ste92z, Ste92-27, Ste92-28, Ste92-29, Ste92-30, Ste92-31, Ste93v, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u]. CPI [SW95b]. CPI-C [SW95b]. CPU [Bar99d, Har94b]. CPUID [Col96c]. CPUs [Sch92b]. Cracking [Ben99c]. Cram [New94a]. Crash [Dun92b]. CRC [Hen97b, Nel92, Wil98c]. CRCs [Kie97b]. Create [Ber96b, Ber96f, Ber97g, Ber97a, Ber97f, Ber97e, Ber97d, Ber98a, Moo98, Tes96b]. created [Tha97]. Creating [Bri99, Dix95, FSW98, Hav98b, JS94, Ram97b, Sip96c]. Creative [Swa99q, Swa95n]. Creator [DDJ99b]. creep [vdW97b]. Creole [Ces96]. Cross [Bro94a, Car97a, Cop95, Dio93b, Dio93a, FH95, Fir93a, Lam95a, NRG^{+}99, SDB^{+}99, JJ91q, Cul94]. Cross-Platform [Cop95, Dio93b, Fir93a, NRG^{+}99, SDB^{+}99, Bro94a, Car97a, Dio93a, FH95, Lam95a, Cul94]. Crosses [Val95h]. Crouch [SWF^{+}99]. Crouch-Echlin [SWF^{+}99]. Cruising [DDJ99b, Lan90]. Cryptanalysis [Sch96g]. CryptoAPI [Boy98c]. Cryptographic [BDP97, DC98, PRB98, Sch95o]. Cryptography [DDJ98a, Eri99j, Fer99, JM97, PKA^{+}98, Smi94, Sti98, Joh97b, Sch92d, Woe97b, vdWS97]. CSORT [Ste90a]. CSplitterWnd [SW96b]. CSplitteWnd [SW97a]. CTrace [Cra90].

Cubic [Cou96]. Cult [BCP^{+}99, Wil97i]. Curmudgeon [Ste92b]. Curmudgery [Gut92]. Current [Soo98b]. CURSES [Ste90l]. Curve [Fer99]. Curves [Gut99, HHW^{+}99, JM92, JM97, Kin90a]. Custom [Bri93, Flo94h, Sch99e, Hav98b]. customize [Zha97]. Customizing [Ana94, Ber98g, Wil96c, Wil96d, WS96a, WS96c]. cut [Kru94a]. Cutting [Eri99c]. CVS [BAC^{+}99, Per99]. CX [Aye90c]. Cyberspace [Bro95f]. cycles [Bra94a]. cyclical [Nel92].

D [Och94b, Sch93l, Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Abr96a, Cla97, Hal93, Jan99b, Koc97, Lew94, Naj95, SL96b, SL96a, SC97, Spi94, Ste91m, Ste91n, Ste91o, Ste91p, Ste91q, Ste91r, Ste91s, Ste91t, Ste91u, Ste92r, Ste92s, Ste92t, Ste92u, Ste92v, Ste92w, Ste92x, Ste92y, Ste92z, Ste92-27, Ste92-28, Ste92-29, Ste92-30, Ste92-31, Ste93v, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u, Ste95l, Ste98f, Wri94]. D-Flat [Ste91m, Ste91n, Ste91o, Ste91p, Ste91q, Ste91r, Ste91s, Ste91t, Ste91u, Ste92r, Ste92s, Ste92t, Ste92u, Ste92v, Ste92w, Ste92x, Ste92y, Ste92z, Ste92-27, Ste92-28, Ste92-29, Ste92-30, Ste92-31, Ste93v, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u, Ste98f, Wri94, Ste95l]. Dancing [Eri98c]. Dangerous [DDJ98h]. d’Art [Flo91c]. Data [Ano91, BAC^{+}99, Ber98d, Ber98c, Bet98, Boy98a, Cha95c, Cho95, Cop98, Die93, FV93b, FV93c, For98b, GE98, Gam96b, Hig97, Jam93, Kis98, Lia90, LR92, Luu94, Mat96, Mül94, Nel91b, Nel96b, Nol90, PF90a, PF90b, Per99, Sch94k, Sch99d, Sch94m, Sea90, Sou99, Sto90, Tes96a, WV91a, Var93d, Wal92b, Wit95b, Bat94, GWL96, Gil92, Hav98a, Kim92, Mye97, Reg90, SC97, Ste97g, Tho91a, WV91b, Cha94]. Data-Compression [Mat96]. Database [Ber97b, Bis96, Cel93, The98b, Dow93, Ell93b, Hig98, KD97, Lip99b, Mud96, Nor94, Obl90, Rod94a, Sav98, Sch96b, Sch98a, Sha93d, Sul94a, Sul97, Bar93b, Bra93, Bro94a, Dem93, Ell93a, FH95, Fel92, Flo93b, Hig97, KR91, Ohl90, Ste93b]. Databases [Reg94, Sul93, Wil94a]. datacentric [Bon93]. Date [BCP^{+}99, MF98]. Dave [Bet93a]. David [Nic96a, One94d]. Davis [Sch93k, Sch93d]. Dawning [Cro98b]. Day [Abr95b, Lar95, SDB^{+}99]. Days [The98b]. DBMS [Gam96b]. DBperl [Sch96b]. DCE [Blo95, Dan95, Yam95, Yam96]. DCTs [Kie99b]. DDA [MSE^{+}99, Cla97]. DDE [New93b, Sha93a]. DDJ [Con99, Ano91, Ano96j, Ano97f, Ano97g, Ano97h, Ano97i, Cus90a, Joh90, Nel91b, Val93a]. DDK [QG97]. DDV [Ber98g]. DDX [Ber98g]. DDX/DDV [Ber98g]. Deadlock [Tei98, Ver97]. deadly [Kis96]. Deal [Eri97l, Sch94n]. deals [Joh97b]. Death [Swa99l]. Debug [Low98b, Sto97]. Debug/Trace [Sto97]. DebugAPI [MSE^{+}99]. Debugger [Dri94, Man92, Pot92, Kno92b, Mil91, Sto97, Tho92]. Debuggers [Wil97g]. Debugging [Hus92, New93a, QTC^{+}99, SJB92, Spe90, Val93b, Wil90b, Ada92, Bar92, Col97c, Per93, Pie92b, Too90]. December [Con99]. Decimal [Mor91b]. decimation [SC97]. Decision [Col96a, Sch99a, Ste97a]. Decoder [Ale99a, Ale99b, Lyp97a]. decomposition [All91]. Defective [Bro96b]. definable [Sch96a]. deflate [Nel97]. Déjà [Bar99c]. Delayed [WS99]. delegate} [Wil99a]. Delphi [Fai96, Bet95c, Bor96, Kau96a, Lis98, Low98a, Wil95c, Wil96c, Wil96g]. Delta [Kno98]. Deltas [Yor95]. Demystifying [Abr90b]. Denver [Swa97y]. Department [Eri98k]. Deploying [Dan95]. Derby [The98b]. describes [Boy97, McC97, Yam96]. describing [Car97c, RC97a]. Description [Boy99]. Design [Ban98b, Che96, CM97, DDJ98g, Fer98, Gro97a, Gro97b, HM95, Jol96, Kau96b, Kum95, Kut92, Lak97, MBK^{+}98, Ngu96, Och94b, PJR^{+}98, Ran99a, Rem97, Rie96, Ros94, She99, SDB^{+}99, Var92, Var94e, Ver96, Wil97j, Coo96, Dic95a, Dic95b, Hag91, Hol97, Hop91, Kap91, Mae92, Nic91a, NA96, RK92, RC97b, Tom96, Ver97, GHJV96]. designed [Len97, Zha97]. Designing [Bon93, Boo94, Bra92, Car97c, CT90, Dru95, GH95a, Gar91, Gou96, JLC98, Man99a, Nic91a, Nog99, PPTT91, Pot92, Rob98b, SW95b, Sin92, Wil97b, vdW96, Che92, JJ91o]. designs [QG97]. Desktop [Var93d]. Detail [CH97b]. detailed [RC97b]. Details [Col98a]. detect [McC97, MK97]. Detecting [Lee93a]. Detection [Hel93, HC90, LHM93, Moo93, Rob95a, Liu90, Sch93c]. Determination [Abr96b, DGR90]. Detrimental [Mar95a]. Developer [Dor95, Flo95a, Sch95b]. developers [Bro94a, Tic97, Tre98]. Developing [Cav95, Mey92, Rod94a, Sch99e, TSR98, Tro91b]. Development [Ana93, BD97, Bar99e, Bec94, Bro95e, Car96b, Cha95d, Cop94, Cul94, Esp97, HQ98, Law95, Mir95, Nor94, Old94, Par95a, QG97, Rem97, Rod95b, Sch93a, Sch94x, Ste93w, Tet94, Wal95a, Wil95b, Wil98e, Wil98f, Wit95b, Bla96, Bro94a, Hat91, Hem96a, Hig97, Kno90, Met96, Pom98, Sch95a, Smi97, Ste96n, Ste96o, Ste96p, Ste96q, Ste96r, vdW97b, Wal97, Eri91a, Flo91b, Val91a]. Device [Kno92a, Krt98, Li94c, One94d, QG97, Rot90, Ten98, Wil98e, Wil98f, JJ91a, JJ92e, JJ92a, JJ92b, Kyl91, Ste97g]. Device-Driver [QG97]. Devices [BAC^{+}99, FN98, Gam96a, Har99, Won99]. DFWrap [Ste95l]. DGI [Mon96]. diagrams [Dru94]. Dialect [Nil98]. Dialin [Eri94a]. Dialing [DDJ98e]. Dialog [Sar92, New93c, Wil96d]. dictionary [Flo90a]. Did [Swa99m]. Diff [Cop95]. Difference [KS97, Kie96b]. Differences [Her98, PSS^{+}98b, Yam96]. Differential [Bri91, Pat90b, PS97b, Sch96g, Syl93]. Dig [Sha99d]. Digital [Boy98c, Deg94, Flo94a, HS94, MM99, Pat90b, RR98, Rus93, Wal95a, Gre92]. dilation [Ada92]. Dime [Eri98b]. Dimensional [Sem96, Str90]. dimensioned [Ros93]. Direct [Ham93, Pie96, Rob96, Wil98b]. Directed [PH95b, PH95a, Coo96, Man94a]. DirectPlay [How96, Mis96, Wil97f]. Disappearing [vdWS97]. Disassembler [Zha97]. Disco [Bar99c]. Discontiguous [Gun98]. Discounts [DDJ98f]. Discovering [GC98]. discrete [WJ92]. Discrimination [BCP^{+}99]. Disk [Wil92b, Cha92, Rea90, RC97a]. disparate [CL97]. Dispatcher [Kre99]. Display [BA96, Ber96a]. displays [RC97a]. Distance [DDJ98f]. distances [Val92d]. Distributed [Blo95, BN97, Dol95, Jen95, Kra91, TL95a, Mor98, Pom96, Pra95, Rou98, Rus96, SM93, Sch93a, Sue95, vdW96, PPTT91, TL95b, TSR98]. Distributing [Kis98]. Distribution [DL98]. Ditch [Har95d]. Dithered [Ste92a]. Ditherer [GS94]. Dive [Wil96i]. diversity [Mae92]. Divide [Val95h]. division [Gra91]. divisor [Gra91]. DLL [Tew97]. DLLs [Sch90d, Sip96b, Sip96a]. DMA [Nol90]. Do [AW99, BA96, Ber96e, Ber96d, Ber96b, Ber96c, Ber97b, Ber97j, Ber97g, Ber97f, Ber97l, Ber97e, Ber97d, Ber97k, Ber97h, Ber98a, Ber98d, Cri98, DDJ98f, GS98, Gri99, Gur95, Moo98, Pit98a, Ses98, Swa98b, Wil99a, Ber96a, Ber96f, Ber97i, Ber97c, Ber98b, Ber98c, Cop94, WW99]. Dobb [Eri95a, Eri96m, Eri97a, Eri98a, Eri97b, Val92b]. doc [Ait96a]. document [Ste96m]. Documentation [Ait96a, Dav93a, Dav93b, FHW97, Sch93k, Sch93d]. Documenting [Dav93a, Dav93b, Sch93k, Sch93d]. Documents [Wil93d]. Does [Eas99, MGL^{+}98, Pur98, Ran99a, SW97d, Ber96a, Mye91a, Mye97, Wal97]. Doing [Swa98r, Val92a, Yam96]. Dollars [Eri94a]. Domain [Tro98]. Donald [Woe96b]. Don’t [DDJ99f, Eri91a, Swa98b, Bak97b, Sto97]. Door [Eri94m]. Doorknob [Ste99b]. DOS-extender [Wil92a]. DosQProcStatus [Fol94]. Down [Eri97k, JJ91p, van98b, Kno93a]. DPMI [Hew92, One92]. Dr [Eri96m, Sha99d]. Dr. [Eri95a, Eri97a, Eri97b, Eri98a, Sha98i, Sha98g, Sha98c, Sha98f, Sha98d, Sha98e, Sha98b, Sha98h, Sha99e, Sha99l, Sha99b, Sha99k, Sha99f, Sha99h, Sha99i, Sha99g, Sha99c, Sha99j, Val92b]. Draft [Bar99c]. Drag [Ber96c, Pur98]. Drag-and-Drop [Ber96c, Pur98]. dramatically [McC97]. Drawing [Kin90a]. Drawings [Jan99b]. Dreams [Swa98j]. Drive [The98b, Eri91a]. Driven [For95, Joh97a, SG97, Bar93b]. Driver [MSE^{+}99, QG97, Ten98, Wil98e, Wil98f, Kno92a, Rea90]. Drivers [Krt98, Li94c, McC95, Soo98b, Wel94, JJ92e, JJ92a, JJ92b, Kyl91, QG97, One94d]. Drives [DDJ98a]. Drop [Ber96c, Pur98]. DSL [The98b]. DSP [Bit91, Che92, Cro94b, Lea94, ME96, McR96c, Rei94c]. Dual [SW97b, Ban98a, Med94]. Dual-UI [Med94]. Dummies [Ste96l]. Duntemann [Flo94h]. Duplication [Wil92b]. DVI [Gre92]. Dylan [Bet92, BH94, Ces96]. Dynamic [Arn96, Bol97, Cre98, Gro97b, Hig97, Lin94c, Rou98, Sar92, Sha93b, Syc90, Yu96, Car96a]. Dynamically [Ber97c, Kla99].

e-Book [Swa99k, DDJ99b]. E-books [Bar99c]. E-Commerce [Asl98, DDJ98d]. e-mail [Ber96d, Flo94c, AH97, Obl90, Ohl90, Sch94h]. E-memes [Gut94]. each [RC97b]. early [PSS98a]. ease [QG97]. easier [Mye98]. easily [Zha97]. easy [Hel98, Kla99, Ros97, Wil98b]. Eavesdropping [Kno93b, Kno93a]. Ecco [Sha99d, Sha98i, Sha98g, Sha98c, Sha98f, Sha98d, Sha98e, Sha98b, Sha98h, Sha99e, Sha99l, Sha99b, Sha99k, Sha99f, Sha99h, Sha99i, Sha99g, Sha99c, Sha99j]. EccoFan [PKA^{+}98]. Echlin [SWF^{+}99]. EchoNets [Gut94]. Economics [MMV94]. Ed [DDJ98f, MGL^{+}98, MBK^{+}98]. edge [Ban98a]. editing [Rod91]. Edition [Hla96a]. Editor [Bek97b, Cra97, Fly96, Hen97b, Ste99c, Ste99e, WL97, Wil96i, TSR98, AB97, Bak97a]. Editorial [Eri90a, Eri90b, Eri90c, Eri90d, Eri90e, Eri90f, Eri90g, Eri90h, Eri90i, Eri90j, Eri90k, Eri90l, Eri91b, Eri91c, Eri91d, Eri91e, Eri91f, Eri91g, Eri91h, Eri91i, Eri91j, Eri91k, Eri91l, Eri91m, Eri92a, Eri92b, Eri92c, Eri92d, Eri92e, Eri92f, Eri92g, Eri92h, Eri92i, Eri92j, Eri92k, Eri92l, Eri93a, Eri93b, Eri93c, Eri93d, Eri93e, Eri93f, Eri93g, Eri93h, Eri93i, Eri93j, Eri93k, Eri93l, Eri94l, Eri94m, Eri94b, Eri94c, Eri94d, Eri94e, Eri94f, Eri94h, Eri94i, Eri94j, Eri94k, Eri95b, Eri95c, Eri95d, Eri95f, Eri95e, Eri95g, Eri95h, Eri95i, Eri95j, Eri95k, Eri95l, Eri95m, Eri96l, Eri96a, Eri96b, Eri96c, Eri96d, Eri96e, Eri96f, Eri96g, Eri96h, Eri96i, Eri96j, Eri96k, Eri97c, Eri97e, Eri97d, Eri97f, Eri97g, Eri97i, Eri97h, Eri97l, Eri97j, Eri97k, Eri97m, Eri97n]. Editorial [Eri98b, Eri98c, Eri98d, Eri98e, Eri98f, Eri98g, Eri98h, Eri98i, Eri98k, Eri98j, Eri98l, Eri99a, Eri99j, Eri99b, Eri99c, Eri99d, Eri99f, Eri99g, Eri99k, Eri99h, Eri99i, Eri99e, Flo91a, Flo93d, Flo94e, Flo94f, Flo94g, Flo95b, Flo95c, Flo95d, Flo95e, Flo96a, Flo96b, Flo96c, Flo96d, Eri94g, Kie96a, Val92c]. Editors [Kin99, Val93f]. Eds [BCP^{+}99]. Effect [Dix95, SWF^{+}99]. effective [McC97]. effectiveness [Cop94]. Effects [Wit97b]. Efficiency [Och94b]. Efficient [GH99, GM95, PB98b, Kis97, NV90, Pan99]. Efficiently [Duv91]. Egos [Swa98p]. Eiffel [How93]. Einstein [SvA^{+}98]. Electrochemical [DDJ99d]. Electronic [BEK99, Cox92, DDJ98c, FSW98, Koo94b, SDB^{+}99, WEM^{+}99]. element [Ver93]. Elements [AH97, GHJV96]. Ellipses [Kie94a]. Elliptic [Fer99, JM97]. Elliptic-Curve [Fer99]. Else [Bro96c]. Emacspeak [Ram97a]. embed [Esp97]. Embeddable [Sad99]. Embedded [Bar99e, Bek97a, BLE96, CP99, DDJ98a, Dru95, dFIC96, Fly97, GS97, HQ98, Kau96b, Kre99, Kut92, LHB^{+}98, Mel98, MC98, Ric98, Rob98a, RK92, Tra90, Yue97, Cro93a, DM91, Low92, Per93, PR91]. Embedded-System [Kut92]. Embedded-Systems [Dru95, Per93]. Emerges [Har96a]. EMM [Oka94]. Employer [Bro95a]. Empty [Mye97]. Emulating [Tro94]. Emulation [Col97c, Moc94, Tro91b]. Emulator [ABCE99, Hav99b, Tew97]. emulators [Bar92]. Enabled [Won99]. Enabling [Pan99, Ram97a, MK97]. Encapsulating [Sch90c]. encoded [Lyp97a]. Encrypt [Ber98c]. Encrypted [AW99]. Encrypting [SS99]. Encryption [DDJ98c, DDJ98f, DDJ98i, KR96, MGL^{+}98, Riv95, SvA^{+}98, Sch93b, Sch95k, Sch95l, Sch98b, Smi93b, Sto90, Sch94f]. End [DDJ98c, IR99, Ste97f]. End-User [IR99]. Endian [Gil94b, Gil94c, Gil94a, Gil95a, Gil95b, Rog95c]. Endian-Neutral [Gil94a, Gil94b, Gil94c]. ends [RC97a]. Enemy [Sha99j, Sha99j]. Engine [Ell93a, Ell93b, EJ95, Kie99d, Kin90b, Rog95c, Sem95, Sem96, Ang99, Avi92, NV90, HCKS99]. Engineering [Bar99d, GPVV98, Wal95a]. Engineers [DDJ99c, Wil98d]. Engines [Bar99a, Mar90, Swa99g, Fel92]. Enhanced [Col98d, GS97]. Enhancing [RCS91a, Hat91, RCS91b]. Ensure [GS98]. Entropy [Tho91a, Wit97a]. Entropy-Coding [Wit97a]. enums [Eas99]. Envelope [Ste99k]. Environment [Ana93, Czu90, Gab92, Gor93, Low95, Hat91, PPTT91, QG97, Ste96n, Ste96o, Ste96p, Ste96q, Ste96r]. Environments [Gou96, McG98b, Par95a, SM93, Sch93a]. EPROM [Moc94]. Equation [Med94].

Equations [KS97]. Era [DDJ98c]. Ergonomic [DDJ98c]. Erickson [Eri97g]. Errata [Con99]. Error [Can94b, Can94a, Dou90, HC90, Lyp97a, Lyp97b, Mak90, McM90, Kod92, Liu90, Sto97]. Error-control [Lyp97a]. Error-Recovery [Can94b, Can94a]. error-spotting [Sto97]. Errors [Nel93b, Spe90, Nel93a, vdW97b]. Escrow [DDJ98i, Mar94b]. essential [Swa97y]. Ethernet [LS99]. Eurocrypt [Sch94i, Sch95n]. Eva [Woe97a]. Evaluating [DDJ99f, Mat96]. Even [The98b, Sha98d, Bak97b, Len97]. Event [Bar93b, Ber96f, For95, Gig97, SG97, Uhl99, Ste97g, WJ92]. Event-Based [Uhl99]. Event-Driven [For95, Bar93b]. every [Boy97, RC97b]. Everything [Nel90]. Evolution [Con94b, Flo91b, PS97b]. examine [MK97, SW96b]. examines [GHJV96, Kis97, Nel97, vdWS97, vdW97b, Wal97, Wil97g, Yam96]. Examining [Ang99, Ano96c, Bat96, Bug98, Cam96, Cop94, Cro94b, Dav99, Ent90a, Fai96, Fir93b, Flo92, Gar96, Haa95, Hem96a, Hij96, Hla96a, Kan94, Lad90a, Law97, Low98b, MK97, Nee99, One94a, Par95b, Pie93a, Rat93, Ric91, RC96a, Rod95a, Rod94b, RC95, RC96b, RC97a, Sch91a, Sch96c, Sch95b, Sch98c, Sch96i, Sch90e, Sch93c, dS95, Soo98b, Tha97, Vil96, vdW97a, Wil96h, Wil98b, Yam93, Yam94, da 95]. Excellence [Eri95a, Eri96m, Eri97a, Eri98a]. Exception [Big98, Hla94, Hla95c, Pit99, Ste99i, Sal92]. Exception-Handling [Big98]. exceptionally [Boy97]. exchanges [Mir92]. Exchanging [Hav98a]. Exclusion [For98a, Mar95a]. Executable [Sho96, Rod94b]. Executing [AW99, Wei93]. Executive [Jus92]. existential [Flo90a]. expand [Nel97]. expanded [Gin91]. expect [Ste97f]. experience [Yam96]. Expert [Kin91b, Wei96b, Ano96a, Wei96a]. Explorer [Tuc99, Wil96d]. Exploring [Cog93]. Expo [Swa95a]. Exponential [Gun98, Sta98a]. exponentiation [Smi94]. Export [DDJ98f]. Express [Hij96]. Expression [HCKS99]. Expressions [Ber99, BF97, BCP^{+}99, KP99, dSW98]. Extended [Dru94, Gri97b, Tha97, Gut94, Huf93c, Tha97]. Extender [Wil90d, Wil90c, Wil92a]. Extenders [Nor99]. Extending [Bak93, Bra93, But99, Flo95f, Jol93a, HW96, Hus94, Jol93b, Mer94, Mis90, Mon96, Ngu96, Nic92, Rob97, Sch98a, Sue95, Sul94b, Var93a, Var93c, Win99]. extensibility [Kla99]. Extensible [Boy99, dFIC96]. Extension [Aye90c, Cod94a, GH95a, Sad99, Fra91, Swa90a]. Extension/Audio [Cod94a]. Extensions [Bea98, Col96d, Col96e, DR97, Ell90, Eri90m, Lic92, MM98b, Wil97c, Bau92b, Jer92, Ste92p, Col98a]. extents [DGR90]. extract [NN97]. Extrema [Cou96].

F00F [Col98c]. Fabrication [DDJ99d]. Face [DDJ98d, Sch91b]. Facemail [Swa98r]. Facilities [Moo95]. Facility [Ger93]. Fact [Col98b]. Factory [Sha97]. Fair [Eri97d, Sha99e]. Faking [New93b]. family [Kle91b]. Famous [Eri98i]. Fanning [Val91a]. Fantasy [Col98b]. Far [Wil99c]. Farey [Ben96]. Fast [MSE^{+}99, BH96, Cam90, Cla97, DDJ98f, Eri98d, GS97, Hei96, Kie99b, Nel96a, NK98, Ric98, Tho91b, Web96, Boy97, Cod92, DKR97, Mai91, PS97b, Ste97g]. FastCGI [DR97]. Faster [Dob95, MFCM96, Pic91, SC97, Gul95]. Father [DDJ99d, Spe97]. Faults [Sch90h, Sch90i]. faxes [Pic91]. FDIV [Coe95]. Fear [Eri98e]. features [Ham95b, McB97]. Fedorov [Sch93l]. Feedback [Hen97b]. ferreting [vdW97b]. Festival [Swa99t]. FFTs [Dob95]. Fibonacci [Boy97]. Ficl [Sad99, SWF^{+}99]. Fiction [Col98b]. Field [BEK99]. Field-Tested [BEK99]. fields [Gil92].

Fighting [DDJ98h]. File [AW99, BT93, But93, Cha95a, Cha95b, Cop98, Dav93a, Dav93b, FR93, Flo94a, Gis93, GM97, Huf93b, Huf93a, Ken93, Kie98, Lus94, Mau93, Nel92, NN97, Rea91, RC95, Sae95a, Sch93k, Sch93d, Sch93l, Ste98g, SWF^{+}99, Tor95, Wee90, Chu91, Flo93b, Gar91, Len97, Lor96, Mor91a, Mur96, Rod94b]. File-streaming [Sae95a]. Filemon [RC97a]. Files [BT93, Cou94a, Gra94, Hla95a, Hla95b, Lia95, Pop96, Row94a, Sch97a, You98, BDE93, Cou94b, NN97, Nel97, Rei93b, Rei93c, Rod91]. Filesystem [JJ92f, Lom98, RC97a, JJ91t]. Filesystems [Len97]. Filling [Gut99]. Filter [CJ94]. Filtering [Byl99, Sta94a]. Filters [NRG^{+}99, Sta94b, Che92, Pro92]. Fin [Swa99f]. final [JJ92d]. find [Kis97]. Finding [Kim92, Man99a, Nel93a, Nel93b, PSS98a, Val92d]. Fingerprint [DDJ98i]. Finite [Car97c, Her98, Hof97, PSS^{+}98b, Ver93]. finite-element [Ver93]. Finite-state [Hof97]. Fire [Swa91a]. Fish [DDJ99f]. Fissures [DDJ98f]. fit [Soo98b]. Five [JW92, Nic91a, WJ92]. Fix [MSE^{+}99]. Flaks [Swa97h]. Flames [Swa90m, Swa90n, Swa90o, Swa90p, Swa90q, Swa90r, Swa90s, Swa90t, Swa90u, Swa90v, Swa90w, Swa90x, Swa91n, Swa91o, Swa91p, Swa91q, Swa91r, Swa91s, Swa91t, Swa91u, Swa91v, Swa91w, Swa91x, Swa91y, Swa92m, Swa92n, Swa92o, Swa92p, Swa92q, Swa92r, Swa92s, Swa92t, Swa92u, Swa92v, Swa92w, Swa92x, Swa92y, Swa93n, Swa93o, Swa93p, Swa93q, Swa93r, Swa93s, Swa93t, Swa93u, Swa93v, Swa93w, Swa93x, Swa94n, Swa94o, Swa94p, Swa94q, Swa94r, Swa94s, Swa94t, Swa94u, Swa94v, Swa94x, Swa94w, Swa94y, Swa95o, Swa95q, Swa95p, Swa95r, Swa95s, Swa95t, Swa95u, Swa95v, Swa95w, Swa95x, Swa95y, Swa95z, Swa96x, Swa96m, Swa96n, Swa96o, Swa96p, Swa96v, Swa96w, Swa97t, Swa97u, Swa97v, Swa97w, Swa97m, Swa97n, Swa97o, Swa97p, Swa97q, Swa97r, Swa97s, Swa98q, Swa98r, Swa98p, Swa98s, Swa98n, Swa98t, Swa98u, Swa98v, Swa98w, Swa98o, Swa98m]. Flames [Swa99k, Swa99q, Swa99l, Swa99r, Swa99s, Swa99m, Swa99n, Swa99o, Swa99t, Swa99p, Swa99u, Val91a, Swa96q, Swa96r, Swa96s, Swa96t, Swa96u]. Flanagan [Nic96a]. Flash [CS95, Dip95, Gis93, Tor95]. Flashes [Swa98d]. Flashsort1 [Neu98]. Flat [Ste91m, Ste91n, Ste91o, Ste91p, Ste91q, Ste91r, Ste91s, Ste91t, Ste91u, Ste92r, Ste92s, Ste92t, Ste92u, Ste92v, Ste92w, Ste92x, Ste92y, Ste92z, Ste92-27, Ste92-28, Ste92-29, Ste92-30, Ste92-31, Ste93v, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u, Ste98f, Wri94, Ste95l]. Flats [Sha99f]. Flattery [Swa97l]. Fletcher [Kod92]. flexibility [Hoe98]. Flexible [Ise95, Mel98, Mye97, Wea97]. Flic [Ken93]. Flicker [Mar95b]. Flicker-Free [Mar95b]. Flip [Swa99r]. Floating [Mot93, Wil93a]. Floating-Point [Mot93, Wil93a]. Flow [Ber98f, Kau96b]. Fluid [Swa95n]. Flushed [Bar99d]. Fly [Shm94]. focuses [Car97c]. focusing [Swa97d]. Following [Eri98f]. Font [WB95]. Foo [Swa98r, Swa98s]. Fool [Eri99e]. Footprint [Nil98]. Form [Swa97l, Xia95]. Format [BAC^{+}99, Cha95a, Cha95b, Cro95a, Dav93a, Dav93b, FR93, JJ92f, Ken93, Lus94, Mau93, NN97, Per99, Sch93k, Sch93d, Sch93l, Ste98g, Mur96, For98b]. Formats [Cop98, Flo94a, GM97, Ber96a, Rod94b]. Formed [BEK99]. Former [Bro95a]. Forms [Boy99, Dav95c]. Formulate [Amb99]. Forté [Hij96]. Forth [BCP^{+}99, Dri94, Ste98a, Woe91a, Woe91b, Yue96b]. FORTRAN [Bra90, Ana93, Bak95, Bra91, Ham95b, Jus92, Kef92, Vai93, Vel97]. Fortrans [Ham93]. FORTREX [Jus92]. Forum [Kae94]. Foundation [BEK99]. four [Soo98b]. FoxPro [Bra93, Bro94a, Gol96]. Fractal [Esp97, MFCM96, Esp97]. fractal-imaging [Esp97]. Fractals [Oli91].

Fraction [Ple95, Kod92]. fractional [Mor91b]. Fractions [Ben96]. fragmentation [Har93b]. Fragments [Swa98k]. Frame [BH95, Sta95a, Tes96a]. Frame-Relay [Sta95a]. Frames [Abr95c, Tes96a]. Framework [Boo94, Car97b, Hou98, Yam95, Fir93b]. Frameworks [Val92e, Val95a]. Francisco [Bar99c]. Free [DDJ98e, Eri97b, Hug95, Keh94, Mar95b]. Freenets [Hug95]. Fresco [Lin94d]. Friends [DDJ98f]. FS [Rea91]. FSF [DDJ99f]. FTL [CS95]. FTP [Hig96]. Fugit [Har96c]. Full [Kie99c, TMP^{+}99]. Full-Text [Kie99c, TMP^{+}99]. Function [BDP97, DC98, Kis96, McC97]. Functional [Sch99a]. functionality [Sto97]. Functions [Jen97, Ros93, Sch91c]. Funny [Eri97e]. Fusion [Hig97]. Future [Dun91a, Eri97h, Gur95, Sta98b, Tre95, ML97]. Fuzzy [Rot98, TFS94, Vig93, Vio93].

G [Sch93l]. Galaxy [Dol95]. Game [Dor95, Gou96, JS94, Sha98i, Ste95n, Wal95a, vdW96]. Games [Cro94c, Fin95b, Hol96, How96, Pla93, Rat95]. Gaming [Sem96]. Gamma [Dor99]. gap [Gri97a, Mir92]. Garbage [CR92, PB98a, Spe97, Jac90]. Gas [Bar99d]. Gateway [DDJ99c]. Gathering [Swa98p]. Gauze [Swa99s]. Gbit [NK98]. Gbit/sec [NK98]. GDI [Shi94]. GeezerWare [The98b]. GEF [Big98]. Gems [Abr92a]. General [Ber96f, Big98, Sch90h, Sch90i, WC99, Gar91, Mor91d]. General-Purpose [Ber96f, Gar91]. Generating [Ait96a, Cal95b, Krt94, Rei94a]. Generation [Iye95, Jac90, Lin94d, Sta94b, Val94f, Tha97]. Generator [Bal90, Xia95, Mai91, Sch92b, Vog91]. Generic [Dav91, EJ95, Ive90, New94b, Nor93, Zig95, Aus97]. Genetic [Cel93, Mor91e, Pri94, Spi93]. Geographic [CS94, Var93d]. Geometry [Kno95]. George [SW96b]. Geospatial [Gut99]. Get [DDJ98a]. Gets [Bar99a, The98b, Val95e, Mye98]. Getting [Bri90, Fly97, Let91, Rei93a, Sha98d, Swa90a, van98b]. Ghost [Swa99s]. Giant [Val94f]. GIF [Lia95]. Gigabit [LS99]. gigabytes [Wil90a]. Gives [Bar99d]. Giving [Eri97f]. Glassner [Abr92a]. Glenn [Woe99]. Global [Jon98, Li94a, Oka94]. Globalization [Van94]. Globally [Mül94, Sch94m]. Globbing [HHW^{+}99]. GNAT [Smy97]. GNU [Smy97]. Go [DDJ98d, Gul95, The98b]. Goal [Coo96]. Goal-directed [Coo96]. Goes [Bar99c, DDJ98e, Eri97g, ML97]. Going [DDJ98c]. good [SC97]. Gopher [GK94]. Gosselink [GS94]. GOST [Sch95l]. Gotchas [Mis96]. GP [Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s]. GPIB [Mor91c, Mor91d]. Graph [All91, PH95b, PH95a]. Graphic [Wit95a, Cro95a]. Graphical [Bek97a, Shi99, WV91a, WV91b]. Graphics [Abr91b, Abr91c, Abr91d, Abr91e, Abr91f, Abr91g, Abr91h, Abr91i, Abr91j, Abr91k, Abr91l, Abr92a, Abr92b, Abr92c, Abr92d, Abr92e, Abr92f, Abr92g, Abr92h, Abr92i, Abr92j, Abr92k, Abr93a, Abr93b, Car97b, Coh92, Hed98, Jan99a, Lus97, NRG^{+}99, Pro92, Ste99j, Str90, SC97]. Graphing [Ber98f]. Graphs [FSW99, Man94a]. Gray [Ste92a]. Great [DDJ99b, Ste98i, Val95h, Jam92]. Greatness [Ste97j]. Greg [Wil97g]. Gregorian [Mey93]. Grep [Pit98b]. Grepping [HHW^{+}99]. Gresham [Har96a]. Grip [Sch94n]. Group [DDJ99d, Sip96d, Dib96]. Grow [DDJ98d]. Guessing [Pla93]. GUI [Ana93, Kep94, Mey92, Nic91a, Sch95a, Tet94, Val92f, Wal97, Wil98c]. Guided [CKS98]. Guidelines [Gro97a]. GUIS [Sch96a, Bra91, Rod94a]. Gutenberg [DDJ99f]. Guy [Swa98g].

H [Och94b]. H1 [DDJ99b]. H1-B [DDJ99b]. Hacking [WBB^{+}97]. Hah [QTC^{+}99]. Hairy [Ste96l]. HAL [Wil97j]. Hamilton [Ric91]. hand [Sto97]. Handbook [NFP96]. Handle [Bri90, Pit99, Bet91a, Len97]. handle-based [Bet91a]. Handler [New94b, Sal92]. Handlers [Hla94, Hla95c]. Handling [Bet94d, Big98, Mak90, Pit99, Ste99i, Wei95]. Handprinting [Val92b, Avi92, Val93a]. Happen [Val94d]. Hard [Har95e, MGL^{+}98, SDB^{+}99]. Hardware [Fri90a, Fri90b, Col97c, QG97, RK92, Tom96]. Harmed [Wil98d]. harness [Wal97]. Harvest [Dan96]. Hash [BDP97, Jen97, Sta94c, Sch91c]. Hashing [Bin96, Pes96, Sta99]. Hate [HHW^{+}99]. HDF [For98b]. Head [DDJ99a]. Headed [Har94b]. Heap [Oua93]. Heaps [Boy97]. Heat [Val94a]. heavyweight [Woe91b]. Help [Bug98, Eri98g, Hla95b, Lor96, PKA^{+}98, Ste94p, NV90, Ken94, Ste97f]. Herb [Ste99d]. Here [Boy97, Val94b, Val95b, Sto97]. Herman [Ste96m]. Heuristic [Mar95a]. Heuristics [Rie96]. Hewlett [Aus97]. Hidden [Abr96c, Tes96a]. Hidden-Frame [Tes96a]. Hidden-Surface [Abr96c]. Hierarchical [Cra99, Höw99, For98b]. Hierarchies [Gro97a]. Hierarchy [Col99]. High [Eri94a, Har95d, Her98, Hof97, Jol92b, JJ94, MPV95, Rod92, Shi99, Car97c, Nic93b, Smy97]. high-level [Car97c]. High-Performance [MPV95, Shi99, Nic93b]. High-Precision [Her98]. high-quality [Smy97]. High-resolution [Rod92]. High-Speed [JJ94, Hof97, Jol92b]. High-Tech [Eri94a, Har95d]. Higher [The98b]. Highlight [Bar99c]. Hilbert [HHW^{+}99]. Hill [Swa98c]. HIPPI [Nic93b, JJ94]. History [BAC^{+}99, Kim99i, Kin93]. Hit [Tre94]. Hits [Lyn95]. HMAC [Sta99]. Hoff [McC96b]. Hogs [DDJ99f]. Home [Lyn95]. Homegrown [Wil90b]. Honors [DDJ99f]. Hook [One92]. Hooked [McC96b]. Hooking [RC97b, RC96a, RC96b]. Hoping [Kis97]. Horizontally [Fai93]. horse [Ste97e]. Host [Gad96, Car97c, Goa99]. Hot [Swa97v, Swa98d, Hed98]. HotJava [Rod95c]. House [DDJ99d]. Housecleaning [Swa98s]. HP [The98b, Yam96]. HP-UX [Yam96]. HTML [Bug98, Con94b, Dav95c, Dav95b, Fly96, Hig96, Tou99, McA94, Swa96l]. HTMLParser [Tuc99]. HTTP [Mon96]. HTTPsync [Cav99]. Huffman [PB98b]. Hybernation [Swa94a]. Hype [Bek97b]. hyper [GGL^{+}95]. hyper-time [GGL^{+}95]. HyperCard [Ell93a, Ell93b]. Hypercubes [HO92, HO93]. Hypertext [Ges90, Joh90, Kin90b, NV90].

I-nodes [Len97]. I/O [BT93, FY96, HS94, Huf93b, Huf93a, Kno94, Rob96, GS97]. I2C [Kah92]. i860 [Coh92]. Ian [Och94c]. IBM [SRA^{+}94, DDJ99g]. IC [DDJ98i]. ICE [Col97b]. ICs [The98b]. ID3 [Col96a, Mon97]. IDEA [Sch93b]. ideal [vdW97b]. Identification [Ban98b]. Identifying [Rid95]. IDL [FHW97]. IDLDOC [FHW97]. Idle [Nic92]. IEEE [Ale99b, Ale99a]. if [Ste97e]. II [And90b, AB97, Bak97a, Col97a, Dav93b, DY99b, Eri97b, Fri90b, HO93, JJ91c, JJ92c, Kry92b, PF90b, Sch90i, Sch93d, Tro91b, Wil90d, Wil98f, Wu97]. Illiacs [Swa97h]. Illuminates [Wil94d]. illustrate [QG97]. I’m [Eri98l]. IMA [Kie97a]. Image [MSE^{+}99, BA96, Cla97, Cro95b, Das93, Kli90, Lin94a, Lin95a, Lin95b, Pat97, Pig99, Sul93, Wal95b, Bri91]. Image-processing [Pat97]. Images [Dav95b, MM99, McC96a, Ste92a, Wit95a, Ano96i, Jan91, Ram97b]. Imaging [DR93, Ram97b, Esp97]. Imake [Hus94]. Implement [Ber96c, Ses98, Wil99a, Boy97, Swa97y]. Implementation [App97, Pap95, Hop91, Kis96, RC97b, Ste92r, Ste92s, Ste92t, Ste92u, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u]. implementations [Wil98a]. implemented [Boy97, Car97c]. Implementing [BW96, Bev96, Blu95, Day92, Dwy95, Fin95b, Hel98, Jar90, JM92, Kar90, Kie99b, LP90, Mar95b, Mey99, Mor91c, Och94c, PSS98a, Pra95, Sha97, Tes96a, Wel95, Law97]. Implicit [Lus97]. Import [Oka94, Pro92]. important [Woe97b, Woe97c]. Imposing [All91]. Improve [Ros97]. improved [Chu92a]. improvements [QTC^{+}99]. Improving [DRG90, Kie96b, Krt96, Lip99a, Mir92]. In-Circuit [Col97c, Bar92]. Includes [Kru96]. Incremental [Cav99, HC99, Nac97]. Incrementally [Man99a]. Independence [NRG^{+}99]. Independent [BC99, Cal96, Kis96]. Index [Cus90a]. Indexed [Krt95]. Indexing [Cel93, Ste97h, Sul93]. Inferno [SWF^{+}99]. Infinite [Ham95a]. Information [CS94, DY99a, DY99b, Fel93, Gol90, War91, RC97b, Osw93]. Infrared [DDJ98b, Mus99]. Infrastructure [Dan95]. Inheritance [Car92, Kau96a, Ler96, Sha99g]. initial [JJ91t, JJ91v]. initialization [BDE93]. Inner [BF97, SvA^{+}98, RC97a]. ins [CH97b]. Inslaw [Eri97b]. Installing [Wil95a]. InstallShield [Hla96a]. Instance [Mül94, Sch94m]. Instant [Sch90e]. Instant-C [Sch90e]. Instantiating [Wil96b]. Instantiation [Czu90]. instruction [Woe92b, Zha97]. instrumentation [Mor91d]. instruments [And94a]. Insulator [DDJ98g]. Insurance [Eri97b]. Integer [Hei96, Duv91, Gra91]. Integers [Gri97b, Ham95a]. Integrated [The98b, Ste96n, Ste96o, Ste96p, Ste96q, Ste96r]. Integrating [Var93d]. Integration [PJR^{+}98]. Integrator [Con95]. integrity [Hel98]. Intel [Jol95c, BL95, Col97d, Col98a, Zha97]. Intellectual [Mar94b, RR98, Sma94]. Intelligence [Swa94a, Swa98q]. Intelligent [Gam96a, Kie94b, MC98]. Intelligently [Hir94a, Hir94b]. Inter [BC98, Mec97]. Inter-Applet [BC98, Mec97]. interaction [Pan99]. Interactive [HM95, Koc97, Rat95, Tes96b]. Interapplication [Sta95b]. interchange [Mur96]. Interest [Ano98e, Bar99b, BL99a, BL99b, Ber93a, Ber93b, Ber93c, Ber94a, Ber94b, Ber94c, Ber94d, Ber94e, Ber94f, Ber94g, Ber94h, Ber94i, Ber94k, Ber94j, Ber95a, Ber95c, Ber95b, Ber95d, Ber95e, Ber95f, Ber95g, Ber95h, Ber95i, Ber95j, Ber95k, Ber95l, Ber96g, Ber96h, Ber96i, Ber96j, Ber96l, Cus90b, Cus90c, Cus90d, Cus90e, Cus90f, Cus90g, Cus90h, Cus90i, Cus90j, Cus90k, Cus90l, Cus90m, Cus91a, Cus91b, Cus91c, Cus91d, Cus91e, Kim96a, Kim96b, Kim96c, Kim97a, Kim97b, Kim97c, Kim97d, Kim97e, Kim97f, Kim97g, Kim97h, Kim97i, Kim97j, Kim97k, Kim98b, Kim98c, Kim98d, Kim98e, Kim98f, Kim98g, Kim98h, Kim98i, Kim98j, Kim98k, Kim99h, Kim99a, Kim99b, Kim99c, Kim99d, Kim99e, Kim99f, Kim99g, Zem91a, Zem91b, Zem91c, Zem91d, Zem91e, Zem91f, Zem91g, Zem91h, Zem91i, Zem91j, Zem91k, Zem91l, Zem92a, Zem92b, Zem92c, Zem92d, Zem92e, Zem92f, Zem92g, Zem92h, Zem92i, Zem92j, Zem92k, Zem92l, Zem93a, Zem93b, Zem93c, Zem93d, Ber96k, Ber96m, Ber96n, Ber96o]. interest [Ber96p]. Interface [Ber96c, Ber98b, Bro94b, Ces96, Cod94a, Ent90a, Gad96, Kah92, Oka94, Ram97a, Saw94, She99, SW97c, Sip96d, Tha97, Wil99b, WSM95b, Jam92, KR91, Man92, Ste92-33, Tho92, Wil97e, WSM95a]. Interfaces [Bak97b, Bra92, Can90, SW97b, Wil98a, Ban98a, Ram97a, Tha97]. Interfacing [And94a, Bro96c, Bur95b]. Interim [Ste99c]. Internals [SW97a]. International [Ber97l, Flo94b, Mye98, PJR^{+}98]. Internationalization [Jon98]. Internet [Val94b, AN98, Bet98, Bri99, Bur95a, Cro94c, Eri98l, GW97, Hig97, Hus99, Joh97a, JJ95a, Kie95, Kim99i, Li94a, MMV94, Mar97, McG98a, Nee99, NK98, OK98, Pom96, Qui97, Sar97, Tay98, TSR98, Tuc99, Val95b, Val95d, Wil97c, deC99, van95]. Internetworking [Sta93a]. Interoperability [Dow93]. Interoperable [Bet94c]. Interpreter [Dri94, Sad99, Ste94y, Ste94s, Ste94t, Ste94u, Ste94v, Ste94w, Ste94x, RB97]. Interprocess [Rod94c, Bla96]. Interrupt [DDJ99b, Oes97, Rob98a, Web96, Kno93a]. Interrupts [Kno93b, Kno93a, One92]. Interview [Li94b, Ste92-32, Woe92a, Woe96b]. InterViews [Lin94d]. Intrinsics [Lai91]. introduced [Aus97]. Introducing [TSR98, Val94c]. Introduction [Sch93k, Sch93d, Sch93l, Sch94m, Sch94w, Sch94v]. intuitive [Ste97f]. Invented [DDJ98d]. Invocation [Ber97j]. Invoke [Ber97c]. involved [Kis96]. IP [Boz99, Bur95b, NK98, Val94f, WV96]. IPC [Pro99]. IPv6 [Sta98b]. IPX [Jam92, Ser93]. Ironic [Swa97e]. IRQs [Rid95]. Is-A [Gro97a]. ISAPI [Cro98a, Wil97c]. ISDN [Dav98]. Isn’t [Swa97e]. ISO [JJ92f]. ISO-9660 [JJ92f]. isolation [McC97]. Isometric [Gou96]. Issue [Eri97b]. issues [Wal97]. Itch [Swa99j]. item [Boy97]. Iterated [Esp97]. iterators [Zig96b]. Itineraries [LRT99]. Itself [The98b]. IUknown [SW97c].

J [Wil98b]. J. [Cha95d]. J/Direct [Wil98b]. Jabs [Swa97g]. JAVA [CH97b, Ait96a, Ait96b, AW99, Ang99, App97, BAC^{+}99, Bal99, BC99, Ban98a, Bar99e, BW98, Bek97b, BA96, Ber96e, Ber96d, Ber96b, Ber96a, Ber96f, Ber96c, Ber97i, Ber97b, Ber97j, Ber97g, Ber97a, Ber97f, Ber97l, Ber97e, Ber97d, Ber97k, Ber97h, Ber97c, Ber98a, Ber98d, Ber98b, Ber98c, Ber99, Bla96, Bou97, Boy98b, BCP^{+}99, Bri99, CL97, Chr99, CM97, Coo98b, CH97a, Cra97, Cri98, DDJ98f, The98b, DDJ98a, DDJ99f, DDJ99d, Del97, Eas99, For98a, Gat97, Gig97, GS98, Gri97b, Gri99, HHW^{+}99, Hav98a, Hem96a, Hem96b, Hen97b, HL99, Jan99a, Jon98, Kay96, Kie99d, Lip99b, Loe99, MGL^{+}98, MM99, Mec97, Met96, Moo98, Mor98, NRG^{+}99, NN97, Nic96a, NFP96, Nil98, Nil97, PSS98a, PH99, PB98a, Pit98a, Pit98b, Pit99, PJR^{+}98, Pur98, Ran99a, Ran99b, Rem97, Rou98, Sch99b, Ses98]. Java [Shi99, Sho96, Soo98a, Sta98c, Ste96l, Ste99e, Sul97, Swa97g, Swa97f, Swa97i, Swa98e, Tei98, Tre98, Tre99a, Tre99b, TMP^{+}99, TMH99, Tym96, Val95e, Ver96, Ver97, Wal97, Whi96, WL97, Wil96i, Wil97e, Wil98b, Wil99a, Wil99b, WBB^{+}97, You97, You98, van95, van98b, McC96b, Nic96b]. JavaBean [Ber97d, Kin99]. javadoc [Ait96a]. JavaScript [Ste97i, Tes96a, Tes96b, Tic97]. JavaServer [Tre99b]. Jbed [TMH99]. JDBC [Soo98b]. JDK [HL99]. Jeff [Flo94h]. Jerry [Swa99e]. Jfactory [Met96]. JFC [Joh99]. Jini [Swa98j, Won99]. Jive [BAC^{+}99, Ste96l, Ste99e]. JNI [Coo98b]. job [HQ98]. Jobs [Swa99c]. John [Boy97, Swa97y]. Joined [Swa99d]. Joke [Wil99c]. Jonathan [Wil97g]. Jose [Har95d]. Journal [Eri95a, Eri97a, Eri98a, Eri96m]. Journaling [RC96c]. JPEG [Lin95a, Lin95b]. JPEG-Like [Lin95b, Lin95a]. Jperl [Bal99]. JPython [Ang99]. JScript [But99]. Judge [DDJ99b]. Julian [Mey93]. Jumps [Bar99c]. Just [DDJ99b, Swa97u, WC99, Swa96y, Tre98, Nic96b].

K-tree [Bat94]. K. [Sch94m, Sch94v]. Kahan [Woe97c]. Kaleida [Val94h].

Kannel [GPVV98]. Keep [Bar99a]. Keeping [BEK99, Wil98a]. Kermit [And90a, And90b, Kie96b]. Kernel [JJ91p, Kot99, Wel95, JJ91s, JJ91b, JJ91c, JJ91a, JJ92m, Len97]. Kertesy [SG97]. Key [DDJ98i, Sta99, DKR97, Sch92d, Smi93b, Tho91a, You98]. Keyboard [BAC^{+}99]. Keys [Cal95b]. Kicking [Ste97i]. Kids [Swa99e]. Kill [Bar99a]. Killer [DDJ99f, Qui97]. killers [Qui97]. Kind [BCP^{+}99, Swa99k, Wil99e]. Kit [Aye90c, Bet94a, Rod95b, Sch95b, WBB^{+}97, Esp97]. KJPrice [PJR^{+}98]. knotty [Nic91a]. Know [DDJ98b, Rei93a, Swa99m, Mye91a]. Knuth [Woe96b]. Kudos [DDJ98e, DDJ98g, DDJ99e, SvA^{+}98].

LabObjX [Wil96h]. laboratory [And94a]. Lambda [Jaf93]. LAN [Len97, She93]. Landauer [ML97]. Landscape [Eri91a, Flo91b, Val91a]. Lane [Bar99d, DDJ98f]. Language [Abr90a, Amb99, Ber98f, Bet94b, BH94, Boy99, dFIC96, Fle93, Gut92, Har94a, How93, Hyd90, Jer93, Jer94, JJ91q, Man93, McG98c, NRG^{+}99, NFP96, Ous94, Pat90a, Pla93, PJR^{+}98, Rie99, Sad99, Sha94, Sum91a, Sum91b, Sum91c, Sum91d, Sum92a, Sum92b, Sum92c, Sum92d, Sum92e, Sum92f, Sum92g, Sum92h, Tem94, Val93c, Var93a, Var93c, WW99, Wil97g, WS90, Bak97b, Bet91c, Bet93a, Flo90b, Fra91, Lea93, Mae92, McS92, Nor91, Omo93, Rec99, Ski91, Wil92c]. Languages [Flo93a, Flo93c, KP99, Kin93, Mel98, Sha99a, Val91b]. LANS [Nic93b]. Lap [Wil92a]. Large [Bau92a, Lak97, Pla92, Wil97j, Tho91b]. Large-Scale [Lak97, Wil97j]. Larry [DDJ99f, Kim98a]. Laser [Sha99h, Low92]. LaserJet [CJ94]. Last [Ste98c, Swa99m]. Late [Cre98, TMP^{+}99]. Later [Sch95k]. Latin [Val94d]. launches [McG93]. Law [Eri99g]. Layer [Sar97]. Layered [RC95]. Layout [Ber96b]. LC [NK98]. LC-tries [NK98]. LDAP [Hus99, Nee99, Sch98c]. leading [SC97]. Leap [Bar99a, Val94f]. Learn [Wil96e]. Learned [Eri98h]. Learning [Cro98a, DL99, SNC97]. Legacy [Swa98n, Wil97j, Bak97b]. Legal [Bro95b]. Lego [Eri99f]. LegOS [Nog99]. Leinecker [Hel93, Moo93]. Leo [Swa98f]. Leopurd [Mir96]. Less [Shi92]. Lessons [Eri98h, Wil96e]. Let [DDJ98e, DDJ98f, Eri98i, Hol96, Swa99f]. Letter [AB97, Bak97a, Bek97b, Cra97, Fly96, Hen97b, Wil96i, Bau90b, Can95, Dri94, Fer95, Ran94, Rus93]. Letters [Ano95b, Ano95c, Ano95a, Ano95d, Ano96d, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano98c, Ano98d, MSE^{+}99, BAC^{+}99, BCP^{+}99, Coo98b, HHW^{+}99, LHB^{+}98, MGL^{+}98, MBK^{+}98, NRG^{+}99, PKA^{+}98, PSS^{+}98b, PJR^{+}98, QTC^{+}99, SvA^{+}98, SDB^{+}99, SWF^{+}99, TKG^{+}99, Tei98, TMP^{+}99, WEM^{+}99, WL97, Ano96e, Ano96f]. Level [CP99, Sip96d, Car97c]. Levels [JW92, WJ92]. Lex [Gor96, Sha99a]. Liability [Bro96b]. Liana [Kru94b, Val93c]. Liars [Sha98c]. Libraries [Bak97b, JLC98, Syc90, Val92e, Ham95b]. Library [Big98, Clo94, Dur97, Dut95, Gar96, Gib94, Hed98, Hla96b, Jol96, Koo94b, Mus93, Nel97, Nor93, Rog95b, SW97f, dSW98, Ste95n, Ste98j, Vil96, Wei93, Wil93c, Wri99, Chu92a, Esp97, Hyd92, McS92, Mey92, Ozi91, Sie91, Yam96, Aus97, Ent90a, Kef95, Vil95]. License [Ste94a]. Licensing [Bar99d]. Life [DDJ98f, Ste98f]. Light [Bar99d]. Lightweight [Fin95a, PH99]. like [Len97, Lin95a, Lin95b]. limitation [Len97]. Linden [Nic96b]. Line [DRG90, DRG96, Jan99b, TKG^{+}99, Whi96, Kis97, Kyl91]. Line-Segment [DRG96]. Linear [BH95, Sch96g, VP96]. Link [Syc90, Len97]. Linked [Boy98b, Hes93]. Linking [Cre98, HC99, KR91, Lin94c, McG98c, Sha93b]. links [Lor96]. Linus [Flo94d, Li94b]. Linux [DDJ98h, DDJ98a, Eri97b, GE98, Len97, Pro99, Swa98v, Tay99, Wel95]. Lippman [vdWS97]. LISP [Swa96y, Chu92a, Ozi91, Swa90a]. Lisp-style [Chu92a, Ozi91]. List [Pro94, Swa98b]. List-Box [Pro94]. Listboxes [Fai93]. Lists [Boy98b, Hes93, Sch94l]. Literate [Hen97a]. Little [Sha99a, Swa98g, Swa99g, Val91b]. Lives [Abr90a, DDJ98d, Mau93]. Lizards [Swa98f]. Loadable [Wel95]. loader [Mor91a]. Loading [Kyl91]. Loathing [Eri98e]. Locales [Mye98]. Locally [Mül94, Sch94m]. Location [Nel90]. Lofti [Woe94b]. Logarithm [DeV96]. logging [Cra90]. Logic [Cra99, Rot98, Tre95, Vig93, Vio93, TFS94]. Logins [BEK99]. Long [Swa99i]. Look [Bet95c, Eri94p, ML97]. Looking [Dun91a, Swa96l]. looks [Swa97d, vdWS97]. Loop [Let96]. Loops [BF97, SvA^{+}98]. Losers [Val94i]. Lotfi [Woe94a]. Lotus [Ste92-33]. Love [DDJ99g]. Low [CP99, Smy97]. low-cost [Smy97]. Low-Level [CP99]. LUA [dFIC96]. LUC [Smi93b]. LUCA [Gri97a]. Lunch [DDJ98f]. Lurking [Swa99n]. LZW [Reg90].

M [Och94b, Sch93l, Shu96]. Mach [Eri99b]. MacHack [Red95a]. Machine [DL99, Sch94w, Sch94v, Zyt94a, Zyt94b]. Machines [Car97c, Hof97, Len97]. Macintosh [Cop96, Esp97, Gas90]. Macro [Dut95, McS92, Rod91]. Macros [Rog95a, Ski91, Gil95a, Gil95b]. Macs [Swa97h]. MacWorld [Swa95a]. Made [Har95c]. Magic [Swa99h]. Mail [Con94a, Duf95, Kie95, Lu94, Sch91a, Sch94h, Ste97l, Ber96d, Flo94c, AH97, Obl90, Ohl90]. Maintaining [Hel98, SC97]. MainWin [Fra99]. Make [DDJ99f, Val94g, Len97]. Maker [Wil96f]. makes [Esp97, Kie96b, Kla99, Tew97]. Making [AJ90, Aye91, DDJ98f, Dah93, Hag91, Sar97, Wes94, Wil98b, Gul95]. manage [Gin91]. Management [App94, Col97e, Col97d, Col97f, Kno95, Sta93b, Sul94a, Sul97, Bet91a, BDE93, Bra93, Dou90, Har91, Sie91, Ste93b]. Manager [Ber96b, Ber97f, Mül94, Sae95b, Sch94m, She93, Yue94, Dib96, Flo94c]. Managing [Car96a, Gam96b, PF90a, PF90b, Ver96]. manifesto [Kap91]. Manipulators [Cha93]. Mantra [Eri98c]. Many [DDJ99a]. Mapcraft [Sha98e]. MAPI [Sta95b, Tha97]. MapMan [Hla95a]. mapped [BT93, Huf93b, Huf93a]. Mapping [Spi94, Var93d, Bet93b]. Mark [Nel97, RC97a, RC97b, Wil97g]. Market [Val94d]. Markets [FSW98]. Markov [Yu96]. Masking [Kie99a]. MASM [Sch91b]. matching [CL97]. Mates [Sha98f]. Math [AB97, Bak97a, Col97a, DDJ98e, Sto94, Wil93a]. Mathematica [KS97, Mae92]. mathematics [KS97]. MATLAB [Wea97, WW99]. Matrices [Iye95, Duv91]. Matrix [Sto94]. matter [Ste97f]. Maximal [Van98c]. Maximizing [Bun94]. Mazes [Nay93]. MB [Col96e]. MC68882 [McG94a]. MC68HC08 [GM95]. McCarthy [Ste97j]. Me [Eri98l, Flo91d, Har95a, Ste99d, Swa98h]. measure [Cop94]. Measuring [Har93b, DDJ99a]. mechanisms [For98a]. Medals [DDJ98b]. Media [DDJ98d, Eri98c, Ber96a, Gri97a, Wil97g]. Median [Byl99, Kru94a, NRG^{+}99]. Median-cut [Kru94a]. medium [Woe91b]. medium-heavyweight [Woe91b]. Meeting [Koo94a]. Meets [DDJ98g, Hew92, Hem96a, Lad91]. Member [Mye97]. memcpy [Fly97]. memes [Gut94]. Memorable [Swa98o]. Memory [App94, BT93, Blu90, Bol97, Bri90, CS95, Dip95, Fri94, GH99, Huf93b, Huf93a, Lam95b, Moe90, Nel93b, Pat97, PB98b, Ric98, Sae95b, Sch90c, Spe90, Bet91a, Gin91, Ham93, Hir92c, Mye97, Nel93a]. Memory-constrained [Pat97]. Memory-Efficient [PB98b]. Memory-mapped [BT93, Huf93b, Huf93a]. Men [Val95f]. Menus [Rol92, Hop91]. mesh [SC97]. Message [Bet94d, Lam95b, Ste92-33, DDJ99b, Arn96, Cra90, Dou90, Sta99, Ger93]. Message-Handling [Bet94d]. Messages [SS99, Joh97b, Ste97g]. Messaging [DDJ98c, Pie93b, Tha97]. Metacard [PJR^{+}98]. Metafiles [Del97]. metaprograms [VP96]. Method [Ber97j, Nic92, Wil97e]. Methodology [HM95]. Metrics [BD97]. Mewel [Ste91v]. MFC [Bev95, Cul95, HW96, Hoe98, Ler96, LHB^{+}98, Sch98a, Sha99a, SW96b, SW97d, SW97e, Val95a, WS96a, WS96c, Yam93]. Michael [Gul95, Swa97d, Woe96a, Yam96]. MICO [Pud98]. Microcontroller [Fly97, Gad97, Van91, McR93]. Microcontrollers [RK92]. Microkernel [BBB94, SG97]. Microprocessor [DDJ98d, Tre93]. Microprocessors [Dun91a, Tre94]. Microsoft [Bar99c, Boy98c, Bug98, Con94a, DDJ99g, Dav99, Lad90b, Len97, PF90a, PF90b, Sch90b, Sch91a, Sch98c, Sch94n, Ste97h, Tor95, Wil98b, Wil99a]. Middleware [Car96b]. MIDI [Ste97g, Ste97k, Ste98g, Wri96]. MIDIFile [Ste98g]. Might [DDJ98f]. Migrating [Nes94]. Migration [Ste98e, Ste98i]. Mike [Ste99f]. Milestone [Val94e]. Milestones [Val95f]. MIME [Kie95]. Mimicking [For98a]. Mindstorms [BAC^{+}99, Wen99]. Mine [DDJ98h]. Mini [SW97e]. Minimal [Pes96, Mil91]. Minimum [Die93, Boy97]. miniport [QG97]. Minis [Bra90]. missing [JJ92m, JJ92c]. mix [Bak97b]. Mixed [DDJ98d, Nor91, WS90]. Mixed-Language [WS90, Nor91]. Mixing [CL97, Loy92]. ML [Var91]. MMPC [SS99]. MMURTL [Bur94]. MMX [Fom99]. mod [Wil92a]. Mode [Col97b, Col98a, Col98d, Dud90, Fri90a, Fri90b, Bar92, Chu92b, Kno92b, Loy92, Col97e, Col97d, Col97f]. Model [Gig97, Hol97, Rob95b, vdWS97]. Modeless [New93c]. Modeling [DDJ99f, Smi97, SWF^{+}99, Var93b, Var93e, Wea97, Nel91a, Var91]. Models [CL97, DV99, Jan99b, SS93, Wit95b, Gol90]. Modern [App97]. modify [Zha97]. Modula [AJ90, Fre95, Har92, Har94a, Kin91a]. Modula-2 [AJ90, Kin91a]. Modula-3 [Fre95, Har92, Har94a]. Modulation [Kno98]. module [McC97, Tew97]. Modules [Wel95]. Money [DDJ98c, Eri98f, DDJ99a, Val95b]. Monitor [Kno94, Sta94a]. Monitoring [GWL96, Kra91, Oes97, Kno92a, Kno93a]. monitors [RC97a]. Monopoles [Sha99i]. month [Swa97d, Wil97g]. Moore [Eri99g, Woe92a]. Moret [Och94b]. Morphing [And94b, Hal93, Lew94]. most [Kis97, SW96b]. Motif [Hus92, Kno95]. Motion [BAC^{+}99, Har99, Mar95b, Wil94e, Wit97b]. Motion-Tracking [BAC^{+}99, Har99]. Motorola [Soj96, Woe93]. Mouth [DDJ98c]. Move [AJ90]. Movement [GF99]. Moving [Ait96b, HL99, Maz92, Ram97a, Swa99o]. MP3 [Bar99a]. Mr. [Eri97g]. MS [Tro94, Wil94c]. MS-DOS [Tro94, Wil94c]. much [Mye97]. MUD [Cro94c]. Müller [Sch94m]. Multi [Ger93]. Multi-app [Ger93]. Multicasting [Qui97]. Multicolumn [Pro94]. Multilevel [Bev96]. Multimedia [Car97a, Kru94b, May95, Mus93, Rat94, Reg94]. Multiplatform [Gra94]. Multiplayer [JS94]. Multiple [Car92, Kal92, KR96, Kau96a, Ler96, PF90a, PF90b, Rog95b, RK92, SS99]. Multiple-precision [Kal92]. Multiplexing [McM90]. Multiprocessing [Aye90a, Aye90c, NV94].

Multiprogramming [JJ91b, JJ91c]. multitasker [Woe91b]. Multitasking [Gib94, Kof95, Nog99, Vai93, Yue96b, JJ91b, JJ91c, McC91, Tar91]. multithreaded [Hig97, Ver97]. Multithreading [Cro98b, Esp96]. Multitool [Ste93w]. Multiuser [Kry92a, Kry92b]. Munging [Fer95]. Music [Rat95, SDB^{+}99, WEM^{+}99]. must [Bak97b]. Mutual [For98a]. MVC [Aye90b]. My [Ber97f, Eri91a, Sha99j, Swa98b, Ber98b, Eri99f]. Myopic [SWF^{+}99].

NAFTA [Sma94]. Named [She93]. Names [SWF^{+}99]. Namespaces [MBK^{+}98, Pen94]. Nanoelectromechanical [DDJ98e]. Nanomedicine [DDJ98f]. Nathan [Mye97]. National [DDJ98b, DDJ99c, Koo94a]. Native [Gri97b, Nor96, Wil97e, Wil98b]. NC [The98b]. ncpfs [Len97]. NCSA [FK98]. need [Sto97, Swa96y, Wal97]. Needs [Eri96l]. Negotiator [Eri99k]. Net [Val95e]. NetBEUI [Bra94c]. NetRexx [Tay98]. Nets [Kli91, Law90, FV93a, Var91]. Netscape [Ber96a, DDJ98i, Dav95a, GW96]. Netware [Sch97b, Kra91, Rei93b, Sch96h, Ser93, Szc93, Cre98]. Network [Bra94c, Cro95a, Czu90, The98b, Jam93, Lu95, ML93, Sch99e, Sta93b, Var95a, Var95b, Won99, Yue94, FY96, Jam92, JJ95b, Met96, Qui97, TWH93, Yor96]. Network-Enabled [Won99]. Networking [Ano96i, Bet95b, BH96, CS94, Gam96a, Hov94, Kle91a, McC96a, Sha93c, Jol92b]. Networks [Cho99, JJ94, Kar93, Kli90, Sta95a, Var93b, Var93e]. Neural [Czu90, FV93a, Kar93, Kli90, Kli91, Law90, TWH93]. neural-network [TWH93]. Neutral [Gil94a, Gil94b, Gil94c]. Never [TMP^{+}99]. newly [Rie96]. News [Ano96j, Ano97g, Ano97h, Ano97i, Bar99c, BEK99, Bar99a, Bar99d, DDJ97c, DDJ97a, DDJ97b, DDJ97d, DDJ98c, DDJ98e, DDJ98f, The98b, DDJ98h, DDJ98g, DDJ98a, DDJ98d, The98a, DDJ98b, DDJ98i, DDJ99b, DDJ99f, DDJ99e, DDJ99g, DDJ99c, DDJ99d, Eri97b, DDJ99a, Ano97f, Krt96]. Next [Lin94d, Ste98i, Val94f]. nfs [Len97]. Niche [Cho99]. Night [SDB^{+}99]. Nimmerics [Sha98g]. Ninth [Dor95]. NLM [Day92]. NLM-Based [Day92]. NLMs [Cav95]. No [Bar99d, DDJ98f, DDJ99b, Wil98d, Ste97f]. Nod [The98b]. nodes [Len97]. noisy [Kim92]. Non [Li94c, Tro94]. Non-DOS [Tro94]. Non-SCSI [Li94c]. Nonsecret [DDJ98i]. NOS [Swa97u]. Notation [Cha95c, Cho95, Cha94]. Notes [DDJ98h, Abr92a]. Notifiers [Pom98]. Novell [Jam92, Kra91, Szc93]. now-classic [GHJV96]. NP [Och94b, Kis97, Sch94j]. NP-complete [Kis97]. NP-completeness [Sch94j]. NSBD [DL98]. NT [QG97, Bar99d, BT93, Bra94b, Fra99, Hig97, Hov94, Let96, Rob96, Rob98a, RC97a, RC97b, Sch96h, Sch99d, Sip97, Ste98l, Ten98, Tro97, Tro98, Vai93, Wil94c, Wil98b, Yam96]. NT-Style [Wil94c]. NTRegmon [RC97b]. Number [Can90, Kal93, Mai91, SC97]. Numbers [GC98, Ple95, Plu94, Swa98o]. Numeric [Fri95, Let91]. Numerical [Jer92, Lea94, KS97, Woe97c]. Nutmeg [Gar96]. Nutshell [Nic96a]. NVOs [Rob95c].

O [McC96b, SRA^{+}94, BT93, FY96, GS97, HS94, Huf93b, Huf93a, Kno94, Rob96]. Oberon [Mar94a, Tem94]. OBJ [Sie91]. Object [BD97, Bet94b, Bla96, BB90, Cal95a, CL97, Dan96, Ent90b, Fer98, Flo91c, Flo93a, Flo93c, GHJV96, GM97, Har91, Hav98a, HG95, Hyd90, Kau96b, Kum95, Moo95, SHR^{+}99, Shi92, Sho96, Smi97, Ste93b, Too90, Val94a, Ver93, vdWS97, Wil94a, Bet91c, DGR90, Fel92, Flo90b, McB97, McC97, McS92, Mil91, Rie96, Won92]. Object-Oriented [Bet94b, Cal95a, Fer98, Flo93a, Flo93c, HG95, Hyd90, Kum95, Moo95, BD97, Bla96, GHJV96, Har91, Kau96b, Shi92, Smi97, Ste93b, Too90, Ver93, Bet91c, Fel92, Flo90b, McB97, McS92, Mil91, Rie96, Won92]. ObjecTime [Smi97]. Objective [Gro97b]. Objective-C [Gro97b]. Objects [Ber97g, Ber97c, Bet95b, Bet98, Bev98, BN97, Cox92, GH95a, JK95, Lad90c, Lew94, Lip99b, Lom98, MBK^{+}98, Pom96, Pra95, Rus96, dS95, Sou99, Ste92-34, Wee90, da 95, Bet94c, Car96a, KR91, Sip96c]. Observations [GH95b, Swa97i]. Observer [GH95b]. occurs [RC97b]. Octrees [Cla96a]. ODBC [Hig97, McC95, Nor96]. ODBC-compliant [Hig97]. Odd [WC99]. Off [DDJ98f, Ban98a, Dur97, MK97, Row97, Wal97]. off-the-shelf [MK97, Row97]. Oft [WS99]. Oft-Delayed [WS99]. Old [Abr94, DDJ98g, Ste98f]. OLE [Fai95, LaP95, Rod95b, WS96d]. OLE2 [Cou94a, Cou94b]. Oleg [Kis97]. Oliver [Och94c]. Olympiad [DDJ98b]. Omniheurist [Sha98i, Sha98g, Sha98c, Sha98f, Sha98d, Sha98e, Sha98b, Sha98h, Sha99e, Sha99d, Sha99l, Sha99b, Sha99k, Sha99f, Sha99h, Sha99i, Sha99g, Sha99c, Sha99j]. Once [Eri97m, Gar91]. One [Abr95d, Ive90, Sch91c, Sch95k, Car97c, SW96b, Ste97g]. One-Pass [Ive90]. One-way [Sch91c]. Online [Bar99c, Bet94a, BCP^{+}99, The98b, DDJ99e, GW97, Koo94a, MGL^{+}98, MBK^{+}98, Row97]. OO [Bet93a]. OODCE [RZ96]. OOPSLA [Eri95n]. Op [BCP^{+}99, MGL^{+}98, MBK^{+}98]. Op-Ed [MGL^{+}98, MBK^{+}98]. Op/Eds [BCP^{+}99]. Open [Con99, DDJ99b, Eri99h, Lis98, Ste92-33, Ste99f, RC97a, Wil96d, Con98, DDJ99c, GF99, NRG^{+}99, Ste99f]. OpenDoc [OH95]. OpenGL [DDJ98f, Fos95a, Fos96, Ram97b]. Opening [Sch90g]. Operating [Bur94, Kau95, McR96b, Nog99, Var94e, Woo99, Hav98b, Jen95, McR96a].

Operational [CM95]. Operations

[Gil95a, Gil95b, Rog95a, Ber99]. Operator [Mey99, dSW98]. OPROMs [RT97]. optical [Gar91]. Optimal [DGR90]. Optimization [Abr94, Fom99, Mye97, Sch90a, Sch90b, Sto95, Abr91a, PS97b, Gul95]. Optimizations [CKS98, Fri95]. Optimizing [Bis96, Fly97, Gra91, Lad90b, McG94a, NAM96, Ros95b, Sch94a, Sto94]. OPTLINK [Pie93a]. Oracle [WSM95b, WSM95a]. Order [DDJ99e, Smi93a, All91]. Ordering [All90]. Oriented [Bet94b, Cal95a, Fer98, Flo93a, Flo93c, HG95, Hyd90, Kum95, Moo95, BD97, Bet91c, Bla96, Fel92, Flo90b, GHJV96, Har91, Hof97, Kau96b, Law97, McB97, McS92, Mil91, Shi92, Smi97, Ste93b, Too90, Ver93, Rie96, Won92]. original [SC97]. originates [RC97a]. OS-9 [Dib96]. OS/2 [And90a, And90b, Bar93a, BDE93, Fol94, Gig93b, Gig93a, Kan94, Mak90, Old94, Pro94, Ric91, Rod94b, Rod94c, Rot90, Sch90g, SRA^{+}94, Wil94e, Wri94]. OSF [Blo95]. OSF/DCE [Blo95]. OSI [CT90, Sta93b]. Other [Abr96a, Swa98c, Swa98e, Swa98o]. Our [Car97c, MK97, QG97, Soo98b]. out-of-range [Lee93a]. overloading [dSW98]. Overturned [Bro95c]. Overview [Val91c]. Own [Ber97f, Bet94b, Bur94, Cro94a, Jam93, JS94, Rog95c, Sem96, Tay99, Wil90d, Wil92b, Yue94, Avi92, Bet91c, Flo90b, Kno92b, Wil90c, Zha97]. Oyster [Ben99c].

P [Och94b]. P. [Sch93k, Sch93d]. P6 [Val95g]. package [Rod91]. Packard [Aus97]. Packet [Sta94a, Cod94b]. Page [Col96e, Lyn95, Swa99u, Tes96a]. Pages [MSE^{+}99, Joh97a, SW98c, Tes96b, Tre99b, WS96d]. Paging [Col96d]. Pair [Sch95o]. Pairing [Ran94]. palette [Bet93b].

Palettes [Cog93]. PalmOS [Win99]. PalmPilot [Ber98b, Chr99, Mus99]. Panama [DC98]. Paper [BEK99]. Paradigm [Aye90b, Mae92]. Paradigms [Swa90a, Swa90b, Swa90c, Swa90d, Swa90e, Swa90f, Swa90g, Swa90h, Swa90i, Swa90j, Swa90k, Swa90l, Swa91b, Swa91c, Swa91d, Swa91e, Swa91f, Swa91g, Swa91h, Swa91i, Swa91j, Swa91k, Swa91l, Swa91m, Swa92a, Swa92b, Swa92c, Swa92d, Swa92e, Swa92f, Swa92g, Swa92h, Swa92i, Swa92j, Swa92k, Swa92l, Swa93b, Swa93c, Swa93d, Swa93e, Swa93f, Swa93g, Swa93h, Swa93i, Swa93j, Swa93l, Swa93k, Swa93m, Swa94b, Swa94c, Swa94d, Swa94e, Swa94f, Swa94g, Swa94h, Swa94i, Swa94j, Swa94l, Swa94k, Swa94m, Swa95n, Swa95c, Swa95d, Swa95e, Swa95g, Swa95f, Swa95h, Swa95i, Swa95j, Swa95k, Swa95l, Swa95m, Swa96l, Swa96k, Swa96a, Swa96b, Swa96c, Swa96i, Swa96j, Swa97k, Swa97e, Swa97g, Swa97f, Swa97d, Swa97h, Swa97i, Swa97l, Swa97j, Swa97a, Swa97b, Swa97c, Swa98b, Swa98a, Swa98c, Swa98d, Swa98e, Swa98f, Swa98k, Swa98l, Swa98g]. Paradigms [Swa98h, Swa98j, Swa98i, Swa99b, Swa99c, Swa99d, Swa99e, Swa99i, Swa99f, Swa99g, Swa99h, Swa99j, Swa99a, Swa96d, Swa96e, Swa96f, Swa96g, Swa96h]. Parallel [Bau92b, Che92, Ell90, Gad96, GS97, Pan99, Sch99a, dSW98, TKG^{+}99]. Parametric [Zig90]. Paranoia [Swa98t]. Parasol [Jer93, Jer94]. parse [MK97]. Parsers [Mir96]. parses [Joh97a]. Parsing [Ber99, BCP^{+}99, EJ95, Ste98g]. Part [And90a, And90b, Cha95a, Cha95b, Dav93a, Dav93b, DY99a, DY99b, Fri90a, Fri90b, Gil94c, HO92, HO93, Hir94b, Kry92a, Kry92b, Lin95a, Lin95b, OH95, PF90a, PF90b, Sch90h, Sch90i, Sch93k, Sch93d, Sch94w, Sch94v, Tro91a, Tro91b, Wil90c, Wil90d, Wil98e, Wil98f, Woe94b, Woe94a, Gil94b, FN98, Zyt94a, Zyt94b]. Partial [Fre95]. Particular [WC99]. Partitioning [Alm96, McC91]. Partitions [Cel94]. parts [Row97]. PASCAL [Lad90c, Flo92, FNL91b, FNL91a, Gin91]. Pass [Ive90]. Passes [DDJ99d]. passing [Arn96]. Passport [MGL^{+}98]. Password [Ber98c, Pop96, Sta94b]. Past [Eri97h, Swa97d]. Patch [SRA^{+}94]. Patch-O-Rama [SRA^{+}94]. Patent [Aha95, DDJ99c]. Patents [Bau90b, Bro95d, DDJ99f, The90, Val94g, MMB^{+}90]. Path [Ste97j, New94a]. Pattern [Ban98b, GH96, Ngu96, Pit98b, Swa98h, Van97, Ver96]. Patterns [Bec94, GHJV96, Gro97b, HG95, Rem97, NA96, Wil96b]. Paul [Flo94h, Kis96]. Pay [DDJ98f]. PC [And94b, Bau93, DDJ99c, Fri94, Ham93, Ham95b, HS94, JJ91v, Rat93, SvA^{+}98]. PCMCIA [Mil94]. PCs [Bra90, NV94, DDJ99a]. PCX [Chu91]. Peace [Wil96f]. Pearls [Ben99c]. Peer [TSR98]. Peer-to-Peer [TSR98]. PenPoint [AH91, Val91c]. Pens [DDJ99f, DDJ99a]. Pentium [Abr94, AB97, Bak97a, BL95, Coe95, Col96d, Col96e, Col97a, Col97e, Col97b, Col98c, Col98d, Dah93, Fri95, Moo93, Sch94a, Sto94, SKD93, Val95h]. People [War91]. Perfect [Pes96]. Performance [BL95, BW98, Bun94, Cal97, Cro93b, Cro93a, Dun93d, Fly97, Fri95, Har94b, KR96, Kie96b, Krt96, Lip99a, MPV95, New93d, Nor96, PRB98, Shi99, Tre94, Tym96, Bet91b, Gri97b, Mot91, NAM96, Nic93b]. periodic [Yam94]. Peripheral [FN98]. Peripherals [Hir94a, Hir94b]. Perl [Bal99, DDJ98a, DDJ99b, Kie98, Kie99c, Ler97, Sha93c, Sha94, Val95i, Zha97]. Perl-Win32 [Ler97]. PerLDAP [Nee99]. Permutation [Iye95]. Persistence [Cha96, Gab92, SHR^{+}99, Tic97]. Persistent [Ber97g, Bri99, Lad90c, Ste92-34, Ste98j]. Personal [DDJ98e, Hir92a, Hir92b, Hir92c]. PersonalJava [DY99a, DY99b]. Perusing [Wil97k]. PETE [HCKS99]. Peter [Nic96b, vdWS97]. Petri [Var91]. Phar [Wil92a]. Phil [ML97]. Philosophical [Swa98k]. Phone [Li94a]. Photon [Var95c]. Physical [FSW99]. Pick [Can90]. Pick-A-Number [Can90]. picoPERC [Nil98]. Picture [DDJ99g, Mur96]. pie [Hop91, Rol92]. pieces [JJ92m, JJ92c]. PIF [Mau93]. pile [McR93]. pile-ups [McR93]. Pin [DDJ99a]. Pioneers [DDJ98e]. Pipe [She93]. Pipeline [Val95g]. Pipes [Cop96, Ler97, Wil94b]. Pique [Swa97w]. Pixel [All90]. PKZIP [Nel97]. Plain [WC99]. Plan [PPTT91]. Planned [DDJ99c]. Planning [Eri98k]. PlateWare [DDJ98g]. Platform [Cop95, Dio93b, Fir93a, NRG^{+}99, SDB^{+}99, Bro94a, Car97a, Dio93a, FH95, Kis96, Lam95a, Cul94]. platform-independent [Kis96]. Play [DDJ98d, Eri97d, RJ95, RT97]. Playback [Dav98, RC96c, Sym94]. Plays [Eri97i]. Please [Ste99k]. PLoP [Vli94]. Plug [Fir93a, Hoe98, LHB^{+}98, RJ95, RT97]. Plug-and-play [RT97]. Plug-In [Fir93a, LHB^{+}98, Hoe98]. Plus [Bau90b, Mud96]. PNG [Cro95a]. Point [Mot93, Wil93a]. Pointer [Pap95, Hel98]. Pointer-based [Pap95]. Pointers [Mey99, Sch90a]. Points [SW98a, Kis97]. Politics [Swa98u]. poll [Fag98a]. Polygon [SC97]. polygonal [Nic91b]. polygons [SC97]. Polymorphic [Jol94, Vos94]. Polynomial [Var93b, Var93e]. Pooling [Sae95b]. Pools [Cal97]. POP3 [Duf95]. Popularity [Cla95]. popularized [Nel97]. Port [Gad96, GS97, Mus99, Rid95, Rob96, Wil93b]. Portability [Hir92b, Hof95a, Keh94, Luu94, Ros94, You98]. Portable [Bra92, Car97b, Gor93, Hil95, Kof95, Tew97, Van93, Wei93, Wei95, WB95, Cal96, Kis96, Mey92, Nic91a, Tro91a, Tro91b, Cro95a, HCKS99]. Porting [Bau93, Bra90, Die99, Dud90, Fra99, Gla91, Han93, Huf93c, JJ91s, JJ91b, JJ91c, JJ91a, JJ91t, JJ91q, JJ91d, JJ91o, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ91r, JJ91u, JJ91p, JJ91v, JJ92e, JJ92a, JJ92b, JJ92d, JJ92m, JJ92c, JJ92g, JJ92h, JJ92i, JJ92j, JJ92k, JJ92l, Kap95, Mat95, One94b, SDB^{+}99, Wri94, Wri99, Sax92a, Yam96]. ports [Yam96]. POSIX.4 [Jol95a]. possible [Esp97, Len97, Ste97g, Tew97]. Post [Val95a]. Post-MFC [Val95a]. Postman [Lu94]. Postmortem [Pie92b]. Poughkeepsie [Eri97m]. POV [Lin94b]. POV-Ray [Lin94b]. Power [Bit91, DDJ99f, Eri97i, Flo94h, Swa98v, Ton90, CJ94, Duv91, Hoe98, Ric91]. PowerBASIC [Sch95b, Ton90]. PowerBuilder [NA96, Rob95c, Van97]. powerful [Col97c, Law97, Man99a, Sto97, WW99]. PowerMac [Kap95]. PowerPC [Fer95, Gil95a, Gil95b, MPV95, Ros95b, SW95c]. PowerTV [Woo99]. PPD [Val91c]. PPTP [DDJ98c]. Practical [BR96, Fla96, JJ91d, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ92g, JJ92h, JJ92i, JJ92j, JJ92k, JJ92l, Kum95, JJ91o, Sha97]. Practices [Val95c]. Praise [Eri98i]. Precision [Gri97b, Her98, Mot93, Rog95b, Kal92]. predicting [FV93a]. Preemptive [Yue96b]. Preprocessing [Arn98]. Present [Eri97h, QG97, RC97a, RC97b]. presents [Car97c, Len97, Ste97g, Sto97, Wal97, Zha97]. previous [Tha97]. previous-generation [Tha97]. Primitives [Fos96]. Principles [PRB98, Nic91a, Sha93d]. Print [Ber97h, CJ94]. Printer [GS97, Tro97, CJ94]. Printers [Kra91, Low92]. printf [Mis90]. Printing [Lu95, You92, Pic91]. Priority [Nel96c]. Private [New93b]. Pro [You97, Col96d]. Problem [GR98, Van98c, Man99a]. problem-solving [Man99a]. Problems [Pea95, Kis97, Kno93a, Nic91a]. Procedures [Nor99, Ste97f]. Process [Cal95c, Col99, Cop94, HHW^{+}99, Hav98b, Hav99a, PJR^{+}98, Swa99q, WEM^{+}99, Dib96, QG97]. Processes [Wei93, Bau90a]. Processing [Coo98a, How98, Kli90, TL95a, Web96, Car97c, Das93, Gal91, McR93, Pat97, TL95b]. Processor [Col96d, Col97b, Dut95, LHM93, May95, SKD93, Val95g, Hel93, Moo93]. product [McG93]. Production [Wil98d]. Professional [Ent90b, WBB^{+}97]. Profile [CKS98, BDE93]. Profile-Guided [CKS98]. Profiler [Boo92, Wil99b]. Profilers [Arm93, van98a]. Profiling [New93d]. Program [Ber96e, Lip99a, Man94b, MK97, Sch94x, Wil92b, Won92, Joh97a, McC97, Mye91a, RC97a, Ste96m, Ste97g, Yam94]. Programmable [IR99, Tre95]. Programmer [Ams99, Ang99, Ber96p, BF97, Ber98e, BDT97, CKS98, Cha95d, Cla96b, Cro98b, Cro98a, DDJ98h, DDJ98a, Dun92a, Dun93a, Dun93b, Dun93c, Dun95, Eri93m, Eri93n, Eri93o, Eri94n, Eri94o, Eri95o, Eri97m, Fer98, FV97, Flo94h, Fra90, Gab96, Gra98, GG95b, Gri96a, Gri96b, GF99, Gul95, HHW^{+}99, Hed98, HT98, JL98, Joh99, Jol92a, Jol93c, Jol95b, Kim97l, Kim99i, Lis98, LS99, McB97, McC96b, McG98b, McR96c, Mit96, ML97, Och94a, One94c, Pfe96, Rei94b, Rot98, Rot99, Rou98, Sal96, Sch99b, Sch92e, Sch92f, She99, SDB^{+}99, dSW98, Soo98b, Sta96, Sta98b, Ste93x, Ste93y, Ste94q, Ste94r, Ste95m, Swa95b, Tay99, Val93d, Val93e, Val95j, Var94a, Var94c, Var94d, Var95d, vdWS97, Wea97, Wil97i, Wil97j, Wil97h, Wil97g, Wil97k, Wil98d, Wil98c, Wil99d, WC99, Wil99c, WS99, Wil99e]. Programmer [You96, You97, van98a]. Programmers [BR96, Joh97b, Aus97, Bak97b]. Programming [Abr91b, Abr91c, Abr91d, Abr91e, Abr91f, Abr91g, Abr91h, Abr91i, Abr91j, Abr91k, Abr91l, Abr92b, Abr92c, Abr92d, Abr92e, Abr92f, Abr92g, Abr92h, Abr92i, Abr92j, Abr92k, Abr93a, Abr93b, Ste94j, Amb99, AN98, Bar93a, Ben99c, BW91, Bet98, BH94, Bri99, Bro94b, Cal95a, DDJ98h, Dun90a, Dun90b, Dun90c, Dun90d, Dun90e, Dun90f, Dun90g, Dun90h, Dun90i, Dun90j, Dun90k, Dun90l, Dun91b, Dun91c, Dun91d, Dun91e, Dun91f, Dun91g, Dun91h, Dun91i, Dun91j, Dun91k, Dun91l, Dun91m, Dun92c, Dun92d, Dun92e, Dun92f, Dun92g, Dun92h, Dun92i, Dun92j, Dun92k, Dun92l, Dun92m, Dun92n, Dun93e, Dun93f, Dun93g, Dun93h, Eri95a, Eri97a, Fai95, Fle93, Flo91b, Flo91d, Flo94h, Fos95a, Fos96, Fre92, FNL91b, Gab92, Gol96, GW97, Gor93, GM95, Gut92, Har92, Har94a]. Programming [Hen97a, How90, How93, How96, Hus99, Hyd90, JL98, Jer93, Jer94, Joh97a, Jol95a, JK95, Kah92, Kef95, Kim95, Kin93, LaO98, Laf99, Man93, Mar90, May95, MPV95, McG98a, Mir94, Moo92, Naj95, Nee99, Nic93a, NFP96, NK98, Och94c, Ous94, Pla93, Rad98, Rod95c, Rod95d, Ros97, Sat90, Saw94, Sha94, Shu96, Sme98, Ste93a, Ste90a, Ste90l, Ste90b, Ste90c, Ste90d, Ste90e, Ste90f, Ste90g, Ste90h, Ste90i, Ste90j, Ste90k, Ste91a, Ste91b, Ste91c, Ste91d, Ste91e, Ste91f, Ste91g, Ste91h, Ste91i, Ste91j, Ste91k, Ste91l, Ste92c, Ste92d, Ste92e, Ste92f, Ste92g, Ste92h, Ste92i, Ste92j, Ste92k, Ste92l, Ste92m, Ste92n, Ste92o, Ste93c, Ste93d, Ste93e, Ste93f, Ste93g, Ste93h, Ste93i, Ste93j, Ste93k, Ste93m, Ste93l, Ste93n, Ste94b, Ste94c, Ste94d, Ste94e, Ste94f, Ste94g]. Programming [Ste94h, Ste94i, Ste94k, Ste94l, Ste94m, Ste94n, Ste94o, Ste95l, Ste95a, Ste95b, Ste95c, Ste95d, Ste95e, Ste95f, Ste95g, Ste95h, Ste95i, Ste95j, Ste95k, Ste96a, Ste96b, Ste96c, Ste96d, Ste96f, Ste96i, Ste96j, Ste96k, Ste97g, Ste97j, Ste97l, Ste97h, Ste97i, Ste97k, Ste97e, Ste97f, Ste97b, Ste97c, Ste97d, Ste98d, Ste98e, Ste98b, Ste98c, Ste98f, Ste98g, Ste98i, Ste98j, Ste98h, Ste98k, Ste98l, Ste99j, Ste99a, Ste99b, Ste99c, Ste99k, Ste99d, Ste99e, Ste99f, Ste99g, Ste99i, Ste99h, SKD93, Swa90a, Swa90b, Swa90c, Swa90d, Swa90e, Swa90f, Swa90g, Swa90h, Swa90i, Swa90j, Swa90k, Swa90l, Swa91b, Swa91c, Swa91d, Swa91e, Swa91f, Swa91g, Swa91h, Swa91i, Swa91j, Swa91k, Swa91l, Swa91m, Swa92a, Swa92b, Swa92c, Swa92d, Swa92e, Swa92f, Swa92g, Swa92h, Swa92i, Swa92j, Swa92k, Swa92l]. Programming [Swa93b, Swa93c, Swa93d, Swa93e, Swa93f, Swa93g, Swa93h, Swa93i, Swa93j, Swa93l, Swa93k, Swa93m, Swa94b, Swa94c, Swa94d, Swa94e, Swa94f, Swa94g, Swa94h, Swa94i, Swa94j, Swa94l, Swa94k, Swa94m, Swa95n, Swa95c, Swa95d, Swa95e, Swa95g, Swa95f, Swa95h, Swa95i, Swa95j, Swa95k, Swa95l, Swa95m, Swa96l, Swa96k, Swa96a, Swa96b, Swa96c, Swa96d, Swa96e, Swa96f, Swa96g, Swa96h, Swa96i, Swa96j, Swa97k, Swa97e, Swa97g, Swa97f, Swa97d, Swa97h, Swa97i, Swa97l, Swa97j, Swa97a, Swa97b, Swa97c, Swa98b, Swa98a, Swa98c, Swa98d, Swa98e, Swa98f, Swa98k, Swa98l, Swa98g, Swa98h, Swa98j, Swa98i, Swa99b, Swa99c, Swa99d, Swa99e, Swa99i, Swa99f, Swa99g, Swa99h, Swa99j, Swa99a, Tay98, Tem94, Tha97, TSR98, Val93c, Van91, Wal92a, Wil92a, Wil97c, Woe93, WS90, van95, Aus97, Bar93b, Dem93, FH95, FNL91a]. programming [Hem96a, Mae92, Mye98, Nor91, Omo93, Sch92a, Ser93, dSW98, Ste96m, Ste96l, Ste96e, Ste96g, Ste96h, Wil92c, Zig95, Eri96m, Eri98a, Wil97g]. Programs [BC99, Bra90, CR92, Dud90, Gul95, Wil95a, FY96, Sax92a, Wes94]. Prograph [Sme98]. progresses [vdW97b]. Project [FV97, Joh90, Ste94s, Ste94t, Ste94u, Ste94v, Ste94w, Ste94x, Ste99c, DDJ99f]. Projects [JL98, Ste98f]. Prolog [Hov94, Mer94]. Property [Kin99, Mar94b, RR98, SW98c, Sma94, WS96d, Boy97]. Proposing [DAM96, Tea91]. PropView [Kle91b]. Protected [Bar92, Dud90, Fri90a, Fri90b, Chu92b, Kno92b, Loy92]. Protected-mode [Bar92, Chu92b, Kno92b, Loy92]. Protection [Bro95d, Mar94b, Sch90h, Sch90i]. Protocol [GPVV98, Man99b, Szc93, Bra94c, Fre92, Gri97a, Len97, Mar97, Sar97, Sch96h, Sch97b]. Protocols [Asl98, Bol97, Boz99, Bur95a, Cho99, Sch99e, Val95d, deC99, Jol94]. Prototyping [Dem93, Lea93]. proverbial [Ste97e]. Provided [TMP^{+}99]. Provider [Tre99a]. provides [Hig97, Mae92]. Proving [LHB^{+}98, Sav98]. Provisions [Sma94]. Proxies [Lip99b]. Proxy [Lea93]. PSCs [DDJ98e]. Pseudo [Dri94, HC99, Sch92b, Mai91]. Pseudo-code [Dri94]. Pseudo-Incremental [HC99]. Pseudo-random [Sch92b]. PT [Yam94]. PTF [Yam96]. Pthreads [Yam96]. PTL [PSS^{+}98b, Ste98d]. Public [DDJ98b, Smi93b, Sch92d]. Public-key [Smi93b, Sch92d]. Publishing [Hig97]. Pure [Wil99a]. Purpose [Ber96f, Mel98, Gar91, Mor91d]. Push [Ber97a, Eri97b]. Pushy [Har95a]. put [Ste97e]. puts [Ang99]. Putting [DDJ98c, Smi93a]. Puzzle [Swa97t, Swa98w]. Python [Ang99, How98, Kla99, McG98a, McG98b].

Q [Sum91a, Sum91b, Sum91c, Sum91d, Sum92a, Sum92b, Sum92c, Sum92d, Sum92e, Sum92f, Sum92g, Sum92h]. Q&A [AW99, BC99, BA96, Ber96e, Ber96d, Ber96b, Ber96a, Ber96f, Ber96c, Ber97i, Ber97b, Ber97j, Ber97g, Ber97a, Ber97f, Ber97l, Ber97e, Ber97d, Ber97k, Ber97h, Ber97c, Ber98a, Ber98d, Ber98b, Ber98c, Cri98, Eas99, GS98, Gri99, HL99, Moo98, Pit98a, Pit99, Pur98, Ran99a, Ran99b, Ses98, Soo98a, Tre99b, Wil99a, Wil99b]. QEMM [Bro94b]. QEMM-386 [Bro94b]. QMOOD [BD97]. QNX [Var95c]. QPI [Bro94b]. quadtrees [Das93]. Quake [Abr96b, Abr96c]. Quality [Pfl98, Smy97].

Quantization [Cla96a, Kru94a]. Quantized [Yor95]. Quantum [Bar99a, DDJ98d]. QUANTUMdsp [Mir94]. Queries [PSS^{+}98b]. Querying [Sav98]. questions [Val91b]. Queue [Cri98]. Queues [Lam95b, Nel96c]. Quick [Bet95c, Nic91b, Wil93b]. Quicksort [BS98a, QTC^{+}99]. QuickTime [Ful92, Wal92a, Car97a]. QuickTime/AVK [Ful92]. QuickWin [Wil93b]. Quincy [Ste94y, Ste94s, Ste94t, Ste94u, Ste94v, Ste94w, Ste94x, Ste96n, Ste96o, Ste96p, Ste96q, Ste96r, Ste98f].

R. [Hel93, Moo93]. Radio [Bur95a, Eri98l]. Radix [BS98a, Cal95b]. RAID [JW92, WJ92]. Rainer [Sto97]. raising [Duv91]. RAM [Rea90]. Rama [SRA^{+}94]. Random [Eri97j, Plu94, Mai91, Sch92b]. Randomness [GW96]. range [Lee93a]. Rapid [Wit97a]. Rattle [Swa99l]. RaveKit [Car97b]. Ray [Lin94b, Lyk90, Man94b, Rus99, Sem95, Zud96, Lin94b, Man94b]. Ray-Casting [Sem95]. Ray-Tracing [Man94b]. RC5 [Riv95]. re [Gol90, Ste92r, Ste92s, Ste92t, Ste92u, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u]. re-implementation [Ste92r, Ste92s, Ste92t, Ste92u, Ste93o, Ste93p, Ste93q, Ste93r, Ste93s, Ste93t, Ste93u]. re-use [Gol90]. reactive [Dru94]. Readers [Swa97v]. Reading [Eri97b, Lia95]. Readings [Eri97j]. ready [van98b]. Real [Bar99a, BW98, Bek97a, BBB94, Bun94, Coo98a, DDJ99d, Eri97l, FV93c, Fer94, HHW^{+}99, HT98, Jam90, Kar90, Kau95, Kot99, Krt98, Lus97, MGL^{+}98, MM98b, McR96b, ME96, New93a, Nol90, PB98a, Ple95, Pot92, QTC^{+}99, SG97, Smi97, Smi98, Ste98a, Swa98v, Tet94, TMH99, Wil96h, Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Gal91, GWL96, GGL^{+}95, Jen95, Loy92, McR96a, Oli91, Roa91, Sha93a, SJB92, vdW97b, van98b, Pro99]. real- [Loy92].

Real-Time [BW98, BBB94, Bun94, Coo98a, DDJ99d, HT98, Jam90, Kar90, Kau95, Kot99, Krt98, MGL^{+}98, MM98b, McR96b, ME96, New93a, Nol90, PB98a, Pot92, QTC^{+}99, Smi98, Ste98a, Tet94, TMH99, Bar99a, Fer94, Lus97, Smi97, Wil96h, Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Gal91, Jen95, McR96a, Roa91, Sha93a, SJB92, Pro99]. Realistic [Krt94, SC97]. realities [Gut94]. Reality [Bek97b, Fos95b]. Really [DDJ98a]. Rebel [Flo95g]. Rebol [Swa99d]. Reboot [Eri97f]. receiver [Hav98a]. Recognition [Kar93, Val92b, Avi92, Val93a]. Reconciling [Gal91]. Reconfigurable [Kla99]. Recording [Dav98]. Recovery [Can94b, Can94a]. Rectangle [Van98c]. Recursion [Ben98, TKG^{+}99]. Recursive [Jan91, Pic93]. recycled [Bet93a]. Recycling [DDJ99c, SvA^{+}98, Wil96g]. Red [Sch92c, Sha98a, Swa97g]. Red-Black [Sha98a, Sch92c]. Redirector [Har93a]. Redo [Bev96, Ste98k, Van97]. reduce [McC97, SC97]. redundancy [Nel92]. Redux [DDJ98c, Ste98k]. Reed [Lyp97b]. Reeves [Woe99]. Reexamining [JS92]. Reference [Abr95c]. references [Lee93a]. Referential [Kin90b]. referred [McC97]. Reflection [Tre98]. Reflects [Ste92b]. Regions [SNH93]. Registering [Bev98, MBK^{+}98]. Registry [PKA^{+}98, RC97b]. Regression [McC97]. Regular [BF97, KP99]. Reinforcement [SNC97]. Reinventing [Fos92]. Rejected [Eri97b]. Relational [Sch96b]. relationship [Ste97e]. Relationships [Cha95c, Cho95]. relativistic [Ada92]. Relay [Sta95a]. Released [DDJ98a, Aus97, Rie96]. Reliable [Hol97, Ros97]. Remote [Ber97c, Lu95, Sar95, Sta94a, Tro91a, Tro91b, Tuc98, Ber97j]. Removal [Abr96c]. Removing [FY96]. Renderer [Abr95d]. Rendering [Jan99a, Jan99b, Kie94a]. Replace [Kef92]. Replacing [Nic92].

Report [Jae91, Sak92, Woe91a, Hig98]. Reporting [Hig98]. request [RC97a]. requirements [Mye97, vdW97b]. rerunning [McC97]. Research [DDJ98c, LMWR^{+}96, JJ91r]. residence [And91]. resident [Mir92]. resident/sub [Mir92]. Resizable [Boy98a]. resolution [Rod92]. resource [WBB^{+}97]. Response [Cra97, MBK^{+}98]. Rest [Swa98l, McC97]. Restarting [BLE96]. Restoring [DDJ99g, Mye91b]. Results [The98b, Nel91b, Ste91w]. Retargetable [Yue97]. Rethinking [App94, Ste97e]. Retrieval [Krt95, Ste90l]. Retro [Ste98h]. Reusable [GHJV96, Gri97a, HG95, JK95, Lin96, Sha97]. Reuse [Hen97a]. Revealed [SW97e]. Revelation [Fre95]. Review [Aye90c, Ber90, Gul95, MBK^{+}98, Nic96a, NFP96, Nic96b, Och94b, Och94c, One94d, Flo94h, McC96b, Wil92b, Yam93]. Reviews [Cha95d]. Revisited [Bin96, DRG96, Sch97b, Swa91a, Reg90]. Revisiting [GS97]. Revolution [WS99]. REXX [Gig93b, Gig93a, Sul94b]. Reznik [Val94c]. Rhealstone [Kar90]. RIFF [Hub94]. Right [DDJ98b, HQ98, Swa96y, Eri99e]. Rights [RR98]. Ring [Yue94]. RINGO [Shm94]. RIPEMD [BDP97]. RIPEMD-160 [BDP97]. RISC [Bun94, Cro93a, Mar90, Moo92]. Risk [Ams99]. River [Eri97k]. Rivest [Woe97b]. RMI [Ses98]. Road [Val94e]. Robot [Eri99f, Car97c]. robot-control [Car97c]. Robotics [TKG^{+}99, Wen99]. Robots [Car97c, Har95b]. Rocket [Har95c]. Rogatkin [Sch93l]. RogueWave [Sch96c]. Role [DDJ98d, JJ95b]. Role-based [JJ95b]. Roll [Flo90b, Wil90c, Wil90d]. Rolling [Eri97k]. ROM [Ale99a, Ale99b, Jol93a, Har93a, Jol93b, Li94c, Ste96l]. Rome [DDJ98b]. ROMS [ML93, Ano96i, McC96a]. Ron [vdWS97, vdW97b, Woe97b]. Room [BCP^{+}99]. Root [Hei96, Boy97, JJ91t]. Rooting [Swa98g]. Rosenberg [Wil97g]. Rosetta [Sha99k]. routines [Nel97]. Routing [NK98, Sta93a]. RPC [Sip97]. RS [Cro94a]. RS-232 [Cro94a]. RSA [Sch94k, Smi93b]. RSX [Dur97]. RTFHelp [Hla95b]. RTMK [BBB94]. Rules [Abr95d, Ber99, Ros97]. Ruling [Bro95c, Eri97b]. Run [DDJ99g, Gri99, Nel93b, RJ95, Swa97v, Bev98, Nel93a]. Run-Time [RJ95, Nel93b, Nel93a]. Running [GW97, Kim96d, Mye91a]. runs [Hig97]. Runtime [Jus92].

S [McC96b, Cal95a, DDJ99b, Sto90].

S-CODER [Sto90]. Saddle [Eri95n]. Safe [Har92]. Safety [DDJ99d]. sage [Ste90a]. Said [Eri99e]. same [Nel97]. sample [QG97]. San [Bar99c, Har95d]. Santa [Swa97k]. Sather [Omo93]. Saving [Mye91b]. Say [DDJ99b, Swa97u]. Says [Bar99d]. SBI [RB97]. Scale [Lak97, Ste92a, Wil97j]. Scaling [MSE^{+}99, Cla97, Pic91]. scavenging [Jac90]. scheduler [Pie92a]. scheduling [Fer94, Kis97]. Scheme [Lea93]. Scheme-based [Lea93]. Schemes [LHM93, Sti98, Hel93, Moo93]. School [The98b]. Schools [DDJ98f]. Science [DDJ98f, Eri99j, Har95c, Tre93, Wil97i]. Scientific [Vel97]. Scot [SW96b]. Screen [Con91, Wei95]. screens [Mye91b]. Scripting [Bea98, Goa99, McG98b, Rie99, SvA^{+}98, Ste97i, WW99, Law97]. Scripts [Goa99]. ScriptX [Rez94a, Rez94b, Val94c]. Scrollable [Fai93]. Scrolling [Ste99e]. SCSI [Har93a, Li94c, QG97, Saw94, Tew97]. Sculpting [Woe92a]. SDK [Dav99, Hla96a, Rod95b, Sch91a]. SDMI [Bar99a]. Seamless [Hir92b]. Search [Bar99a, BS98b, Cam90, Hof96, Kie99d, Val95c]. Searches [Mud96, Nel96a, PKA^{+}98, Wil90e]. Searching [Boy98b, DDJ98f, Kie99c, Ste97a, TMP^{+}99]. Season [Eri99h]. sec [NK98]. Second [Eri94p, Swa98w]. Secret [DDJ98i]. Secrets [Col97f]. sector [JJ91r]. Secure [BEK99, GS98, Smi93b, Sar97, Sch92d, Sta94c]. Securely [Ber98d]. Security [Ber97f, Bou97, Boz99, JJ95a, KR96, Mar97, Rob95b, Sch94h, Sch94k, Sch95m, Sho96, Wil97f, Boy98c, JJ95b, Woe97b]. Segment [DRG90, DRG96]. Segments [PF90a, PF90b]. select [Fag98a]. Self [AW99, Bev98, Kin90b, Lia90, MBK^{+}98]. Self-Adjusting [Lia90]. Self-Executing [AW99]. Self-Referential [Kin90b]. Self-Registering [Bev98, MBK^{+}98]. Semiconductors [Mal96]. Send [Ber96d]. sender [Hav98a]. sending [Ste97g]. sensor [FNL91a]. separate [Wil98a]. sequence [Sch92b]. Sequential [Cal95b, Wil90e]. Serial [Rid95, Wel94, Yor96]. Serialization [Bev95, Cul95, Hav98a, SHR^{+}99]. Server [Bar99d, Bet94d, Bis96, Cal97, Car96b, Day92, Gam96b, GK94, Kru96, Law95, SvA^{+}98, Wit95b, Bla96, Cal96, Hig97, LaP95]. Server-Side [Kru96, SvA^{+}98]. Servers [Kla99, Mon96, New93b, SW95b, Uhl99, Val95l, Len97]. Servet [Ber97k]. Service [DDJ99c, Hig97, RC96a, RC96b]. Services [Hou99, RJ95, Ber96p, Bra94b]. Servicing [Bro96c]. Servlets [Pit98a]. Set [DDJ98h, Woe92b, Zha97]. sets [Pla92]. Settled [Eri97b]. Setup [One94a]. Setups [Ste96l]. Seven [Swa99j, Kis96]. Seven-Year [Swa99j]. Severe [Wic99]. Seybold [Bar99c]. SGI [Aus97, Ram97b]. sGs [GK94]. SHA [Sta94c]. Shaio [McC96b]. shaped [Sip96c]. shapes [Kin90a]. Shapiro [Och94b]. Shared [Fri94, Lam95b]. shares [Kis96]. Shareware [DDJ98g, Ste99g]. Sharing [Hir94a, Hir94b, Tes96a]. Sharp [Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s]. shelf [MK97, Row97]. Shell [Gig93a, Ric91, Bar93a, Gig93b]. Shelves [Eri96l]. Shines [Swa95a]. Shirts [Swa96x]. shopping [BW96]. Short [DDJ99b, DAM96]. Shortage [DDJ98h, DDJ98a]. shortcomings [Rec99]. Shots [Wu97]. should [Kis96]. Show [Swa98h, RC97b]. shows [Kis96, Mye97]. Shuttles [Sha99h]. Side [Kru96, SvA^{+}98, Swa99r, Tic97]. sidebar [Abr94, BT93, Bet95a, Clo94, Con94b, Dah93, Gil95a, Gil95b, Hel93, Hov94, Lin94d, Mis96, Moo93, Sym94, Tre94, Val94c, Wal95a]. SIGNA [JJ94]. Signal [McG94b]. Signatures [Boy98c]. Signed [Ber97e, Ber98a, Bri99]. significance [Kim92]. Signs [Eri97b]. Silicon [DDJ98g, Val95k, Woe92a]. Silicon-On-Insulator [DDJ98g]. Simple [DDJ98g, GK94, Har95c, Bet91a, PS97b, WW99, Tha97]. simplifies [Smi97]. Simplifying [Nee99, Pom98, Sch95a, Wil95b]. Simulated [DDJ99d, Rec99, TKG^{+}99]. Simulating [FSW99, HO92, HO93, Swa97y, Wic99]. Simulation [Cra99, Hof95a, Kut92, QG97, Rez94a, Rez94b, Val94c, Var93a, Var95a, Var93c, Var95b, vdW96, Swa97y, Var91, WJ92]. Simulations [IR99]. SimulEdit [TSR98]. Sincerest [Swa97l]. Single [Cro95b, Mae92]. Single-Image [Cro95b]. sins [Kis96]. site [Sto97]. Sixth [Ste99h]. Size [Col96e]. Sizes [PKA^{+}98]. Sizing [Val92e, Val92f]. Skip [Sch94l]. SLA [Koo94b]. Slippery [Wal95b]. Slow [Bar99d, Nac97]. slugging [Dun93d]. Small [Eri99i, Har96b, Nil98, Val94f, Wil99e, Yue97, Ste97f, Ano96b, Hov94, NRG^{+}99, RB97, Rie99, Yue96a]. Small-Footprint [Nil98]. Smaller [DDJ98a, DDJ98a]. Smallest [Bar99d]. Smalltalk [Alm96, Aye90b, Aye90a, Aye91, DM91]. Smalltalk/V [Aye90b, Aye90a]. Smart [Con98, Con99, DDJ98e, DDJ98d, DDJ98b, DDJ99f, DDJ99d, Mey99, Row97].

SMARTDrive [Cha92]. SmartSockets [Cam96]. SMB [Len97]. smbfs [Len97]. SMIL [Bar99c]. Smith [Swa97g]. Smoothing [Sta98a]. smoothly [vdW97b]. Snippets [Sto92a, Sto92b, Sto92c, Sto92d, Sto93]. SNMP [Sta94a]. Snooper [Jam93]. Snow [DDJ98e]. Sockets [CPRB93]. SoftRAM [RCS96]. Software [Aha95, Bak94, BEK99, Bar99d, Bec94, Bra94a, Bro95b, Bro95d, Bro95e, Bro96b, Bro96c, Cha95d, Che96, Cop94, DDJ98e, DDJ99f, Die93, Die99, DL98, Eri91a, FV97, Flo91b, HG95, Hug95, JL98, Keh94, KP99, The90, Lak97, McC91, Mil94, MMB^{+}90, Pfl98, Rod95b, Sch94n, Val91a, Val94g, Van94, Wil94e, Wil97j, Wil97f, Col97c, Coo96, Fra91, Gar91, Gil94b, Gil94c, Gil94a, Gol90, Hag91, Har91, Hol97, JJ91o, Jon98, Kap91, Kis96, Kno93a, Law97, NK98, Ros97, Row97, Smi97, Ver93, GHJV96, ML97]. Solid [Fal98]. Solomon [Lyp97b]. solutions [Wea97]. Solver [Med94]. Solving [GR98, Nay93, Man99a]. SOM [OH95]. Some [Swa97i, Abr92a]. Someone [Bro96c]. Something [Abr94, SvA^{+}98]. Sometimes [Swa96y, Bak97b]. Son [DDJ98g]. Sophisticated [Arn98, Nel93a]. sort [Mau93, Ste90a]. Sorting [BS98a, BT93, Boy98b, DDJ99e, PS97a, Tho91b]. Sound [Coo98a, Cro94a, MGL^{+}98, Rat95, Wal95a, Wal92b, Yor95]. Sounding [Dur97]. Source [Boo92, Mar94b, van98a, MK97, Vog91, GF99, NRG^{+}99, Ste99f]. Source-Code [Mar94b, van98a]. Space [Gut99]. Space-Filling [Gut99]. Spam [Swa99t]. SPARCstation [BL95]. Spare [Eri98b]. Spark [Kot99]. Spatial [LR92]. Spec [Con94b, DDJ99c, DDJ98a]. Special [Dix95, Mel98]. Special-Effect [Dix95]. Special-Purpose [Mel98]. specific [Ste92p]. Specification [Hub94, Mon98, RT97, WB95, Haa95, JJ91o]. Specifications [CM95]. Specifying [Ber99].

Speech [Dav99, Deg94, Eri97b, LMWR^{+}96, Ram97a, Row94b]. Speech-Enabling [Ram97a]. Speed [JJ94, Kim96d, Hof97, Jol92b, Let91, NK98]. Speeding [DDJ99f, Mud96, Jam92, Reg90]. Speedy [Ton91]. SPICE [Wri99]. SPIFF [Mur96]. Spin [Hol97]. Splay [Cla92]. Spline [Cou96]. Splines [LP90]. splitting [Let96]. spotting [Sto97]. spreadsheets [Swa97d]. Spring [DDJ98d]. Sprites [Gru95]. Spying [Bur93]. SQL [Ber97b, Bet94d, Lam99, Nor93, Nor96, Rei93b, Sip96d]. Square [DKR97, Hei96]. SSH [Sar97]. Stack [Van97]. stacks [Fre92]. staff [Ano96j, Ano97f, Ano97g, Ano97h, Ano97i]. Stagnation [Swa96x]. Stalking [Sch90h, Sch90i]. standalone [JJ91u]. Standard [Aus97, DAM96, Ste98g, Vil95, Hyd92, Ste97f, Tea91, Aus97, DDJ98a, Jae91, Jae95, Kef95, Sak92, Ste98c, Ste98i]. Standardization [PJR^{+}98, Jae93]. standardize [Tha97]. Standards [Cal95c, DDJ98b, Jol93b, Jol93a]. Stanford [Kae94]. Stanley [vdWS97]. Starbuch [McC96b]. Startup [Val95k]. StarView [Fir93b]. State [Car97c, BW91, Dru94, Hof97]. static [MK97]. Station [Smi98]. Statistical [Mot91, Nel91a]. Statistics [Aha95]. Status [Jae91, Rea91, Sak92, Woe91a, SW95a]. Steeps [Sha99f]. Steganography [Joh97b]. Step [Val94f, JJ92d]. Stepping [Koo94b]. Stereograms [Cro95b]. Stereoscopic [DR93, Ram97b]. Steve [Swa99c]. Still [Fos92, Mur96]. Stinkin [Eri96l]. STL [LS99, Mye97, Nel96c, Rot99, Sha97, Sha98a, Zig95, Zig96a, Zig96b]. Stocking [Swa97j, Swa98i]. Stone [DDJ98h]. storage [Gar91]. Store [AW99, HL99]. Stored [Nor99]. Story [Abr95d, DDJ98i]. Strange [Eri94l]. Strategies [Bro94a, GR98, Hes93]. Strategy [Kum95, Kie96b, PS97b]. stream [Ste97g]. Streaming [Moo98, Sae95a]. streams [Jam92]. String [Hil95, Nel96a, Tea91, Tho91b, Val92d]. Strings [BS98a]. Stripped [JJ91p]. Stripped-Down [JJ91p]. Stronger [SvA^{+}98]. Stroustrup [Ste92q, Ste92-32]. structure [Sha97]. Structured [Dun90a, Dun90b, Dun90c, Dun90d, Dun90e, Dun90f, Dun90g, Dun90h, Dun90i, Dun90j, Dun90k, Dun90l, Dun91b, Dun91c, Dun91d, Dun91e, Dun91f, Dun91g, Dun91h, Dun91i, Dun91j, Dun91k, Dun91l, Dun91m, Dun92c, Dun92d, Dun92e, Dun92f, Dun92g, Dun92h, Dun92i, Dun92j, Dun92k, Dun92l, Dun92m, Dun92n, Dun93e, Dun93f, Dun93g, Dun93h]. Structures [Boy98a, Lia90, Sou99, Bat94]. study [RK92]. Stuff [Eri98i, Har95e]. Stuffers [Swa97j, Swa98i]. Style [AH97, Wil90b, Chu92a, Ozi91, Wil94c]. styles [Bak97b]. sub [Mir92]. Subclassing [Kle91c, LR96]. subobjects [Mye97]. subroutines [Zha97]. subtree [Boy97]. Subway [Sha98h]. Success [Bar99d]. Suffix [Nel96a]. Suite [Bha98, Oro98]. Suits [DDJ99c]. Sultan [Sha99l]. Summer [Swa99m]. Sun [BL95, Swa98j]. Super [Eri96l, How90]. Supercharging [Wil90e]. Superclassing [LR96]. Supercomputers [DDJ98e]. Supercomputing [Fri94, Hir92b, Hir92a, Hir92c]. Superdistribution [Cox92]. Supply [DDJ99b]. Support [Eas99, Ran99a, Ste97k, Bak97b, Ber96a, JJ91q, Sto97]. Supported [Sch90d]. Supporting [But99]. Surface [Abr96b, Abr96c, SL96b, SL96a]. Surfaces [Lus97]. Surfeit [Val95l]. surprises [Sha93d]. surprising [Mae92]. Swaine [Swa90m, Swa90n, Swa90o, Swa90p, Swa90q, Swa90r, Swa90s, Swa90t, Swa90u, Swa90v, Swa90w, Swa90x, Swa91n, Swa91o, Swa91p, Swa91q, Swa91r, Swa91s, Swa91t, Swa91u, Swa91v, Swa91w, Swa91x, Swa91y, Swa92m, Swa92n, Swa92o, Swa92p, Swa92q, Swa92r, Swa92s, Swa92t, Swa92u, Swa92v, Swa92w, Swa92x, Swa92y, Swa93n, Swa93o, Swa93p, Swa93q, Swa93r, Swa93s, Swa93t, Swa93u, Swa93v, Swa93w, Swa93x, Swa94n, Swa94o, Swa94p, Swa94q, Swa94r, Swa94s, Swa94t, Swa94u, Swa94v, Swa94x, Swa94w, Swa94y, Swa95o, Swa95q, Swa95p, Swa95r, Swa95s, Swa95t, Swa95u, Swa95v, Swa95w, Swa95x, Swa95y, Swa95z, Swa96x, Swa96m, Swa96n, Swa96o, Swa96p, Swa96q, Swa96r, Swa96s, Swa96t, Swa96u, Swa96v, Swa96w, Swa97t, Swa97u, Swa97v, Swa97w, Swa97m, Swa97n, Swa97o, Swa97p, Swa97q, Swa97r, Swa97s, Swa98q, Swa98r, Swa98p, Swa98s, Swa98n, Swa98t]. Swaine [Swa98u, Swa98v, Swa98w, Swa98o, Swa98m, Swa99k, Swa99q, Swa99l, Swa99r, Swa99s, Swa99m, Swa99n, Swa99o, Swa99t, Swa99p, Swa99u, WL97]. Swapping [BB90]. Swedes [Sha99e]. SWIG [Bea98]. Swing [Soo98a]. Symantec [Ano96c, Hem96a, Rod95a, Yam94]. symbol [SW95a]. Symbolic [KS97, PJR^{+}98]. Symbols [Hla95a]. Symera [FK98]. Symmetric [NV94]. Synchronization [Chr99, For98a]. Synthesis [Row94b]. synthesizer [TWH93]. Sysadmin [Tay99]. System [Bet94a, Bur94, RCS91a, Col97e, Cro94a, Esp97, Fag98b, GE98, Ges90, Kau95, Kut92, TL95a, Lu94, MGL^{+}98, Mar94a, Mel98, Nog99, Obl90, Pie93b, Pud98, RK92, RC97b, Sch99d, Tor95, Var94e, Wei96b, Woo99, Yam96, Ano96a, Car97c, Dic95a, Dic95b, Gar91, Hav98b, Hig97, JJ91u, NV90, Ohl90, Pom98, TL95b, Wea97, Wei96a, Col97d, Col97f, RCS91b, RC95, Ste91v, Str90, Swa97y]. System-call [RC97b]. Systems [Bar99e, Bek97a, Blu90, BLE96, CS94, CP99, DDJ98a, Dru95, Fly97, FSW99, GS97, Gis93, HQ98, Ise95, JS94, Kau96b, Kin91b, Kre99, Krt98, Kry92a, McR96b, Mel98, MC98, New93a, PB98a, Rat94, Ric98, Rob98b, Rob98a, Rou98, Tra90, Tro94, TMH99, Var93b, Var93e, Yue97, Cro93a, Dem93, Dic95b, Dru94, DM91, Jen95, Kry92b, McR96a, Per93, PR91, SJB92, Ste93b, Tha97, Esp97, Vil96]. table [Yam94]. tables [BW91]. Take [Pea95, PSS^{+}98b]. takes [Wal97]. Taking [Ams99, And91, Ban98a, Eri98j]. Talent [DDJ98f]. Taming [TMP^{+}99]. Tapping [Eri94m]. TARGA [Bor96]. Task [Kre99]. Tasks [Fin95a]. Taste [Bet92]. TAWK [Law97]. Taxman [DDJ98e]. Tcl [Arn98, Bek97b, DDJ98e, Gra98, Gra97, McG98b, Ous94, Sta98c, Uhl99]. Tcl/Tk [Bek97b, McG98b]. Tclblend [Sta98c]. TCP [Bur95b, WV96]. TCP/IP [WV96, Bur95b]. Tea [DDJ99a]. Teaching [Ste99h]. TeamworX [Ste97j]. Tecate [Koc97]. Tech [Eri94a, Har95d, Ste97k]. Technical [Val91c, Car97c]. Technique [Tes96a]. Techniques [App97, Pea95, Sym94, Man99a]. Technology [DDJ98e, Fom99, Laf99, Gar91]. Technomadic [Rob98b]. Telephony [deC99]. Tell [Kli91]. Template [MSE^{+}99, Aus97, HCKS99, How98, Joh97a, Kef95, SW97f, Ste98j, VP96]. Template-Driver [MSE^{+}99]. Templates [BR93, Rei93d, Ste99i, Wil92d, But93, Joh97a, Wil96a]. Temporal [DV99]. Tempus [Har96c]. Ten [Fos92, Ros97]. Terminal [Con98, Con99, Tro91b]. terminals [Tro91a, Tro91b]. Terminology [QTC^{+}99]. Ternary [BS98b, PKA^{+}98]. Terrain [Krt94]. Territory [Sha98i]. Terse [Kau95]. Tessellation [LR92]. Test [Ran99b, Ang99, CT90, Wal97]. Testability [Wil97b]. Testboard [Kut92]. Tested [BEK99]. Testers [vdW97b]. Testing [Bak94, Bet91b, DDJ99f, HHW^{+}99, Ise95, Kep94, McC97, Nor96, Sel99, SDB^{+}99, vdW97b, Wal97, Bra94a]. tests [McC97, Ste97g]. Text [Gat97, Kie99c, Krt95, Mon97, TMP^{+}99, Val93f, Mir92, Ste90l, TSR98]. TEXTSRCH [Ste90l]. Texture [Spi94]. TGI [Mon96]. Theatrix [Ste95n]. Theme [DDJ99g]. Theorem [LHB^{+}98, Sav98]. There [BCP^{+}99, Eri97m]. these [Car97c].

Thielen [One94d]. Thin [Tou99]. Thing [BCP^{+}99, Val92a]. Things [DDJ99e, WW99]. Think [Mül94, Sch94m]. Thinking [BEK99, DDJ98h]. Thomas [ML97]. thorniest [Nel93a]. those [vdW97b]. though [Bak97b, Len97]. Thoughts [Abr96a]. thousand [Ste90l]. Thread [Cal97, Pan99, Rod95d, TKG^{+}99]. Threads [Ber96f, Cri98, For95, Lee93b, Wil94c, Yam95, Yam96, Kan94, Pan99]. Three [BS98a, Car96b, Str90, JJ91v]. Three-Dimensional [Str90]. Three-tier [Car96b]. Three-Way [BS98a]. Threshold [Sti98]. Thumbs [Eri98d]. Thunking [Can95, Pie96]. TI [May95]. tier [Car96b]. Tightened [Sch94n]. Time [Ada92, BW98, Bek97a, BBB94, Bun94, Coo98a, DDJ99d, Eri94p, Eri97n, FV93c, HT98, Jam90, Kar90, Kau95, Kim99i, Kot99, Krt98, MGL^{+}98, MM98b, McR96b, ME96, New93a, Nol90, PB98a, Pot92, QTC^{+}99, RJ95, SG97, Smi98, Ste98a, Ste99h, Swa99p, Tet94, TMH99, Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Arn98, Bar99a, Bev98, Fer94, Gal91, GWL96, GGL^{+}95, Jen95, Lus97, McR96a, Nel93a, Nel93b, Roa91, Sha93a, SJB92, Smi97, Ste97e, Wil96h, Pro99]. Timed [Sta92]. Timing [DDJ99b, McR95, Rod92]. Tiny [Bet94b, Kau95, Yue96b, Bet91c]. Tk [Bek97b, McG98b]. TLT30G [Kis98]. Today [Bro95d, Val94b]. Token [Yue94]. Token-Ring [Yue94]. Tony [Zha97]. Too [Wil99c, Mye97]. Tool [Bha98, Hig98, Sto97, Col97c, HQ98, Hig97, Law97, Nel93a, RC97b, Smi97, Ste91w, Swa96y, Tre98, Wea97]. Toolbox [Bat96, Gas90]. Toolchest [Ang99, CKS98, Hed98, HT98, Joh99, Lis98, McB97, Rou98, Sch99b, She99, dSW98, Soo98b, Wea97, van98a]. ToolHelp [Sax92b]. ToolKit [Vil96, Cam96, Fir93a, Lai91, Fra99, Lin94b, Nic91a, Sar94, Bug98, Hem96b, One94a, Sch96i]. Toolkits [Ten98, Val92f]. Tools [Arm95, DDJ99f, Lis98, Old94, Per93, Var94b, Wit95b, Fra90, JJ91q, MK97, vdW97b]. Tools.h [Sch96c]. Topological [PS97a]. TopSpeed [Lan90]. Topview [Mau93]. Torvalds [Flo94d, Li94b]. Tough [DDJ99g]. TPU [Soj96]. Trace [Ger93, Sto97]. tracebacks [Dun92b]. Tracer [Rus99]. Tracing [Fag98b, Lyk90, Man94b, Zud96, Lin94b]. Track [BEK99]. Trackball [Swa99n]. Tracking [BAC^{+}99, Har99, Lyn95, Tro98, Kno93a]. Trade [Pat90a]. Trademark [Bro95f, Eri97b]. Trail [Eri98e]. training [GGL^{+}95]. Trains [Sha99l]. Transactions [Tou99]. Transfer [Ber98d]. transform [Cod92, Cod94b, Nel96b]. Translator [Smy97]. Transparent [AN98]. Transport [Sar97, Cal96]. transport-independent [Cal96]. Transportability [Dow93]. Traveling [Eri97n]. Treaps [Nil97]. Tree [Abr95a, Bat94, Mud96]. Trees [BS98b, Col96a, Dwy95, Höw99, Nel96a, Sch99a, Sch99c, Sha98a, Ste97a, Bet93b, Boy97, Cla92, JS92, Sch92c]. Trends [DDJ98f, Gut92, Var94e]. trick [Hof97]. Tricks [Pat90a]. tries [NK98]. Triumphant [Har96a]. Trouble [ML97]. troublesome [MK97]. TrueType [Rei93a]. Truly [Plu94]. truncation [Kru92]. Truth [Bar99c]. Tuning [Ben99b, Ell93a, Ell93b, Tym96, Dun93d, Sha93d]. Tunnel [DDJ99b]. Turbo [Eck90, Flo92, FNL91b, FNL91a, Gin91, Lad90c, Nic92, Pea91, Tho92]. Turkey [NRG^{+}99]. Turn [Eri98l, Eri99e]. Tutor [DDJ99g]. Tutorial [Mae92]. TWAIN [Lin94a]. Two [Abr95d, Eri98d, Sem96, Kis97]. Two-Dimensional [Sem96]. Twofish [Sch98b]. Type [Wal92b].

U.S. [Aha95]. Ubiquitous [DDJ99d]. UC [DDJ99c, SNH93]. UCR [Hyd92]. UDFs [Nor99]. UDP [Red95b]. UI [AH91, Gat97, Med94, Sip96c]. UML [DV99, MM98b, Wil98c]. underlying [Tha97]. Understanding [Col96e, Hus99, Kie97b, Mor91d, Sta93b, Ser93]. Undo [Bev96, Ste98k, Van97]. Undo/Redo [Bev96, Ste98k, Van97]. Undocumented [Chu92b, Col96b, Col96c, Col96d, Col96e, Col97e, Col97c, Col97d, Col97b, Col97f, Col98a, Col98d, Col98e, Fol94, Ham95b, Sch96h, Sch93k, SNH93, Sch93e, Sch93f, Sch93g, Sch93h, Sch93l, Sch93i, Sch93j, Sch94w, Sch94v, SRA^{+}94, Sch94o, Sch94p, Sch94q, Sch94r, Sch94s, Sch94t, Sch94u, Sch95p, She93, SW96a, SW96b, SW97a, SW97f, SW97c, SW97b, SW97d, SW97e, SW98c, SW98b, SW98a, WS96b, WS96c, WS96d]. unexpected [McC97]. Unicode [FV97, Nes94, Van94]. Unification [PH95b, PH95a]. Unit [McC97]. units [McC97]. Universal [Man92]. University [Smy97]. UNIX [CJ94, Dun92b, Fra99, Gal91, Gla91, HO93, JL98, JJ91s, JJ91b, JJ91c, JJ91a, JJ91t, JJ91q, JJ91d, JJ91o, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ91r, JJ91p, JJ91v, JJ92e, JJ92a, JJ92b, JJ92d, JJ92g, JJ92h, JJ92i, JJ92j, JJ92k, JJ92l, Len97, Rea91, Sha93b, Wel94, Bau90a, HO92, JJ91u, JJ92m, JJ92c, Lu94, Rea90, Ric91]. UnixWare [Rod94c, Rod95d]. Unlimited [Van97]. unpredictability [Sch99a]. unravel [Nic91a]. Unraveling [Sch90b]. Unscrambling [DDJ98c]. Unsharp [Kie99a]. Untangling [CPRB93, Cha92, Law90, Sch92d]. Untrusted [Gri99]. Update [Cav99, Jae95, LHB^{+}98, MGL^{+}98, DDJ99a, SWF^{+}99, TMP^{+}99, Jae93, TFS94]. updates [Ste97f]. Updating [Yam94]. Upgrading [Bro96c]. ups [McR93]. URL [BAC^{+}99]. Usability [Jol96]. Usage [Tro98]. Use [Ber97j, Pit98a, Bak97b, Ber97i, Gol90, Mye97, QG97]. used [Nel97]. useful [Sto97]. Usenet [Krt96]. USENIX [DDJ98h]. User [Bra92, IR99, KR91, Ram97a, RC97a, Sch96a]. user-definable [Sch96a]. Users [Ste97f]. uses [DKR97, RC97b, Swa97y]. Using [BD97, BH95, Car92, Con94a, Dav95a, Dwy95, DR97, Fag98a, Fra99, Gad96, Iye95, Kis98, Kru96, Lin94a, McC96a, NN97, Nol90, One92, Roa91, RZ96, Rog95b, Sta95b, Str90, Tay98, Tes96a, Tes96b, Tuc99, Yor95, Ano96i, Bar92, BW91, Bet93b, Bra94c, But93, Cla96a, Das93, Duf95, Gre92, Hig97, LaP95, Nel92, Nel97, Nic91b, Rei93d, Row97, Swa96y, Tho91b]. usually [Boy97]. utilities [JJ91v]. Utility [Cav99, Pit98b, Rea91, Ser93, Osw93]. UX [Yam96].

V [Aye90b, Aye90a]. v86 [Col98d]. Valley [Swa91a, Val95k]. Vaporware [DDJ98f]. Variables [Low95]. variably [Ros93]. Variation [DDJ99g]. VB [Wil98c]. VBE [Haa95]. VBSCRIPT [Lam99]. VBX [Red95b]. VCR [Row97]. Vectors [Blu95]. VerCheck [GC98, MGL^{+}98]. Verification [Cro93b, Cro93a, Nel92]. Version [GC98, HHW^{+}99, Hav99a, NRG^{+}99, WEM^{+}99]. versions [Hav98b]. Versus [Ben96, MGL^{+}98, Vel97, Bar99d, BL95, Bek97b, CS95, Yam96]. Very [JJ94, Lin94c]. VESA [BH95, Cod94a, Eri90m, Haa95]. VGA [Eri90m, Abr90b, Bet93b, How90, Mye91b]. VHDL [Tom96]. Via [McG94b, Boy97, Hig97, MK97]. Video [Flo94a, Kel95, May95, Row97, Wil94e, Gre92, Tho92]. View [Shi92, Swa99i]. Viewer [Bor96, Chu91, Flo93b, Ste96m]. Views [Ano96j, Ano97g, Ano97h, Ano97i, Bar99c, BEK99, Bar99a, Bar99d, Ber90, DDJ97c, DDJ97a, DDJ97d, DDJ98c, DDJ98e, DDJ98f, The98b, DDJ98h, DDJ98g, DDJ98a, DDJ98d, The98a, DDJ98b, DDJ98i, DDJ99b, DDJ99f, DDJ99e, DDJ99g, DDJ99c, DDJ99d, Eri97b, Ros95a, DDJ99a, DDJ97b, Gol90, Ano97f, Hed98]. village [Jon98]. Vintage [DDJ98h]. Virtual [Bri90, DDJ99f, Fos95b, One94d, Rei93d, Sch94w, Sch94v, Wil93c, Zud96, Zyt94a, Zyt94b, Hir92c, Col98a]. Virtual-array [Rei93d]. Visas [DDJ99b]. Visible [Abr96b, Ste91w]. Visible-Surface [Abr96b]. VisiCorpse [Swa97g]. Vision [Nic92]. Visions [Woe94b, Woe94a]. Visitor [Ngu96, Pit98b]. Visual [Amb99, FV93b, Hem96a, JLC98, JK95, Laf99, Naj95, PKA^{+}98, Sme98, Ste97h, Sti98, Var93a, Var93c, Var94b, Cul94, Esp96, Flo95f, Gol96, LaO98, LaP95, Lic92, Lom98, Nor94, Rob97, Sar95, Sch94x, dS95, Ste91w, Wen99, Wil97d, Xia95, da 95]. Visualization [WV91a, Var92, WV91b]. Visualizer [BW98]. Visualizing [FV93c]. Visually [Dru95, Wil95c]. Viterbi [Lyp97a]. VM [Wil99b]. VME [Col98a, Col98e]. Voice [Dav98]. Volker [Len97]. Volume [Och94b]. Voronoi [LR92]. VRML [NAM96, Ste95o]. VT100 [Tro91b]. VTOOLSD [Mir95]. Vu [Bar99c]. VWinL [Wil93c]. VXD [Mir95, RC96a, RC96b]. VXDs [Mat95, Bro96a, Shm94].

wait [Nac97]. Waiting [Val94d]. Walking [Rot90]. Wall [Tre94, DDJ99f, Kim98a]. walls [GG95a]. Wanted [Eri98g, PKA^{+}98]. War [Swa97w, Swa98e]. Warp [Kim96d]. Warren [GHJV96]. Wars [Bro95f, Col96c, Val94a, Val95m, DAM96]. Was [Eri97m, Tha97]. Washington [DDJ98f, Eri97g]. Watchdogs [Oes97]. Watcom [Sch90f]. Wave [LMWR^{+}96]. Waveform [Row94a]. Wavelet [Cod93, Cod92, Cod94b]. Wavelets [Pig99]. Way [BS98a, Hel98, Sch91c, WW99]. Wayner [vdWS97]. WDM [Ale99b]. Weapon [DDJ98h]. Weapons [Swa98e]. Weather [Smi98, Wic99]. Web [MSE^{+}99, Bar99d, BW96, Cal96, DDJ98e, DDJ98g, DAM96, Dut96, DR97, Gol96, Hig97, Hig98, Hou99, Joh97a, Law95, Lor96, MR99, McA94, MC98, Rem97, Sel99, Sho96, Ste95o, Tes96a, Tes96b, Tic97, Val95m]. Web-Based [MC98]. Web-Executable [Sho96]. Webl [MR99]. Week [Lar95, DDJ99c]. Weighing [KR96]. Wer [Wil98d]. were [Aus97]. Wesley [Cha95d]. West [DDJ98c]. WFC [Joh99]. Wheel [Fos92]. wheelchair [McR93]. Wheeler [Nel96b]. Where [DDJ98c, DDJ98d, Har94b, Val95b]. while [SC97, vdWS97]. Who [DDJ98d, Eri94m, Eri96l]. Whose [Swa99u]. Wide [Dut96, Law95, McA94, Ste95o]. Widen [Val95m]. Widgets [Hus92, Aye91]. WIDL [Mon98]. Will [Bar99a, DDJ98d, DDJ99g, Kef92]. William [vdWS97, Woe97c]. Win32 [Rad98, For98a, Han93, Ler97, Low98b, QTC^{+}99, Sch97b, Ste97k, Tew97, vdW97a, van98a]. WinASPI [Tew97]. Winbatch [Sch96i]. WinCE [MSE^{+}99, QTC^{+}99]. Window [Ana94, Hem96b, PKA^{+}98, RCS91b, Ste91v, Str90, Wil93c, RCS91a]. windowing [Kie96b]. Windows [Flo94h, One94d, Sch93k, Sch93d, Sch93l, Sch94w, Sch94v, SRA^{+}94, Ana95, Bak93, Bar99d, BW91, Boy98c, Bra94b, Bri93, Bro96a, CPRB93, Can95, Chu91, Chu92b, Cog93, Con91, Cou94b, Dav93a, Dav93b, Del97, Die99, Esp97, FR93, Fin95b, Flo91d, Flo92, Fos95b, FY96, Ger93, Goa99, Hav98b, Hav99b, Hig97, Hla94, Hla95a, Hla95b, Hla95c, Hov94, Kel95, Ken94, Kle91a, Kle91b, Kle91c, Kno94, Kno90, Lad91, Lee93b, Let96, Lic92, Lin94c, LR96, Low95, LHB^{+}98, Mat95, Mir92, Mül94, Mus93, Mye91a, New93c, New94a, Nor91, Oka94, One92, One94a, One94b, PF90a, PF90b, Pie92a, Pie93a, Pie93b, Pie96, Pro92, QG97, Rad98, Rob96, Rob98a, Rol92, Ros95a, RC95, RC96c, RC97a, RC97b, Sal92, Sax92a, Sax92b, Sch98a, Sch92a, Sch96h, Sch97a, Sch99d, Sch93c]. Windows [SNH93, Sch94m, Sch95p, Sha93a, She93, Shi94, Sip96b, Sip96a, Sip97, Ste91w, Ste93w, Ste94p, Ste97f, Ste98l, Ten98, Tro97, Tro98, Tuc98, Vai93, Val93b, Van93, Web96, Wil94b, Wil95a, Wil95b, Wil96g, Wil94d, WV96, Wil97e, Wil98b, Wil98e, Wil98f, Yam96, You92, Zyt94a, Zyt94b]. winds [FV93a]. WinG [Kel95]. WinHelp [Bur93, Kis96]. WinMock [Hla96b]. Winner [DDJ99g, QTC^{+}99]. Winners [Val94i]. WinSock [Duf95, BH96]. WINTHERE [Mye91a]. Wire [Mar95a]. Wireless [Die93, Man99b]. Wisdom [DDJ99a]. WiT [Höw99]. without [Len97, Mus99, Mye97, Rei93b, Smi94]. Wizdom [Val95i]. WNet [Low98a]. Woes [DDJ98i, Ver97]. wonders [Ste97e]. Woodman [Wil97g]. Woodruff [One94d]. Wookie [ABCE99]. Word [DDJ99g]. Wordsworth [Cha95d]. Work [Ber96f, MGL^{+}98, Pur98, KS97, Wil97g]. workaround [Len97]. Working [DDJ99d, Rec99]. workings [RC97a]. Workplace [Gig93b, Gig93a, Bar93a]. works [Hig97]. Workshop [Sch94g]. Workstation [Par95a]. Workstations [Bau93]. World [Abr95b, Bar99d, Law95, Nic93a, Och94c, Rus93, Ste98e, Ste95o, Val95a, Oli91, vdW97b, Dut96, McA94]. Worlds [Pic93, NAM96]. WorldToolKit [Fos95b]. Worry [Flo91d]. wrap [Val93a]. wrap-up [Val93a]. WRAPI [Sar94]. wrapper [Ste95l]. Writable [Woe92b]. Write [Ber96e, Ber97l, Ber97k, Gar91]. write-once [Gar91]. Writers [DDJ99f]. Writing [Bri93, Cal96, Kin99, Kru94b, Li94c, McC95, Mil94, One94d, Sch96a, Shi99, Tew97, Van93, Wel94, Wil97c, Kis96, QG97]. Written [Zha97, Ver96, Wal97]. Wrong [SDB^{+}99, Val92a]. WSNETWRK [Clo94].

X [Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s, Bau92a, RCS91a, Lai91, RCS91b, Str90]. X-arRAY [Bau92a]. X-Sharp [Abr92l, Abr92m, Abr92n, Abr92o, Abr92p, Abr92q, Abr92r, Abr92s]. X-Window [RCS91a]. X.509 [Tre99c]. x86 [Zha97]. Xalloc [Gin91]. Xchg [Ste97g]. XDE [Fra99]. XFDL [Boy99]. XLink [McG98c]. Xlisp [Swa97y]. XML [Bar99c, Ham99, McG98a, McG98c]. XYModem [Kie94b].

Y2K [BCP^{+}99, DDJ98h, DDJ98b, DDJ99g, Eri97b, Eri98e, GR98, MGL^{+}98, PSS^{+}98b, Ste99j, Swa98o, TKG^{+}99]. Yacc [Gor96, Kin91b, Sha99a]. Year [MBK^{+}98, Sch95k, Swa99j, MSE^{+}99, Bha98, MF98]. Years [Fos92]. York [Smy97].

Z [Cha95d]. Z80180 [Kal93]. Zen [Gul95, Tra90]. Zero [Bur95b]. Zero-copy [Bur95b]. Zinc [Ent90a]. Zip [NN97]. zlib [Nel97, Nel97]. Zortech [Lad90a]. Zytaruk [Sch94w, Sch94v].

References

Atkinson-Barr:1997:LEP

[AB97] Martin Atkinson-Barr. Letter to the Editor: Pentium II math bug. Dr. Dobb’s Journal of Software Tools, 22(10):10, October 1997. CODEN DDJOEB. ISSN 1044-789X. Identifies himself as the “Mr. X” cited in [Col97a], and provides more the background on the discovery of the Pentium FIST (floating-point to integer store) instruction.

Anderson:1999:WE

[ABCE99] Kalle Anderson, Jason Buttron, Paul Clarke, and Matt Enwald. Wookie: A 68HC11 emulator. Dr. Dobb’s Journal of Software Tools, 24(3):50, 52–55, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/thrsim11.
zip;   http://www.ddj.com/
ftp/1999/1999_03/wookcode.
zip;   http://www.ddj.com/
ftp/1999/1999_03/wookdemo.
zip;   http://www.ddj.com/
ftp/1999/1999_03/wookexe.
zip;   http://www.ddj.com/
ftp/1999/1999_03/wookie.txt.

Abrash:1990:ALL

[Abr90a] Michael Abrash. Assembly language lives! Dr. Dobb’s Journal of Software Tools, 15(3):16– 26, 94, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1990:DV

[Abr90b] Michael Abrash. Demystifying 16-bit VGA. Dr. Dobb’s Journal of Software Tools, 15(5):70–81, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:O

[Abr91a] Michael Abrash. 80*86 optimization. Dr. Dobb’s Journal of Software Tools, 16(3):16–18, 20, 22, 24, 26, 73, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPa

[Abr91b] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(2): 153–??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPb

[Abr91c] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(3):129–??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPc

[Abr91d] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(4):133–??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPd

[Abr91e] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(5):131–??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPe

[Abr91f] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(6):139–??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPf

[Abr91g] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(7):133–??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPg

[Abr91h] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(8): 165–??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPh

[Abr91i] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(9): ??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPi

[Abr91j] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPj

[Abr91k] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(11): ??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1991:GPk

[Abr91l] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 16(12): ??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GGA

[Abr92a]  Michael Abrash.  Graphics

Gems, by Andrew Glassner [some notes on]. Dr. Dobb’s Journal of Software Tools, 17 (7):137–??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPa

[Abr92b] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPb

[Abr92c] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPc

[Abr92d] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPd

[Abr92e] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPe

[Abr92f] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPf

[Abr92g] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPg

[Abr92h] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(9):141– ??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPh

[Abr92i] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(10): 151–??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPi

[Abr92j] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(11):171– ??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:GPj

[Abr92k] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 17(12):143– ??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRa

[Abr92l] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRb

[Abr92m] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRc

[Abr92n] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRd

[Abr92o] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRe

[Abr92p] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRf

[Abr92q] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRg

[Abr92r] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(9):??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1992:SDRh

[Abr92s] Michael Abrash. ‘X-sharp’: 3-D real-time animation in GP. Dr. Dobb’s Journal of Software Tools, 17(10):??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1993:GPa

[Abr93a] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 18(1):

127–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1993:GPb

[Abr93b] Michael Abrash. Graphics programming. Dr. Dobb’s Journal of Software Tools, 18(2): 127–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1994:POS

[Abr94] Michael Abrash. Pentium optimization: Something old something new [sidebar]. Dr. Dobb’s Journal of Software Tools, 19 (1):40–??, January 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Abrash:1995:CBT

[Abr95a] Michael Abrash. Compiling a BSP tree. Dr. Dobb’s Journal of Software Tools, 20(??):49– ??, July/August 1995. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1995:DWC

[Abr95b] Michael Abrash. The day the world changed. Dr. Dobb’s Journal of Software Tools, 20(??):48– ??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1995:FR

[Abr95c] Michael Abrash. Frames of reference. Dr. Dobb’s Journal of Software Tools, 20(??):46–??, September/October 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Abrash:1995:OST

[Abr95d] Michael Abrash. One story two rules and a BSP renderer. Dr. Dobb’s Journal of Software Tools, 20(??):47–??, November/December 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Abrash:1996:DCO

[Abr96a] Michael Abrash. 3-D clipping and other thoughts. Dr. Dobb’s Journal of Software Tools, 21 (??):43–??, March/April 1996. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1996:IQV

[Abr96b] Michael Abrash. Inside quake:

Visible-surface determination. Dr. Dobb’s Journal of Software Tools, 21(??):41–??, January/February 1996. CODEN DDJOEB. ISSN 1044-789X.

Abrash:1996:QHS

[Abr96c]  Michael Abrash.  Quake’s

hidden-surface removal. Dr. Dobb’s Journal of Software Tools, 21(??):48–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Adams:1992:TDR

[Ada92] Edwards N. Adams, III. Time dilation and relativistic debugging. Dr. Dobb’s Journal of Software Tools, 17(11):80–81, 84–85, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Alden:1991:PU

[AH91] Roland Alden and Tony Hoeber.

The PenPoint UI. Dr. Dobb’s Journal of Software Tools, 16 (12):108–109, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Angell:1997:EMS

[AH97] David Angell and Brent Heslop. Elements of E-Mail style. Dr. Dobb’s Journal of Software Tools, 22(5):124–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Aharonian:1995:USP

[Aha95] Gregory Aharonian. 1994 U.S. software patent statistics. Dr. Dobb’s Developer Update, 2(4): 2–3, April 1995. CODEN ???? ISSN 1079-8595.

Aitken:1996:AGJ

[Ait96a] Gary Aitken. Automatically generating Java documentation: javadoc and the doc comment. Dr. Dobb’s Journal of Software Tools, 21(7):42, 44, 46–49, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Aitken:1996:MCJ

[Ait96b] Gary Aitken. Moving from C++ to Java. Dr. Dobb’s Journal of Software Tools, 21(3):52, 54–56, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Auping:1990:MMM

[AJ90] Judy V. Auping and J. C. Johnston. Making the move to Modula-2. Dr. Dobb’s Journal of Software Tools, 15(9):16– ??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Alexander:1999:ICR

[Ale99a] William F. Alexander. IEEE 1394 configuration ROM decoder. Dr. Dobb’s Journal of Software Tools, 24(8):66, 68, 70–73, August 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_08/
1394rom.txt;   http://www.
ddj.com/ftp/1999/1999_08/
1394rom.zip.

Alexander:1999:WIC

[Ale99b] William F. Alexander. A WDM IEEE 1394 configuration ROM decoder. Dr. Dobb’s Journal of Software Tools, 24(12): 64, 66, 68, 70–71, December 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
12/1394wdm.txt;   http://
www.ddj.com/ftp/1999/1999_
12/1394wdm.zip.

Allen:1990:POA

[All90] Norton T. Allen. A pixel ordering algorithm. Dr. Dobb’s Journal of Software Tools, 15 (6):56–61, 116–117, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Allburn:1991:GDI

[All91] Edward Allburn. Graph decomposition. imposing order on chaos. Dr. Dobb’s Journal of Software Tools, 16(1):88, 90–92, 94–96, 118–120, 122, 124, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Almarode:1996:PAS

[Alm96]  Jay Almarode.  Partition-

ing applications in Smalltalk. Dr. Dobb’s Journal of Software Tools, 21(??):23–??, January/February 1996. CODEN DDJOEB. ISSN 1044-789X.

Ambler:1999:FVP

[Amb99] Allen Ambler. The Formulate visual programming language. Dr. Dobb’s Journal of Software Tools, 24(8):21–22, 24, 26, 28, August 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL 9908b/9908b.htm;
                                         http://designlab.ukans.edu/
                                        ~ambler/ddj/;   http://www.
                                        ddj.com/ftp/1999/1999_08/
                                        formulate.txt.

Amsterdam:1999:PBT

[Ams99] Jonathan Amsterdam. Programmer’s bookshelf: Taking a risk. Dr. Dobb’s Journal of Software Tools, 24(3):125–126, March 1999. CODEN DDJOEB. ISSN 1044-789X.

Armstrong:1998:IPT

[AN98] Tom Armstrong and Mark Nelson. Internet programming: Transparent ATL controls. Dr. Dobb’s Journal of Software Tools, 23(3):72, 74– 76, 105–106, March 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_03/
                                        trans.txt;   http://www.ddj.
                                        com/ftp/1998/1998_03/trans.
                                        zip.

Anantharaman:1993:GEF

[Ana93] Vinod Anantharaman. A GUI environment for Fortran development. Dr. Dobb’s Journal of Software Tools, 18(6):104, 106– 108, 110–111, June 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Anantharaman:1994:CWB

[Ana94] Vinod Anantharaman. Customizing window behavior. Dr. Dobb’s Journal of Software Tools, 19(??):50–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Anantharaman:1995:WCC

[Ana95] Vinod Anantharaman. Windows 95 common controls. Dr. Dobb’s Journal of Software Tools, 20(5):52, 54–56, 58, 60–63, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1990:KOPa

[And90a] Brian R. Anderson. Kermit for OS/2: Part I. Dr. Dobb’s Journal of Software Tools, 15(9):70– ??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1990:KOPb

[And90b] Brian R. Anderson. Kermit for OS/2: Part II. Dr. Dobb’s Journal of Software Tools, 15(10): 72–??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Andrews:1991:TRC

[And91] R. Bradley Andrews. Taking up residence with Coderunner. Dr. Dobb’s Journal of Software Tools, 16(6):74, 78, 104, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1994:ILI

[And94a] Brian R. Anderson. Interfacing laboratory instruments. Dr. Dobb’s Journal of Software Tools, 19(13):62, 64, 66, 68, 97– 98, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Anderson:1994:MYP

[And94b] Scott Anderson. Morphing on your PC. Dr. Dobb’s Journal of Software Tools, 19(??):4– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Angell:1999:PTE

[Ang99] Kirby W. Angell. Programmer’s toolchest: Examining JPython: A Java test engine puts Python to the test. Dr. Dobb’s Journal of Software Tools, 24(4):78, 81–83, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
04/jpython.txt;   http://
www.ddj.com/ftp/1999/1999_
04/jpython.zip.

Anonymous:1991:DDC

[Ano91] Anonymous. The DDJ data compression contest. Dr. Dobb’s Journal of Software Tools, 16 (2):52–??, February 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Anonymous:1995:LYc

[Ano95a] Anonymous. Letters — by you. Dr. Dobb’s Journal of Software Tools, 20(10):10–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1995:LYa

[Ano95b] Anonymous. Letters — you.

Dr. Dobb’s Journal of Software Tools, 20(8):10–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1995:LYb

[Ano95c] Anonymous. Letters — you.

Dr. Dobb’s Journal of Software Tools, 20(9):8–??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1995:LYd

[Ano95d] Anonymous. Letters — you.

Dr. Dobb’s Journal of Software Tools, 20(11):10–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:BES

[Ano96a] Anonymous. Building an expert system. Dr. Dobb’s Journal of Software Tools, 21(8):??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:CSC

[Ano96b] Anonymous. Concurrent Small C. Dr. Dobb’s Journal of Software Tools, 21(8):??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:ESC

[Ano96c] Anonymous. Examining Symantec’s Café. Dr. Dobb’s Journal of Software Tools, 21(8):??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:LYa

[Ano96d] Anonymous. Letters — you. Dr. Dobb’s Journal of Software Tools, 21(4):10–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:LYb

[Ano96e] Anonymous. Letters — you. Dr. Dobb’s Journal of Software Tools, 21(8):8–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:LYc

[Ano96f] Anonymous. Letters — you. Dr. Dobb’s Journal of Software Tools, 21(9):8–??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:LYd

[Ano96g] Anonymous. Letters — you. Dr. Dobb’s Journal of Software Tools, 21(10):8–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:LYe

[Ano96h] Anonymous. Letters — you. Dr. Dobb’s Journal of Software Tools, 21(11):8–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:NCU

[Ano96i] Anonymous. Networking CD-ROMs using cached CD images. Dr. Dobb’s Journal of Software Tools, 21(8):??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:NVD

[Ano96j] Anonymous. News and views — DDJ staff. Dr. Dobb’s Journal of Software Tools, 21(11): 12–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:LYa

[Ano97a] Anonymous. Letters — by you.

Dr. Dobb’s Journal of Software Tools, 22(5):10–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:LYb

[Ano97b] Anonymous. Letters — by you.

Dr. Dobb’s Journal of Software Tools, 22(6):10–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:LYc

[Ano97c] Anonymous. Letters — by you. Dr. Dobb’s Journal of Software Tools, 22(8):10, 12, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:LYd

[Ano97d] Anonymous. Letters — by you.

Dr. Dobb’s Journal of Software Tools, 22(10):10, 12, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:LYe

[Ano97e] Anonymous. Letters — by you.

Dr. Dobb’s Journal of Software Tools, 22(12):10, 12, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:NVDa

[Ano97f]  Anonymous.  NEWS and

VIEWS — by the DDJ staff. Dr. Dobb’s Journal of Software Tools, 22(1):12–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:NVDb

[Ano97g] Anonymous. News and views — by the DDJ staff. Dr. Dobb’s Journal of Software Tools, 22 (2):16–??, February 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:NVDc

[Ano97h] Anonymous. News and views by the DDJ staff. Dr. Dobb’s Journal of Software Tools, 22 (7):16–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:NVDd

[Ano97i] Anonymous. News and views by the DDJ staff. Dr. Dobb’s Journal of Software Tools, 22 (8):16, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1998:La

[Ano98a] Anonymous. Letters. Dr. Dobb’s Journal of Software Tools, 23 (1):10, 12, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Anonymous:1998:Lb

[Ano98b] Anonymous. Letters. Dr. Dobb’s Journal of Software Tools, 23 (2):10, 12, February 1998.

Anonymous:1998:Lc

[Ano98c] Anonymous. Letters. Dr. Dobb’s Journal of Software Tools, 23 (5):10, 12, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1998:Ld

[Ano98d] Anonymous. Letters. Dr. Dobb’s Journal of Software Tools, 23 (9):10, 12, September 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Anonymous:1998:I

[Ano98e] Anonymous. Of interest. Dr.

Dobb’s Journal of Software Tools, 23(6):130–131, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Applegate:1994:RMM

[App94]  Arthur D. Applegate.  Re-

thinking memory management. Dr. Dobb’s Journal of Software Tools, 19(6):52, 54–55, 96–97, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Appel:1997:MCI

[App97]  Andrew W. Appel.  Mod-

ern Compiler Implementation in Java: Basic Techniques. Cambridge University Press, Cambridge, UK, February 1997. ISBN 0-521-58387-X (hardback), 0-521-58654-2 (paperback). x + 398 pp. LCCN QA76.73.J38A66 1997. US$74.95 (hardback), US$29.95 (paperback). See [Wil97i].

Armistead:1993:AAP

[Arm93] Mike Armistead. All about profilers. Dr. Dobb’s Journal of Software Tools, 18(11):24– ??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Armbruster:1995:CCT

[Arm95] Jeffrey L. Armbruster. Comparing CASE tools. Dr. Dobb’s Journal of Software Tools, 20 (6):76–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Arnaud:1996:DMP

[Arn96] Pierre Arnaud. Dynamic message passing in C++. Dr. Dobb’s Journal of Software Tools, 21 (8):46, 48, 50–51, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Arney:1998:CPT

[Arn98] Jonathan S. Arney. C preprocessing with Tcl: Sophisticated compile-time preprocessing. Dr. Dobb’s Journal of Software Tools, 23(8):46, 48– 49, 91, August 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1998/1998_08/cpdq.
txt;   http://www.ddj.com/
ftp/1998/1998_08/cpdq.zip.

Aslam:1998:PC

[Asl98] Taimur Aslam. Protocols for Ecommerce. Dr. Dobb’s Journal of Software Tools, 23(12):52, 54– 56, 58, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Austern:1997:SST

[Aus97] Matthew H. Austern. The SGI Standard Template Library: C++ programmers were introduced to generic programming in 1994, when Hewlett– Packard released the Standard Template Library. Dr. Dobb’s Journal of Software Tools, 22 (8):18–20, 22–24, 26, 27, 90, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Avitzur:1992:YOH

[Avi92] Ron Avitzur. Your own handprinting recognition engine. Dr. Dobb’s Journal of Software Tools, 17(4):32, 34, 36–37, 103, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Angel:1999:JQH

[AW99] Dave Angel and Andy Wilson.

Java Q&A: How do I store a Java app in a self-executing encrypted file? Dr. Dobb’s Journal of Software Tools, 24(2):115– 116, 118, 120–121, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jqa299.txt;
 http://www.ddj.com/ftp/1999/
1999_02/jqa299.zip.

Ayers:1990:MS

[Aye90a]  Kenneth E. Ayers.  Multi-

processing with Smalltalk/V. Dr. Dobb’s Journal of Software Tools, 15(5):82–90, 114, 118–120, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Ayers:1990:MPS

[Aye90b] Kenneth E. Ayers. The MVC paradigm in Smalltalk/V. Dr. Dobb’s Journal of Software Tools, 15(11):168, 170, 172–174, 175, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Ayers:1990:RCM

[Aye90c] Kenneth E. Ayers. Review of CX multiprocessing extension kit. Dr. Dobb’s Journal of Software Tools, 15(5):82–??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Ayers:1991:MSW

[Aye91]  Kenneth E. Ayers.  Mak-

ing Smalltalk with Widgets. Dr. Dobb’s Journal of Software Tools, 16(5):64, 66, 68–70, 98, 100–102, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHa

[BA96] Cliff Berg and Steve Alexander. Java Q&A: How do I display an image? Dr. Dobb’s Journal of Software Tools, 21(5):103, 117, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Babcock:1999:LCH

[BAC^{+}99] Dave Babcock, Peter T. Anderson, Norman S. Clerman, Ralph Hempel, Steve Grappone, Jim Cloughley, Kevin Johnson, John Bartley, and Cesar A. Gonzalez Perez. Letters: Computer history; motion-tracking devices; asleep at the keyboard?; mindstorms; Java jive; analyzing algorithms; the CVS data format URL. Dr. Dobb’s Journal of Software Tools, 24(7):12, 14, July 1999. CODEN DDJOEB. ISSN 1044-789X.

Baker:1993:EWB

[Bak93] Steven Baker. Extending Windows to 32 bits. Dr. Dobb’s Journal of Software Tools, 18 (1):18–20, 22, 24, 88, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Baksaas:1994:CST

[Bak94] Birger Baksaas. Computeraided software testing. Dr. Dobb’s Journal of Software Tools, 19(2):36, 38, 78, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Baker:1995:CF

[Bak95] Steven Baker. Complying with Fortran 90. Dr. Dobb’s Journal of Software Tools, 20(1):68, 70, 72, 74–76, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Baker:1997:LEP

[Bak97a] Louis Baker. Letter to the Editor: Pentium II math bug. Dr. Dobb’s Journal of Software Tools, 22(10):10, October 1997. CODEN DDJOEB. ISSN 1044-789X. Comments on the Ariane 5 missile failure mentioned in [Col97a].

Baker:1997:CIC

[Bak97b] Larry E. Baker, Jr. C++ interfaces for C-language libraries: C++ programmers must sometimes use legacy C support libraries, even though C and C++ styles don’t always mix. Dr. Dobb’s Journal of Software Tools, 22(8):34, 36, 37, 90, 91, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Baldwin:1990:ACC

[Bal90] Wahhab Baldwin. Awk as a C code generator. Dr. Dobb’s Journal of Software Tools, 15 (8):36–??, August 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Balamurugan:1999:JAP

[Bal99] S. Balamurugan. Jperl: Accessing Perl from Java. Dr. Dobb’s Journal of Software Tools, 24(2):103–105, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/jperl.txt;
                                         http://www.ddj.com/ftp/1999/
                                        1999_02/jperl.zip.

Bandes:1998:JCA

[Ban98a] Kenneth Bandes. Java and COM automation: Taking the edge off dual interfaces. Dr. Dobb’s Journal of Software Tools, 23(1):28, 30, 32, 34–35, 98–99, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Bansiya:1998:ADP

[Ban98b] Jagdish Bansiya. Automating design pattern identification. Dr. Dobb’s Journal of Software Tools, 23(6):20–22, 24, 26, 28, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Barron:1992:PDU

[Bar92] Tovey Barron. Protected-mode debugging using in-circuit emulators. Dr. Dobb’s Journal of Software Tools, 17(2):46, 48, 50, 52–53, 55–56, 58, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Barnum:1993:POW

[Bar93a] Joel Barnum. Programming for the OS/2 2.0 Workplace Shell. Dr. Dobb’s Journal of Software Tools, 18(4):72–73, 94–97, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Bartels:1993:EDP

[Bar93b]  Dirk Bartels.  Event-driven

database programming in C++. Dr. Dobb’s Journal of Software Tools, 18(4 (supplement)): 42S–8, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Baran:1999:NVQ

[Bar99a] Nicholas Baran. News and views: A quantum leap; will SDMI kill MP3?; search engines can’t keep up; real-time CORBA gets real. Dr. Dobb’s Journal of Software Tools, 24 (10):18, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Baran:1999:Ia

[Bar99b] Nicholas Baran. Of interest.

Dr. Dobb’s Journal of Software Tools, 24(10):134–135, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Baran:1999:NVB

[Bar99c] Nick Baran. News and views:

E-books highlight Seybold San Francisco; Microsoft jumps on the XML bandwagon; truth and consequences goes online; disco Déjà vu; SMIL Boston draft announced. Dr. Dobb’s Journal of Software Tools, 24(12): 18, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Baran:1999:NVW

[Bar99d] Nick Baran. News and views: The world’s smallest Web server; CPU flushed with success; light in the slow lane; Windows CE gives you gas; the ACM says no to software engineering licensing; NT versus CE. Dr. Dobb’s Journal of Software Tools, 24(11):18, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Barnes:1999:ESJ

[Bar99e] Darryl Barnes. Embedded systems: Java Card application development. Dr. Dobb’s Journal of Software Tools, 24(2): 72, 74, 76–78, 80, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jcard.txt.

Bates:1994:KCD

[Bat94] Rodney Bates. K-tree container data structures. Dr. Dobb’s Journal of Software Tools, 19 (10):26–28, 30, 32, 34, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Bates:1996:ECT

[Bat96] Rodney M. Bates. Examining the Cocktail toolbox. Dr. Dobb’s Journal of Software Tools, 21 (3):78, 80–82, 95–96, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Bauer:1990:CBP

[Bau90a] Barr E. Bauer. Controlling background processes under Unix. Dr. Dobb’s Journal of Software Tools, 15(12):16–18, 20, 82–84, 86–88, 90, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Bauer:1990:PP

[Bau90b] Barr E. Bauer. A plus for patents [letter]. Dr. Dobb’s Journal of Software Tools, 15 (7):8–??, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Bauer:1992:ALA

[Bau92a]  Barr E. Bauer.  Accessing

large arrays with X-arRAY. Dr. Dobb’s Journal of Software Tools, 17(6):74–??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Bauer:1992:PCE

[Bau92b] Barr E. Bauer. Parallel C extensions. Dr. Dobb’s Journal of Software Tools, 17(8):110, 112– 114, 124, 127, August 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Bauer:1993:PWP

[Bau93] Barr E. Bauer. Porting from workstations to PC’s. Dr. Dobb’s Journal of Software Tools, 18(9):68, 70, 72–73, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Bottorff:1990:OS

[BB90] Jan Bottorff and Jim Bolland. Object swapping. Dr. Dobb’s Journal of Software Tools, 15 (5):48–56, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Bortolotti:1994:RRM

[BBB94] J. F. Bortolotti, P. Bernard, and E. Bouchet. RTMK: A real-time microkernel. Dr. Dobb’s Journal of Software Tools, 19(5):70, 72, 74, 76, 105–106, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Ball:1998:CIA

[BC98] Steve Ball and John Miller Crawford. Channels for interapplet communication. Dr. Dobb’s Journal of Software Tools, 23(9):18–20, 22, 91, September 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/1998/1998_09/channel.
txt;   http://www.ddj.com/
ftp/1998/1998_09/channel.
zip.

Ball:1999:JQJ

[BC99] Steve Ball and John Miller Crawford. Java Q&A: Are Java applets independent programs? Dr. Dobb’s Journal of Software Tools, 24(4):101–105, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1999/1999_04/
jqa499.txt;   http://www.
ddj.com/ftp/1999/1999_04/
webcad.zip.

Braun:1999:LTA

[BCP^{+}99] Jay Braun, David Cox, Christophe Ponsard, Reg Charney, Torken Danielsson, Christian Lanctot, John D. Kent, and Steve Furlong. Letters: There’s always room for Forth; online op/eds; parsing expressions in Java; age discrimination; A better date?; more Y2K; it’s a cult kind of thing. Dr. Dobb’s Journal of Software Tools, 24 (3):10, 12, March 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm.

Bansiya:1997:AMO

[BD97] Jagdish Bansiya and Carl Davis.

Automated metrics and objectoriented development: Using QMOOD++ for object-oriented metrics. Dr. Dobb’s Journal of Software Tools, 22(12):42, 44– 48, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Blain:1993:OIF

[BDE93] Derrel Blain, Kurt Delimon, and Jeff English. OS/2 2.x initialization files and profile management. Dr. Dobb’s Journal of Software Tools, 18(5):32, 34, 36, 38, 84, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Bosselaers:1997:RCH

[BDP97] Antoon Bosselaers, Hans Dogbbertin, and Bart Preneel. The RIPEMD-160 cryptographic hash function. Dr. Dobb’s Journal of Software Tools, 22(1):24, 26, 28, 78, 80, January 1997.

CODEN DDJOEB. ISSN 1044-789X.

Brath:1997:PB

[BDT97] Richard Brath, Ray Duncan, and Edward R. Tufte. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 22 (6):133–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Beazley:1998:SAC

[Bea98] David Beazley. SWIG and automated C/C++ scripting extensions. Dr. Dobb’s Journal of Software Tools, 23(2):30, 32, 34– 36, 100, February 1998.

Beck:1994:PSD

[Bec94] Kent Beck. Patterns and software development. Dr. Dobb’s Journal of Software Tools, 19 (2):18–20, February 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Beker:1997:GER

[Bek97a] Harry Beker. Graphical embedded real time systems. Dr. Dobb’s Journal of Software Tools, 22(4):54–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Beker:1997:LEJ

[Bek97b] Harry Beker. Letter to the editor: Java hype versus Tcl/Tk reality. Dr. Dobb’s Journal of Software Tools, 22(1):8, 10, January 1997. CODEN DDJOEB. ISSN 1044-789X. Argues that Tcl/Tk provides a much more advanced windowing interface than Java currently does.

Baran:1999:NVK

[BEK99] Nick Baran, Jonathan Erickson, and Eugene Eric Kim. News and views: Keeping track and field-tested thinking and Apache Software Foundation formed and more better secure logins and electronic paper. Dr. Dobb’s Journal of Software Tools, 24 (9):16, September 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Bender:1996:CFV

[Ben96] Andreas O. Bender. Algorithm alley: Continued fractions versus Farey fractions. Dr. Dobb’s Journal of Software Tools, 21 (5):99–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Bentley:1998:AAC

[Ben98] Jon Bentley. Algorithm alley: The cost of recursion. Dr. Dobb’s Journal of Software Tools, 23(6):111–114, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Bentley:1999:AAA

[Ben99a] Jon Bentley. Algorithm alley: Analysis of algorithms. Dr. Dobb’s Journal of Software Tools, 24(4):107–110, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9904o/9904o.
htm;   http://www.ddj.com/
ftp/1999/1999_04/aa499.zip.

Bentley:1999:AAC

[Ben99b] Jon Bentley. Algorithm alley: Code tuning in context.

Dr. Dobb’s Journal of Software Tools, 24(5):125–126, 128, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/9905e/9905m/9905m.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_05/aa599.zip.

Bentley:1999:AAP

[Ben99c] Jon Bentley. Algorithm alley: Programming pearls: Cracking the oyster. Dr. Dobb’s Journal of Software Tools, 24 (12):117–119, December 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        12/aa1299.txt.

Bergman:1990:RCT

[Ber90] Noel J. Bergman. Review of C talk/views from CNS. Dr. Dobb’s Journal of Software Tools, 15(2):66–74, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Berg:1993:Ia

[Ber93a]  Monica Berg.  Of interest.

Dr. Dobb’s Journal of Software Tools, 18(7):146–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Berg:1993:Ib

[Ber93b] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 18(12):??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Berg:1993:Ic

[Ber93c] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 18(13):138–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ia

[Ber94a] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(1):138–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ib

[Ber94b] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 19(2):130–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ic

[Ber94c] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(3):162–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Id

[Ber94d] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 19(4):148–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ie

[Ber94e] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 19(5):148–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:If

[Ber94f] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 19(6):158–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ig

[Ber94g] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(7):148–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ih

[Ber94h] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(8):148–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ii

[Ber94i] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(10):??, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ik

[Ber94j] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(13):142–??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1994:Ij

[Ber94k] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 19(15):134–??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ia

[Ber95a] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(1):150–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ic

[Ber95b] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 20(3):??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ib

[Ber95c] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(3):148–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Id

[Ber95d] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 20(4):150–??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ie

[Ber95e] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 20(5):??, May 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Berg:1995:If

[Ber95f] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 20(6):142–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ig

[Ber95g] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 20(7):142–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ih

[Ber95h] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(8):150–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ii

[Ber95i] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(9):??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ij

[Ber95j] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(10):164–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Ik

[Ber95k] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(11):150–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1995:Il

[Ber95l] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 20(12):150–??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHe

[Ber96a] C. Berg. Java Q&A: How do I display media formats that Netscape does not support? Dr. Dobb’s Journal of Software Tools, 21(10):111, 114– 115, 129, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHd

[Ber96b] Cliff Berg. Java Q&A: How do I create a layout manager? Dr. Dobb’s Journal of Software Tools, 21(9):101–??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHg

[Ber96c] Cliff Berg. Java Q&A: How do I implement a drag-and-drop interface? Dr. Dobb’s Journal of Software Tools, 21(12):110– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHc

[Ber96d] Cliff Berg. Java Q&A: How do I send e-mail from a Java applet? Dr. Dobb’s Journal of Software Tools, 21(8): 111–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHb

[Ber96e] Cliff Berg. Java Q&A: How do I write a chat program? Dr. Dobb’s Journal of Software Tools, 21(6):105–106, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHf

[Ber96f] Cliff Berg. Java Q&A: How do threads work and how can I create a general-purpose event? Dr. Dobb’s Journal of Software Tools, 21(11):111–??, November 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ia

[Ber96g] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(1):140–??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ib

[Ber96h] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(2):142–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ic

[Ber96i] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(3):133–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Id

[Ber96j] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(4):134–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ie

[Ber96k] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(5):118–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:If

[Ber96l] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(6):??, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ig

[Ber96m] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(7):118–??, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ih

[Ber96n] Monica E. Berg. Of interest. Dr. Dobb’s Journal of Software Tools, 21(8):134–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:Ii

[Ber96o] Monica E. Berg. Of interest.

Dr. Dobb’s Journal of Software Tools, 21(9):119–??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:PSI

[Ber96p] Monica E. Berg. Programmer’s services of interest. Dr. Dobb’s Journal of Software Tools, 21 (6):126–??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHe

[Ber97a] Cliff Berg. Java Q&A: How can I create a push Java channel? Dr. Dobb’s Journal of Software Tools, 22(5):99, 101–103, 118, 122, 123, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHb

[Ber97b] Cliff Berg. Java Q&A: How do I access a SQL database from an applet? Dr. Dobb’s Journal of Software Tools, 22(2):103, 105, 113, 114, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHl

[Ber97c] Cliff Berg. Java Q&A: How do I browse and dynamically invoke remote objects. Dr. Dobb’s Journal of Software Tools, 22 (12):121–123, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHi

[Ber97d] Cliff Berg. Java Q&A: How do I create a JavaBean? Dr. Dobb’s Journal of Software Tools, 22 (9):102–104, 114, 115, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHh

[Ber97e] Cliff Berg. Java Q&A: How do I create a signed applet? Dr. Dobb’s Journal of Software Tools, 22(8):109–111, 122, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHf

[Ber97f] Cliff Berg. Java Q&A: How do I create my own security manager? Dr. Dobb’s Journal of Software Tools, 22(6):115, 117–120, 132, June 1997. CO-DEN DDJOEB. ISSN 1044-789X. See letter to the editor [Bou97] pointing out a flaw in the security manager.

Berg:1997:JQHd

[Ber97g] Cliff Berg. Java Q&A: How do I create persistent Java objects? Dr. Dobb’s Journal of Software Tools, 22(4):98–101, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHk

[Ber97h] Cliff Berg. Java Q&A: How do I print in Java? Dr. Dobb’s Journal of Software Tools, 22(11):

110–114, November 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHa

[Ber97i] Cliff Berg. Java Q&A: How do I use CORBA from Java? Dr. Dobb’s Journal of Software Tools, 22(1):100, 102, 103, 114, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHc

[Ber97j] Cliff Berg. Java Q&A: How do I use Java Remote Method Invocation from an applet? Dr. Dobb’s Journal of Software Tools, 22(3):101–103, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHj

[Ber97k] Cliff Berg. Java Q&A: How do I write a Java servet? Dr. Dobb’s Journal of Software Tools, 22 (10):121–123, 130, 131, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHg

[Ber97l] Cliff Berg. Java Q&A: How do I write an international application? Dr. Dobb’s Journal of Software Tools, 22(7):105, 106, 116, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1998:JQHa

[Ber98a] Cliff Berg. Java Q&A: How do I create a signed Castanet channel? Dr. Dobb’s Journal of Software Tools, 23(1):121–124, 129, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Berg:1998:JQHc

[Ber98b] Cliff Berg. Java Q&A: How do I interface Java to my PalmPilot. Dr. Dobb’s Journal of Software Tools, 23(4):118, 120, 122, 139, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        04/jqa498.txt.

Berg:1998:JQHd

[Ber98c]  Cliff Berg.  Java Q&A:

How do I password encrypt data? Dr. Dobb’s Journal of Software Tools, 23(8):107– 109, 117, August 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_08/
                                        jqa898.txt;   http://www.
                                        ddj.com/ftp/1998/1998_08/
                                        jqa898.zip.

Berg:1998:JQHb

[Ber98d] Cliff Berg. Java Q&A: How do I transfer data securely? Dr. Dobb’s Journal of Software Tools, 23(2):119–121, February 1998.

Bernecky:1998:PB

[Ber98e] Robert Bernecky. Programmer’s bookshelf: Computer architecture. Dr. Dobb’s Journal of Software Tools, 23(4):141–143, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Beroset:1998:ALC

[Ber98f] Edward J. Beroset. Assemblylanguage control-flow graphing. Dr. Dobb’s Journal of Software Tools, 23(9):58, 60, 62, 64, 66, 68, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
09/asmflow.txt.

Bertron:1998:CDD

[Ber98g] Jean-Denis Bertron. Customizing DDX/DDV. Dr. Dobb’s Journal of Software Tools, 23 (5):105–107, 137, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1996/1996_
03/mirc/mirc.htm;   http:/
/www.ddj.com/ftp/1998/1998_
05/ddxddv.zip.

Berg:1999:PEJ

[Ber99] Cliff Berg. Parsing expressions in Java: Specifying rules for common operations. Dr. Dobb’s Journal of Software Tools, 24 (1):50, 52–53, 56–58, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ddj/ftp/1999/1999_
01/jparser.txt;   http://
www.ddj.com/ddj/ftp/1999/
1999_01/jparser.zip.

Betz:1991:SHM

[Bet91a] David Betz. A simple handlebased memory management. Dr. Dobb’s Journal of Software Tools, 16(12):64–66, 151, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Betz:1991:TCC

[Bet91b] David Betz. Testing C compiler performance. Dr. Dobb’s Journal of Software Tools, 16(8):80, 82, 84, 87–88, 90, 93, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Betz:1991:YOT

[Bet91c] David Betz. Your own tiny object-oriented language. Dr. Dobb’s Journal of Software Tools, 16(9):26, 28, 30, 32–33, 86, 88–89, September 1991. CO-DEN DDJOEB. ISSN 1044-789X. Reprinted in [Bet94b].

Betz:1992:TD

[Bet92] David Betz. A taste of Dylan. Dr. Dobb’s Journal of Software Tools, 17(10):50–52, 57, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Betz:1993:DRO

[Bet93a] David Betz. Dave’s recycled OO language. Dr. Dobb’s Journal of Software Tools, 18(11): 74–78, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Betz:1993:VPM

[Bet93b] Mark Betz. VGA palette mapping using BSP trees. Dr. Dobb’s Journal of Software Tools, 18(7):28, 30, 32, 34, 36, 94, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Betz:1994:OCS

[Bet94a] David Betz. An online conferencing system construction kit. Dr. Dobb’s Journal of Software Tools, 19(??):26–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Betz:1994:YOT

[Bet94b] David Betz. Your own tiny object-oriented language. Dr. Dobb’s Journal of Software Tools, 19(13):26–??, November 1994. CODEN DDJOEB. ISSN 1044-789X. Reprint of [Bet91c].

Betz:1994:IO

[Bet94c] Mark Betz. Interoperable objects. Dr. Dobb’s Journal of Software Tools, 19(11):18–20, 24–26, 28, 32, 34, 36–39, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Betz:1994:SSM

[Bet94d] Mark Betz. An SQL server message-handling class. Dr. Dobb’s Journal of Software Tools, 19(13):18–20, 22, 24, 90, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Betz:1995:CB

[Bet95a] David Betz. Callable Bob [sidebar]. Dr. Dobb’s Journal of Software Tools, 20(5):66–??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Betz:1995:NOC

[Bet95b] Mark Betz. Networking objects with CORBA. Dr. Dobb’s Journal of Software Tools, 20 (11):18–20, 22, 25–26, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Betz:1995:QLA

[Bet95c] Mark Betz. A quick look at Borland’s Delphi 95. Dr. Dobb’s Developer Update, 2(1):2–??, January 1995. CODEN ???? ISSN 1079-8595.

Betz:1998:ADO

[Bet98]  Mark Betz.  Internet pro-

gramming: Active data objects and ASP. Dr. Dobb’s Journal of Software Tools, 23 (5):88–91, 111–112, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
05/adoasp.txt.

Beveridge:1995:IMS

[Bev95] Jim Beveridge. Inside MFC serialization. Dr. Dobb’s Journal of Software Tools, 20(10):62, 64, 66–67, 122–123, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Beveridge:1996:IMU

[Bev96] Jim Beveridge. Implementing multilevel undo/redo. Dr. Dobb’s Journal of Software Tools, 21(2):64–66, 100, 102– 104, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Beveridge:1998:SRO

[Bev98] Jim Beveridge. Self-registering objects in C++: Registering new classes at run time. Dr. Dobb’s Journal of Software Tools, 23(8):38, 40–41, 90, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
08/register.txt;   http://
www.ddj.com/ftp/1998/1998_
08/register.zip.

Bernecky:1997:PBI

[BF97] Robert Bernecky and Michael Fitzpatrick. Programmer’s bookshelf: Inner loops and regular expressions. Dr. Dobb’s Journal of Software Tools, 22 (12):141, 143, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Bowen:1994:DPL

[BH94] Tamme D. Bowen and Kelly M.

Hall. The Dylan programming language. Dr. Dobb’s Journal of Software Tools, 19(??):54–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Bennett:1995:UVB

[BH95] Kendall Bennett and Brian Hook. Using the VESA BIOS 2.0 linear frame buffer. Dr. Dobb’s Journal of Software Tools, 20(??):35–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1996:FNW

[BH96] Derek Brown and Martin Hall.

Fast networking with Winsock 2.0. Dr. Dobb’s Journal of Software Tools, 21(2):76, 78, 105– 109, 111, February 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Bhattacharyya:1998:YTS

[Bha98] Dev Bhattacharyya. A Year 2000 tool suite. Dr. Dobb’s Journal of Software Tools, 23 (5):36, 38, May 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

                                        ddj.com/ftp/1998/1998_05/
                                        javay2k.zip.

Bigby:1998:GGE

[Big98] Bruce W. Bigby. The GEF general exception-handling library. Dr. Dobb’s Journal of Software Tools, 23(11):60, 62, 64, 66, November 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1998/1998_11/GEF.
                                        TXT;   http://www.ddj.com/
                                        ftp/1998/1998_11/gef1.ZIP.

Binstock:1996:HR

[Bin96] Andrew Binstock. Hashing revisited. Dr. Dobb’s Journal of Software Tools, 21(4):24, 26, 28, 30, 33, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Bist:1996:OCS

[Bis96] Gary Bist. Optimizing client/ server database applications. Dr. Dobb’s Journal of Software Tools, 21(11):24–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Bittman:1991:APD

[Bit91] James Bittman. Adding the power of DSP to your applications. Dr. Dobb’s Journal of Software Tools, 16(5):30, 32, 34, 90–96, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Barlock:1995:PCS

[BL95] Thomas W. Barlock and David S. Lindsay. Performance comparison: Sun SPARCstation 5 versus Intel Pentium 90. Dr. Dobb’s Developer Update, 2(11):

2–3, November 1995. CODEN ???? ISSN 1079-8595.

Baran:1999:Ib

[BL99a] Nick Baran and Amy Lincicum.

Of interest. Dr. Dobb’s Journal of Software Tools, 24(11): 132, 135–136, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Baran:1999:Ic

[BL99b] Nick Baran and Amy Lincicum.

Of interest. Dr. Dobb’s Journal of Software Tools, 24(12):134– 135, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Blaha:1996:OIC

[Bla96] Stephen Blaha. Object-oriented interprocess communication: Client/server development in C++ and Java. Dr. Dobb’s Journal of Software Tools, 21 (8):24, 26, 28, 30–31, 89–92, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Bredberg:1996:RES

[BLE96] Fredrik Bredberg, Ola Liljedahl, and Bengt Eliasson. Restarting embedded systems. Dr. Dobb’s Journal of Software Tools, 21 (11):56–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Bloomer:1995:DCO

[Blo95]  John Bloomer.  Distributed

computing and the OSF/DCE. Dr. Dobb’s Journal of Software Tools, 20(2):18–20, 22, 24, 26, 28, 30, 96, 98–100, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Blum:1990:BAM

[Blu90] Adam Blum. Bidirectional associative memory systems in C++. Dr. Dobb’s Journal of Software Tools, 15(4):16–20, 24– 26, 84–86, 88, April 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Blustein:1995:IBV

[Blu95] James Blustein. Implementing bit vectors in C. Dr. Dobb’s Journal of Software Tools, 20 (8):42, 44, 46, 96, 98–100, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1997:DAO

[BN97] Marc H. Brown and Marc A. Najork. Distributed active objects. Dr. Dobb’s Journal of Software Tools, 22(3):34–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Bollella:1997:DMA

[Bol97]  Gregory Bollella.  Dynamic

memory allocation and communication protocols. Dr. Dobb’s Journal of Software Tools, 22 (3):60–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Bonneau:1993:DCD

[Bon93] Paul Bonneau. Designing complex datacentric applications. Dr. Dobb’s Journal of Software Tools, 18(3):40, 42, 102, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Boone:1992:SCP

[Boo92] Keith W. Boone. A source code profiler. Dr. Dobb’s Journal of Software Tools, 17(10):98–104, 128–131, October 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Booch:1994:DAF

[Boo94] Grady Booch. Designing an application framework. Dr. Dobb’s Journal of Software Tools, 19 (2):24, 26, 28, 30, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Born:1996:TVB

[Bor96] Günter Born. A TARGA viewer in Borland Delphi. Dr. Dobb’s Journal of Software Tools, 21 (1):90–??, January 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Boukanov:1997:JS

[Bou97] Igor Boukanov. Java security. Dr. Dobb’s Journal of Software Tools, 22(9):10, September 1997. CODEN DDJOEB. ISSN 1044-789X. Points out a flaw in the security manager described in [Ber97f].

Boyer:1997:AAH

[Boy97] John Boyer. Algorithm alley — heaps are usually implemented via binary trees, with the property that for every subtree, the root is the minimum item. here, John describes how to implement exceptionally fast ‘Fibonacci’ heaps. Dr. Dobb’s Journal of Software Tools, 22 (1):106–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Boyer:1998:AAR

[Boy98a] John Boyer. Algorithm alley: Resizable data structures. Dr. Dobb’s Journal of Software Tools, 23(1):115–116, 118, 129, January 1998. CODEN DDJOEB. ISSN 1044-789X. Discusses some deficiencies of the Java library hash table support, and compares it with his algorithm and that used in the C++ Standard Template Library. Also compares the Jenkins hash function [Jen97] with the one proposed in this paper.

Boyer:1998:AAS

[Boy98b] John Boyer. Algorithm alley: Sorting and searching linked lists in Java. Dr. Dobb’s Journal of Software Tools, 23(5):126– 129, 137–138, May 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_05/
                                        aa0598.txt;   http://www.
                                        ddj.com/ftp/1998/1998_05/
                                        aa0598.zip.

Boyer:1998:DSM

[Boy98c] John Boyer. Digital signatures with the Microsoft CryptoAPI: Adding security to Windows applications. Dr. Dobb’s Journal of Software Tools, 23(6):80, 82–85, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Boyer:1999:XEF

[Boy99] John Boyer. XFDL: The extensible forms description language. Dr. Dobb’s Journal of Software Tools, 24(12):86, 88– 89, 91–92, December 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_12/xfdl.
txt.

Bozoki:1999:ISP

[Boz99] Eva Bozoki. IP security protocols. Dr. Dobb’s Journal of Software Tools, 24(12):42, 44, 46–49, 51–52, 54–55, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Becker:1993:CT

[BR93] Peter Becker and D. Reilly.

C++ templates. Dr. Dobb’s Journal of Software Tools, 18 (8):44, 46–51, 102–103, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Binstock:1996:PAC

[BR96] Andrew Binstock and John Rex.

Practical algorithms for C programmers. Dr. Dobb’s Journal of Software Tools, 21(2): 139–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Bradberry:1990:PFP

[Bra90] John L. Bradberry. Porting FORTRAN programs from minis to PCs. Dr. Dobb’s Journal of Software Tools, 15(9):26– 36, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Bradberry:1991:FG

[Bra91] John L. Bradberry. Fortran and GUIs. Dr. Dobb’s Journal of Software Tools, 16(6):16–18, 20, 22, 25, 101–102, June 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Bradberry:1992:DPU

[Bra92] John L. Bradberry. Designing portable user interfaces. Dr. Dobb’s Journal of Software Tools, 17(11):58, 60–62, 130– 132, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Brachman:1993:EFD

[Bra93] M. L. Brachman. Extending FoxPro (database management). Dr. Dobb’s Journal of Software Tools, 18(4 (supplement)):35S–8, April 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Bradley:1994:STC

[Bra94a] N. Scott Bradley. Software testing cycles. Dr. Dobb’s Journal of Software Tools, 19(2):95– ??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Brain:1994:IWN

[Bra94b] Marshall Brain. Inside Windows NT services. Dr. Dobb’s Journal of Software Tools, 19(5):48– 51, 100–101, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Brain:1994:NCU

[Bra94c] Marshall Brain. Network communications using the NetBEUI protocol. Dr. Dobb’s Journal of Software Tools, 19(11):82, 84, 86–87, 110–111, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Bright:1990:GHV

[Bri90] Walter Bright. Getting a handle on virtual memory. Dr. Dobb’s Journal of Software Tools, 15 (5):40–46, 110, May 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Bridges:1991:DIC

[Bri91] John Bridges. Differential image compression. Dr. Dobb’s Journal of Software Tools, 16(2):38, 40–42, 44, 46, 48, 51, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Brindle:1993:WWC

[Bri93] Dan Brindle. Writing Windows custom controls. Dr. Dobb’s Journal of Software Tools, 18 (10):64–??, October 1993. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Brigner:1999:IPC

[Bri99] Paul Brigner. Internet programming: Creating signed, persistent Java applets. Dr. Dobb’s Journal of Software Tools, 24 (2):82, 84–88, February 1999. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/persist.
txt;   http://www.ddj.com/
ftp/1999/1999_02/persist.
zip.

Brown:1994:CDD

[Bro94a] J. Randolph Brown. Crossplatform database development. strategies for FoxPro developers. Dr. Dobb’s Journal of Software Tools, 19(6):84, 88, 90–91, 102, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Brown:1994:QQP

[Bro94b] Ralf Brown. QPI: the QEMM-386 programming interface. Dr. Dobb’s Journal of Software Tools, 19(7):123–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Brown:1995:CYF

[Bro95a]  Marc E. Brown.  Compet-

ing with your former employer. Dr. Dobb’s Journal of Software Tools, 20(??):61– ??, July/August 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1995:CSC

[Bro95b]  Marc E. Brown.  Copying

software concepts can be legal. Dr. Dobb’s Journal of Software Tools, 20(??):61–??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1995:CRA

[Bro95c] Marc E. Brown. Copyright ruling against Borland overturned. Dr. Dobb’s Developer Update, 2 (5):2–4, May 1995. CODEN ???? ISSN 1079-8595.

Brown:1995:PBP

[Bro95d] Marc E. Brown. Patents: Best protection for software today? Dr. Dobb’s Journal of Software Tools, 20(??):55–??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1995:SDC

[Bro95e] Marc E. Brown. Software development contracts. Dr. Dobb’s Journal of Software Tools, 20 (??):61–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1995:TWC

[Bro95f] Marc E. Brown. Trademark wars in cyberspace. Dr. Dobb’s Journal of Software Tools, 20(??):57–??, November/December 1995. CODEN DDJOEB. ISSN 1044-789X.

Brown:1996:BVW

[Bro96a] John Forrest Brown. Building VxDs in Windows 95. Dr. Dobb’s Journal of Software Tools, 21(??):29–??, March/ April 1996. CODEN DDJOEB. ISSN 1044-789X.

Brown:1996:LDS

[Bro96b] Marc E. Brown. Liability for defective software. Dr. Dobb’s Journal of Software Tools, 21 (??):61–??, March/April 1996. CODEN DDJOEB. ISSN 1044-789X.

Brown:1996:SUI

[Bro96c] Marc E. Brown. Servicing upgrading or interfacing with someone else’s software. Dr. Dobb’s Journal of Software Tools, 21(??):61–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Bentley:1998:AA

[BS98a] Jon Bentley and Robert Sedgewick.

Algorithm alley: Sorting strings with three-way radix quicksort. Dr. Dobb’s Journal of Software Tools, 23(11):133– 134, 136–138, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
11/shas/shas.htm;   http:/
/www.ddj.com/ftp/1998/1998_
11/AA118.TXT;   http://www.
ddj.com/ftp/1998/1998_11/
AA118.ZIP.

Bentley:1998:TST

[BS98b] Jon Bentley and Robert Sedgewick. Ternary search trees. Dr. Dobb’s Journal of Software Tools, 23 (4):20–22, 24–25, April 1998. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.cs.
princeton.edu/~rs/strings/
;   http://www.ddj.com/ftp/
1998/1998_04/ternary.txt;
 http://www.ddj.com/ftp/1998/
1998_04/ternary.zip. All of

the source code described in this article is available at http:/

/www.cs.princeton.edu/~rs/
strings/.

Bergman-Terrell:1993:SFN

[BT93] Eric Bergman-Terrell. Sorting files with NT’s memory-mapped file I/O [sidebar]. Dr. Dobb’s Journal of Software Tools, 18 (??):16–??, ???? 1993. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Bugg:1998:EMH

[Bug98] Keith Bugg. Examining the Microsoft HTML Help Toolkit.

Dr. Dobb’s Journal of Software Tools, 23(10):98, 100, 102– 103, October 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/1998/1998_10/htmlhelp.
txt;   http://www.ddj.com/
ftp/1998/1998_10/htmlhelp.
zip.

Bunnell:1994:MPR

[Bun94] Mitchell Bunnell. Maximizing performance of real-time RISC applications. Dr. Dobb’s Journal of Software Tools, 19(1): 54, 56, 58, 60, 62, 64, 90, 94, 96, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Burk:1993:SW

[Bur93] Ronald L. Burk. Spying on WinHelp. Dr. Dobb’s Journal of Software Tools, 18(6):153–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Burgess:1994:MYO

[Bur94] Richard Burgess. MMURTL:

Your own 32-Bit operating system. Dr. Dobb’s Journal of Software Tools, 19(5):38, 40, 44–45, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Burcham:1995:PIR

[Bur95a] Robert Burcham. Protocols for Internet radio. Dr. Dobb’s Developer Update, 2(7):1, 3, 4, July 1995. CODEN ???? ISSN 1079-8595.

Burd:1995:ZIT

[Bur95b] Dana Burd. Zero-copy interfacing to TCP/IP. Dr. Dobb’s Journal of Software Tools, 20 (9):68, 70, 72, 74, 76, 78, 106, 108–110, September 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Butterfield:1993:FCU

[But93] Timothy Butterfield. File conversion using C++ templates. Dr. Dobb’s Journal of Software Tools, 18(3):26, 28, 98, 100, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Butcher:1999:EJS

[But99]  Paul Butcher.  Extending

JScript: Supporting constructors and arrays in JScript. Dr. Dobb’s Journal of Software Tools, 24(1):42, 44, 46, 48–49, January 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/ftp/1999/1999_01/
jstest.txt;   http://www.
ddj.com/ddj/ftp/1999/1999_
01/jstest.zip.

Bertrand:1991:PWU

[BW91] Michael A. Bertrand and William R. Welch. Programming Windows using state tables. Dr. Dobb’s Journal of Software Tools, 16(12 (supplement)):39–46, December 1991. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Baron:1996:IWS

[BW96] Chris Baron and Bob Weil. Implementing a Web shopping cart. Dr. Dobb’s Journal of Software Tools, 21(9):64, 66, 68–69, 83–85, September 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Barton:1998:RTP

[BW98] John J. Barton and John Whaley. A real-time performance visualizer for Java. Dr. Dobb’s Journal of Software Tools, 23 (3):44, 46–48, 105, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
03/visualz.txt;   http://
www.ddj.com/ftp/1998/1998_
03/visualz.zip.

Bylsma:1999:AAM

[Byl99] Wesley Bylsma. Algorithm alley: Median filtering. Dr. Dobb’s Journal of Software Tools, 24(10):119–121, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/aa1099.txt.

Calaway:1995:OP

[Cal95a]  Richard Calaway.  Object-

oriented programming in S. Dr. Dobb’s Journal of Software Tools, 20(10):44, 46, 48–49, 51, 113–115, October 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Callahan:1995:GSK

[Cal95b] Gene Callahan. Generating sequential keys in an arbitrary radix. Dr. Dobb’s Journal of Software Tools, 20(12):131– ??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Calvert:1995:ASP

[Cal95c] Kevin Calvert. ATASPI and the standards process. Dr. Dobb’s Developer Update, 2(9): 2, September 1995. CODEN ???? ISSN 1079-8595.

Calcote:1996:WPT

[Cal96]  John Calcote.  Writing a

portable transport-independent Web server. Dr. Dobb’s Journal of Software Tools, 21(5):60, 62, 64–65, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Calcote:1997:TPS

[Cal97] John Calcote. Thread pools and server performance. Dr. Dobb’s Journal of Software Tools, 22 (7):60–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Campise:1990:FS

[Cam90] Leon Campise. Fast search. Dr. Dobb’s Journal of Software Tools, 15(9):42–??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Campbell:1996:EST

[Cam96] Douglass L. Campbell. Examining the SmartSockets toolkit. Dr. Dobb’s Journal of Software Tools, 21(7):74, 76–77, 84, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Canup:1990:PNI

[Can90] Bob Canup. Pick-A-number interfaces. Dr. Dobb’s Journal of Software Tools, 15(2):38– 42, February 1990. CODEN DDJOEB. ISSN 1044-789X.

DeCanne:1994:EC

[Can94a] B. De Canne. Error-recovery codes. Dr. Dobb’s Journal of Software Tools, 19(15):44, 46, 48, 50, 92–93, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Canne:1994:ERC

[Can94b] Bart De Canne. Error-recovery codes. Dr. Dobb’s Journal of Software Tools, 19(15):44– ??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Canup:1995:TW

[Can95] Bob Canup. Thunking and Windows 95 [letter]. Dr. Dobb’s Journal of Software Tools, 20 (11):16–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Cargill:1992:UMI

[Car92] Tom Cargill. Using multiple inheritance in C++. Dr. Dobb’s Journal of Software Tools, 17 (12 (special supplement)):48s– 51s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Cargill:1996:MDO

[Car96a] Tom Cargill. Managing dynamic objects in C++. Dr. Dobb’s Journal of Software Tools, 21 (6):16–18, 22, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Carone:1996:MTC

[Car96b] Timothy E. Carone. Middleware and three-tier client/server development. Dr. Dobb’s Journal of Software Tools, 21(11):

16–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Carolan:1997:QCP

[Car97a]  Mark Carolan.  Quicktime

and cross-platform multimedia. Dr. Dobb’s Journal of Software Tools, 22(4):70–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Carolan:1997:RPG

[Car97b] Mark Carolan. RaveKit: A portable graphics framework. Dr. Dobb’s Journal of Software Tools, 22(7):30–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Carter:1997:RFS

[Car97c] Everett F. Carter, Jr. Robots and finite state machines — designing and building autonomous robots presents a host of technical challenges. our author focuses on one of these challenges, describing the high-level processing he implemented in designing a robot-control system. Dr. Dobb’s Journal of Software Tools, 22(2):50–??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Cave:1995:DCN

[Cav95] W. Dale Cave. Developing C++ NLMs. Dr. Dobb’s Journal of Software Tools, 20(8):90, 92, 94– 95, 107–109, August 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Cavalier:1999:HIU

[Cav99] Forrest J. Cavalier, III. The HTTPsync incremental update utility. Dr. Dobb’s Journal of Software Tools, 24(7):52, 54, 56, July 1999. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
articles/1999/9907/9907e/
9907e.htm;   http://www.ddj.
com/ftp/1999/1999_07/httpsync.
txt;   http://www.ddj.com/
ftp/1999/1999_07/httpsync.
zip.

Corbit:1996:AAM

[CC96] Dann Corbit and Rafael Collantes-Bellido. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 21(12):118–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Celko:1993:GAD

[Cel93] Joseph Celko. Genetic algorithms and database indexing. Dr. Dobb’s Journal of Software Tools, 18(4):30–32, 34, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Celko:1994:P

[Cel94] Joseph Celko. Partitions. Dr.

Dobb’s Journal of Software Tools, 19(13):116–117, 140– 141, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Cessna:1996:DCI

[Ces96] Edward Cessna. Dylan’s Creole interface. Dr. Dobb’s Journal of Software Tools, 21(3):18, 20, 22, 24, 26, 90–91, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Cornell:1997:CJ

[CH97a] Gary Cornell and Cay S. Horstmann. Core Java. SunSoft Press Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, second edition, January 1997. ISBN 0-13-596891-7. xxxii + 766 pp. LCCN QA76.73.J38C67 1997. US$44.95. URL http:

//www.sun.com/books/books/
Cornell/Cornell.html. Avail-

able in German translation [CH97b]. See review in [You97]. Includes CD-ROM.

Cornell:1997:JBI

[CH97b] Gary Cornell and Cay S. Horstmann. JAVA bis ins Detail. SunSoft Press Java series. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-087-0. 648 pp. LCCN ???? 99,80 DM.

URL http://www.emedia.de/
bin/bookshop?show=4553&id=.

German translation of [CH97a]. Includes CD-ROM.

Chappell:1992:USD

[Cha92] Geoff Chappell. Untangling SMARTDrive (disk caching). Dr. Dobb’s Journal of Software Tools, 17(1):64, 66, 68, 71–72, 90–92, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Charney:1993:CMA

[Cha93] Reginald B. Charney. C++ manipulators and applicators. Dr. Dobb’s Journal of Software Tools, 18(11):150–155, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Charney:1994:DAN

[Cha94] Reginald B. Charney. Data Attribute Notation in C++. Dr. Dobb’s Journal of Software Tools, 19(8):139, 144–151, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Charlap:1995:BFFa

[Cha95a] David Charlap. The BMP file format. part 1. Dr. Dobb’s Journal of Software Tools, 20(3):44, 46–48, 50, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Charlap:1995:BFFb

[Cha95b] David Charlap. The BMP file format. part 2. Dr. Dobb’s Journal of Software Tools, 20(4):34, 36–38, 40, 42, 92–93, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Charney:1995:DAN

[Cha95c] Reginald B. Charney. Data attribute notation relationships. Dr. Dobb’s Journal of Software Tools, 20(1):106–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Charney:1995:PBR

[Cha95d] Reginald B. Charney. Programmer’s bookshelf: Reviews: Software Development with Z, by J. B. Wordsworth [Addison-Wesley 1992]. Dr. Dobb’s Journal of Software Tools, 20(6): 125–126, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Channon:1996:PC

[Cha96] David Channon. Persistence for C++. Dr. Dobb’s Journal of Software Tools, 21(10): 46–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Chen:1992:PDD

[Che92] Daniel Chen. Parallel DSP for designing adaptive filters. Dr. Dobb’s Journal of Software Tools, 17(1):16–18, 20, 22, 74– 77, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Cheung:1996:ASA

[Che96] Derek Cheung. ATM software analysis and design. Dr. Dobb’s Journal of Software Tools, 21 (10):70–??, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Chorney:1995:DAN

[Cho95]  R. B. Chorney.  Data at-

tribute notation relationships. Dr. Dobb’s Journal of Software Tools, 20(1):106–107, 145, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Chocholac:1999:PNA

[Cho99]  Jaromir Chocholac.  Proto-

cols for niche area networks. Dr. Dobb’s Journal of Software Tools, 24(7):68, 70, 72– 74, 76, 78, July 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1999/1999_07/nan.
                                        txt.

Christ:1999:JSP

[Chr99] Tilo Christ. Java, synchronization, and the Palmpilot. Dr. Dobb’s Journal of Software Tools, 24(7):58, 60–64, 66, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
07/syncbuild.txt;   http:/
/www.ddj.com/ftp/1999/1999_
07/syncbuild.zip.

Chui:1991:CPF

[Chu91] Paul Chui. A C++ PCX file viewer for Windows 3. Dr. Dobb’s Journal of Software Tools, 16(7):62, 64, 66, 69, 71, 73, 97–102, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Chubb:1992:ILL

[Chu92a]  Douglas Chubb.  An im-

proved Lisp-style library for C. Dr. Dobb’s Journal of Software Tools, 17(9):52, 54, 56–57, 112, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Chui:1992:UPW

[Chu92b] Paul Chui. Undocumented DOS from protected-mode Windows 3. Dr. Dobb’s Journal of Software Tools, 17(2):70, 72–74, 76, 102–103, February 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Covington:1994:PFU

[CJ94] Michael A. Covington and Mark Juric. A print filter for UNIX: More power for your LaserJet 4M printer. Dr. Dobb’s Journal of Software Tools, 19(10):64–65, 67, 95–96, September 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Carleton:1998:PTP

[CKS98] Gary Carleton, Knud Kirkegaard, and David Sehr. Programmer’s toolchest: Profile-guided optimizations. Dr. Dobb’s Journal of Software Tools, 23(5):98, 100–103, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Chauvet:1997:OMJ

[CL97] Jean-Marie Chauvet and Marc Lerman. Object models and Java: Mixing and matching disparate object models. Dr. Dobb’s Journal of Software Tools, 22(12):20–22, 243, 99, 100, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Clark:1992:ST

[Cla92]  Dean Clark.  Splay trees.

Dr. Dobb’s Journal of Software Tools, 17(12):56, 58, 60, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Clark:1995:PA

[Cla95] Dean Clark. Popularity algorithm. Dr. Dobb’s Journal of Software Tools, 20(7):121–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Clark:1996:CQU

[Cla96a] Dean Clark. Color quantization using octrees. Dr. Dobb’s Journal of Software Tools, 21(1):54– 57, 102–104, January 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Clark:1996:PB

[Cla96b]  Dean Clark.  Programmer’s

bookshelf. Dr. Dobb’s Journal of Software Tools, 21(5):113–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Clark:1997:DDA

[Cla97] Dean Clark. A 2-D DDA algorithm for fast image scaling. Dr. Dobb’s Journal of Software Tools, 22(4):46–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Clouden:1994:WL

[Clo94]  Mark Clouden.  The WS-

NETWRK library [sidebar]. Dr. Dobb’s Journal of Software Tools, 19(??):50–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Coats:1995:COS

[CM95] Mark Coats and Terry Mellon.

Constructing operational specifications. Dr. Dobb’s Journal of Software Tools, 20(6):18–20, 22, 24, 26, 28, 30, 33, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Coad:1997:JDB

[CM97] Peter Coad and Mark Mayfield.

Java Design: Building Better Apps and Applets. Yourdon Press Computing Series. Yourdon Press, Upper Saddle River, NJ, USA, January 1, 1997. ISBN 0-13-271149-4. xiii + 238 pp. LCCN QA76.73.J38C65 1996. US$39.95. See [Wil97i].

Cody:1992:FWT

[Cod92] Mac A. Cody. The fast wavelet transform. Dr. Dobb’s Journal of Software Tools, 17(4):16–18, 20, 24, 26, 28, 100–101, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Cody:1993:WA

[Cod93] Mac A. Cody. A wavelet analyzer. Dr. Dobb’s Journal of Software Tools, 18(4):44–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Cody:1994:VBE

[Cod94a] Doug Cody. The VESA BIOS extension/audio interface. Dr. Dobb’s Journal of Software Tools, 19(??):58–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Cody:1994:WPT

[Cod94b] Mac A. Cody. The wavelet packet transform. Dr. Dobb’s Journal of Software Tools, 19 (4):44–46, 50, 52, 54, 100, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Coe:1995:IPF

[Coe95] Tim Coe. Inside the Pentium FDIV bug. Dr. Dobb’s Journal of Software Tools, 20(4):129–??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Cogswell:1993:EWP

[Cog93] Jeffrey M. Cogswell. Exploring Windows palettes. Dr. Dobb’s Journal of Software Tools, 18 (5):125–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Cohen:1992:IGC

[Coh92] Debra Cohen. The i860 as A graphics controller. Dr. Dobb’s Journal of Software Tools, 17 (7):64, 66, 68, 70, 72–73, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Colin:1996:AAB

[Col96a] Andrew Colin. Algorithm alley: Building decision trees with the ID3 algorithm. Dr. Dobb’s Journal of Software Tools, 21 (6):107–109, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Collins:1996:UC

[Col96b] Robert R. Collins. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 21(9): 112–??, September 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Collins:1996:UCC

[Col96c] Robert R. Collins. Undocumented corner: CPUID algorithm wars. Dr. Dobb’s Journal of Software Tools, 21(11): 119, 122–125, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Collins:1996:UCP

[Col96d] Robert R. Collins. Undocumented corner: Paging extensions for the Pentium Pro processor. Dr. Dobb’s Journal of Software Tools, 21(7):107–??, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Collins:1996:UCU

[Col96e] Robert R. Collins. Undocumented corner: Understanding Pentium’s 4-MB page size extensions. Dr. Dobb’s Journal of Software Tools, 21(5):105–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Collins:1997:IPI

[Col97a] Robert C. Collins. Inside the Pentium II math bug. Dr. Dobb’s Journal of Software Tools, 22(8):52, 55–57, August 1997. CODEN DDJOEB. ISSN 1044-789X. See letters [AB97, Bak97a].

Collins:1997:UCIb

[Col97b] Robert R. Collins. Undocumented corner: ICE mode and the Pentium processor. Dr. Dobb’s Journal of Software Tools, 22(11):121–123, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Collins:1997:UCCb

[Col97c] Robert R. Collins. Undocumented corner: In-circuit emulation: A powerful hardware tool for software debugging. Dr. Dobb’s Journal of Software Tools, 22(7):113–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Collins:1997:UCIa

[Col97d] Robert R. Collins. Undocumented corner: Intel’s System Management Mode. Dr. Dobb’s Journal of Software Tools, 22 (1):110–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Collins:1997:UCCa

[Col97e] Robert R. Collins. Undocumented corner: The caveats of Pentium System Management Mode. Dr. Dobb’s Journal of Software Tools, 22(5):109–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Collins:1997:UCS

[Col97f] Robert R. Collins. Undocumented corner: The secrets of System Management Mode. Dr. Dobb’s Journal of Software Tools, 22(3):109–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Collins:1998:UC

[Col98a] Robert L. Collins. Undocumented corner: Details of Intel’s Virtual Mode Extensions (VME). Dr. Dobb’s Journal of Software Tools, 23(3):133– 135, 138, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Collins:1998:BFF

[Col98b] Robert R. Collins. Benchmarks: Fact, fiction, or fantasy? Dr. Dobb’s Journal of Software Tools, 23(3):30, 32, 34–37, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
03/benchmk.zip.

Collins:1998:PFB

[Col98c] Robert R. Collins. The Pentium F00F bug. Dr. Dobb’s Journal of Software Tools, 23(5):62, 64–66, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
05/f00fbug.txt;   http://
www.ddj.com/ftp/1998/1998_
05/f00fbug.zip.

Collins:1998:UCP

[Col98d] Robert R. Collins. Undocumented corner: The Pentium’s enhanced v86 mode. Dr. Dobb’s Journal of Software Tools, 23 (1):126–128, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Collins:1998:UCV

[Col98e] Robert R. Collins. Undocumented corner: VME: Coming out of the cold. Dr. Dobb’s Journal of Software Tools, 23 (5):130–133, 138, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
05/shas/shas.htm;   http:/
/www.ddj.com/ftp/1998/1998_
05/ud598.txt.

Colin:1999:AAA

[Col99] Andrew Colin. Algorithm alley: The analytic hierarchy process. Dr. Dobb’s Journal of Software Tools, 24(2):123–125, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/aa299.txt.

Conger:1991:SCW

[Con91] James Conger. Screen capturing for Windows 3.0. Dr. Dobb’s Journal of Software Tools, 16 (2):85–87, February 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Conger:1994:UMM

[Con94a] James Conger. Using the Microsoft mail API. Dr. Dobb’s Journal of Software Tools, 19 (8):72–76, 98, 100, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Connolly:1994:EHS

[Con94b]  Dan Connolly.  Evolution

of the HTML spec [sidebar]. Dr. Dobb’s Journal of Software Tools, 19(15):20, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Conway:1995:CIC

[Con95] Darrel J. Conway. A C++ integrator class. Dr. Dobb’s Journal of Software Tools, 20(12):52, 54– 56, 58, 106–108, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Conklin:1998:SCO

[Con98] Edward K. Conklin. Smart cards and the Open Terminal Architecture. Dr. Dobb’s Journal of Software Tools, 23(12): 70, 72, 74, 76, 78, 80, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
12/../../../ftp/1998/1998_
12/sc_ota.txt. See errata

[Con99].

Conklin:1999:ESC

[Con99] Edward K. Conklin. Errata: “Smart Cards and the Open Terminal Architecture” (DDJ, December 1998). Dr. Dobb’s Journal of Software Tools, 24 (2):18, February 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/1999/9902/9902toc.htm.

See [Con98].

Cooper:1996:GSD

[Coo96] Alan Cooper. Goal-directed software design. Dr. Dobb’s Journal of Software Tools, 21 (9):16–18, 20, 22, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Cook:1998:RTS

[Coo98a] Randall Cook. Real-time sound processing. Dr. Dobb’s Journal of Software Tools, 23(10): 50, 52–55, 110–112, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        10/rtsound.txt;   http://
                                        www.ddj.com/ftp/1998/1998_
                                        10/rtsound.zip.

Cooper:1998:LJJ

[Coo98b] Greg Cooper. Letters: Java JNI. Dr. Dobb’s Journal of Software Tools, 23(1):10, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Coplien:1994:ESD

[Cop94] James O. Coplien. Examining the software development process: How do you measure the effectiveness of a development process? Dr. Dobb’s Journal of Software Tools, 19(11):88, 90, 92, 94–97, October 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Coppieters:1995:CBD

[Cop95] Kristiaan Coppieters. A crossplatform binary diff. Dr. Dobb’s Journal of Software Tools, 20 (5):32, 35–36, May 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Coppieters:1996:PM

[Cop96] Kristiaan Coppieters. Pipes for Macintosh. Dr. Dobb’s Journal of Software Tools, 21(5):26, 28, 30, 78–80, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Copp:1998:FFA

[Cop98] Lee R. Copp. File formats and automotive data acquisition. Dr. Dobb’s Journal of Software Tools, 23(11):84, 86, 88–90, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1998/1998_11/DFR.
TXT;   http://www.ddj.com/
ftp/1998/1998_11/DFR.ZIP.

Corbit:1996:AA

[Cor96] Dann Corbit. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 21(10):117–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Cousins:1994:OF

[Cou94a] Billy Cousins. OLE2 and .INI files. Dr. Dobb’s Journal of Software Tools, 19(11):98–100, 112, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Cousins:1994:OWI

[Cou94b] Billy Cousins. OLE2 and Windows .INI files. Dr. Dobb’s Journal of Software Tools, 19(13): 98–??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Courtney:1996:CSE

[Cou96] Mike J. Courtney. A cubic spline extrema algorithm. Dr. Dobb’s Journal of Software Tools, 21 (4):34, 36, 38, 40, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Cox:1992:SEO

[Cox92] Brad Cox. Superdistribution and electronic objects. Dr. Dobb’s Journal of Software Tools, 17(10):44–48, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Cunningham:1999:ESL

[CP99] Tom Cunningham and Chad Peckham. Embedded systems: Low-level APIs for embedded systems. Dr. Dobb’s Journal of Software Tools, 24 (3):68, 70, 72, 74–77, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_03/mcore.txt.

Calbaum:1993:UWS

[CPRB93] Mike Calbaum, Frank Porcaro, Mark Ruegsegger, and Bruce Backman. Untangling the Windows Sockets API. Dr. Dobb’s Journal of Software Tools, 18 (2):66, 68, 70–71, 96, 98– 99, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Carlini:1992:GCC

[CR92] Giuliano Carlini and Susan Rendina. Garbage collection for C programs. Dr. Dobb’s Journal of Software Tools, 17(11):46, 48–49, 128–129, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Cramer:1990:CML

[Cra90] William D. Cramer. CTrace:

a message logging class. Dr. Dobb’s Journal of Software Tools, 15(11):44, 46, 48, 52, 55, 116, 118, 120, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Cravit:1997:LEJ

[Cra97] Matthew Cravit. Letter to the editor: Java response. Dr. Dobb’s Journal of Software Tools, 22(2):10, 12, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Craig:1999:HLS

[Cra99] Donald C. Craig. Hierarchical logic simulation. Dr. Dobb’s Journal of Software Tools, 24 (3):32, 34, 36–38, March 1999. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/digitcl.
txt;   http://www.ddj.com/

                                        ftp/1999/1999_03/digitcl.
                                        zip.

Creighton:1998:DLL

[Cre98] N. Thomas Creighton. Dynamic linking and late binding for Netware. Dr. Dobb’s Journal of Software Tools, 23(9):26, 28, 30– 31, 91–92, September 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1998/1998_09/dnl.
                                        txt;   http://www.ddj.com/
                                        ftp/1998/1998_09/dnl.zip.

Criscolo:1998:JQ

[Cri98] Mike Criscolo. Java Q&A: How do I queue Java threads? Dr. Dobb’s Journal of Software Tools, 23(10):127–129, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_10/
                                        jqa108.txt;   http://www.
                                        ddj.com/ftp/1998/1998_10/
                                        jqa108.zip.

Crooks:1993:PVC

[Cro93a] R. Crooks. Performance verification: cache, RISC, and embedded systems. Dr. Dobb’s Journal of Software Tools, 18(12):46, 48, 50, 52, 56, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Crooks:1993:PV

[Cro93b] Roger Crooks. Performance verification. Dr. Dobb’s Journal of Software Tools, 18(11):46– ??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Cronin:1994:BYO

[Cro94a] Dennis Cronin. Build your own RS-232 sound system. Dr. Dobb’s Journal of Software Tools, 19(??):70–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Cronin:1994:EAD

[Cro94b] Dennis Cronin. Examining audio DSP algorithms. Dr. Dobb’s Journal of Software Tools, 19 (7):78, 80, 82, 84, 104, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Cronin:1994:MGI

[Cro94c] Dennis Cronin. MUD games on the Internet. Dr. Dobb’s Journal of Software Tools, 19(??):22–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Crocker:1995:PPN

[Cro95a] Lee Daniel Crocker. PNG: The Portable Network Graphic format. Dr. Dobb’s Journal of Software Tools, 20(7):36, 38, 40, 42, 44, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Cronin:1995:SS

[Cro95b] Dennis Cronin. Single-image stereograms. Dr. Dobb’s Journal of Software Tools, 20(7):18–20, 22, 24, 27, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Cromwell:1998:PBL

[Cro98a] Jeff Cromwell. Programmer’s bookshelf: Learning ISAPI. Dr. Dobb’s Journal of Software Tools, 23(3):141, 143, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Cromwell:1998:PBD

[Cro98b] Jeff Cromwell. Programmer’s bookshelf: The dawning of the age of multithreading. Dr. Dobb’s Journal of Software Tools, 23(9):127, 129, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Civille:1994:CNG

[CS94] Richard Civille and Renée E. Sieber. Civic networking with geographic information systems. Dr. Dobb’s Journal of Software Tools, 19(??):68–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Culcasi:1995:FMF

[CS95] Kevin Culcasi and Nick Schwartzman. Flash memory: FTL versus ATA. Dr. Dobb’s Developer Update, 2(12):2, December 1995. CODEN ???? ISSN 1079-8595.

Crocker:1990:DOT

[CT90] Kenneth L. Crocker and Michael T. Thompson. Designing an OSI test bed. Dr. Dobb’s Journal of Software Tools, 15(12):24, 28, 30, 32, 34–36 92–93, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Cullens:1994:CDV

[Cul94] Chane Cullens. Cross-Platform development with Visual C++. Dr. Dobb’s Journal of Software Tools, 19(3):64, 68, 70, 72–73, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Cullens:1995:SM

[Cul95] Chane Cullens. Serialization and MFC. Dr. Dobb’s Journal of Software Tools, 20(4):76–??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:DI

[Cus90a] Janna Custer. DDJ index for 1989. Dr. Dobb’s Journal of Software Tools, 15(4):65A, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ia

[Cus90b] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 15(1):??, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ib

[Cus90c] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 15(2):??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ic

[Cus90d] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 15(3):??, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Id

[Cus90e] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 15(4):??, April 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ie

[Cus90f] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 15(5):??, May 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1990:If

[Cus90g] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 15(6):??, June 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ig

[Cus90h] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 15(7):??, July 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ih

[Cus90i] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 15(8):??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ii

[Cus90j] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 15(9):??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ij

[Cus90k] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 15(10):??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Ik

[Cus90l] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 15(11):??, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1990:Il

[Cus90m] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 15(12):??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Custer:1991:Ia

[Cus91a] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 16(1):??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Custer:1991:Ib

[Cus91b] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 16(2):??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Custer:1991:Ic

[Cus91c] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 16(3):??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Custer:1991:Id

[Cus91d] Janna Custer. Of interest.

Dr. Dobb’s Journal of Software Tools, 16(4):??, April 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Custer:1991:Ie

[Cus91e] Janna Custer. Of interest. Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Czuchry:1990:NNI

[Czu90] Andrew J. Czuchry, Jr. A neural network instantiation environment. Dr. Dobb’s Journal of Software Tools, 15(4):28–36, 93, 95–96, 98, April 1990. CODEN DDJOEB. ISSN 1044-789X.

daSilva:1995:ECO

[da 95] Rod da Silva. Examining CAvisual objects. Dr. Dobb’s Journal of Software Tools, 20(10): 90–92, 95–96, 98, 124, 126– 127, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Dahms:1993:MCP

[Dah93] John Dahms. Making compilers Pentium aware [sidebar]. Dr. Dobb’s Journal of Software Tools, 18(6):38–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Doyle:1996:PSW

[DAM96] Michael Doyle, Cheong Ang, and David Martin. Proposing a standard Web API: Short circuiting the API wars. Dr. Dobb’s Journal of Software Tools, 21(2):18–20, 22, 24, 26, 91, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Danahy:1995:DDI

[Dan95] Jack Danahy. Deploying DCE as an infrastructure. Dr. Dobb’s Journal of Software Tools, 20 (2):88, 90–93, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Danzig:1996:HOC

[Dan96] Peter B. Danzig. The Harvest object cache. Dr. Dobb’s Journal of Software Tools, 21(4):70, 72–74, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Dash:1993:IPU

[Das93] Raj Kumar Dash. Image processing using quadtrees. Dr. Dobb’s Journal of Software Tools, 18(7):44, 46, 48–49, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Davidson:1991:GCC

[Dav91]  Andrew Davidson.  Generic

container classes in C++. Dr. Dobb’s Journal of Software Tools, 16(8):50, 52, 124– 125, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Davis:1993:DDWa

[Dav93a] Pete Davis. Documenting documentation: The Windows .HLP file format, part I. Dr. Dobb’s Journal of Software Tools, 18 (9):119–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Davis:1993:DDWb

[Dav93b] Pete Davis. Documenting documentation: The Windows .HLP file format, part II. Dr. Dobb’s Journal of Software Tools, 18 (10):143–??, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Davison:1995:AUN

[Dav95a] Andrew Davison. Animation using the Netscape browser. Dr. Dobb’s Journal of Software Tools, 20(??):13–??, November/December 1995. CODEN DDJOEB. ISSN 1044-789X.

Davison:1995:CIH

[Dav95b] Andrew Davison. Clickable images in HTML. Dr. Dobb’s Journal of Software Tools, 20(9):18– 20, 22, 24, 27, 98–100, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Davison:1995:CHF

[Dav95c] Andrew Davison. Coding with HTML forms. Dr. Dobb’s Journal of Software Tools, 20(6):70, 72–75, 106–109, June 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Davies:1998:VRP

[Dav98] Martyn Davies. Voice recording and playback with ISDN. Dr. Dobb’s Journal of Software Tools, 23(9):42, 44, 46, 93, September 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_09/
                                        isdnrec.txt;   http://www.
                                        ddj.com/ftp/1998/1998_09/
                                        isdnrec.zip.

Davis:1999:EMS

[Dav99] Peter Davis. Examining Microsoft’s speech SDK. Dr. Dobb’s Journal of Software Tools, 24(7):86, 88–90, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1999/1999_07/
spsdk.txt;   http://www.ddj.
com/ftp/1999/1999_07/spsdk.
zip.

Day:1992:INC

[Day92] Michael Day. Implementing NLM-Based client/server architectures. Dr. Dobb’s Journal of Software Tools, 17(10):78– 84, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Daemen:1998:PCF

[DC98] Joan Daemen and Craig Clapp.

The Panama cryptographic function. Dr. Dobb’s Journal of Software Tools, 23(12): 42, 44, 46, 48–49, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
12/../../../ftp/1998/1998_
12/panama.zip.

DDJStaff:1997:NVa

[DDJ97a] DDJ Staff. News and views.

Dr. Dobb’s Journal of Software Tools, 22(5):16–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1997:NVb

[DDJ97b] DDJ Staff. News and views.

Dr. Dobb’s Journal of Software Tools, 22(6):16–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1997:NVD

[DDJ97c] DDJ Staff. News and views.

Dr. Dobb’s Journal of Software Tools, 22(10):16, October 1997.

CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1997:NVc

[DDJ97d] DDJ Staff. News and views. Dr. Dobb’s Journal of Software Tools, 22(12):18, December 1997. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVS

[DDJ98a] DDJ Staff. News and views: A Standard Linux? cryptography contest; drives get smaller and Smaller; Perl conference; really embedded systems; programmer shortage?; Beowulf: Linux clustering; Java SPEC released. Dr. Dobb’s Journal of Software Tools, 23(11):16, November 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVb

[DDJ98b] DDJ Staff. News and views: Computing olympiad; the public’s right to know; smart cards; more Y2K; why Rome burns; biocomputing; new infrared standards; national medals awarded. Dr. Dobb’s Journal of Software Tools, 23(3):18, March 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVG

[DDJ98c] DDJ Staff. News and views: Going west; end of an era; electronic messaging available; putting your money where your mouth is; Amiga redux; unscrambling encryption; ergonomic research; PPTP bug. Dr. Dobb’s Journal of Software Tools, 23(8):16, August 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVY

[DDJ98d] DDJ Staff. News and views:

In your face; who invented the microprocessor?; mixed media; but where will they go for spring break?; quantum computing lives; E-commerce continues to grow…; … but will smart cards play a role? Dr. Dobb’s Journal of Software Tools, 23(12):18, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

microcomputerhistory.com/.

DDJStaff:1998:NVK

[DDJ98e] DDJ Staff. News and views: Kudos for free software pioneers; PSCs: Personal supercomputers; smart dialing; let it snow…; math for the Web; the taxman changes; advances in nanoelectromechanical technology; Tcl goes it alone. Dr. Dobb’s Journal of Software Tools, 23(5):18, May 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVN

[DDJ98f] DDJ Staff. News and views:

New trends in vaporware; distance ed might pay off; life in the fast lane; making friends in Washington; news on OpenGL 1.2; no discounts for schools; let’s do lunch; encryption export challenge?; Java fissures; searching for talent in science; nanomedicine. Dr. Dobb’s Journal of Software Tools, 23(6):18, June 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVPb

[DDJ98g] DDJ Staff. News and views: Plateware; it’s that simple; shareware kudos; son of AARD; old meets new; silicon-on-insulator breakthrough; Web design competition. Dr. Dobb’s Journal of Software Tools, 23 (10):16, October 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ddj/1998/1998_10/
                                        index.htm.

DDJStaff:1998:NVPa

[DDJ98h] DDJ Staff. News and views: Programming contests announced; Y2K set in stone; USENIX notes; thinking about clusters; “it’s mine, all mine!”; vintage computers; fighting the programmer shortage; Linux: A dangerous weapon. Dr. Dobb’s Journal of Software Tools, 23 (9):16, September 1998. CO-DEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVc

[DDJ98i] DDJ Staff. News and views: The secret story of nonsecret encryption; Netscape news; key escrow woes; fingerprint IC. Dr. Dobb’s Journal of Software Tools, 23 (4):18, April 1998. CODEN DDJOEB. ISSN 1044-789X. Discusses a claim by the British GHCQ agency to have invented public-key cryptography in unpublished classified work prior to the RSA and Diffie–Hellman publications.

Staff:1999:NVC

[DDJ99a] DDJ Staff. News and views:

Conventional wisdom; 2B update; coffee, tea, or PCs; … or your money back; how many pens on the head of a pin; measuring up. Dr. Dobb’s Journal of Software Tools, 24(4):16, April 1999. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1999:NVJ

[DDJ99b] DDJ Staff. News and views:

Just say no to carpal tunnel; ACM award for S creator; cruising with Perl; you can’t judge an open E-book …; H1-B visas in short supply; we interrupt you with this message…;

great timing; www.waitandsee.

org. Dr. Dobb’s Journal of Software Tools, 24(8):18, August 1999. CODEN DDJOEB. ISSN 1044-789X.

DDJstaff:1999:NVN

[DDJ99c] DDJ staff. News and views: National Engineers Week at UC Berkeley; recycling that PC; patent suits; Open Service Gateway spec planned. Dr. Dobb’s Journal of Software Tools, 24 (5):16, May 1999. CODEN DDJOEB. ISSN 1044-789X.

DDJstaff:1999:NVR

[DDJ99d] DDJ staff. News and views:

Real-time Java working group; simulated safety; A house of smart cards; father of ubiquitous computing passes away; electrochemical fabrication. Dr. Dobb’s Journal of Software Tools, 24(7):18, July 1999. CO-DEN DDJOEB. ISSN 1044-789X.

DDJStaff:1999:NVSb

[DDJ99e] DDJ staff. News and views: Sorting things out; congratulations are in order; more kudos; cheaper chips; online astronomy. Dr. Dobb’s Journal of Software Tools, 24(3):16, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.

htm. New world record in sorting: Datamation’s benchmark of one-million 100-byte records read, sorted, and written in 1.18 seconds using a cluster of 16 off-the-shelf PCs, each with two Pentium processors, two hard disk, and Windows NT 4.0.

DDJStaff:1999:NVSa

[DDJ99f] DDJ Staff. News and views: Speeding up 3D modeling; Project Gutenberg; FSF honors Larry Wall; smart pens don’t make smart writers; power hogs; virtual fish: Java’s killer app?; evaluating testing tools; software patents con. Dr. Dobb’s Journal of Software Tools, 24 (1):18, January 1999. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1999:NVT

[DDJ99g] DDJ staff. News and views: Tough love; restoring an IBM 1620; picture this; and the winner is …; but will it run Microsoft Word?; computer tutor; variation on a Y2K theme. Dr. Dobb’s Journal of Software Tools, 24(2):18, February 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm.

deCarmo:1999:ITP

[deC99] Linden deCarmo. Internet telephony protocols. Dr. Dobb’s Journal of Software Tools, 24 (7):30, 32, 34–36, 38–39, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/articles/1999/
9907/9907c/9907c.htm.

Deel:1994:ABC

[Dee94] Ernie F. Deel. Adaptive block coding. Dr. Dobb’s Journal of Software Tools, 19(3):127–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Degener:1994:DSC

[Deg94] Jutta Degener. Digital speech compression. Dr. Dobb’s Journal of Software Tools, 19(15):30, 32, 34, 88–89, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Delessio:1997:CWM

[Del97]  Carmen Delessio.  Convert-

ing Windows Metafiles to Java. Dr. Dobb’s Journal of Software Tools, 22(5):34, 36, 37, 74, 76, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Dempsey:1993:PPD

[Dem93] Miles Dempsey. Prototyping and programming database systems. Dr. Dobb’s Journal of Software Tools, 18(4 (supplement)):19S–23S, April 1993. CODEN DDJOEB. ISSN 1044-789X.

DeVos:1996:CLA

[DeV96] John K. DeVos. A compact logarithm algorithm. Dr. Dobb’s Journal of Software Tools, 21 (3):117–??, March 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Figueiredo:1996:LEE

[dFIC96] Luiz Henrique de Figueiredo, Roberto Ierusalimschy, and Waldemar Celes. LUA: An extensible embedded language. Dr. Dobb’s Journal of Software Tools, 21(12):26–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Duvanenko:1990:ODO

[DGR90] Victor J. Duvanenko, Ronald S. Gyurcsik, and W. E. Robbins. Optimal determination of object extents. Dr. Dobb’s Journal of Software Tools, 15(10): 58, 60, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Dibble:1996:PGM

[Dib96]  Peter C. Dibble.  A pro-

cess group manager for OS-9. Dr. Dobb’s Journal of Software Tools, 21(9):52, 54, 56, 58, 60– 62, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Dickerson:1995:SD

[Dic95a] Willard J. Dickerson. 68HC05based system design. Dr. Dobb’s Journal of Software Tools, 20 (8):??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Dickerson:1995:SDA

[Dic95b] Willard J. Dickerson. 68HC05based system design (antilock brake systems). Dr. Dobb’s Journal of Software Tools, 20 (8):68, 70, 72, 74, 76, 78, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Diem:1993:WDM

[Die93] Darrell Diem. Wireless data and minimum airtime software. Dr. Dobb’s Journal of Software Tools, 18(13):90, 92, 94–96, 102– 103, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Diener:1999:PCS

[Die99] Oliver Diener. Porting communications software to Windows CE. Dr. Dobb’s Journal of Software Tools, 24(9): 56, 58–60, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/portwce.txt.

Dion:1993:CPC

[Dio93a] Pierre J. Dion. Cross-platform compression. Dr. Dobb’s Journal of Software Tools, 18(12): 32–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Dion:1993:CC

[Dio93b] Pierre J. Dion. Cross-platform compression. Dr. Dobb’s Journal of Software Tools, 18(13): 32, 34, 36–37, December 1993.

CODEN DDJOEB. ISSN 1044-789X.

Dipert:1995:IFM

[Dip95] Brian L. Dipert. Inside flash memory. Dr. Dobb’s Journal of Software Tools, 20(10):68, 70, 72, 74, 76, 78, 80, 82–83, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Dixit:1995:CSE

[Dix95] Saurabh Dixit. Creating specialeffect bitmaps. Dr. Dobb’s Journal of Software Tools, 20(??):41– ??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Daemen:1997:BCS

[DKR97] Joan Daemen, Lars R. Knudsen, and Vincent Rijmen. The block cipher square algorithm: A fast block cipher that uses a 128-bit key. Dr. Dobb’s Journal of Software Tools, 22(10):54, 56–57, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Dykstra:1998:NSD

[DL98] Dave Dykstra and Katherine Lato. NSBD and software distribution. Dr. Dobb’s Journal of Software Tools, 23(9):84, 87– 88, 97, September 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1998/1998_09/nsbd.
txt;   http://www.ddj.com/
ftp/1998/1998_09/nsbdsrc.
zip.

Ding:1999:MLA

[DL99] Zhimin Ding and Li Liu. Machine learning and agent-based computing. Dr. Dobb’s Journal of Software Tools, 24(11):88, 90, 92, 94, 96, November 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_11/mle.
txt;   http://www.ddj.com/
ftp/1999/1999_11/mle.zip.

Duimovich:1991:SES

[DM91] John Duimovich and Mike Milinkovich. Smalltalk and embedded systems. Dr. Dobb’s Journal of Software Tools, 16(10):86, 88, 90–95, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Dobbe:1995:FF

[Dob95] J. G. G. Dobbe. Faster FFTs.

Dr. Dobb’s Journal of Software Tools, 20(2):125–??, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Dolberg:1995:BDA

[Dol95]  Stan Dolberg.  Building

distributed applications with Galaxy. Dr. Dobb’s Journal of Software Tools, 20(3):80, 82, 84, 86–87, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Dorsey:1995:NCG

[Dor95]  John Dorsey.  Ninth com-

puter game developer’s conference. Dr. Dobb’s Developer Update, 2(6):3, June 1995. CODEN ???? ISSN 1079-8595.

Dorbie:1999:GC

[Dor99] Angus Dorbie. Gamma correction. Dr. Dobb’s Journal of Software Tools, 24(9):30, 32, 34–37, September 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_09/
gamma.txt;   http://www.ddj.
com/ftp/1999/1999_09/gamma.
zip.

Douglas:1990:EMM

[Dou90] Rohan T. Douglas. Error message management. Dr. Dobb’s Journal of Software Tools, 15 (1):48, 51, 108, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Dowgiallo:1993:DIA

[Dow93] Edward Dowgiallo. Database interoperability and application transportability. Dr. Dobb’s Journal of Software Tools, 18 (13):38, 40, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Duvanenko:1993:ASI

[DR93] Victor J. Duvanenko and W. E. Robbins. Algorithms for stereoscopic imaging. Dr. Dobb’s Journal of Software Tools, 18 (4):18–20, 22, 24, 26, 76, 78, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Dybiec:1997:WEA

[DR97] Scott Dybiec and Philip Rousselle. Web extensions and applications using FastCGI. Dr. Dobb’s Journal of Software Tools, 22(6):88–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Duvanenko:1990:ILS

[DRG90] Victor J. Duvanenko, W. E.

Robbins, and Ronald S. Gyurcsik. Improving line segment clipping. Dr. Dobb’s Journal of Software Tools, 15(7):36, 38, 40, 42, 44– 45, 98, 100, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Duvanenko:1996:LCR

[DRG96] Victor J. Duvanenko, W. E.

Robbins, and R. S. Gyurcsik. Line-segment clipping revisited. Dr. Dobb’s Journal of Software Tools, 21(1):107–110, 116, 117, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Drissel:1994:FIC

[Dri94] Warren E. Drissel. Forth interpreter compiler and debugger in pseudo-code [letter]. Dr. Dobb’s Journal of Software Tools, 19 (??):12–??, March or May?? 1994. CODEN DDJOEB. ISSN 1044-789X.

Drusinsky:1994:ESD

[Dru94] Doron Drusinsky. Extended state diagrams and reactive systems. Dr. Dobb’s Journal of Software Tools, 19(11):72, 74, 76, 78, 80, 106–107, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Drusinsky:1995:VDE

[Dru95] Doron Drusinsky. Visually designing embedded-systems applications. Dr. Dobb’s Journal of Software Tools, 20(6):62, 64, 66, 68, 104–106, June 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Silva:1995:ECV

[dS95] Rod da Silva. Examining CA-Visual objects. Dr. Dobb’s Journal of Software Tools, 20(10): 90–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Simone:1998:PTA

[dSW98] Mauricio de Simone and Gregory V. Wilson. Programmer’s toolchest: The Active Expressions library: Parallel programming with C++ operator overloading. Dr. Dobb’s Journal of Software Tools, 23 (8):76, 78–79, 92–94, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_08/
                                        active.txt;   http://www.
                                        ddj.com/ftp/1998/1998_08/
                                        active.zip.

Dudley:1990:PCP

[Dud90] William F. Dudley, Jr. Porting C programs to 80386 protected mode. Dr. Dobb’s Journal of Software Tools, 15(8):16–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Duffy:1995:PMC

[Duf95] Robert A. Duffy. A POP3 mail client using WinSock. Dr. Dobb’s Journal of Software Tools, 20(4):62–66, 103–104, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPa

[Dun90a] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(1):

140–??, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPb

[Dun90b] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(2): 134–??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPc

[Dun90c] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(3): 134–??, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPd

[Dun90d] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(4): 135–??, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPe

[Dun90e] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(5): 141–??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPf

[Dun90f] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(6): 143–??, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPg

[Dun90g] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(7): 139–??, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPh

[Dun90h] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(8): 161–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPi

[Dun90i] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(9): 137–??, September 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPj

[Dun90j] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(10): 139–??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPk

[Dun90k] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(11): 155–??, November 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Duntemann:1990:SPl

[Dun90l] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 15(12): 131–??, December 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Duncan:1991:LFM

[Dun91a] Ray Duncan. Looking into the future of microprocessors. Dr. Dobb’s Journal of Software Tools, 16(2):76–??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPa

[Dun91b] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(1): 140–??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPb

[Dun91c] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(2): 145–??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPc

[Dun91d] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(3): 119–??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPd

[Dun91e] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(4): 125–??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPe

[Dun91f] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(5): 123–??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPf

[Dun91g] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(6): 131–??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPg

[Dun91h] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(7): 123–??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPh

[Dun91i] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(8): 157–??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPi

[Dun91j] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(9): ??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPj

[Dun91k] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(10): ??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPk

[Dun91l] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(11): ??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1991:SPl

[Dun91m] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 16(12): ??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1992:PB

[Dun92a] Ray Duncan. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 17(11):179– ??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Dunham:1992:CTU

[Dun92b] Alan Dunham. Crash tracebacks in UNIX. Dr. Dobb’s Journal of Software Tools, 17(9):80, 82, 84, 86, 89, 113, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPa

[Dun92c] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(1): ??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPb

[Dun92d] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(2): ??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPc

[Dun92e] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPd

[Dun92f] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPe

[Dun92g] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPf

[Dun92h] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPg

[Dun92i] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPh

[Dun92j] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(8): ??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPi

[Dun92k] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(9): 135–??, September 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPj

[Dun92l] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(10): 145–??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPk

[Dun92m] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(11): 165–??, November 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Duntemann:1992:SPl

[Dun92n] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 17(12): 135–??, December 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Duncan:1993:PBa

[Dun93a] Ray Duncan. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18(2): 131–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1993:PBb

[Dun93b] Ray Duncan. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18(4):127–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1993:PBc

[Dun93c] Ray Duncan. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18(8): 143–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Dunlavey:1993:PTS

[Dun93d] Michael R. Dunlavey. Performance tuning: slugging it out! Dr. Dobb’s Journal of Software Tools, 18(12):18–20, 22, 24–26, 90–91, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1993:SPa

[Dun93e] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 18(1): 121–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1993:SPb

[Dun93f] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 18(2): 119–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1993:SPc

[Dun93g] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 18(3): 119–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:1993:SPd

[Dun93h] Jeff Duntemann. Structured programming. Dr. Dobb’s Journal of Software Tools, 18(4): 111–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Duncan:1995:PB

[Dun95] Ray Duncan. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20(12):133– ??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Durham:1997:SRL

[Dur97] Steve Durham. Sounding off with the RSX library. Dr. Dobb’s Journal of Software Tools, 22(7):66–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Dutton:1995:BMP

[Dut95] Brett Dutton. Bob as a macro processor library. Dr. Dobb’s Journal of Software Tools, 20 (5):64, 66–69, 104, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Dutt:1996:CWW

[Dut96] G. Dinesh Dutt. CGI and the World Wide Web. Dr. Dobb’s Journal of Software Tools, 21 (2):42, 44, 46, 48, 50, 92, 94– 96, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Duvanenko:1991:ERM

[Duv91] Victor J. Duvanenko. Efficiently raising matrices to an integer power. Dr. Dobb’s Journal of Software Tools, 16(6):86– 87, 157, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Douglass:1999:TMU

[DV99] Bruce Powel Douglass and Srini Vasan. Temporal models in UML. Dr. Dobb’s Journal of Software Tools, 24(12):74, 76, 78–80, 82, 85, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Dwyer:1995:IUB

[Dwy95] Nathan Dwyer. Implementing and using BSP trees. Dr. Dobb’s Journal of Software Tools, 20 (7):46, 48–49, July 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Davis:1991:CC

[DY91] Warren Davis and Kan Yabumoto. A coprocessor for a coprocessor? Dr. Dobb’s Journal of Software Tools, 16(5):16– 28, 84–88, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Dolvane:1999:PIAa

[DY99a] Jaison Dolvane and Kumanan Yogaratnam. PersonalJava and information appliances, part I. Dr. Dobb’s Journal of Software Tools, 24(1):60, 62, 64–68, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/ftp/1999/
1999_01/pjia1.txt.

Dolvane:1999:PIAb

[DY99b] Jaison Dolvane and Kumanan Yogaratnam. PersonalJava and information appliances, part II. Dr. Dobb’s Journal of Software Tools, 24(2):64, 66–68, 70–71, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ddj/1999/1999_02/9902f/9902f.
htm;   http://www.ddj.com/
ftp/1999/1999_02/pjphone.
txt;   http://www.ddj.com/
ftp/1999/1999_02/pjphone.
zip.

Easton:1999:JQD

[Eas99] Evan Easton. Java Q&A: Does Java support enums? Dr. Dobb’s Journal of Software Tools, 24(10):113–117, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/jqa1099.txt;   http://

www.ddj.com/ftp/1999/1999_
10/jqa1099.zip.

Eckel:1990:CTC

[Eck90] Bruce Eckel. Collections in Turbo C++. Dr. Dobb’s Journal of Software Tools, 15(8): 94–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Esposito:1995:GPE

[EJ95] Todd D. Esposito and Andrew K. Johnson. A generic parsing engine in C++. Dr. Dobb’s Journal of Software Tools, 20(??):37–??, July/ August 1995. CODEN DDJOEB. ISSN 1044-789X.

Ellis:1990:PEC

[Ell90] Graham K. Ellis. Parallel extensions to C. Dr. Dobb’s Journal of Software Tools, 15(8): 70–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Elliot:1993:THD

[Ell93a] J. Elliot. Tuning up Hyper-Card’s database engine. Dr. Dobb’s Journal of Software Tools, 18(4):39S–41S, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Elliott:1993:THD

[Ell93b] Jeff Elliott. Tuning up Hyper-Card’s database engine. Dr. Dobb’s Journal of Software Tools, 18(4 (supplement)):39S, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Entsminger:1990:EZI

[Ent90a] Gary Entsminger. Examining the Zinc Interface Library. Dr. Dobb’s Journal of Software Tools, 15(12):64, 66, 68, 70–71, 101–102, December 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Entsminger:1990:IOP

[Ent90b] Gary Entsminger. Inside object professional. Dr. Dobb’s Journal of Software Tools, 15(9):62– ??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Epstein:1996:CC

[Eps96] David Epstein. Conditional compilation. Dr. Dobb’s Journal of Software Tools, 21(5):44, 46, 80–83, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ea

[Eri90a] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 15(1):6–??, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Eb

[Eri90b] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 15(2):6–??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ec

[Eri90c] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 15(3):6–??, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ed

[Eri90d] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 15(4):6–??, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ee

[Eri90e] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 15(5):6–??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ef

[Eri90f] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 15(6):6–??, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Eg

[Eri90g] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 15(7):6–??, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Eh

[Eri90h] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 15(8):6–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ei

[Eri90i] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 15(9):6–??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ej

[Eri90j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 15(10):6–??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:Ek

[Eri90k] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 15(11):6–??, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1990:El

[Eri90l] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 15(12):6–??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Ericsson:1990:VVB

[Eri90m] Bo Ericsson. VESA VGA BIOS extensions. Dr. Dobb’s Journal of Software Tools, 15(4): 6511–6570, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:BDY

[Eri91a] Jonathan Erickson. Baby don’t you drive my car in “The Changing Landscape of Software Development”. Dr. Dobb’s Journal of Software Tools, 16(1 (supplement)):96S–?, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ea

[Eri91b] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 16(1):6–??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Eb

[Eri91c] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 16(2):6–??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ec

[Eri91d] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 16(3):6–??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ed

[Eri91e] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 16(4):6–??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ee

[Eri91f] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ef

[Eri91g] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 16(6):??, June 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Eg

[Eri91h] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 16(7):??, July 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Eh

[Eri91i] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 16(8):6–??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ei

[Eri91j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 16(9):??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ej

[Eri91k] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:Ek

[Eri91l] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 16(11):??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1991:El

[Eri91m] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 16(12):??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ea

[Eri92a] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Eb

[Eri92b] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ec

[Eri92c] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ed

[Eri92d] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ee

[Eri92e] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ef

[Eri92f] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Eg

[Eri92g] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Eh

[Eri92h] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ei

[Eri92i] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 17(9):8–??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ej

[Eri92j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 17(10):??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:Ek

[Eri92k] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 17(11):8–??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1992:El

[Eri92l] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 17(12):8–??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ea

[Eri93a] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 18(1):8–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Eb

[Eri93b] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 18(2):8–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ec

[Eri93c] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 18(3):8–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ed

[Eri93d] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 18(4):8–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ee

[Eri93e] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 18(5):8–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ef

[Eri93f] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 18(6):8–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Eg

[Eri93g] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 18(7):8–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Eh

[Eri93h] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 18(8):8–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ei

[Eri93i] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 18(9):6–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ej

[Eri93j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 18(10):4–??, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:Ek

[Eri93k] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 18(11):6–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:El

[Eri93l] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 18(12):??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:PBa

[Eri93m] Jonathan Erickson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18 (6):163–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1993:PBb

[Eri93n] Jonathan Erickson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18 (11):157–??, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1993:PBc

[Eri93o] Jonathan Erickson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18 (13):131–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:DHT

[Eri94a] Jonathan Erickson. Dialin’ for high-tech dollars. Dr. Dobb’s Developer Update, 1(10):2, December 1994. CODEN ???? ISSN 1079-8595.

Erickson:1994:Ea

[Eri94b] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(1):6–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Eb

[Eri94c] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(2):6–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Ec

[Eri94d] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(3):6–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Ed

[Eri94e] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 19(4):??, April 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Ee

[Eri94f] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 19(5):??, May 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Jonathan:1994:E

[Eri94g] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 19(5):6–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Ef

[Eri94h] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 19(6):6–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Eg

[Eri94i] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 19(7):6–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Eh

[Eri94j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 19(8):6–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:Ei

[Eri94k] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(11):6–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:ESB

[Eri94l]  Jonathan Erickson.  Edito-

rial: Not so strange bedfellows. Dr. Dobb’s Journal of Software Tools, 19(15):6–??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:EWT

[Eri94m] Jonathan Erickson. Editorial:

Who’s that tapping at your back door? Dr. Dobb’s Journal of Software Tools, 19(13): 6, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:PBa

[Eri94n] Jonathan Erickson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19 (3):141–??, March 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1994:PBb

[Eri94o] Jonathan Erickson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19 (7):133–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1994:TSL

[Eri94p] Jonathan Erickson. Time for a second look. Dr. Dobb’s Developer Update, 1(6):2–??, June 1994. CODEN ???? ISSN 1079-

Erickson:1995:DDJ

[Eri95a] Jonathan Erickson. Dr. Dobb’s Journal Excellence In Programming Awards. Dr. Dobb’s Journal of Software Tools, 20(3): 18–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ea

[Eri95b] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 20(1):6–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Eb

[Eri95c] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 20(2):??, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ec

[Eri95d] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 20(3):6–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ee

[Eri95e] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 20(5):??, May 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ed

[Eri95f] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 20(5):6–??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ef

[Eri95g] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(6):6–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Eg

[Eri95h] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(7):6–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Eh

[Eri95i] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(8):6–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ei

[Eri95j] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(9):??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ej

[Eri95k] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(10):??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:Ek

[Eri95l] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(11):??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:El

[Eri95m] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 20(12):??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1995:OBS

[Eri95n] Jonathan Erickson. OOPSLA ’95: Back in the saddle again. Dr. Dobb’s Developer Update, 2 (12):3, December 1995. CODEN ???? ISSN 1079-8595.

Erickson:1995:PB

[Eri95o] Jonathan Erickson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20 (9):141–??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ea

[Eri96a] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 21(1):??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Eb

[Eri96b] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 21(2):??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ec

[Eri96c] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 21(2):6–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ed

[Eri96d] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(3):??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ee

[Eri96e] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(4):??, April 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ef

[Eri96f] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(5):??, May 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Eg

[Eri96g] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(6):??, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Eh

[Eri96h] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(7):??, July 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ei

[Eri96i] Jonathan Erickson. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(8):??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ej

[Eri96j] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 21(9):??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:Ek

[Eri96k] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 21(12):6, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:EWN

[Eri96l] Jonathan Erickson. Editorial: Who needs your stinkin’ super shelves. Dr. Dobb’s Journal of Software Tools, 21(11): 6, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1996:DDJ

[Eri96m] Jonathan Erickson. Dr. Dobb’s Journal Excellence in Programming awards. Dr. Dobb’s Journal of Software Tools, 21(3): 16–17, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:DDJ

[Eri97a] Jonathan Erickson. Dr. Dobb’s Journal Excellence in Programming Awards. Dr. Dobb’s Journal of Software Tools, 22(5): 18–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:DDN

[Eri97b] Jonathan Erickson. Dr. Dobb’s news and views: Linux trademark issue settled; Inslaw ruling rejected; push over?; reading signs for the blind; free speech I; free speech II; Y2K insurance. Dr. Dobb’s Journal of Software Tools, 22(11): 16, November 1997. URL http:

//www.iplawyers.com/text/

linux.htm. A patent and trademark claim dispute on the name Linux has been resolved in favor of Linus Torvalds and the Linux community. A U.S. District Court Judge has ruled that source code is protected speech under the First Amendment to the U.S. Constitution, and that the U.S. Commerce Department acted illegally in requiring academics to obtains a government license before discussing cryptographic research with scholars on the Internet.

Erickson:1997:E

[Eri97c]  Jonathan Erickson.  Edito-

rial: Browsing for the best buy. Dr. Dobb’s Journal of Software Tools, 22(5):6, May 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1997:EFP

[Eri97d] Jonathan Erickson. Editorial:

Fair play. Dr. Dobb’s Journal of Software Tools, 22(4):6, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:EF

[Eri97e] Jonathan Erickson. Editorial:

Funny, as in “ha ha”. Dr. Dobb’s Journal of Software Tools, 22 (8):6, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:EGI

[Eri97f] Jonathan Erickson. Editorial: Giving it the reboot. Dr. Dobb’s Journal of Software Tools, 22 (11):6, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:EME

[Eri97g] Jonathan Erickson. Editorial: Mr. Erickson goes to Washington. Dr. Dobb’s Journal of Software Tools, 22(1):6, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:EPPb

[Eri97h]  Jonathan Erickson.  Edito-

rial: Past, present, and future. Dr. Dobb’s Journal of Software Tools, 22(3):6, March 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1997:EPPa

[Eri97i] Jonathan Erickson. Editorial: Power plays. Dr. Dobb’s Journal of Software Tools, 22(2): 6, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:ERRa

[Eri97j] Jonathan Erickson. Editorial: Random readings. Dr. Dobb’s Journal of Software Tools, 22 (7):6, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:ERRb

[Eri97k] Jonathan Erickson. Editorial: Rolling down the river — into 1998! Dr. Dobb’s Journal of Software Tools, 22(10):6, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:ERD

[Eri97l] Jonathan Erickson. Editorial:

The real deal. Dr. Dobb’s Journal of Software Tools, 22(12): 6, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:ETO

[Eri97m]  Jonathan Erickson.  Edito-

rial: There once was a programmer from Poughkeepsie. … Dr. Dobb’s Journal of Software Tools, 22(9):6, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:ETT

[Eri97n] Jonathan Erickson. Editorial:

Time traveling. Dr. Dobb’s Journal of Software Tools, 22 (6):6, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:DDJ

[Eri98a] Jonathan Erickson. Dr. Dobb’s Journal Excellence In Programming awards. Dr. Dobb’s Journal of Software Tools, 23(3): 20–21, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:EBC

[Eri98b] Jonathan Erickson. Editorial:

Banker, can you spare a dime? Dr. Dobb’s Journal of Software Tools, 23(6):6, June 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1998:EDM

[Eri98c] Jonathan Erickson. Editorial:

Dancing to the media mantra. Dr. Dobb’s Journal of Software Tools, 23(11):8, November 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:EFC

[Eri98d] Jonathan Erickson. Editorial: Fast, cheaps & two thumbs up. Dr. Dobb’s Journal of Software Tools, 23(9):6, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:EFL

[Eri98e] Jonathan Erickson. Editorial: Fear and loathing on the Y2K trail. Dr. Dobb’s Journal of Software Tools, 23(5):4, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:EFM

[Eri98f] Jonathan Erickson. Editorial: Following the money. Dr. Dobb’s Journal of Software Tools, 23 (12):8–??, December 1998. CO-DEN DDJOEB. ISSN 1044-789X. Extensive comments on the U.S. H-1B visa reform bill, officially known as the “American Competitiveness Act” and “Workforce Improvement and Protection Act of 1998”.

Erickson:1998:EHW

[Eri98g] Jonathan Erickson. Editorial: Help wanted. Dr. Dobb’s Journal of Software Tools, 23(4):6, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:ELL

[Eri98h] Jonathan Erickson. Editorial: Lessons learned. Dr. Dobb’s Journal of Software Tools, 23 (8):6, August 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:ELU

[Eri98i] Jonathan Erickson. Editorial:

Let us now praise famous stuff. Dr. Dobb’s Journal of Software Tools, 23(3):6, March 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1998:ETC

[Eri98j]  Jonathan Erickson.  Edito-

rial: Taking care of business. Dr. Dobb’s Journal of Software Tools, 23(10):8, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
10/index.htm.

Erickson:1998:EPD

[Eri98k]  Jonathan Erickson.  Edito-

rial: The planning department. Dr. Dobb’s Journal of Software Tools, 23(1):6, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1998:EYT

[Eri98l] Jonathan Erickson. Editorial:

You turn me on, I’m an Internet radio. Dr. Dobb’s Journal of Software Tools, 23(2):6, February 1998.

Erickson:1999:E

[Eri99a] Jonathan Erickson. Editorial:

2B or not 2B. Dr. Dobb’s Journal of Software Tools, 24 (1):8, January 1999. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:EBC

[Eri99b]  Jonathan Erickson.  Edito-

rial: Bumper cars at Mach 1. Dr. Dobb’s Journal of Software Tools, 24(3):8, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ddj/1999/1999_03/9903p/9903p.
htm.

Erickson:1999:ECCa

[Eri99c]  Jonathan Erickson.  Edito-

rial: Clear cutting the concept. Dr. Dobb’s Journal of Software Tools, 24(5):8, May 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:1999:ECCb

[Eri99d] Jonathan Erickson. Editorial: Coffee clashing. Dr. Dobb’s Journal of Software Tools, 24 (7):8, July 1999. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
articles/1999/9907/9907o/
9907o.htm.

Erickson:1999:EST

[Eri99e] Jonathan Erickson. Editorial: “I said turn RIGHT, fool!”. Dr. Dobb’s Journal of Software Tools, 24(11):8, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:ELM

[Eri99f] Jonathan Erickson. Editorial: Lego of my robot. Dr. Dobb’s Journal of Software Tools, 24 (4):8, April 1999. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:EML

[Eri99g] Jonathan Erickson. Editorial:

Moore’s law. Dr. Dobb’s Journal of Software Tools, 24(8):8, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:EOS

[Eri99h]  Jonathan Erickson.  Edito-

rial: Open season. Dr. Dobb’s Journal of Software Tools, 24 (2):8, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ddj/1999/1999_02/9902q/9902q.

htm. Comments on the GNU open software approach to software development, compared to the Microsoft model.

Erickson:1999:ESC

[Eri99i]  Jonathan Erickson.  Edito-

rial: Small can be beautiful. Dr. Dobb’s Journal of Software Tools, 24(10):8, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:EAS

[Eri99j] Jonathan Erickson. Editorial:

The art and science of cryptography. Dr. Dobb’s Journal of Software Tools, 24(12): 8, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1999:EN

[Eri99k] Jonathan Erickson. Editorial:

The negotiator. Dr. Dobb’s Journal of Software Tools, 24 (9):8, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Esposito:1996:MVB

[Esp96] Dino Esposito. Multithreading and Visual Basic. Dr. Dobb’s Journal of Software Tools, 21 (12):46–??, December 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Esposito:1997:IIS

[Esp97] Dino Esposito. Inside Iterated Systems’ Fractal Development Kit — the Fractal Development Kit from Iterated Systems is a library that makes it possible for you to embed fractal-imaging capabilities into C/C++ Windows and Macintosh applications. Dr. Dobb’s Journal of Software Tools, 22 (1):74–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Fagan:1998:ACU

[Fag98a] Sean Eric Fagan. Asynchronous communications using select and poll. Dr. Dobb’s Journal of Software Tools, 23(9): 48, 50–51, 93–95, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
09/bsdasync.txt.

Fagan:1998:TBS

[Fag98b] Sean Eric Fagan. Tracing BSD system calls. Dr. Dobb’s Journal of Software Tools, 23(3): 38, 40, 42–43, 105, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1998/1998_03/
truss.txt;   http://www.ddj.

com/ftp/1998/1998_03/truss.
zip.

Faison:1993:HSL

[Fai93] Ted Faison. Horizontally scrollable listboxes. Dr. Dobb’s Journal of Software Tools, 18 (10):47–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Faison:1995:BCO

[Fai95] Ted Faison. Borland C++ 4.5 and OLE 2.0 programming. Dr. Dobb’s Journal of Software Tools, 20(4):68–??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Faison:1996:EBD

[Fai96] Ted Faison. Examining Borland Delphi 2.0. Dr. Dobb’s Journal of Software Tools, 21(6):64, 66–67, 70–71, 87–88, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Faler:1998:BSC

[Fal98] Wes Faler. Building solid code.

Dr. Dobb’s Journal of Software Tools, 23(11):106, 108– 110, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_11/
SOLID.TXT.

Felton:1992:BOD

[Fel92] Sam Felton. BLOBs and objectoriented database engines. Dr. Dobb’s Journal of Software Tools, 17(12):90, 92, 94, 120– 121, 158, December 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Felsenstein:1993:CI

[Fel93] Lee Felsenstein. The commons of information. Dr. Dobb’s Journal of Software Tools, 18(5):18– 20, 22, 24, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Ferrari:1994:RSA

[Fer94] Alberto Daniel Ferrari. Realtime scheduling algorithms. Dr. Dobb’s Journal of Software Tools, 19(15):60, 62, 64, 66, 94, 96, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Ferris:1995:PAM

[Fer95] Jeffrey Ferris. PowerPC address munging [letter]. Dr. Dobb’s Journal of Software Tools, 20 (1):10–??, January 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Fernandez:1998:PBO

[Fer98]  Eduardo Fernandez.  Pro-

grammer’s bookshelf: Objectoriented design. Dr. Dobb’s Journal of Software Tools, 23 (12):135, 137, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Fernandes:1999:ECC

[Fer99] Andrew D. Fernandes. Ellipticcurve cryptography. Dr. Dobb’s Journal of Software Tools, 24 (12):56, 58, 60–63, December 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        12/ellip.zip.

Fairman:1995:CDP

[FH95] William Fairman and Randal Hoff. Cross-platform database programming. Dr. Dobb’s Journal of Software Tools, 20(3):36, 38, 40, 42, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Friedman-Hill:1997:IAD

[FHW97] Ernest J. Friedman-Hill and Robert A. Whiteside. IDLDOC: Automatic documentation for CORBA IDL. Dr. Dobb’s Journal of Software Tools, 22(6): 46–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Finger:1995:LTC

[Fin95a] Jonathan Finger. Lightweight tasks in C. Dr. Dobb’s Journal of Software Tools, 20(5):48, 50, 102, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Finnegan:1995:IGW

[Fin95b] James Finnegan. Implementing games for Windows. Dr. Dobb’s Journal of Software Tools, 20 (??):42–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Firoozye:1993:CPT

[Fir93a]  Ramin Firoozyé.  A cross-

platform plug-in toolkit. Dr. Dobb’s Journal of Software Tools, 18(6):120, 122–125, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Firoozye:1993:ESA

[Fir93b] Ramin Firoozyé. Examining the StarView application framework. Dr. Dobb’s Journal of Software Tools, 18(13):78–80, 82, 84, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Flanigan:1998:NS

[FK98] Pat Flanigan and Jawed Karim. NCSA Symera. Dr. Dobb’s Journal of Software Tools, 23 (11):20–22, 24, 26–27, 32– 33, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_11/
SYMERA.TXT.

Flamig:1996:PAC

[Fla96] Bryan Flamig. Practical algorithms in C++. Dr. Dobb’s Journal of Software Tools, 21 (2):139–??, February 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Fleming:1993:CPL

[Fle93] Jim Fleming. The C+@ programming language. Dr. Dobb’s Journal of Software Tools, 18 (11):24–32, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1990:ED

[Flo90a] Edwin T. Floyd. An existential dictionary. Dr. Dobb’s Journal of Software Tools, 15(11):20, 22, 24, 28, 30, 32, 110–112, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1990:RYO

[Flo90b] Michael A. Floyd. Roll your own object-oriented language. Dr. Dobb’s Journal of Software Tools, 15(11):16–18, 102– 104, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1991:E

[Flo91a]  Michael A. Floyd.  Edito-

rial. Dr. Dobb’s Journal of Software Tools, 16(12 (supplement)):??, December 1991. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Floyd:1991:ECB

[Flo91b] Michael A. Floyd. The evolution of component-based programming in “The Changing Landscape of Software Development”. Dr. Dobb’s Journal of Software Tools, 16(1 (supplement)):96S, January 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1991:OD

[Flo91c] Michael A. Floyd. The object d’art. Dr. Dobb’s Journal of Software Tools, 16(10): 52–??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1991:WMW

[Flo91d] Michael A. Floyd. What? me worry about Windows programming? Dr. Dobb’s Journal of Software Tools, 16(12 (supplement)):4–??, December 1991. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Floyd:1992:ETP

[Flo92] Michael A. Floyd. Examining Turbo Pascal for Windows.

Dr. Dobb’s Journal of Software Tools, 17(11):94, 97, 99– 100, 145–146, 148–149, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1993:COL

[Flo93a]  Michael Floyd.  Comparing

object-oriented languages. Dr. Dobb’s Journal of Software Tools, 18(11):104–118, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1993:BDF

[Flo93b] Michael A. Floyd. Building a database file viewer. Dr. Dobb’s Journal of Software Tools, 18 (4 (supplement)):25S–31S, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1993:COO

[Flo93c] Michael A. Floyd. Comparing object-oriented languages. Dr. Dobb’s Journal of Software Tools, 18(10):104–??, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1993:E

[Flo93d] Michael A. Floyd. Editorial. Dr. Dobb’s Journal of Software Tools, 18(4 (supplement)):??, April 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Florence:1994:DVF

[Flo94a] Mark Florence. Digital video file formats. Dr. Dobb’s Journal of Software Tools, 19(??):10–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1994:BIC

[Flo94b] Michael A. Floyd. Borland International Conference. Dr. Dobb’s Developer Update, 1(7): 2, September 1994. CODEN ???? ISSN 1079-8595.

Floyd:1994:BEM

[Flo94c] Michael A. Floyd. Building an email manager. Dr. Dobb’s Journal of Software Tools, 19(15): 82, 84–85, 100, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1994:CLT

[Flo94d] Michael A. Floyd. A conversation with Linus Torvalds. Dr. Dobb’s Journal of Software Tools, 19(5):27–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1994:Ea

[Flo94e] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(??):??, ???? 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1994:Eb

[Flo94f] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(??):??, ???? 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1994:Ec

[Flo94g] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 19(??):??, ???? 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1994:WPP

[Flo94h] Michael A. Floyd. Programmer’s bookshelf: Windows Programming Power with Custom Controls, by Paul Cilwa and Jeff Duntemann [review]. Dr. Dobb’s Journal of Software Tools, 19 (6):133–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1994:WAW

[Flo94i] Michael A. Floyd. Why ask why. Dr. Dobb’s Journal of Software Tools, 19(??):3–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1995:BDC

[Flo95a] Michael A. Floyd. Borland developer’s conference. Dr. Dobb’s Developer Update, 2(10): 2–3, October 1995. CODEN ???? ISSN 1079-8595.

Floyd:1995:Ea

[Flo95b] Michael A. Floyd. Editorial. Dr. Dobb’s Journal of Software Tools, 20(7):??, July 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1995:Eb

[Flo95c] Michael A. Floyd. Editorial. Dr. Dobb’s Journal of Software Tools, 20(8):??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1995:Ec

[Flo95d] Michael A. Floyd. Editorial. Dr. Dobb’s Journal of Software Tools, 20(9):??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1995:Ed

[Flo95e] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 20(10):??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1995:EVB

[Flo95f] Michael A. Floyd. Extending Visual Basic’s comm control. Dr. Dobb’s Journal of Software Tools, 20(12):44, 46, 48–49, 105– 106, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1995:RA

[Flo95g] Michael A. Floyd. Rebel alliance. Dr. Dobb’s Journal of Software Tools, 20(??):4–??, September/October 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Floyd:1996:Ea

[Flo96a] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(1):??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1996:Eb

[Flo96b] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(2):??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1996:Ec

[Flo96c] Michael A. Floyd. Editorial.

Dr. Dobb’s Journal of Software Tools, 21(3):??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Floyd:1996:Ed

[Flo96d] Michael A. Floyd. Editorial. Dr. Dobb’s Journal of Software Tools, 21(4):??, April 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Flynn:1996:LEA

[Fly96] Peter Flynn. Letter to the editor: Automating HTML conversion. Dr. Dobb’s Journal of Software Tools, 21(10):8, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Flynn:1997:ESO

[Fly97] James Flynn. Embedded systems — optimizing microcontroller performance: Getting around a memcpy bottleneck. Dr. Dobb’s Journal of Software Tools, 22(11):62, 64, 66, 68, 70, 73, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Forte:1998:BPD

[FN98] Derrick B. Forte and Hai T. Nguyen. 68HC05-based peripheral devices: Part I. Dr. Dobb’s Journal of Software Tools, 23 (6):66, 68, 70, 72–73, June 1998. CODEN DDJOEB. ISSN 1044-789X.

FridNielsen:1991:CPT

[FNL91a] L. Frid-Nielsen and A. Lane. Celestial programming with Turbo Pascal (CCD sensor). Dr. Dobb’s Journal of Software Tools, 16(6):80, 82–84, 106–112, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Frid-Nielsen:1991:CPT

[FNL91b] Lars Frid-Nielsen and Alex Lane. Celestial programming with Turbo Pascal. Dr. Dobb’s Journal of Software Tools, 16 (6):80–??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Folger:1994:UD

[Fol94] Troy Folger. Undocumented OS/2: DosQProcStatus. Dr. Dobb’s Journal of Software Tools, 19(8):125–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Fomitchev:1999:MTC

[Fom99]  Max I. Fomitchev.  MMX

technology code optimization. Dr. Dobb’s Journal of Software Tools, 24(9):38, 40, 42– 46, 48, September 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_09/mmx.
txt;   http://www.ddj.com/
ftp/1999/1999_09/mmx.zip.

Ford:1994:AC

[For94] Dan Ford. Associations in C++. Dr. Dobb’s Journal of Software Tools, 19(8):62, 64, 66, 68, 96, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Ford:1995:ETC

[For95] Dan Ford. Event-driven threads in C++. Dr. Dobb’s Journal of Software Tools, 20(6):48–50, 52, 54, 98, 100, 102, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Ford:1998:MES

[For98a] Dan Ford. Mutual exclusion and synchronization in Java: Mimicking Win32 synchronization mechanisms. Dr. Dobb’s Journal of Software Tools, 23(1):62, 64, 66, 68–70, 73–75, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Fortner:1998:HHD

[For98b] Brand Fortner. HDF: The Hierarchical Data Format. Dr. Dobb’s Journal of Software Tools, 23(5):42, 44–48, 109–110, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        05/hdf.txt.

Foshay:1992:TYA

[Fos92] Laird Foshay. Ten years after: Still reinventing the wheel. Dr. Dobb’s Journal of Software Tools, 17(9):16–??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Fosner:1995:PO

[Fos95a] Ron Fosner. Programming with OpenGL. Dr. Dobb’s Journal of Software Tools, 20(7):86, 88, 90– 96, 98–107, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Fosner:1995:VRW

[Fos95b] Ron Fosner. Virtual reality and the WorldToolKit for Windows. Dr. Dobb’s Journal of Software Tools, 20(1):78, 80, 82, 102– 105, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Fosner:1996:POP

[Fos96] Ron Fosner. Programming with OpenGL primitives. Dr. Dobb’s Journal of Software Tools, 21 (??):42–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Fedorov:1993:WRF

[FR93] Alex G. Fedorov and Dmitry M.

Rogatkin. The Windows .RES file format. Dr. Dobb’s Journal of Software Tools, 18(8): 133–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Franz:1990:PTA

[Fra90] Marty Franz. Programmer tools for Actor 3.0. Dr. Dobb’s Journal of Software Tools, 15(11):86, 88–90, 132, 134–135, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Franks:1991:AEL

[Fra91] Neville Franks. Adding an extension language to your software. Dr. Dobb’s Journal of Software Tools, 16(9):34, 36, 38, 40, 42–43, September 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Frazier:1999:PCC

[Fra99] George F. Frazier. Porting C++ code from NT to UNIX: Using the MainWin XDE toolkit. Dr. Dobb’s Journal of Software Tools, 24(4):84, 86–89, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
04/nt2unix.txt;   http://

                                        www.ddj.com/ftp/1999/1999_
                                        04/nt2unix.zip.

Free:1992:PCP

[Fre92] Gordon Free. Programming with communication protocol stacks. Dr. Dobb’s Journal of Software Tools, 17(3):72, 74–75, 77–78, 92, 94, 96, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Freeman:1995:PRM

[Fre95] Steve Freeman. Partial revelation and Modula-3. Dr. Dobb’s Journal of Software Tools, 20 (10):36, 38, 40, 42, 110, 112, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Fried:1990:AHPa

[Fri90a] Stephen Fried. Accessing hardware from 80386 protected mode: Part I. Dr. Dobb’s Journal of Software Tools, 15 (5):92–98, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Fried:1990:AHPb

[Fri90b] Stephen S. Fried. Accessing hardware from 80386 protected mode: Part II. Dr. Dobb’s Journal of Software Tools, 15(6):78– 84, 122–125, June 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Fried:1994:SMP

[Fri94] Stephen S. Fried. Shared memory and PC supercomputing. Dr. Dobb’s Journal of Software Tools, 19(1):18–20, 22, 24, 26, 28, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Fried:1995:PON

[Fri95] Stephen S. Fried. Pentium optimizations and numeric performance. Dr. Dobb’s Journal of Software Tools, 20(1):18–20, 22, 26–29, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Fan:1998:CEM

[FSW98] Ming Fan, Jan Stallaert, and Andrew B. Whinston. Creating electronic markets. Dr. Dobb’s Journal of Software Tools, 23 (11):52, 54–57, November 1998. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1998/1998_11/fbts.
txt.

Frick:1999:SGP

[FSW99] Arne Frick, Georg Sander, and Kathleen Wang. Simulating graphs as physical systems. Dr. Dobb’s Journal of Software Tools, 24(8):58, 60–64, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Fulco:1992:QAC

[Ful92] William Fulco. The QuickTime/ AVK connection. Dr. Dobb’s Journal of Software Tools, 17 (7):28, 30, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Farley:1993:NNP

[FV93a] James F. Farley and Peter D.

Varhol. Neural nets for predicting behavior (of winds). Dr. Dobb’s Journal of Software Tools, 18(2):82, 84, 86, 88, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Farley:1993:VAD

[FV93b] James F. Farley and Peter D. Varhol. A visual approach to data acquisition. Dr. Dobb’s Journal of Software Tools, 18 (5):145–147, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Farley:1993:VDR

[FV93c] James F. Farley and Peter D. Varhol. Visualizing data in real time. Dr. Dobb’s Journal of Software Tools, 18(13):58, 62, 64, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Fitzpatrick:1997:PBS

[FV97] Michael E. Fitzpatrick and Laurence Vanhelsuwé. Programmer’s bookshelf: The software project and Unicode. Dr. Dobb’s Journal of Software Tools, 22 (8):125, 127, August 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Frazier:1996:RBN

[FY96] George F. Frazier and Derek Yenzer. Removing blocking network I/O from Windows programs. Dr. Dobb’s Journal of Software Tools, 21(6):56, 58–60, 86–87, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Gabriel:1992:PPE

[Gab92] Richard P. Gabriel. Persistence in A programming environment. Dr. Dobb’s Journal of Software Tools, 17(12):46, 48, 50, 52, 54– 55, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Gablon:1996:PB

[Gab96] Dean Gablon. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 21(2): 139–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Gadre:1996:UPA

[Gad96] Dhananjay V. Gadre. Using the parallel adapter as a host interface port. Dr. Dobb’s Journal of Software Tools, 21(4):60, 62, 64, 66, 68, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Gadre:1997:AAM

[Gad97]  Dhananjay Ve Gadre.  At-

mel’s AT89C2051 microcontroller. Dr. Dobb’s Journal of Software Tools, 22(7):46–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Gallmeister:1991:RUA

[Gal91] William O. Gallmeister. Reconciling UNIX, Ada, and real-time processing. Dr. Dobb’s Journal of Software Tools, 16(6):56, 58, 60, 62, 64, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Gameiro:1996:NID

[Gam96a] Gil Gameiro. Networking intelligent devices. Dr. Dobb’s Journal of Software Tools, 21(2):68, 70, 72, 74, 104–105, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Gams:1996:CSD

[Gam96b] Richard A. Gams. A client/ server DBMS for managing clinical data. Dr. Dobb’s Journal of Software Tools, 21(??):29–??, January/February 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Garfinkel:1991:DWF

[Gar91]  Simson L. Garfinkel.  De-

signing a write-once file system (a general-purpose optical storage software technology). Dr. Dobb’s Journal of Software Tools, 16(1):78, 80, 82–86, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Garrett:1996:ENC

[Gar96] Doug Garrett. Examining the Nutmeg class library. Dr. Dobb’s Journal of Software Tools, 21 (10):80–??, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Gaspar:1990:MCT

[Gas90]  Don Gaspar.  The Macin-

tosh communications toolbox. Dr. Dobb’s Journal of Software Tools, 15(12):38, 40–41, 44, 94– 96, 98–99, December 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Gathman:1997:TUJ

[Gat97] Stuart D. Gathman. A text UI for the Java AWT. Dr. Dobb’s Journal of Software Tools, 22(9):38, 40–42, 45, 84, 86, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Graham-Cumming:1998:VDC

[GC98] John Graham-Cumming. VerCheck: Discovering component version numbers. Dr. Dobb’s Journal of Software Tools, 23(3):99–102, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
03/vercheck.txt;   http://
www.ddj.com/ftp/1998/1998_
03/vercheck.zip.

Gadre:1998:DAS

[GE98] Dhananjay V. Gadre and Sunu Engineer. A data acquisition system for Linux. Dr. Dobb’s Journal of Software Tools, 23 (2):62, 64–67, 101, February 1998.

Gerencir:1993:MAM

[Ger93] Ivan Gerencir. A multi-app Message Trace Facility for Windows. Dr. Dobb’s Journal of Software Tools, 18(11):88–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Gessner:1990:BHS

[Ges90] Rick Gessner. Building a hypertext system. Dr. Dobb’s Journal of Software Tools, 15(6):22– 28, 31, 33, 86, 88, 90, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1999:PBC

[GF99] Lou Grinzo and Laryn Fernandes. Programmer’s bookshelf: Clarifying the Open Source movement. Dr. Dobb’s Journal of Software Tools, 24 (9):119–120, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1995:BIW

[GG95a] Lou Grinzo and Steve Gallagher. Books from inside the walls. Dr. Dobb’s Journal of Software Tools, 20(7):??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1995:PB

[GG95b] Lou Grinzo and Steve Gallagher. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20(7):137–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Guckenberger:1995:ATH

[GGL^{+}95] Dutch Guckenberger, Liz Guckenberger, Frank Luongo, Kay Stanney, and Jose Sepulveda. Above-real-time training and the hyper-time algorithm. Dr. Dobb’s Journal of Software Tools, 20(4):52, 54, 56, 58, 60– 61, 100, 102, April 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Gamma:1995:DOE

[GH95a] Erich Gamma and Richard Helm. Designing objects for extension. Dr. Dobb’s Journal of Software Tools, 20(??): 56–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Gamma:1995:OO

[GH95b] Erich Gamma and Richard Helm. Observations on observer. Dr. Dobb’s Journal of Software Tools, 20(??):59–??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

Gamma:1996:CP

[GH96] Erich Gamma and Richard Helm. The Courier pattern. Dr. Dobb’s Journal of Software Tools, 21(??):55–??, January/February 1996. CODEN DDJOEB. ISSN 1044-789X.

Gontmakher:1999:AA

[GH99] Sasha Gontmakher and Ilan Horn. Algorithm alley: Efficient memory allocation. Dr. Dobb’s Journal of Software Tools, 24(1):116–119, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/ftp/1999/
1999_01/aa199.txt;   http:
//www.ddj.com/ddj/ftp/1999/
1999_01/aa199.zip.

Gamma:1996:WEN

[GHJV96] Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. Warren examines the now-classic Design Patterns: Elements of Reusable Objectoriented Software. Dr. Dobb’s Journal of Software Tools, 21 (10):134–??, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Gibson:1994:CMC

[Gib94] Ken Gibson. A C++ multitasking class library. Dr. Dobb’s Journal of Software Tools, 19 (5):28, 30, 32, 34, 96–98, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Giguere:1993:RWS

[Gig93a] Eric Giguere. REXX and the OS/2 workplace shell. Dr. Dobb’s Journal of Software Tools, 18(12):44–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Giguere:1993:ROW

[Gig93b] Eric Giguere. REXX and the OS/2 Workplace Shell. Dr. Dobb’s Journal of Software Tools, 18(13):44, 46, 48, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Giguere:1997:JBN

[Gig97] Eric Giguère. Java Beans and the new event model. Dr. Dobb’s Journal of Software Tools, 22 (4):50, 52, 53, 79, 80, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Giles:1992:CDF

[Gil92] Bruce (Bear) Giles. Contouring data fields. Dr. Dobb’s Journal of Software Tools, 17(6):44, 46, 91, 93–95, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Gillig:1994:ES

[Gil94a] James R. Gillig. Endian-neutral software. Dr. Dobb’s Journal of Software Tools, 19(13):62– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Gillig:1994:ENSa

[Gil94b] James R. Gillig. Endian-neutral software, part 1. Dr. Dobb’s Journal of Software Tools, 19 (11):62, 64, 68–70, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Gillig:1994:ENSb

[Gil94c] James R. Gillig. Endian-neutral software. part 2. Dr. Dobb’s Journal of Software Tools, 19 (13):44, 46–49, 51, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Gillig:1995:PBEa

[Gil95a] James R. Gillig. PowerPC biendian capabilities [sidebar] in ‘Bit Operations with C Macros’. Dr. Dobb’s Journal of Software Tools, 20(9):43–??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Gillig:1995:PBEb

[Gil95b] James R. Gillig. PowerPC biendian capabilities [sidebar] in ‘Bit Operations with C Macros’. Dr. Dobb’s Journal of Software Tools, 20(10):43–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Gintis:1991:XEM

[Gin91] Herbert Gintis. Xalloc (an expanded memory manage for Turbo Pascal). Dr. Dobb’s Journal of Software Tools, 16(10):74, 121–122, October 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Gislason:1993:FFS

[Gis93] Drew Gislason. Flash file systems. Dr. Dobb’s Journal of Software Tools, 18(5):54, 56, 58– 59, 94, 96–97, May 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Goldman:1994:SSG

[GK94] Jonny Goldman and Bob Kaehms. sGs: A simple Gopher server. Dr. Dobb’s Journal of Software Tools, 19(??):36–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Glass:1991:PUA

[Gla91] David N. Glass. Porting UNIX applications to DOS. Dr. Dobb’s Journal of Software Tools, 16(11):68, 70–72, 76, 78, 80, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Gray:1995:EMP

[GM95] Rand Gray and Deepak Mulchandani. Efficient MC68HC08 programming. Dr. Dobb’s Journal of Software Tools, 20(5):70, 72, 74, 76, 78, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Gray:1997:OFF

[GM97] Rand Gray and Deepak Mulchandani. Object file formats. Dr. Dobb’s Journal of Software Tools, 22(5):47–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Goalby:1999:SWS

[Goa99] John Goalby. Scripts for Windows Scripting Host. Dr. Dobb’s Journal of Software Tools, 24 (1):86, 88–92, January 1999. CODEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ddj/ftp/1999/1999_01/
                                        wsh.txt;   http://www.ddj.
                                        com/ddj/ftp/1999/1999_01/
                                        wsh.zip.

Goldberg:1990:IMV

[Gol90] Adele Goldberg. Information models, views, and controllers (software re-use). Dr. Dobb’s Journal of Software Tools, 15 (7):54, 56–59, 61, 106–107, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Goley:1996:WPV

[Gol96] George F. Goley IV. Web programming with Visual FoxPro. Dr. Dobb’s Journal of Software Tools, 21(12):80–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Gorman:1993:BPP

[Gor93] Ian E. Gorman. Building A portable programming environment. Dr. Dobb’s Journal of Software Tools, 18(5):76, 78, 80– 81, 106–107, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Gorman:1996:LY

[Gor96] Ian E. Gorman. Lex and yacc.

Dr. Dobb’s Journal of Software Tools, 21(2):86–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Goudie:1996:DIG

[Gou96] Nate Goudie. Designing isometric game environments. Dr. Dobb’s Journal of Software Tools, 21(??):13–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Grano:1998:PEK

[GPVV98] Kari Granö, Jukka Paakki, Jukka Viljamaa, and Antti Viljamaa. Protocol engineering and Kannel. Dr. Dobb’s Journal of Software Tools, 23(4):46, 48, 50, 52–53, 100, April 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/1998/1998_03/../.
./1994/1994_10/drus/drus.
htm;   http://www.ddj.com/
ftp/1998/1998_04/kannel.txt;
 http://www.ddj.com/ftp/1998/
1998_04/kannel1.zip;   http:
//www.ddj.com/ftp/1998/1998_
04/kannel2.zip.

Gothard:1998:SSY

[GR98] William Gothard and Les Rodner. Strategies for solving the Y2K problem. Dr. Dobb’s Journal of Software Tools, 23 (5):26, 28, 30, 32, 109, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
05/straty2k.txt.

Grappel:1991:OID

[Gra91] Robert D. Grappel. Optimizing integer division by a constant divisor. Dr. Dobb’s Journal of Software Tools, 16(2):80, 82–84, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Graf:1994:MIF

[Gra94] Joseph J. Graf. Multiplatform .INI files. Dr. Dobb’s Journal of Software Tools, 19(3):36, 38, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Gray:1997:ATM

[Gra97] Robert S. Gray. Agent Tcl. Dr. Dobb’s Journal of Software Tools, 22(3):18–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Graef:1998:PBC

[Gra98] Gerald L. Graef. Programmer’s bookshelf: From C to Tcl. Dr. Dobb’s Journal of Software Tools, 23(2):133–134, February 1998.

Green:1992:CDV

[Gre92] James L. Green. Capturing digital video using DVI. Dr. Dobb’s Journal of Software Tools, 17 (7):16–18, 20, 22, 24, 90, 92–95, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1996:PBa

[Gri96a]  Lou Grinzo.  Programmer’s

bookshelf. Dr. Dobb’s Journal of Software Tools, 21(3):127–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1996:PBb

[Gri96b]  Lou Grinzo.  Programmer’s

bookshelf. Dr. Dobb’s Journal of Software Tools, 21(9):117– ??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Griffin:1997:LRC

[Gri97a]  Bennett Griffin.  LUCA:

Reusable communication code: Bridging the media and protocol gap. Dr. Dobb’s Journal of Software Tools, 22(10):98, 100–102, 104, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1997:EPN

[Gri97b] Lou Grinzo. Extended precision native integers for Java: When performance counts. Dr. Dobb’s Journal of Software Tools, 22 (11):42, 44–45, 91–92, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1999:JQH

[Gri99] Lou Grinzo. Java Q&A: How do you run untrusted classes? Dr. Dobb’s Journal of Software Tools, 24(5):121–123, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1999/1999_05/
jqa599.txt;   http://www.
ddj.com/ftp/1999/1999_05/
jqa599.zip.

Grosberg:1997:DGH

[Gro97a] John A. Grosberg. Design guidelines for Is-A hierarchies. Dr. Dobb’s Journal of Software Tools, 22(6):36–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Grosso:1997:DDP

[Gro97b] William Grosso. Dynamic design patterns in Objective-C. Dr. Dobb’s Journal of Software Tools, 22(8):38, 40–44, 91– 93, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Gruber:1995:AS

[Gru95] Diana Gruber. Attached sprites. Dr. Dobb’s Journal of Software Tools, 20(??):29–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Gosselink:1994:AAG

[GS94] Pieter Gosselink and Bruce Schneier. Algorithm alley: The Gosselink ditherer. Dr. Dobb’s Journal of Software Tools, 19 (15):117–119, 133, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Gadre:1997:ESE

[GS97] Dhananjay V. Gadre and Larry A. Stein. Embedded systems — the enhanced parallel printer port for fast I/O: Revisiting the parallel printer adapter. Dr. Dobb’s Journal of Software Tools, 22(10):66, 68, 70, 73, 75–76, 78, 80–81, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Golomb:1998:JQH

[GS98] Kenneth Golomb and Thomas Sorgie. Java Q&A: How do I ensure secure communications from a Java applet? Dr. Dobb’s Journal of Software Tools, 23 (6):107–109, 124, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Gulutzan:1995:RZC

[Gul95] Peter Gulutzan. Programmer’s bookshelf: Review of Zen of Code Optimization, by Michael Abrash, in “Making Programs Go Faster”. Dr. Dobb’s Journal of Software Tools, 20(1): 133–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Gunther:1998:AAD

[Gun98] John C. Gunther. Algorithm alley: Discontiguous exponential averaging. Dr. Dobb’s Journal of Software Tools, 23 (9):117–119, 125, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        09/aa998.txt.

Gurley:1995:DBF

[Gur95] George Gurley. Do books have a future? Dr. Dobb’s Developer Update, 2(4):3, April 1995. CO-DEN ???? ISSN 1079-8595.

Guthery:1992:CPL

[Gut92] Scott B. Guthery. A curmudgery on programming language trends. Dr. Dobb’s Journal of Software Tools, 17(12): 104–105, 140–141, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Guthery:1994:EEE

[Gut94] Scott B. Guthery. EchoNets, Ememes, and extended realities. Dr. Dobb’s Journal of Software Tools, 19(4):72, 74, 76, 78, 80, 82, 84, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Gutman:1999:AAS

[Gut99] Ron Gutman. Algorithm alley: Space-filling curves in geospatial applications. Dr. Dobb’s Journal of Software Tools, 24(7):115–118, 120– 121, July 1999. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        articles/1999/9907/9907m/
                                        9907m.htm;   http://www.ddj.
                                        com/ftp/1999/1999_07/aa799.
                                        txt.

Goldberg:1996:RNB

[GW96] Ian Goldberg and David Wagner. Randomness and the Netscape browser. Dr. Dobb’s Journal of Software Tools, 21 (1):66, 68–70, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Gorda:1997:IPB

[GW97] Brent Gorda and Gregory V.

Wilson. Internet programming — building and running online auctions. Dr. Dobb’s Journal of Software Tools, 22(10):84, 86– 88, 91, 104, October 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Garcia:1996:MDR

[GWL96] Tony Garcia, Jeff Woodard, and Craig Liddell. Monitoring data in real time. Dr. Dobb’s Journal of Software Tools, 21(7):56, 58, 60, 62, 64, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Haakenson:1995:EVV

[Haa95] Brad Haakenson. Examining the VESA VBE 2.0 specification. Dr. Dobb’s Journal of Software Tools, 20(7):80–83, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Hagerty:1991:MCS

[Hag91]  Michael Hagerty.  Making

a CASE for software design. Dr. Dobb’s Journal of Software Tools, 16(1):108, 110–112, 114, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Hall:1993:M

[Hal93] Valerie Hall. Morphing in 2-D and 3-D. Dr. Dobb’s Journal of Software Tools, 18(7):18–20, 22, 24, 26, 132, 92–93, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Hamilton:1993:DMA

[Ham93] Kenneth G. Hamilton. Direct memory access from PC Fortrans. Dr. Dobb’s Journal of Software Tools, 18(5):70, 72–74, 102–104, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Hamilton:1995:BAI

[Ham95a] Jeffrey W. Hamilton. Basic arithmetic with infinite integers. Dr. Dobb’s Journal of Software Tools, 20(1):44–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Hamilton:1995:UFP

[Ham95b] Kenneth G. Hamilton. Undocumented features of PC Fortran libraries. Dr. Dobb’s Journal of Software Tools, 20(1):30, 32, 34, 84–85, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Hamstra:1999:XC

[Ham99]  Dirk Hamstra.  XML and

CORBA. Dr. Dobb’s Journal of Software Tools, 24(11): 98, 100, November 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_11/
                                        xmlcorba.txt.

Handsman:1993:PWA

[Han93] Peter Handsman. “porting to the Win32 API”. Dr. Dobb’s Journal of Software Tools, 18 (1):74, 76–78, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Harter:1991:OSC

[Har91] Richard Harter. Object-oriented software configuration management. Dr. Dobb’s Journal of Software Tools, 16(10):36, 38, 40, 42, 44, 46, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Harbison:1992:SPM

[Har92] Sam Harbison. Safe programming with Modula-3. Dr. Dobb’s Journal of Software Tools, 17 (10):88–96, 126–127, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Harper:1993:RSC

[Har93a] Jim Harper. A DOS redirector for SCSI CD-ROM. Dr. Dobb’s Journal of Software Tools, 18 (3):44, 48, 50, 103–105, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Harrington:1993:MF

[Har93b] James Harrington. Measuring fragmentation. Dr. Dobb’s Journal of Software Tools, 18(4):66, 68, 70, 92, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Harbison:1994:MPL

[Har94a] Sam Harbison. The Modula-3 programming language. Dr.

Dobb’s Journal of Software Tools, 19(??):24, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Repeat of Oct92 article.

Hardenbergh:1994:CPW

[Har94b] Hal W. Hardenbergh. CPU performance: Where are we headed? Dr. Dobb’s Journal of Software Tools, 19(1):30, 32, 34, 36, 38, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1995:PM

[Har95a] Hal W. Hardenbergh. Pushy? me?? Dr. Dobb’s Journal of Software Tools, 20(??):53–??, November/December 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1995:RAU

[Har95b] Hal W. Hardenbergh. Robots around us. Dr. Dobb’s Journal of Software Tools, 20(??):53– ??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1995:RSM

[Har95c] Hal W. Hardenbergh. Rocket science made simple. Dr. Dobb’s Journal of Software Tools, 20 (??):53–??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1995:SJH

[Har95d] Hal W. Hardenbergh. San jose’s high-tech ditch. Dr. Dobb’s Journal of Software Tools, 20 (??):53–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1995:SH

[Har95e] Hal W. Hardenbergh. This stuff is hard. Dr. Dobb’s Journal of Software Tools, 20(??):54– ??, July/August 1995. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1996:GET

[Har96a] Hal W. Hardenbergh. Gresham emerges triumphant. Dr. Dobb’s Journal of Software Tools, 21 (??):47–??, January/February 1996. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1996:SC

[Har96b] Hal W. Hardenbergh. Small catastrophes. Dr. Dobb’s Journal of Software Tools, 21(??):53– ??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Hardenbergh:1996:TF

[Har96c] Hal W. Hardenbergh. Tempus fugit. Dr. Dobb’s Journal of Software Tools, 21(??):57– ??, March/April 1996. CODEN DDJOEB. ISSN 1044-789X.

Harrington:1999:CMT

[Har99]  Mike Harrington.  Control-

ling motion-tracking devices. Dr. Dobb’s Journal of Software Tools, 24(3):90, 92–96, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/motion.txt.

Hatchett:1991:EAD

[Hat91]  Steve Hatchett.  Enhancing

the Actor development environment. Dr. Dobb’s Journal of Software Tools, 16(10):100–102, 125–128, October 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Haverlock:1998:OSC

[Hav98a] Kevin Haverlock. Object serialization, C++, and Java: Exchanging data between a sender and a receiver. Dr. Dobb’s Journal of Software Tools, 23(8):32, 34, 36–37, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
08/serial.zip.

Havewala:1998:WCB

[Hav98b] Aspi Havewala. The Windows CE build process: Creating custom versions of the operating system. Dr. Dobb’s Journal of Software Tools, 23(8):50, 52, 54, 56–57, 59, 91–92, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
08/cebuild.txt.

Havewala:1999:VCP

[Hav99a] Aspi Havewala. The version control process. Dr. Dobb’s Journal of Software Tools, 24(5):100, 102, 104–111, May 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Havewala:1999:WCE

[Hav99b] Aspi Havewala. The Windows CE emulator. Dr. Dobb’s Journal of Software Tools, 24(3): 56, 58, 60, 62, 64–65, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/ceemul.txt.

Hsi-Chiu:1990:ACE

[HC90] Liu Hsi-Chiu. Algebraic codes for error detection and correction. Dr. Dobb’s Journal of Software Tools, 15(12):46– ??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Hoffman:1999:PIL

[HC99] William A. Hoffman and Rupert W. Curwen. Pseudoincremental linking for C/C++. Dr. Dobb’s Journal of Software Tools, 24(10):46, 48–50, 52, 54, October 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_10/
linking.txt;   http://www.
ddj.com/ftp/1999/1999_10/
linking.zip.

Haney:1999:PPE

[HCKS99] Scott Haney, James Crotinger, Steve Karmesin, and Stephen Smith. PETE: The Portable Expression Template Engine. Dr. Dobb’s Journal of Software Tools, 24(10):88, 90–92, 94–95, October 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_10/pete-
2_0.sit;   http://www.ddj.
com/ftp/1999/1999_10/pete-
2_0.tgz;   http://www.ddj.
com/ftp/1999/1999_10/pete-
2_0.zip;   http://www.ddj.

                                        com/ftp/1999/1999_10/pete.
                                        txt.

Heddle:1998:PTH

[Hed98] David P. Heddle. Programmer’s toolchest: The Hot Views graphics library. Dr. Dobb’s Journal of Software Tools, 23 (5):92, 94–95, 112–113, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        05/hv.txt.

Heinrich:1996:AAF

[Hei96] Peter Heinrich. Algorithm alley: A fast integer square root. Dr. Dobb’s Journal of Software Tools, 21(4):113–114, 130, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Heller:1993:CDP

[Hel93] Steve Heller. 486 cache detection [sidebar in “Processor Detection Schemes” (R. C. Leinecker)]. Dr. Dobb’s Journal of Software Tools, 18(6):48–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Hellwig:1998:IAM

[Hel98]  Frank Hellwig.  Implement-

ing associations: Maintaining pointer integrity the easy way. Dr. Dobb’s Journal of Software Tools, 23(6):86, 88–91, 94–95, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Hemrajani:1996:ESC

[Hem96a] Anil Hemrajani. Examining Symantec’s Café: Visual development meets Java programming. Dr. Dobb’s Journal of Software Tools, 21(8):78, 80, 82, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Hemrajani:1996:JAW

[Hem96b] Anil Hemrajani. The Java Abstract Window Toolkit. Dr. Dobb’s Journal of Software Tools, 21(9):40, 42, 44, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Hendseth:1997:LPC

[Hen97a]  Sverre Hendseth.  Literate

programming and code reuse. Dr. Dobb’s Journal of Software Tools, 22(6):18–20, 22, 24, 97, 98, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Henson:1997:LEJ

[Hen97b] Scott G. Henson. Letter to the editor: Java and CRC feedback. Dr. Dobb’s Journal of Software Tools, 22(6):10, June 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Herstine:1998:AHP

[Her98] Michael Herstine. Algorithms for high-precision finite differences. Dr. Dobb’s Journal of Software Tools, 23(5):52, 54–58, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Hester:1993:SBL

[Hes93] Garyl Hester. Strategies for better linked lists. Dr. Dobb’s Journal of Software Tools, 18(8):32, 34, 36, 38, 40, 42, 100–101, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Hewett:1992:DMC

[Hew92] Frederick Hewett. DPMI meets C++. Dr. Dobb’s Journal of Software Tools, 17(10):60– 68, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Helm:1995:PRO

[HG95] Richard Helm and E. Gamma. Patterns for reusable objectoriented software. Dr. Dobb’s Journal of Software Tools, 20 (??):56–??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Harrison:1999:LRP

[HHW^{+}99] Brian Harrison, Philip Hendrickson, Murel Warren, Jr., Lee Kamentsky, Ron Gutman, Brenton Hoff, Martin Handwerker, Tom Culliton, and Aspi Havewala. Letters: Real programmer’s hate Cobol; 1984; Hilbert curves; grepping and globbing; testing Java classes; the version control process. Dr. Dobb’s Journal of Software Tools, 24 (9):10, 12, September 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Hightower:1996:HCF

[Hig96] Lauren Hightower. HTML conversion and FTP automation. Dr. Dobb’s Journal of Software Tools, 21(4):82, 84, 86, 94, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Hightower:1997:PDD

[Hig97] Lauren Hightower. Publishing dynamic data on the Internet — Allaire’s Cold Fusion is a development tool that provides access (via the Web) to any database the Web server can access using ODBC. Cold Fusion runs as a multithreaded Windows NT system service and works with any ODBC-compliant database. Dr. Dobb’s Journal of Software Tools, 22(1):70–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Hightower:1998:WRD

[Hig98] Lauren Hightower. The Web Report database reporting tool. Dr. Dobb’s Journal of Software Tools, 23(10):90, 92–93, 95, 113–114, October 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_10/
webrept.txt;   http://www.
ddj.com/ftp/1998/1998_10/
webrept.zip.

Hijazi:1996:EFE

[Hij96] Nabil Hijazi. Examining Forté Express. Dr. Dobb’s Journal of Software Tools, 21(11):86– ??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Hill:1995:PCS

[Hil95] William Hill. A portable C++ string class. Dr. Dobb’s Journal of Software Tools, 20(??):19– ??, July/August 1995. CODEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1992:PS

[Hir92a] Ian Hirschsohn. Personal supercomputing. Dr. Dobb’s Journal of Software Tools, 17(6):16–18, 20, 22, 24, 26–27, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1992:PSS

[Hir92b] Ian Hirschsohn. Personal supercomputing: Seamless portability. Dr. Dobb’s Journal of Software Tools, 17(7):40, 42–44, 46, 48, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1992:PSV

[Hir92c] Ian Hirschsohn. Personal supercomputing: virtual memory, 64bit. Dr. Dobb’s Journal of Software Tools, 17(8):50, 52, 54, 56, 60, 62–63, August 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1994:SP

[Hir94a] Ian Hirschsohn. Sharing peripherals intelligently. Dr. Dobb’s Journal of Software Tools, 19 (13):52–60, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1994:SPI

[Hir94b] Ian Hirschsohn. Sharing peripherals intelligently. part 2. Dr. Dobb’s Journal of Software Tools, 19(15):52, 54, 56, 58, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Hittleman:1999:JQW

[HL99] Kenneth Hittleman and Ted Leung. Java Q&A: What’s in store when moving from JDK 1.1 to JDK 1.2? Dr. Dobb’s Journal of Software Tools, 24(1):

112–115, January 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/ftp/1999/1999_01/
jqa199.txt.

Hlavaty:1994:EHW

[Hla94] Joseph Hlavaty. Exception handlers and Windows applications. Dr. Dobb’s Journal of Software Tools, 19(??):64–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Hlavaty:1995:MBW

[Hla95a]  Joseph Hlavaty.  MapMan:

Building Windows symbols files. Dr. Dobb’s Journal of Software Tools, 20(5):40, 42–47, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Hlavaty:1995:RWH

[Hla95b]  Joseph Hlavaty.  RTFHelp

for Windows help files. Dr. Dobb’s Journal of Software Tools, 20(2):135–136, 138–139, 154, 156, February 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Hlavaty:1995:WAE

[Hla95c]  Joseph Hlavaty.  Windows

apps and exception handlers. Dr. Dobb’s Journal of Software Tools, 20(??):7–??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Hlavaty:1996:EIS

[Hla96a] Joseph Hlavaty. Examining the InstallShield SDK Edition.

Dr. Dobb’s Journal of Software Tools, 21(9):70, 72–74, 76, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Hlavaty:1996:WL

[Hla96b] Joseph Hlavaty. The WinMock library. Dr. Dobb’s Journal of Software Tools, 21(7):66, 68, 70– 73, 83–84, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Herold:1995:IDM

[HM95] Phil Herold and Carla Merrill (or Merill??). Interactive design methodology. Dr. Dobb’s Journal of Software Tools, 20 (6):40, 42, 44–47, 98, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Hamilton:1992:SHU

[HO92] Jeffrey W. Hamilton and Eileen M. Ornsby. Simulating hypercubes in Unix. part I. Dr. Dobb’s Journal of Software Tools, 17(12):72, 74, 76, 108– 110, 112–113, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Hamilton:1993:SHU

[HO93] Jeffrey W. Hamilton and Eileen M. Ormsby. Simulating hypercubes in UNIX. part II. Dr. Dobb’s Journal of Software Tools, 18(1):66, 68, 70, 99–100, 102–105, January 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Hoenig:1998:PCM

[Hoe98] Stefan Hoenig. Plug-in components for MFC: Adding power and flexibility to MFC. Dr. Dobb’s Journal of Software Tools, 23(8):82, 84–86, 94, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
08/mfcplug.txt;   http://
www.ddj.com/ftp/1998/1998_
08/mfcplug.zip.

Hoffman:1995:SCP

[Hof95a] Marc E. Hoffman. Simulation compilation and portability. Dr. Dobb’s Journal of Software Tools, 20(3):52, 54–55, 58–59, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Hofri:1995:AA

[Hof95b] Micha Hofri. Algorithm analysis. Dr. Dobb’s Journal of Software Tools, 20(3):125–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Hofri:1996:BS

[Hof96] Micha Hofri. Binary search.

Dr. Dobb’s Journal of Software Tools, 21(2):135–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Hoff:1997:HSF

[Hof97] Brenton Hoff. High-speed finitestate machines: Byte-oriented applications benefit from an assembly trick. Dr. Dobb’s Journal of Software Tools, 22(11): 54, 56–61, 93, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Holub:1996:LGB

[Hol96] Allen I. Holub. Let the games begin. Dr. Dobb’s Journal of Software Tools, 21(??):63– ??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Holzmann:1997:SMC

[Hol97] Gerard J. Holzmann. Spin model checking: Reliable design of concurrent software. Dr. Dobb’s Journal of Software Tools, 22(10):92, 94, 96– 97, 104, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Hopkins:1991:DIP

[Hop91] Don Hopkins. The design and implementation of pie menus. Dr. Dobb’s Journal of Software Tools, 16(12):16–26, 94, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Houlding:1998:CBF

[Hou98] David Houlding. A CORBA bean framework. Dr. Dobb’s Journal of Software Tools, 23 (11):34, 36, 38–40, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_11/
                                        CORBABeans.exe;   http://
                                        www.ddj.com/ftp/1998/1998_
                                        11/CORBEAN.TXT.

Houlding:1999:AWS

[Hou99] David Houlding. An architecture for Web services. Dr. Dobb’s Journal of Software Tools, 24(7):80, 82, 84–85, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
07/corbean.txt.

Hovel:1994:SPW

[Hov94] David Hovel. Small Prolog and Windows NT networking [sidebar]. Dr. Dobb’s Journal of Software Tools, 19(8):80–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Howard:1990:SVP

[How90] Christopher A. Howard. Super VGA programming. Dr. Dobb’s Journal of Software Tools, 15 (7):16–18, 22, 26, 28, 82, 84– 88, 90–92, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Howard:1993:EPL

[How93] Robert Howard. The Eiffel programming language. Dr. Dobb’s Journal of Software Tools, 18 (11):68–73, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Howard:1996:GPD

[How96] Christopher A. Howard. Games programming with DirectPlay. Dr. Dobb’s Journal of Software Tools, 21(??):7–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Howes:1998:TPC

[How98] Brad Howes. Template processing classes for Python. Dr. Dobb’s Journal of Software Tools, 23(2):38, 40, 42, 44–46, 48, 100, February 1998.

Howing:1999:WHT

[Höw99] Frank Höwing. WiT and hierarchical trees. Dr. Dobb’s Journal of Software Tools, 24 (9):80, 82–86, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/hoptree.txt;   http://
www.ddj.com/ftp/1999/1999_
09/hoptree.zip.

Hair:1998:ESE

[HQ98] Don Hair and Cesar Quiroz. Embedded systems: Embedded development compilers: The right tool for the right job. Dr. Dobb’s Journal of Software Tools, 23(8):60, 62, 64, 66–69, August 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_08/
embedcpp.txt.

Hook:1994:DIP

[HS94] Brian Hook and Dennis Shuman. Digital I/O with the PC. Dr. Dobb’s Journal of Software Tools, 19(4):64, 66, 68, 70, 103, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Huddleston:1998:PTA

[HT98] Charles H. Huddleston and Douglas A. Troy. Programmer’s toolchest: An ActiveX control for real-time computer control. Dr. Dobb’s Journal of Software Tools, 23(2):94, 96–99, 105, February 1998.

Hubbard:1994:IRS

[Hub94] Hamish Hubbard. Inside the RIFF specification. Dr. Dobb’s Journal of Software Tools, 19 (10):38, 40, 42–44, 89–90, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Huffman:1993:MMF

[Huf93a]  Doug Huffman.  Memory-

mapped file I/O. Dr. Dobb’s Journal of Software Tools, 18 (??):14–??, ???? 1993. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Huffman:1993:MFI

[Huf93b]  Doug Huffman.  Memory-

mapped file I/O. Dr. Dobb’s Journal of Software Tools, 18 (10):14–??, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Huffman:1993:PE

[Huf93c] Joe Huffman. Porting from 16-bit to 32-bit extended DOS. Dr. Dobb’s Journal of Software Tools, 18(1):28, 30, 32, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Hughes:1995:FFS

[Hug95] Phil Hughes. Freenets and free software. Dr. Dobb’s Developer Update, 2(2):2–3, February 1995. CODEN ???? ISSN 1079-8595.

Husain:1992:DMW

[Hus92] Kamran Husain. Debugging Motif widgets. Dr. Dobb’s Journal of Software Tools, 17(11):86, 88, 90–91, 135–136, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Husain:1994:EI

[Hus94] Kamran Husain. Extending Imake. Dr. Dobb’s Journal of Software Tools, 19(6):70, 72, 74–76, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Hussain:1999:UL

[Hus99] Basit Hussain. Internet programming: Understanding LDAP. Dr. Dobb’s Journal of Software Tools, 24(3): 78, 80–82, March 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ddj/1999/1999_03/9903h/9903h.
htm;   http://www.ddj.com/
ftp/1999/1999_03/ldap.txt;
 http://www.ddj.com/ftp/1999/
1999_03/ldap.zip.

Hoenig:1996:EM

[HW96] Stefan Hoenig and Scot Wingo. Extending MFC. Dr. Dobb’s Journal of Software Tools, 21 (1):58–59, 62–64, 100–101, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Hyde:1990:OPA

[Hyd90]  Randall L. Hyde.  Object-

oriented programming with assembly language. Dr. Dobb’s Journal of Software Tools, 15 (3):66–73, 110–111, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Hyde:1992:USA

[Hyd92] Randall L. Hyde. The UCR standard ASM library. Dr. Dobb’s Journal of Software Tools, 17(3):16–18, 20, 22, 80, 82, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Ioannidou:1999:EUP

[IR99] Andri Ioannidou and Alexander Repenning. End-user programmable simulations. Dr. Dobb’s Journal of Software Tools, 24(8):40, 42–48, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Isenberg:1995:FTS

[Ise95] Herb Isenberg. Flexible testing systems. Dr. Dobb’s Journal of Software Tools, 20(6):88–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Ives:1990:GOA

[Ive90] William E. Ives. A generic one-pass assembler. Dr. Dobb’s Journal of Software Tools, 15 (9):50–??, September 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Iyer:1995:PGU

[Iye95] Mani G. Iyer. Permutation generation using matrices. Dr. Dobb’s Journal of Software Tools, 20(11):133–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Jackson:1990:GSG

[Jac90]  Frank Jackson.  Generation

scavenging (garbage collection).

Dr. Dobb’s Journal of Software Tools, 15(5):16–23, 26–28, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Jaeschke:1991:SCS

[Jae91] Rex Jaeschke. Standard C: A status report. Dr. Dobb’s Journal of Software Tools, 16(8):16– 18, 20, 22, 24, 26, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Jaeschke:1993:CSU

[Jae93] R. Jaeschke. C/C++ standardization: an update. Dr. Dobb’s Journal of Software Tools, 18 (8):26, 28, 30, August 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Jaeschke:1995:SCU

[Jae95] Rex Jaeschke. Standard C: An update. Dr. Dobb’s Journal of Software Tools, 20(8):28, 30, 32, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Jaffer:1993:ALC

[Jaf93] Aubrey Jaffer. Algebra and the lambda calculus. Dr. Dobb’s Journal of Software Tools, 18 (9):36, 38, 40, 88, September 1993. CODEN DDJOEB. ISSN 1044-789X.

James:1990:RA

[Jam90] Rahner James. Real-time animation. Dr. Dobb’s Journal of Software Tools, 15(1):16–17, 19– 20, 22, 24, 82, 84, 86, 88, 90, 92–93, January 1990. CODEN DDJOEB. ISSN 1044-789X.

James:1992:IGC

[Jam92] Rahner James. IPX the great communicator: speeding up a Novell network with a streams interface. Dr. Dobb’s Journal of Software Tools, 17(5):46, 48, 50, 52, 87–88, May 1992. CODEN DDJOEB. ISSN 1044-789X.

James:1993:YON

[Jam93] Rahner James. Your own network data snooper. Dr. Dobb’s Journal of Software Tools, 18 (5):62, 64–67, 98, 100, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Janke:1991:RI

[Jan91] Steven Janke. Recursive images. Dr. Dobb’s Journal of Software Tools, 16(7):16–18, 20, 22, 74, 76–78, July 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Jannak:1999:JGR

[Jan99a] Torpum Jannak. Java 2 graphics rendering. Dr. Dobb’s Journal of Software Tools, 24(9): 19–20, 22, 24, 26–28, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/java2gra.txt;   http://
www.ddj.com/ftp/1999/1999_
09/java2gra.zip.

Janzen:1999:AAR

[Jan99b] Thomas E. Janzen. Algorithm alley: Rendering line drawings from 3D models. Dr. Dobb’s Journal of Software Tools, 24 (9):109–113, September 1999.

CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_09/
aa999.txt;   http://www.ddj.
com/ftp/1999/1999_09/aa999.
zip.

Jarvis:1990:ICA

[Jar90]  Pitts Jarvis.  Implementing

Cordic algorithms. Dr. Dobb’s Journal of Software Tools, 15 (10):152–154, 56, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Jensen:1995:DRO

[Jen95] E. Douglas Jensen. Distributed real-time operating systems. Dr. Dobb’s Journal of Software Tools, 20(2):32–34, 36, 38, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Jenkins:1997:AAH

[Jen97] Bob Jenkins. Algorithm alley: Hash functions. Dr. Dobb’s Journal of Software Tools, 22 (9):107–109, 115–116, September 1997. CODEN DDJOEB. ISSN 1044-789X. Describes a new hash function which is much better at producing uniform key distributions than others commonly used, yet remains acceptably fast. See [Boy98a] for comparison with a related algorithm.

Jervis:1992:NEC

[Jer92] Robert Jervis. Numerical extensions to C. Dr. Dobb’s Journal of Software Tools, 17(8):26, 28, 30, 32, 34, 39, August 1992.

CODEN DDJOEB. ISSN 1044-789X.

Jervis:1993:PPL

[Jer93] Robert Jervis. The Parasol programming language. Dr. Dobb’s Journal of Software Tools, 18 (11):34–41, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Jervis:1994:PPL

[Jer94] Robert Jervis. The Parasol programming language. Dr. Dobb’s Journal of Software Tools, 19 (??):6–??, October 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBd

[JJ91a] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386. the basic kernel (device autoconfiguration). Dr. Dobb’s Journal of Software Tools, 16(11):46, 48, 50, 52, 54, 56–58, 60, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBb

[JJ91b] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386. the basic kernel multiprogramming and multitasking. I. Dr. Dobb’s Journal of Software Tools, 16 (9):44, 46, 48, 50, 52, 54, 56, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBc

[JJ91c] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386. the basic kernel multiprogramming and multitasking. II. Dr. Dobb’s Journal of Software Tools, 16 (10):62, 64, 66, 68, 70, 72, 118– 120, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPa

[JJ91d] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(1):??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPc

[JJ91e] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(2):??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPd

[JJ91f] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(3):??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPe

[JJ91g] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(4):??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPf

[JJ91h] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPg

[JJ91i] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(6):??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPh

[JJ91j] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(7):??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPi

[JJ91k] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(8):??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPj

[JJ91l] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(9): ??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPk

[JJ91m] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPl

[JJ91n] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 16(11): ??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPb

[JJ91o] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach (designing the software specification). Dr. Dobb’s Journal of Software Tools, 16(1):16– 18, 20, 22–24, 28, 30, 32, 34, 36, 38–40, 42, 46, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUSb

[JJ91p] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A strippeddown kernel. Dr. Dobb’s Journal of Software Tools, 16(7):32, 34– 36, 38, 40, 84–88, 88, 90, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUL

[JJ91q] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: Language tools cross support. Dr. Dobb’s Journal of Software Tools, 16 (4):34, 36–38, 42, 44–45, 90, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUR

[JJ91r] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386 research and the commercial sector. Dr. Dobb’s Journal of Software Tools, 16(6):50, 52, 54, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBa

[JJ91s] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: the basic kernel. Dr. Dobb’s Journal of Software Tools, 16(8):54, 56–60, 64– 65, 126, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUI

[JJ91t] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386. the initial root filesystem. Dr. Dobb’s Journal of Software Tools, 16(5):46, 48, 50, 52–54, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUSa

[JJ91u] William Frederick Jolitz and Lynne Greer Jolitz. Porting Unix to the 386: the standalone system. Dr. Dobb’s Journal of Software Tools, 16(3):38, 40, 42, 44, 47–48, 50, 76–82, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUT

[JJ91v] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: three initial PC utilities. Dr. Dobb’s Journal of Software Tools, 16(2):54, 56, 58–61, 111–112, 114–115, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUDb

[JJ92a] W. F. Jolitz and L. G. Jolitz. Porting UNIX to the 386: device drivers. Dr. Dobb’s Journal of Software Tools, 17(3):44, 46, 48, 50, 90–91, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUDc

[JJ92b] W. F. Jolitz and L. G. Jolitz. Porting UNIX to the 386: device drivers. Dr. Dobb’s Journal of Software Tools, 17(4):64–67, 108–109, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUMb

[JJ92c] W. F. Jolitz and L. G. Jolitz. Porting Unix to the 386 missing pieces. II. Dr. Dobb’s Journal of Software Tools, 17(6):56, 58– 60, 62, 96, 101–104, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUF

[JJ92d] W. F. Jolitz and L. G. Jolitz. Porting UNIX to the 386: the final step. Dr. Dobb’s Journal of Software Tools, 17(7):50, 52, 54, 56, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUDa

[JJ92e] W. F. Jolitz and L. Greer Jolitz. Porting UNIX to the 386 device drivers. Dr. Dobb’s Journal of Software Tools, 17(2):30, 32– 33, 93, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:IIF

[JJ92f] William Frederick Jolitz and Lynne Greer Jolitz. Inside the ISO-9660 filesystem format. Dr. Dobb’s Journal of Software Tools, 17(12):80, 82–84, 89, 114– 115, 118–119, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPa

[JJ92g] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPb

[JJ92h] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPc

[JJ92i] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPd

[JJ92j] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPe

[JJ92k] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPf

[JJ92l] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: A practical approach. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUMa

[JJ92m] William Frederick Jolitz and Lynne Greer Jolitz. Porting Unix to the 386 missing pieces. I. completing the 386BSD kernel. Dr. Dobb’s Journal of Software Tools, 17(5):54, 56, 59– 60, 62, 91, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1994:VHS

[JJ94] Lynne Greer Jolitz and William Frederick Jolitz. Very high-speed networks: HiPPI and SIGNA. Dr. Dobb’s Journal of Software Tools, 19(??):32–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1995:ISB

[JJ95a] William F. Jolitz and Lynne Greer Jolitz. Internet security breach. Dr. Dobb’s Developer Update, 2 (3):3–4, March 1995. CODEN ???? ISSN 1079-8595.

Jolitz:1995:RNS

[JJ95b] William F. Jolitz and Lynne Greer Jolitz. Role-based network security. Dr. Dobb’s Journal of Software Tools, 20(5):80, 82, 84–85, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Jones:1995:VPR

[JK95] Carol Jones and Morgan Kinne.

Visual programming with reusable objects. Dr. Dobb’s Journal of Software Tools, 20(12):96– ??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Jaekl:1998:PBS

[JL98] Chris Jaekl and Paul Lu. Programmer’s bookshelf: Software projects and UNIX programming. Dr. Dobb’s Journal of Software Tools, 23(6):127, 129, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Jolin:1998:DCL

[JLC98] Arthur T. Jolin, David Lavin, and Susan Carpenter. Designing class libraries for visual builders. Dr. Dobb’s Journal of Software Tools, 23(6):38, 40, 42–43, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1992:ICC

[JM92] Stephen P. Johnson and Tom McReynolds. Implementing curves in C++. Dr. Dobb’s Journal of Software Tools, 17 (12 (special supplement)):53s– 55s, 57s–64s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Jurisic:1997:ECC

[JM97] Aleksandar Jurisic and Alfred J. Menezes. Elliptic curves and cryptography. Dr. Dobb’s Journal of Software Tools, 22(4): 26–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1990:DHP

[Joh90] J. Scott Johnson. The DDJ hypertext project. Dr. Dobb’s Journal of Software Tools, 15 (6):16–18, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1993:WCC

[Joh93] R. Colin Johnson. What is cognitive computing? Dr. Dobb’s Journal of Software Tools, 18 (2):18–20, 22, 24, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Johansen:1997:IPT

[Joh97a] Jay Johansen. Internet programming — template driven Web pages: A CGI program that parses templates. Dr. Dobb’s Journal of Software Tools, 22(11):74, 76–78, 93– 94, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1997:SPS

[Joh97b]  Alan Johnson.  Steganogra-

phy for DOS programmers — steganography is a branch of cryptography that deals with concealing messages. Dr. Dobb’s Journal of Software Tools, 22 (1):48–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Johnson:1999:PTC

[Joh99] David M. Johnson. Programmer’s toolchest: Comparing WFC and JFC. Dr. Dobb’s Journal of Software Tools, 24 (2):90, 92–95, February 1999. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/relay.txt;
 http://www.ddj.com/ftp/1999/
1999_02/relay.zip.

Jolitz:1992:PB

[Jol92a] William Jolitz. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 17(9):149– ??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:HN

[Jol92b] William Frederick Jolitz. Highspeed networking. Dr. Dobb’s Journal of Software Tools, 17 (8):86, 88, 90–93, 122, August 1992. CODEN DDJOEB. ISSN 1044-789X.

GreerJolitz:1993:ESC

[Jol93a] L. Greer Jolitz. Extending standards for CD-ROM. Dr. Dobb’s Journal of Software Tools, 18 (7):74, 76, 78–79, 82, 101, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1993:ESC

[Jol93b] Lynne Greer Jolitz. Extending standards for CD-ROM. Dr. Dobb’s Journal of Software Tools, 18(7):74–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1993:PB

[Jol93c] Lynne Greer Jolitz. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18 (9):127–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1994:PP

[Jol94] William Frederick Jolitz. Polymorphic protocols. Dr. Dobb’s Journal of Software Tools, 19 (1):66, 68, 70, 72, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1995:PP

[Jol95a] Lynne Greer Jolitz. POSIX.4 programming. Dr. Dobb’s Developer Update, 2(8):2–3, August 1995. CODEN ???? ISSN 1079-8595.

Jolitz:1995:PB

[Jol95b] Lynne Greer Jolitz. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20 (8):133–??, August 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Jolitz:1995:III

[Jol95c] William F. Jolitz. Inside “intel inside”. Dr. Dobb’s Developer Update, 2(2):3–4, February 1995. CODEN ???? ISSN 1079-8595.

Jolin:1996:UCL

[Jol96] Arthur T. Jolin. Usability and class library design. Dr. Dobb’s Journal of Software Tools, 21 (10):16–??, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Jones:1998:JIA

[Jon98] Carol A. Jones. The Java internationalization API: Global software for the global village. Dr. Dobb’s Journal of Software Tools, 23(1):54, 56–69, 103– 104, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1992:RB

[JS92] Ted Johnson and Dennis Shasha.

Reexamining B-trees. Dr. Dobb’s Journal of Software Tools, 17(1):44, 46–47, January 1992. CODEN DDJOEB. ISSN 1044-789X.

James:1994:CYO

[JS94] Rahner James and Linus Sphinx. Creating your own multiplayer game systems. Dr. Dobb’s Journal of Software Tools, 19(??):56–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Justice:1992:FFR

[Jus92] Harold R. Justice. FORTREX:

A Fortran runtime executive. Dr. Dobb’s Journal of Software Tools, 17(6):82–??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1992:FLR

[JW92] Steve Johnson and M. Wiebel. The five levels of RAID. Dr. Dobb’s Journal of Software Tools, 17(1):24–??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Kaehms:1994:SCF

[Kae94] Bob Kaehms. Stanford computer forum. Dr. Dobb’s Developer Update, 1(10):3, December 1994. CODEN ???? ISSN 1079-8595.

Kahn:1992:PII

[Kah92] Mitchell Kahn. Programming the I2C interface. Dr. Dobb’s Journal of Software Tools, 17 (6):64–??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Kaliski:1992:MAC

[Kal92] Burton S. Kaliski, Jr. Multipleprecision arithmetic in C. Dr. Dobb’s Journal of Software Tools, 17(8):40, 42, 44, 46–48, 116–119, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Kaliski:1993:ZBA

[Kal93] Burton S. Kaliski, Jr. The Z80180 and big-number arithmetic. Dr. Dobb’s Journal of Software Tools, 18(9):50, 52, 54, 56, 58, 90–91, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Kanalakis:1994:EOT

[Kan94] John M. Kanalakis, Jr. Examining OS/2 2.1 threads.

Dr. Dobb’s Journal of Software Tools, 19(1):74, 76, 78– 79, 96, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Kapor:1991:SDM

[Kap91] Mitchell Kapor. A software design manifesto. Dr. Dobb’s Journal of Software Tools, 16(1):62, 64–67, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Kaplan:1995:PP

[Kap95] Paul Kaplan. Porting to the PowerMac. Dr. Dobb’s Journal of Software Tools, 20(??): 7–??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

Kar:1990:IRR

[Kar90] Rabindra P. Kar. Implementing the Rhealstone real-time benchmark. Dr. Dobb’s Journal of Software Tools, 15(4):46–55, 100–104, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Karnofsky:1993:NNC

[Kar93] Ken Karnofsky. Neural networks and character recognition. Dr. Dobb’s Journal of Software Tools, 18(6):96, 98–100, 102–103, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Kauler:1995:TTR

[Kau95]  Barry Kauler.  Terse: A

tiny real-time operating system. Dr. Dobb’s Journal of Software Tools, 20(12):72, 74, 76, 78, 80, 110, 114, 116, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Kaufmann:1996:MID

[Kau96a] Roland Kaufmann. Multiple inheritance for Delphi. Dr. Dobb’s Journal of Software Tools, 21 (10):26–??, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Kauler:1996:OFD

[Kau96b] Barry Kauler. Object-oriented flow design for embedded systems. Dr. Dobb’s Journal of Software Tools, 21(10):54– ??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Kauffmann:1997:AA

[Kau97] Robert F. Kauffmann. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 22(5):105–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Kay:1996:JC

[Kay96] Jeffrey Kay. Java and C++. Dr. Dobb’s Journal of Software Tools, 21(11):8, 10, November 1996. Comments on [Ste96l].

Kasperink:1997:CDC

[KD97] Harold R. Kasperink and John C. Dekker. Concurrent database commands and C++. Dr. Dobb’s Journal of Software Tools, 22(8):84, 86, 88, 89, 98, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Keffer:1992:WCW

[Kef92] Thomas Keffer. Why C++ will replace Fortran. Dr. Dobb’s Journal of Software Tools, 17(12 (special supplement)):39s, 40s, 42s–47s, December 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Keffer:1995:PST

[Kef95] Thomas Keffer. Programming with the Standard Template Library. Dr. Dobb’s Journal of Software Tools, 20(??):7–??, July/August 1995. CODEN DDJOEB. ISSN 1044-789X.

Kehoe:1994:PFS

[Keh94] Brendan Kehoe. Portability and free software. Dr. Dobb’s Developer Update, 1(11):2–??, October 1994. CODEN ???? ISSN 1079-8595.

Kelly:1995:VWW

[Kel95] Christopher Kelly. Video for Windows and WinG. Dr. Dobb’s Journal of Software Tools, 20 (??):19–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Kent:1993:FFF

[Ken93] Jim Kent. The Flic file format. Dr. Dobb’s Journal of Software Tools, 18(3):18–20, 22, 92, 94, 96, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Kent:1994:AAW

[Ken94] Peter Kent. Adding animation to Windows Help. Dr. Dobb’s Journal of Software Tools, 19 (15):74, 76, 78, 80, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Kepple:1994:BAG

[Kep94] Laurence R. Kepple. The black art of GUI testing. Dr. Dobb’s Journal of Software Tools, 19 (2):40–?? (or 42–??), February 1994. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1994:AAR

[Kie94a] Tim Kientzle. Algorithm alley: Rendering circles and ellipses. Dr. Dobb’s Journal of Software Tools, 19(7):119–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1994:IX

[Kie94b] Tim Kientzle. Intelligent XY-Modem. Dr. Dobb’s Journal of Software Tools, 19(15):36, 38, 40, 42, 90–91, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1995:MIM

[Kie95] Tim Kientzle. MIME and Internet mail. Dr. Dobb’s Journal of Software Tools, 20(9):54, 56, 58, 60, 104, 106, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1996:E

[Kie96a] Tim Kientzle. Editorial. Dr. Dobb’s Journal of Software Tools, 21(??):??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1996:IKP

[Kie96b] Tim Kientzle. Improving Kermit performance: A windowing strategy makes all the difference. Dr. Dobb’s Journal of Software Tools, 21(2):28, 30, 32, 34, 36, 38, 91–92, February 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Kientzle:1997:AAI

[Kie97a] Tim Kientzle. Algorithm alley: Inside IMA ADPCM. Dr. Dobb’s Journal of Software Tools, 22(11):116, 118, 124, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1997:AAU

[Kie97b] Tim Kientzle. Algorithm alley: Understanding CRCs. Dr. Dobb’s Journal of Software Tools, 22(4):103–107, 114, 115, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Kientzle:1998:AFC

[Kie98] Tim Kientzle. Automatic file conversions with Perl. Dr. Dobb’s Journal of Software Tools, 23(2):56, 58, 101, February 1998.

Kientzle:1999:AAA

[Kie99a] Tim Kientzle. Algorithm alley: Adaptive unsharp masking. Dr. Dobb’s Journal of Software Tools, 24(11):119–121, November 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
11/aa1199.zip.

Kientzle:1999:AAI

[Kie99b] Tim Kientzle. Algorithm alley: Implementing fast DCTs. Dr. Dobb’s Journal of Software Tools, 24(3):115–119, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/aa399.txt;

 http://www.ddj.com/ftp/1999/
1999_03/aa399.zip.

Kientzle:1999:FTS

[Kie99c] Tim Kientzle. Full-text searching in Perl. Dr. Dobb’s Journal of Software Tools, 24(1): 34, 38–41, January 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ddj/ftp/1999/1999_
01/perlsrch.txt;   http://
www.ddj.com/ddj/ftp/1999/
1999_01/perlsrch.zip.

Kientzle:1999:JAS

[Kie99d] Tim Kientzle. A Java applet search engine. Dr. Dobb’s Journal of Software Tools, 24 (2):32, 36, 38–40, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jsearch.
txt;   http://www.ddj.com/
ftp/1999/1999_02/jsearch.
zip.

Kimbrell:1992:FSN

[Kim92] Roy E. Kimbrell. Finding significance in noisy data. Dr. Dobb’s Journal of Software Tools, 17 (6):30, 32, 37–38, 40–42, 90, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Kim:1995:PCC

[Kim95] Eugene E. Kim. Programming CGI in C. Dr. Dobb’s Journal of Software Tools, 20(??):21– ??, November/December 1995. CODEN DDJOEB. ISSN 1044-789X.

Kim:1996:Ia

[Kim96a] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 21(10):137–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Kim:1996:Ib

[Kim96b] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 21(11):130–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Kim:1996:Ic

[Kim96c] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 21(12):137–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Kim:1996:RAW

[Kim96d] Eugene Eric Kim. Running at Warp speed. Dr. Dobb’s Journal of Software Tools, 21(12):12, December 1996. Brief discussion of Java products under IBM OS/2.

Kim:1997:Ia

[Kim97a] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(1):118–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ib

[Kim97b] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(2):119–??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ic

[Kim97c] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(3):119–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Id

[Kim97d] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(4):119–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ie

[Kim97e] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(5):126–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:If

[Kim97f] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(6):138–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ig

[Kim97g] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(7):120–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ih

[Kim97h] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 22(8):129–130, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ii

[Kim97i] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 22(10):137–138, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ij

[Kim97j] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 22(11):129–130, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:Ik

[Kim97k] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 22(12):145–146, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1997:PB

[Kim97l] Eugene Eric Kim. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 22 (4):116–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:CLW

[Kim98a] Eugene Eric Kim. A conversation with Larry Wall. Dr. Dobb’s Journal of Software Tools, 23 (2):20–22, 24, 28, February 1998.

Kim:1998:Ia

[Kim98b] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(1):137–138, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:Ib

[Kim98c] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(2):137–138, February 1998.

Kim:1998:Ic

[Kim98d] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(3):145–146, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:Id

[Kim98e] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(4):145–146, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:Ie

[Kim98f] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(5):145–146, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:If

[Kim98g] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(8):122–123, August 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:Ig

[Kim98h] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(9):130–131, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:Ih

[Kim98i] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 23(10):146–147, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
10/index.htm.

Kim:1998:Ii

[Kim98j] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 23(11):146–147, November 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1998:Ij

[Kim98k] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 23(12):138, 139, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Kim:1999:Ia

[Kim99a] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(1):134–135, January 1999. CODEN DDJOEB. ISSN 1044-789X.

Kim:1999:Ib

[Kim99b] Eugene Eric Kim. Of interest.

Dr. Dobb’s Journal of Software Tools, 24(2):142–143, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ddj/1999/1999_02/9902u/9902u.
htm.

Kim:1999:Ic

[Kim99c] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(3):134–135, March 1999. CODEN DDJOEB. ISSN 1044-789X.

Kim:1999:Id

[Kim99d] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(4):126–127, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/1999/9904u/9904u.
                                        htm.

Kim:1999:Ie

[Kim99e] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(5):142–143, May 1999. CODEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/9905s/9905s.htm.

Kim:1999:If

[Kim99f] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(7):134–135, July 1999. CODEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/9907s/9907s.htm.

Kim:1999:Ig

[Kim99g] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(8):126–127, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/articles/1999/
                                        9908/9908t/9908t.htm.

Kim:1999:I

[Kim99h] Eugene Eric Kim. Of interest. Dr. Dobb’s Journal of Software Tools, 24(9):126–127, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Kim:1999:PBI

[Kim99i] Eugene Eric Kim. Programmer’s bookshelf: Internet time as history. Dr. Dobb’s Journal of Software Tools, 24(12):127–128, 130, December 1999. CODEN DDJOEB. ISSN 1044-789X.

King:1990:DCS

[Kin90a] Todd King. Drawing character shapes with Bézier curves. Dr. Dobb’s Journal of Software Tools, 15(7):46, 48, 50, 52, 102, 104, July 1990. CODEN DDJOEB. ISSN 1044-789X.

King:1990:SHE

[Kin90b] Todd King. A self-referential hypertext engine. Dr. Dobb’s Journal of Software Tools, 15(6):34– 38, 92–95, June 1990. CODEN DDJOEB. ISSN 1044-789X.

King:1991:WNM

[Kin91a] K. N. King. What’s new with Modula-2? Dr. Dobb’s Journal of Software Tools, 16(6):42, 44, 46, 48–49, June 1991. CODEN DDJOEB. ISSN 1044-789X.

King:1991:YES

[Kin91b] Todd King. YACC for expert systems. Dr. Dobb’s Journal of Software Tools, 16(2):88– ??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

King:1993:HPL

[Kin93] K. N. King. The history of programming languages. Dr. Dobb’s Journal of Software Tools, 18(8):18–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Kinne:1999:WJP

[Kin99] Morgan Kinne. Writing Javabean property editors. Dr. Dobb’s Journal of Software Tools, 24(2):52, 54–56, 58, 60, 62, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/propedit.
txt;   http://www.ddj.com/
ftp/1999/1999_02/propedit.
zip.

Kissel:1996:IWA

[Kis96] Paul Kissel. Inside the Winhelp() API function. Paul shows what’s involved in building a platform-independent implementation of the WinHelp() function. he then shares “seven deadly sins” you should avoid when writing portable software. Dr. Dobb’s Journal of Software Tools, 21(12):22–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Kiselyov:1997:AAH

[Kis97] Oleg Kiselyov. Algorithm alley — hoping to find the most efficient line between two points, Oleg examines scheduling algorithms and NP-complete problems. Dr. Dobb’s Journal of Software Tools, 22(2):107– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Kiselyov:1998:DDU

[Kis98] Oleg Kiselyov. Distributing data using TLT30G. Dr.

Dobb’s Journal of Software Tools, 23(9):34, 36, 38, 40, 92–93, September 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_09/
lapnet.txt;   http://www.
ddj.com/ftp/1998/1998_09/
lapnet.zip.

Klatchko:1999:DRS

[Kla99] Ron Klatchko. Dynamically reconfigurable servers: Python’s extensibility makes it easy. Dr. Dobb’s Journal of Software Tools, 24(1):80, 82–84, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ddj/ftp/1999/1999_
01/cprofile.zip.

Klein:1991:NW

[Kle91a] Mike Klein. Networking with Windows 3.0. Dr. Dobb’s Journal of Software Tools, 16(3):64– 66, 68, 90, 92, 94–98, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Klein:1991:PWF

[Kle91b] Mike Klein. PropView: a Windows family browser. Dr. Dobb’s Journal of Software Tools, 16(12 (supplement)):47–59, December 1991. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Klein:1991:SAW

[Kle91c] Mike Klein. Subclassing applications (Windows 3). Dr. Dobb’s Journal of Software Tools, 16(12 (supplement)):19–27, December 1991. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Klimasauskas:1990:NNI

[Kli90] Casimir C. “Casey” Klimasauskas. Neural networks and image processing. Dr. Dobb’s Journal of Software Tools, 15(4):77–82, 114, 116, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Klimasauskas:1991:NNT

[Kli91] Casimir C. “Casey” Klimasauskas. Neural nets tell why. Dr. Dobb’s Journal of Software Tools, 16 (4):16, 18, 20, 22, 24, 78, 80, 82–84, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Knowles:1990:WAD

[Kno90]  Walter Knowles.  Windows

3.0 application development. Dr. Dobb’s Journal of Software Tools, 15(11):92, 94, 96, 99, 100, 136, 138–139, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Knoblaugh:1992:DDM

[Kno92a] Rick Knoblaugh. Device driver monitoring. Dr. Dobb’s Journal of Software Tools, 17(3):52, 54, 57, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Knoblaugh:1992:YOP

[Kno92b] Rick Knoblaugh. Your own protected-mode debugger. Dr. Dobb’s Journal of Software Tools, 17(9):38, 40–41, 108– 109, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Knoblaugh:1993:EIT

[Kno93a]  R. Knoblaugh.  Eavesdrop-

ping on interrupts: tracking down software problems by monitoring interrupt activity. Dr. Dobb’s Journal of Software Tools, 18(12):42, 44, 94, 96, 98, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Knoblaugh:1993:EI

[Kno93b] Rick Knoblaugh. Eavesdropping on interrupts. Dr. Dobb’s Journal of Software Tools, 18(11): 42–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Knoblaugh:1994:WM

[Kno94] Rick Knoblaugh. A Windows I/O monitor. Dr. Dobb’s Journal of Software Tools, 19(??): 42–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Knowlton:1995:MGM

[Kno95] Scott Knowlton. Motif geometry management. Dr. Dobb’s Developer Update, 2(3):2, March 1995. CODEN ???? ISSN 1079-8595.

Knott:1998:AAA

[Kno98] Gary D. Knott. Algorithm alley:

Adaptive delta modulation: Algorithms for audio compression. Dr. Dobb’s Journal of Software Tools, 23(4):124, 126, 128, 139, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/aa498.txt.

Kochevar:1997:TID

[Koc97] Peter D. Kochevar. Tecate and interactive 3-D. Dr. Dobb’s Journal of Software Tools, 22 (7):72–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Kodis:1992:FCE

[Kod92] John Kodis. Fletcher’s checksum error correction at a fraction of the cost. Dr. Dobb’s Journal of Software Tools, 17 (5):32, 34, 36, 38, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Kofoed:1995:PMC

[Kof95] Stig Kofoed. Portable multitasking in C++. Dr. Dobb’s Journal of Software Tools, 20 (11):70, 72, 74, 76, 78, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Koolish:1994:NOM

[Koo94a] Ruth Koolish. National Online Meeting ’94. Dr. Dobb’s Developer Update, 1(8):3–??, August 1994. CODEN ???? ISSN 1079-8595.

Koolish:1994:SSE

[Koo94b] Ruth Koolish. SLA: Stepping into the electronic library. Dr. Dobb’s Developer Update, 1(11): 2–??, October 1994. CODEN ???? ISSN 1079-8595.

Kotlarsky:1999:SRT

[Kot99] Anatoly Kotlarsky. The spark real-time kernel. Dr. Dobb’s Journal of Software Tools, 24 (5):80, 82, 84–86, May 1999.

CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/spark.txt.

Kernighan:1999:REL

[KP99] Brian W. Kernighan and Rob Pike. Regular expressions: Languages, algorithms, software. Dr. Dobb’s Journal of Software Tools, 24(4):19–22, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1999/1999_04/
regexp.txt;   http://www.
ddj.com/ftp/1999/1999_04/
regexp.zip.

Kwang:1991:LUI

[KR91] Eng-Kee Kwang and Christopher Rosebrugh. Linking user interface and database objects. Dr. Dobb’s Journal of Software Tools, 16(12):44–52, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Kaliski:1996:MEW

[KR96] Burton S. Kaliski, Jr. and M. J. B. Robshaw. Algorithm alley: Multiple encryption: Weighing security and performance. Dr. Dobb’s Journal of Software Tools, 21(1):123, 124, 126, 127, January 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Krammes:1991:MDP

[Kra91] V. James Krammes. Monitoring distributed printers under Novell NetWare. Dr. Dobb’s Journal of Software Tools, 16(11):82, 84–86, 100, 102, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Kreymborg:1999:TDE

[Kre99] Ron Kreymborg. A task dispatcher for embedded systems. Dr. Dobb’s Journal of Software Tools, 24(8):74, 76–80, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        08/dispatch.txt;   http://
                                        www.ddj.com/ftp/1999/1999_
                                        08/dispatch.zip.

Krten:1994:GRT

[Krt94] Robert Krten. Generating realistic terrain. Dr. Dobb’s Journal of Software Tools, 19(7):26, 28, 88, 90–91, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Krten:1995:ITR

[Krt95] Robert Krten. Indexed text retrieval. Dr. Dobb’s Journal of Software Tools, 20(11):54, 56, 59–61, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Krten:1996:IUN

[Krt96]  Robert Krten.  Improving

Usenet News performance. Dr. Dobb’s Journal of Software Tools, 21(5):66, 68–70, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Krten:1998:DDR

[Krt98] Robert Krten. Device drivers and real-time systems. Dr. Dobb’s Journal of Software Tools, 23(10):34, 36, 38– 39, October 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_10/
rtosdd.txt;   http://www.
ddj.com/ftp/1998/1998_10/
rtosdd.zip.

Kruger:1992:BTC

[Kru92] Anton Kruger. Block truncation compression. Dr. Dobb’s Journal of Software Tools, 17(4):48, 50, 53–55, 104, 106, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Kruger:1994:MCQ

[Kru94a] Anton Kruger. Median-cut color quantization. Dr. Dobb’s Journal of Software Tools, 19(10):46, 48, 50, 52, 54, 91–92, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Krupansky:1994:WMA

[Kru94b] Jack Krupansky. Writing a multimedia app in liana. Dr. Dobb’s Journal of Software Tools, 19 (??):54–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Kruse:1996:USI

[Kru96] Matt Kruse. Using server-side includes. Dr. Dobb’s Journal of Software Tools, 21(2):52, 54, 56, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Kryszak:1992:MCSa

[Kry92a] Richard Kryszak. Multiuser DOS for control systems. part I. Dr. Dobb’s Journal of Software Tools, 17(4):70, 73–74, 76, 78– 79, 100–111, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Kryszak:1992:MCSb

[Kry92b] Richard Kryszak. Multiuser DOS for control systems. part II. building an application. Dr. Dobb’s Journal of Software Tools, 17(5):70, 72, 75, 95– 96, 98, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Knapp:1997:DEC

[KS97] Robert Knapp and Mark Sofroniou. Difference equations and chaos in Mathematica: Symbolic and numerical mathematics at work. Dr. Dobb’s Journal of Software Tools, 22(11):84, 85, 88–90, 95, 98, 99, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Kumar:1995:PSO

[Kum95] Kanchan Kumar. A practical strategy for object-oriented design. Dr. Dobb’s Journal of Software Tools, 20(6):34, 36, 38–39, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Kutter:1992:STE

[Kut92] Michael Kutter. Simulation and testboard for embedded-system design. Dr. Dobb’s Journal of Software Tools, 17(12):62, 66– 68, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Kyle:1991:LDD

[Kyl91]  Jim Kyle.  Loading device

drivers from the DOS command line. Dr. Dobb’s Journal of Software Tools, 16(11):30–34, 39–42, 90–98, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Ladd:1990:EZC

[Lad90a] Scott Robert Ladd. Examining Zortech C++ 2.0. Dr. Dobb’s Journal of Software Tools, 15 (1):64, 66, 69, 71, 73, 112, 114, 117, 119, January 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Ladd:1990:OMC

[Lad90b] Scott Robert Ladd. Optimizing with Microsoft C 6.0. Dr. Dobb’s Journal of Software Tools, 15 (8):84–??, August 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Ladd:1990:POT

[Lad90c] Scott Robert Ladd. Persistent objects in Turbo PASCAL. Dr. Dobb’s Journal of Software Tools, 15(9):36–??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Ladd:1991:WMC

[Lad91] Scott Robert Ladd. Windows meets C++. Dr. Dobb’s Journal of Software Tools, 16(12 (supplement)):29–36, December 1991. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Lafever:1999:VPA

[Laf99] Dave Lafever. Visual programming and assistive technology. Dr. Dobb’s Journal of Software Tools, 24(8):32, 34–36, 38, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Lainhart:1991:IXT

[Lai91] Todd Lainhart. Intrinsics of the X toolkit. Dr. Dobb’s Journal of Software Tools, 16(2):94– ??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Lakos:1997:LSC

[Lak97] John Lakos. Large-scale C++ software design. Dr. Dobb’s Journal of Software Tools, 22 (5):124–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Lam:1995:CCC

[Lam95a] Richard B. Lam. Cross-platform communication classes. Dr. Dobb’s Journal of Software Tools, 20(3):20, 22, 24, 26, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Lam:1995:SMM

[Lam95b] Richard B. Lam. Shared memory and message queues. Dr. Dobb’s Journal of Software Tools, 20(5):28, 30, 97, 100–101, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Lambert:1999:VSC

[Lam99] John Donovan Lambert. VB-SCRIPT and SQL calendars. Dr. Dobb’s Journal of Software Tools, 24(5):40, 42, 46–48, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1999/1999_05/
                                        vbsql.txt;   http://www.ddj.
                                        com/ftp/1999/1999_05/vbsql.
                                        zip.

Lane:1990:CTC

[Lan90] Alex Lane. Cruising with Top-Speed (C compiler). Dr. Dobb’s Journal of Software Tools, 15 (4):72–75, 109–113, April 1990. CODEN DDJOEB. ISSN 1044-789X.

LaOr:1998:CPV

[LaO98] Ofer LaOr. CGI programming and Visual Basic. Dr. Dobb’s Journal of Software Tools, 23 (6):74, 77–79, 93–94, June 1998. CODEN DDJOEB. ISSN 1044-789X.

LaPlante:1995:BOS

[LaP95] John LaPlante. Building an OLE server using Visual C++ 2.0. Dr. Dobb’s Journal of Software Tools, 20(2):82, 84, 86–87, 112–113, February 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Larsen:1995:CDW

[Lar95] Kim S. Larsen. Computing the day of the week. Dr. Dobb’s Journal of Software Tools, 20 (4):125–??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Lawrence:1990:UNN

[Law90] Jeannette “Jet” Lawrence. Untangling neural nets. Dr. Dobb’s Journal of Software Tools, 15 (4):38–44, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Lawless:1995:CSD

[Law95] Jim Lawless. Client/server development and the world wide Web. Dr. Dobb’s Journal of Software Tools, 20(??):35–??, November/December 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Lawless:1997:ETC

[Law97] James K. Lawless. Examining the TAWK compiler: A powerful scripting tool for implementing batch-oriented software. Dr. Dobb’s Journal of Software Tools, 22(5):60, 62, 64, 77, 79, 80, May 1997. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.tasoft.com/
~thompson.

Leavenworth:1993:PSP

[Lea93] Burt Leavenworth. Proxy: a Scheme-based prototyping language. Dr. Dobb’s Journal of Software Tools, 18(3):86, 88, 90, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Leary:1994:NCD

[Lea94] Kevin Leary. Numerical C and DSP. Dr. Dobb’s Journal of Software Tools, 19(8):18– 24, 90, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Lee:1993:DOR

[Lee93a] Chan Y. Lee. Detecting outof-range references. Dr. Dobb’s Journal of Software Tools, 18 (6):50, 52, 54, 56, 128, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Lee:1993:TW

[Lee93b] David Lee. Threads for Windows 3. Dr. Dobb’s Journal of Software Tools, 18(10):84–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Lendecke:1997:UFI

[Len97] Volker Lendecke. UNIX filesystems without I-nodes — the Linux kernel nfs, smbfs, and ncpfs filesystems make it possible to link Linux machines to file servers across a LAN — even though Microsoft’s SMB protocol is not designed to handle UNIX clients like Linux. Volker presents a workaround for this limitation. Dr. Dobb’s Journal of Software Tools, 22(2):60– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Leroy:1996:MIM

[Ler96] Jean-Louis Leroy. Multiple inheritance for MFC 4.0. Dr. Dobb’s Journal of Software Tools, 21(1):84–??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Leroy:1997:PCP

[Ler97] Jean-Louis Leroy. Pipes, channels, and Perl-Win32. Dr. Dobb’s Journal of Software Tools, 22(4):74–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Letcher:1991:GNC

[Let91] John H. Letcher. Getting numeric coprocessors up to speed. Dr. Dobb’s Journal of Software Tools, 16(5):36, 38, 40, 42, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Letner:1996:LSU

[Let96] Charles Letner. Loop splitting under Windows NT. Dr. Dobb’s Journal of Software Tools, 21 (7):50, 52, 54–55, 82–83, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Lewis:1994:MOC

[Lew94] Glenn M. Lewis. Morphing 3-D objects in C++. Dr. Dobb’s Journal of Software Tools, 19 (7):18–20, 22, 86–87, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Lynch:1998:LMP

[LHB^{+}98] Joshua Lynch, Russ Heeschen, Antonio Bigazzi, Bill Wood, Andy Thomas-Cramer, and Niall Murphy. Letters: MFC plug-in components; embedded compilers; theorem proving update; Windows CE. Dr. Dobb’s Journal of Software Tools, 23(10):12, 14, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ddj/1998/1998_
                                        10/index.htm.

Leinecker:1993:PDS

[LHM93] Richard C. Leinecker, S. Heller, and R. Moote. Processor detection schemes. Dr. Dobb’s Journal of Software Tools, 18(6):46, 48–49, 126–127, June 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Li:1994:BIG

[Li94a] Sing Li. Building an Internet global phone. Dr. Dobb’s Journal of Software Tools, 19(??): 46–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Li:1994:ILT

[Li94b] Sing Li. An interview with Linus Torvalds. Dr. Dobb’s Journal of Software Tools, 19(5):26–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Li:1994:WNC

[Li94c] Sing Li. Writing Non-SCSI CD-ROM device drivers. Dr. Dobb’s Journal of Software Tools, 19 (3):102, 104–111, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Liao:1990:SAD

[Lia90] Andrew M. Liao. Self-adjusting data structures. Dr. Dobb’s Journal of Software Tools, 15 (2):44–57, February 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Liaw:1995:RGF

[Lia95] Wilson MacGyver Liaw. Reading GIF files. Dr. Dobb’s Journal of Software Tools, 20(2):56, 58, 60, 103, 106–107, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Lichtman:1992:VBW

[Lic92] Moshe Lichtman. Visual Basic and Windows 3.1 extensions. Dr. Dobb’s Journal of Software Tools, 17(5):80, 82, 84–85, 106, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1994:IAU

[Lin94a] Craig A. Lindley. Image acquisition using TWAIN. Dr. Dobb’s Journal of Software Tools, 19 (10):76, 78–80, 98–100, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1994:RTP

[Lin94b] Craig A. Lindley. Ray tracing and the POV-Ray toolkit. Dr. Dobb’s Journal of Software Tools, 19(7):68, 70, 72, 74, 76, 103, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1994:VDL

[Lin94c] Craig A. Lindley. Very dynamic linking in Windows. Dr. Dobb’s Journal of Software Tools, 19 (??):12–??, ???? 1994. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Linton:1994:FNG

[Lin94d] Mark Linton. Fresco: The next generation InterViews [sidebar]. Dr. Dobb’s Journal of Software Tools, 19(6):49–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1995:JICa

[Lin95a]  Craig A. Lindley.  JPEG-

like image compression. part 1. Dr. Dobb’s Journal of Software Tools, 20(7):50, 52, 54–58, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1995:JICb

[Lin95b]  Craig A. Lindley.  JPEG-

Like image compression. part 2.

Dr. Dobb’s Journal of Software Tools, 20(8):62, 64–66, 102– 105, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Linenbach:1996:RBA

[Lin96] Terris Linenbach. Reusable binary associations in C++. Dr. Dobb’s Journal of Software Tools, 21(6):34, 36–41, 84, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Lippman:1999:ICP

[Lip99a]  Stanley Lippman.  Improv-

ing C++ program performance. Dr. Dobb’s Journal of Software Tools, 24(10):40, 42, 44–45, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/cppperf.txt.

Lipton:1999:JPD

[Lip99b] Paul Lipton. Java proxies for database objects. Dr. Dobb’s Journal of Software Tools, 24 (5):34, 36–39, May 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_05/
jproxy.txt.

Lischner:1998:PTD

[Lis98] Ray Lischner. Programmer’s toolchest: The Delphi Open Tools API. Dr. Dobb’s Journal of Software Tools, 23(2):88, 90– 92, 104–105, February 1998.

Liu:1990:ACE

[Liu90] Hsi-Chiu Liu. Algebraic codes for error detection and correction. Dr. Dobb’s Journal of Software Tools, 15(12):46, 48, 50– 51, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Lutter:1996:SRW

[LMWR^{+}96] Petra Lutter, Michael Muller-Wernhart, Jurgen Ramharter, Frank Rattay, and Peter Slowik. Speech research with Wave GL. Dr. Dobb’s Journal of Software Tools, 21(11):50–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Loeb:1999:JA

[Loe99] Bill Loeb. The Java 2D API. Dr. Dobb’s Journal of Software Tools, 24(2):44, 46–49, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/j2d.txt.

Lomax:1998:VBF

[Lom98] Paul Lomax. Visual Basic 6 filesystem objects. Dr. Dobb’s Journal of Software Tools, 23 (10):104, 106–107, 114–115, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
10/vb6file.txt.

Lord:1996:HFW

[Lor96] Robert Lord. Help file Web links. Dr. Dobb’s Journal of Software Tools, 21(7):78, 80– 81, 84–85, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Lowell:1992:AEC

[Low92] Robert Lowell. The Am29000 as an embedded controller (for laser printers). Dr. Dobb’s Journal of Software Tools, 17(3):58, 60–62, 64, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Lowrey:1995:EVW

[Low95] John Fritz Lowrey. Environment variables and Windows 3.1. Dr. Dobb’s Journal of Software Tools, 20(10):84, 86, 88– 89, 123–124, October 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Lowrey:1998:DWA

[Low98a] Fritz Lowrey. Delphi 4 and the WNet API. Dr. Dobb’s Journal of Software Tools, 23(12):102, 104–106, 107, December 1998. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/1998/1998_12/../..
/../ftp/1998/1998_12/wnet.
txt;   http://www.ddj.com/
ddj/1998/1998_12/../../../
ftp/1998/1998_12/wnet.zip.

Lowrey:1998:EWD

[Low98b] Fritz Lowrey. Examining the Win32 Debug API. Dr. Dobb’s Journal of Software Tools, 23 (11):98, 100–102, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1998/1998_11/
QDBUG.TXT;   http://www.ddj.
com/ftp/1998/1998_11/QDBUG.
ZIP.

Loynd:1992:MRP

[Loy92] Kerry Loynd. Mixing realand protected-mode code. Dr. Dobb’s Journal of Software Tools, 17(2):24, 26, 28, 82, 84, 86, 88, 90, February 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Lauzzana:1990:IBS

[LP90] Raymond G. Lauzzana and Denise E. M. Penrose. Implementing bicubic splines. Dr. Dobb’s Journal of Software Tools, 15(8):48–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Lukatela:1992:SDV

[LR92] Hrvoje Lukatela and John Russell. Spatial data and the Voronoi tessellation. Dr. Dobb’s Journal of Software Tools, 17 (12):18–20, 24–25, 96, 98– 100, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Locke:1996:WSS

[LR96] Jonathan Locke and Jeffrey Richter. Windows 95 subclassing and superclassing. Dr. Dobb’s Journal of Software Tools, 21(??):35–??, March/ April 1996. CODEN DDJOEB. ISSN 1044-789X.

Lentini:1999:AI

[LRT99] Russell P. Lentini, Goutham P. Rao, and Jon N. Thies. Agent itineraries. Dr. Dobb’s Journal of Software Tools, 24(5): 60, 62, 64, 66, 68, 70, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/agent.txt.

Lu:1999:PB

[LS99] Paul Lu and William Stallings.

Programmer’s bookshelf: STL & gigabit Ethernet. Dr. Dobb’s Journal of Software Tools, 24 (1):127, 129, January 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Lu:1994:PBU

[Lu94] Zongnan H. Lu. Postman: A bridge to the Unix mail system. Dr. Dobb’s Journal of Software Tools, 19(7):64, 66, 101–102, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Lu:1995:RNP

[Lu95] Zongnan H. Lu. Remote network printing. Dr. Dobb’s Journal of Software Tools, 20(1):62, 64, 66, 100–101, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Luse:1994:BFF

[Lus94] Marv Luse. The BMP file format. Dr. Dobb’s Journal of Software Tools, 19(10):18–20, 22, 82, 84, 85, September 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Lussier:1997:ISR

[Lus97] Kyle Lussier. Implicit surfaces and real-time graphics. Dr. Dobb’s Journal of Software Tools, 22(7):18–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Luu:1994:BP

[Luu94] José Luu. Binary-data portability. Dr. Dobb’s Journal of Software Tools, 19(3):18–20, 22, 24, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Lyke:1990:RT

[Lyk90] Daniel Lyke. Ray tracing. Dr. Dobb’s Journal of Software Tools, 15(9):152–??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Lynnworth:1995:THP

[Lyn95] Ann Lynnworth. Tracking home page hits. Dr. Dobb’s Journal of Software Tools, 20(??):27–??, November/December 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Lyppens:1997:CEC

[Lyp97a] Hugo Lyppens. Convolutional error-control codes: An encoded and a Viterbi decoder in C++. Dr. Dobb’s Journal of Software Tools, 22(11):46, 48–53, 92– 93, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Lyppens:1997:REC

[Lyp97b] Hugo Lyppens. Reed–Solomon error correction. Dr. Dobb’s Journal of Software Tools, 22 (1):30–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Maeder:1992:DMP

[Mae92] Roman E. Maeder. The design of the Mathematica programming language: a single paradigm provides surprising diversity. (tutorial). Dr. Dobb’s Journal of Software Tools, 17 (4):86–88, 90, 92, 94, 97, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Maher:1992:CB

[Mah92] Tim Maher. A C++ beautifier.

Dr. Dobb’s Journal of Software Tools, 17(12 (special supplement)):23s–27s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Maier:1991:FPR

[Mai91] William L. Maier. A fast pseudo random number generator. Dr. Dobb’s Journal of Software Tools, 16(5):152, 154–157, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Mak:1990:HOE

[Mak90] Nico Mak. Handling OS/2 error codes. Dr. Dobb’s Journal of Software Tools, 15(8): 134–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Malone:1996:BS

[Mal96] Michael S. Malone. Beyond semiconductors. Dr. Dobb’s Developer Update, 3(1):1, 3, January 1996. CODEN ???? ISSN 1079-8595.

Mann:1992:UDI

[Man92] Daniel Mann. The Universal Debugger Interface. Dr. Dobb’s Journal of Software Tools, 17 (9):58, 60, 62, 64, 66, 68, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Mann:1993:BPL

[Man93] Steve Mann. The Beta programming language. Dr. Dobb’s Journal of Software Tools, 18 (11):56–66, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Mangano:1994:ADG

[Man94a] Salvatore R. Mangano. Algorithms for directed graphs. Dr. Dobb’s Journal of Software Tools, 19(4):92, 94, 103, 106– 7, 147, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Mangen:1994:RRP

[Man94b] Alain Mangen. Ray: A ray-tracing program in C++. Dr. Dobb’s Journal of Software Tools, 19(7):40–43, 96, 98, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Manber:1999:DAI

[Man99a] Udi Manber. Designing algorithms incrementally: Finding powerful problem-solving techniques. Dr. Dobb’s Journal of Software Tools, 24(4):26, 28–31, April 1999. CODEN DDJOEB. ISSN 1044-789X.

Mann:1999:WAP

[Man99b] Steve Mann. The wireless application protocol. Dr. Dobb’s Journal of Software Tools, 24 (10):56, 58–60, 62–64, 66, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/wap.txt;   http://www.
ddj.com/ftp/1999/1999_10/
wap.zip.

Margulis:1990:PRE

[Mar90] Neil Margulis. Programming RISC engines. Dr. Dobb’s Journal of Software Tools, 15 (2):116–117, 128–132, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Marais:1994:OS

[Mar94a] Johannes L. Marais. Oberon System 3. Dr. Dobb’s Journal of Software Tools, 19(11): 42–50, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Martin:1994:SCE

[Mar94b] Michelle San Martin. Sourcecode escrow for intellectual property protection. Dr. Dobb’s Developer Update, 1(7): 2, September 1994. CODEN ???? ISSN 1079-8595.

Martino:1995:DWE

[Mar95a] Paul J. Martino. The detrimental wire exclusion heuristic. Dr. Dobb’s Journal of Software Tools, 20(4):28, 30, 32–33, 88, 90–91, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Martino:1995:IFM

[Mar95b] Paul J. Martino. Implementing flicker-free motion. Dr. Dobb’s Journal of Software Tools, 20(9):88–89, 114, 116– 117, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Markham:1997:ISP

[Mar97] Tom Markham. Internet Security Protocol. Dr. Dobb’s Journal of Software Tools, 22 (6):70–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Matthews:1995:PVW

[Mat95] Don Matthews. Porting VXDs from Windows 3.1 to Windows 95. Dr. Dobb’s Journal of Software Tools, 20(11):64, 66– 68, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Mathews:1996:EDA

[Mat96]  G. Jason Mathews.  Eval-

uating data-compression algorithms. Dr. Dobb’s Journal of Software Tools, 21(1):50– 53, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Maurice:1993:PFF

[Mau93] Michael P. Maurice. The PIF file format or topview (sort of) lives! Dr. Dobb’s Journal of Software Tools, 18(7):127–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

May:1995:PTM

[May95] William D. May. Programming TI’s multimedia video processor. Dr. Dobb’s Journal of Software Tools, 20(11):46, 48, 50– 52, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Mazur:1992:MAC

[Maz92] Beth Mazur. Moving from assembly to C. Dr. Dobb’s Journal of Software Tools, 17(8):72, 74, 76, 78, 80, 82, 84, 120, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Morita:1998:LSR

[MBK^{+}98] Aki Morita, John Burkhardt, Steve Kelleher, Mike Kelley, Jeff Feeley, Byron Miller, and Ted Velkoff. Letters: Self-registering objects; C++ namespaces; online op-ed; year 2000; book review response; design by contract. Dr. Dobb’s Journal of Software Tools, 23(11):12, 14, November 1998. CODEN DDJOEB. ISSN 1044-789X.

Milligan:1998:ESB

[MC98] Tom Milligan and Steve Coffin. Embedded systems: Building intelligent Web-based control systems. Dr. Dobb’s Journal of Software Tools, 23(3):56, 58, 60, 62, 64, 66, 68, 70–71, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
03/control.txt.

McArthur:1994:WWW

[McA94] Douglas C. McArthur. World Wide Web and HTML. Dr. Dobb’s Journal of Software Tools, 19(15):18–20, 22, 24, 26, 86, December 1994. CODEN DDJOEB. ISSN 1044-789X.

McBride:1997:PTA

[McB97] Blake McBride. Programmer’s toolchest — advanced objectoriented features for C/C++. Dr. Dobb’s Journal of Software Tools, 22(8):78, 80, 82, August 1997. CODEN DDJOEB. ISSN 1044-789X.

McCracken:1991:SPM

[McC91] David McCracken. Software partitioning for multitasking communication. Dr. Dobb’s Journal of Software Tools, 16 (9):64, 66, 68, 70, 72, 74–75, 96, September 1991. CODEN DDJOEB. ISSN 1044-789X.

McCarthy:1995:WOD

[McC95] Dennis R. McCarthy. Writing ODBC drivers. Dr. Dobb’s Journal of Software Tools, 20(11): 88–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

McCoy:1996:NCR

[McC96a] John H. McCoy. Networking CD-ROMs using cached CD images. Dr. Dobb’s Journal of Software Tools, 21(8):74, 76–77, 96–99, August 1996. CODEN DDJOEB. ISSN 1044-789X.

McCoy:1996:PBH

[McC96b] John H. McCoy. Programmer’s bookshelf: Hooked on Java, by A. van Hoff, S. Shaio, and O. Starbuch [review]. Dr. Dobb’s Journal of Software Tools, 21 (4):128–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

McCartju:1997:URT

[McC97] Adrian McCartju. Unit and regression testing — testing a function, module, or object in isolation from the rest of the program–referred to as “unit testing” coupled with rerunning tests to detect unexpected changes in behavior (“regression testing”) can dramatically reduce your bug counts. Adrian describes how to build effective units tests and automate regression testing. Dr. Dobb’s Journal of Software Tools, 22(2):18– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

McGary:1993:APL

[McG93] Diane McGary. The art of product launches. Dr. Dobb’s Journal of Software Tools, 18(11): 80–83, October 1993. CODEN DDJOEB. ISSN 1044-789X.

McGrath:1994:OMC

[McG94a] Gary McGrath. Optimizing MC68882 code. Dr. Dobb’s Journal of Software Tools, 19 (6):58, 60, 62, 64, 66, 98–99, June 1994. CODEN DDJOEB. ISSN 1044-789X.

McGrath:1994:SAB

[McG94b] Gary McGrath. Signal analysis via the bootstrap. Dr. Dobb’s Journal of Software Tools, 19 (2):48, 50, 52, 54–57, 81– 82, February 1994. CODEN DDJOEB. ISSN 1044-789X.

McGrath:1998:IPX

[McG98a] Sean McGrath. Internet programming: XML programming in Python. Dr. Dobb’s Journal of Software Tools, 23(2):82, 84– 87, 101–104, February 1998.

McGrath:1998:PBT

[McG98b] Sean McGrath. Programmer’s bookshelf: The Tcl/Tk and Python scripting environments. Dr. Dobb’s Journal of Software Tools, 23(10):143, 145, October 1998. CODEN DDJOEB.

ISSN 1044-789X. URL http://

                                        www.ddj.com/ddj/1998/1998_
                                        10/index.htm.

McGrath:1998:XXL

[McG98c] Sean McGrath. XLink: The XML linking language. Dr. Dobb’s Journal of Software Tools, 23(12):94, 96–101, December 1998. CODEN DDJOEB. ISSN 1044-789X.

McMahon:1990:MEC

[McM90] William J. McMahon. Multiplexing error codes. Dr. Dobb’s Journal of Software Tools, 15 (2):60–65, February 1990. CO-DEN DDJOEB. ISSN 1044-789X.

McRae:1993:AMP

[McR93] Eric McRae. Avoiding microcontroller processing pile-ups (wheelchair). Dr. Dobb’s Journal of Software Tools, 18(11): 84–92, October 1993. CODEN DDJOEB. ISSN 1044-789X.

McRae:1995:T

[McR95] Eric McRae. Timing for the 68332. Dr. Dobb’s Journal of Software Tools, 20(1):50, 52, 54, 56, 58, 60, January 1995. CO-DEN DDJOEB. ISSN 1044-789X.

McRae:1996:BRO

[McR96a] Eric McRae. Benchmarking real-time operating systems. Dr. Dobb’s Journal of Software Tools, 21(5):??, May 1996. CO-DEN DDJOEB. ISSN 1044-789X.

McRae:1996:BRT

[McR96b] Eric McRae. Benchmarking real-time operating systems. Dr. Dobb’s Journal of Software Tools, 21(5):48, 50, 52, 54, 56– 58, 83–86, May 1996. CODEN DDJOEB. ISSN 1044-789X.

McRae:1996:PBB

[McR96c] Eric McRae. Programmer’s bookshelf: Benchmarks and DSP algorithms. Dr. Dobb’s Journal of Software Tools, 21 (11):128, November 1996. CO-DEN DDJOEB. ISSN 1044-789X.

McSwain:1992:OAL

[McS92]  Donald J. McSwain.  An

object-oriented assembly language macro library. Dr. Dobb’s Journal of Software Tools, 17 (3):24, 26, 30, 32, 34, 84–87, March 1992. CODEN DDJOEB. ISSN 1044-789X.

McRae:1996:CAR

[ME96] Eric McRae and Paul M. Embree. C algorithms for realtime DSP. Dr. Dobb’s Journal of Software Tools, 21(11):128– ??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Meckler:1997:JIA

[Mec97] Andrew Meckler. Java and inter-applet communication. Dr. Dobb’s Journal of Software Tools, 22(10):46, 48, 50–53, 103, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Medwin:1994:DCE

[Med94] Larry Medwin. A Dual-UI constrain equation solver in C++. Dr. Dobb’s Journal of Software Tools, 19(6):44, 46, 48– 50, 93–94, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Mello:1998:ESF

[Mel98] Russell W. Mello. Embedded systems: Flexible system control and special-purpose languages. Dr. Dobb’s Journal of Software Tools, 23(2):68, 70, 72, 74, 76– 80, February 1998.

Merritt:1994:ECP

[Mer94] Dennis Merritt. Extending C with Prolog. Dr. Dobb’s Journal of Software Tools, 19(8): 78–82, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Metz:1996:JJN

[Met96] Eldon Metz. Java, Jfactory, and network development. Dr. Dobb’s Journal of Software Tools, 21(9):78, 80–82, 85– 87, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Meyer:1992:DPC

[Mey92] Andreas Meyer. Developing a portable C++ GUI class library. Dr. Dobb’s Journal of Software Tools, 17(11):102, 105– 107, 109, November 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Meyer:1993:JGC

[Mey93] Peter J. G. Meyer. Julian and Gregorian calendars. Dr. Dobb’s Journal of Software Tools, 18 (3):152, 154, 156, 158–159, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Meyers:1999:IOS

[Mey99] Scott Meyers. Implementing operator ->* for smart pointers. Dr. Dobb’s Journal of Software Tools, 24(10):32, 34–39, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/smartptr.txt;   http://
www.ddj.com/ftp/1999/1999_
10/smartptr.zip.

Moore:1998:DCY

[MF98] Robert L. Moore and D. Gregory Foley. Date compression and Year 2000 challenges. Dr. Dobb’s Journal of Software Tools, 23(5):20–22, 24, 109, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
05/y2kcomp.txt;   http://
www.ddj.com/ftp/1998/1998_
05/y2kcomp.zip.

McGregor:1996:FFC

[MFCM96] D. R. McGregor, R. J. Fryer, P. Cockshott, and P. Murray. Faster fractal compression. Dr. Dobb’s Journal of Software Tools, 21(1):34, 36, 38–40, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Mackenzie:1998:LPS

[MGL^{+}98] Don Mackenzie, Andrew J.

Gryc, Graziano Lo Russo, Gary Clouse, C. J. Hinke, Bruce E.

Hogman, Thomas Fleischer, and John Graham-Cumming. Letters: The passport system does work; real-time sound; C++ versus Java; online op-ed; hard encryption; Y2K; VerCheck update. Dr. Dobb’s Journal of Software Tools, 23(12):12, 16– 17, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Miller:1991:MOD

[Mil91] William M. Miller. A minimal object-oriented debugger for C++. Dr. Dobb’s Journal of Software Tools, 16(10):16–18, 20, 23, 110, 112, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Miles:1994:WPS

[Mil94] Troy A. Miles. Writing PCM-CIA software. Dr. Dobb’s Journal of Software Tools, 19(6):150, 152, 154, 156–157, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Mirho:1992:BGR

[Mir92] Charles Albert Mirho. Bridging the gap with resident/sub - /C improving text exchanges between DOS and Windows applications. Dr. Dobb’s Journal of Software Tools, 17(5):76, 79, 100, 102, 104, May 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Mirho:1994:PQ

[Mir94] Charles Albert Mirho. Programming the QUANTUMdsp. Dr. Dobb’s Journal of Software Tools, 19(??):42–??, ???? 1994.

CODEN DDJOEB. ISSN 1044-789X.

Mirho:1995:VVD

[Mir95] Charles Albert Mirho. VTOOLSD for VXD development. Dr. Dobb’s Journal of Software Tools, 20(12):90–??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Mirchandani:1996:BPL

[Mir96] Thor Mirchandani. Building parsers with Leopurd. Dr. Dobb’s Journal of Software Tools, 21(3):30, 32–34, 91–93, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Mischel:1990:EP

[Mis90]  James Mischel.  Extending

printf(). Dr. Dobb’s Journal of Software Tools, 15(8): 60–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Mischel:1996:DG

[Mis96]  James Mischel.  DirectPlay

gotchas [sidebar]. Dr. Dobb’s Journal of Software Tools, 21 (??):10–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1996:PB

[Mit96] Phil Mitchell. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 21(8): 131–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Meyers:1997:ECP

[MK97] Scott Meyers and Martin Klaus.

Examining C++ program analyzers — our authors examine a bevy of off-the-shelf tools that parse and analyze C++ source code, enabling you to detect troublesome C++ code via static analysis. Dr. Dobb’s Journal of Software Tools, 22(2):68– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

McCoy:1993:NAC

[ML93] John H. McCoy and Wuhsiung Lu. Network access to CD-ROMS. Dr. Dobb’s Journal of Software Tools, 18(8):72, 74, 78– 80, 113, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1997:PBP

[ML97] Phil Mitchell and Derek Leebaert. Programmer’s bookshelf — Phil goes back to the future in his look at Thomas Landauer’s The Trouble With Computers and The Future of Software. Dr. Dobb’s Journal of Software Tools, 22(2):117–??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

McCormick:1998:A

[MM98a] Hays W. McCormick and Raphael Malveaux. AntiPatterns. Dr. Dobb’s Journal of Software Tools, 23(6):44, 46, 48–50, June 1998. CODEN DDJOEB. ISSN 1044-789X.

McLaughlin:1998:RTE

[MM98b] Michael J. McLaughlin and Alan Moore. Real-time extensions to UML. Dr. Dobb’s Journal of Software Tools, 23(12):82, 84, 86–93, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Martin:1999:JDI

[MM99] David H. Martin and Johnny Martin. Java and digital images. Dr. Dobb’s Journal of Software Tools, 24(5):72, 74–76, 78–79, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/snapshot.txt;   http://
www.ddj.com/ftp/1999/1999_
05/snapshot.zip.

Minsky:1990:SP

[MMB^{+}90] M. Minsky, J. McCarthy, R. S. Boyer, R. Stallman, and S. Garfinkel. Software patents. Dr. Dobb’s Journal of Software Tools, 15(11):56, 58, 62, 65–67, 70–73, November 1990. CODEN DDJOEB. ISSN 1044-789X.

MacKie-Mason:1994:EI

[MMV94] Jeffrey K. MacKie-Mason and Hal Varian. The economics of the Internet. Dr. Dobb’s Journal of Software Tools, 19(??):6–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Mockridge:1994:EE

[Moc94] David Mockridge. EPROM emulation. Dr. Dobb’s Journal of Software Tools, 19(10):56, 58, 60–63, 93–94, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Moeser:1990:MC

[Moe90] Robert A. Moeser. A memory controller. Dr. Dobb’s Journal of Software Tools, 15(5):58–68, 111–113, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Montefusco:1996:EHS

[Mon96] Andrew Montefusco. Extending HTTP servers with DGI and TGI. Dr. Dobb’s Journal of Software Tools, 21(12): 68–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Monson:1997:AAC

[Mon97] Lynn Monson. Algorithm alley; classifying text with ID3 and C4.5. Dr. Dobb’s Journal of Software Tools, 22(10):117–119, 128–130, October 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Monson:1998:WS

[Mon98] Lynn Monson. The WIDL specification. Dr. Dobb’s Journal of Software Tools, 23(11):92, 94–96, November 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1998/1998_11/WIDL.
                                        TXT;   http://www.ddj.com/
                                        ftp/1998/1998_11/WIDL.ZIP.

Moore:1992:PRP

[Moo92] David L. Moore. Programming the 29050 (RISC programming). Dr. Dobb’s Journal of Software Tools, 17(1):34, 36, 38, 40, 42, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Moote:1993:PDP

[Moo93] Robert Moote. Pentium detection [sidebar in “Processor Detection Schemes” (R. C. Leinecker)]. Dr. Dobb’s Journal of Software Tools, 18(6):49–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Moore:1995:OFA

[Moo95]  David L. Moore.  Object-

oriented facilities in Ada 95. Dr. Dobb’s Journal of Software Tools, 20(10):28, 30, 32, 34– 35, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Moore:1998:JQH

[Moo98] L. Richard Moore. Java Q&A:

How do I create a streaming audio Java applet? Dr. Dobb’s Journal of Software Tools, 23 (5):122–124, May 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_05/
jqa598.txt;   http://www.
ddj.com/ftp/1998/1998_05/
jqa598.zip.

Morgan:1991:CFL

[Mor91a] Don Morgan. A COFF file loader for the 34010. Dr. Dobb’s Journal of Software Tools, 16 (7):42, 44, 46, 93–95, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Morgan:1991:DFC

[Mor91b] Don Morgan. Decimal fractional conversion. Dr. Dobb’s Journal of Software Tools, 16(8): 76, 78, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Morgan:1991:IG

[Mor91c] Don Morgan. Implementing the GPIB. Dr. Dobb’s Journal of Software Tools, 16(5):56, 58– 60, 96, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Morgan:1991:UGG

[Mor91d] Don Morgan. Understanding the GPIB (general purpose instrumentation bus). Dr. Dobb’s Journal of Software Tools, 16 (4):48, 50, 52–53, 92, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Morrow:1991:GA

[Mor91e] Michael Morrow. Genetic algorithms. Dr. Dobb’s Journal of Software Tools, 16(4):26, 28, 30, 32, 86, 88–89, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Morgan:1998:BDA

[Mor98] Bryan Morgan. Building distributed applications with Java and CORBA. Dr. Dobb’s Journal of Software Tools, 23 (4):94, 96–99, 104–105, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/corbajav.txt;   http://
www.ddj.com/ftp/1998/1998_
04/corbajav.zip.

Motteler:1991:SPA

[Mot91] Fred Motteler. Statistical performance analysis. Dr. Dobb’s Journal of Software Tools, 16 (12):68–73, 76, 100–104, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Motteler:1993:APF

[Mot93] Frederick C. Motteler. Arbitrary precision floating-point arithmetic. Dr. Dobb’s Journal of Software Tools, 18(9):28, 30, 32, 34, 84, 86–87, September 1993. CODEN DDJOEB. ISSN 1044-789X.

McClanahan:1995:HPP

[MPV95] Kip McClanahan, Mike Phillip, and Mark VandenBrink. Highperformance programming for the PowerPC. Dr. Dobb’s Journal of Software Tools, 20(??):29– ??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

McRae:1996:B

[MR96] Grace McRae and Eric Richard.

The BenchmarkBook. Dr. Dobb’s Journal of Software Tools, 21(11):128–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Marais:1999:AWW

[MR99] Hannes Marais and Tom Rodeheffer. Automating the Web with Webl. Dr. Dobb’s Journal of Software Tools, 24(1):20– 23, 26–27, January 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/ftp/1999/1999_01/
webl.txt.

Anonymous:1999:L

[MSE^{+}99] Roger McFarlane, Costa Siourbas, Nassie Evyatar, Joseph Fenton, Justin Gale, Robert Stafford, Sam Holden, and Oleg Kiselyov. Letters: WinCE’s CeRapiInvoke API; Year 2000 fix; DebugAPI; template-driver Web pages; DDA and fast image scaling. Dr. Dobb’s Journal of Software Tools, 24(1):10, 12, 16–17, January 1999. CODEN DDJOEB. ISSN 1044-789X.

Mudd:1996:SCP

[Mud96]  John Mudd.  Speeding up

C-tree Plus database searches. Dr. Dobb’s Journal of Software Tools, 21(4):76, 78–80, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Mueller:1994:TGA

[Mül94] Klaus Müller. Think globally act locally: Inside the Windows instance data manager. Dr. Dobb’s Journal of Software Tools, 19(4):125–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Murray:1996:SSP

[Mur96] James D. Murray. SPIFF: Still picture interchange file format. Dr. Dobb’s Journal of Software Tools, 21(7):34, 36, 38, 40–41, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Musser:1993:MCL

[Mus93]  John Musser.  A multime-

dia class library for Windows. Dr. Dobb’s Journal of Software Tools, 18(7):84, 86, 88–90, 102– 103, 106, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Musgrove:1999:PIP

[Mus99] A. J. Musgrove. The Palmpilot’s infrared port: Communication without cables. Dr. Dobb’s Journal of Software Tools, 24(4):52, 56, 58–61, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
04/batlship.txt;   http://
www.ddj.com/ftp/1999/1999_
04/batlship.zip.

Myers:1991:WDY

[Mye91a] B. Myers. WINTHERE (does your program know when Windows is running?). Dr. Dobb’s Journal of Software Tools, 16 (1):116–117, January 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Myers:1991:SRV

[Mye91b] Ben Myers. Saving and restoring VGA screens. Dr. Dobb’s Journal of Software Tools, 16(7):24, 26, 28, 30, 79–82, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Myers:1997:EMC

[Mye97] Nathan Myers. The empty member C++ optimization: The STL is flexible but does it use too much memory? Nathan shows how the STL — and you — can use empty subobjects without bloating your data requirements. Dr. Dobb’s Journal of Software Tools, 22(8):30, 32, 90, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Myers:1998:CLI

[Mye98]  Nathan Myers.  C++ lo-

cales: International programming gets easier. Dr. Dobb’s Journal of Software Tools, 23 (8):42, 44–45, 90–91, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        08/locales.txt.

Nielsen:1996:ADP

[NA96] Mark Nielsen and Nick Abdo. Applying design patterns to PowerBuilder. Dr. Dobb’s Journal of Software Tools, 21(6):78, 80, 82, 91–92, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Nackman:1997:CIC

[Nac97] Lee R. Nackman. CodeStore and incremental C++: Why wait for slow builds. Dr. Dobb’s Journal of Software Tools, 22(12):92–95, 107, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Najork:1995:VP

[Naj95] Marc Najork. Visual programming in 3-D. Dr. Dobb’s Journal of Software Tools, 20(12):18–20, 22, 24, 28, 31, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Nadeau:1996:OPV

[NAM96] David R. Nadeau, Andrea L. Ames, and John L. Moreland. Optimizing the performance of VRML worlds. Dr. Dobb’s Journal of Software Tools, 21(7):16– 18, 20, 22, 24, July 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Nayfeh:1993:CAS

[Nay93] Basem A. Nayfeh. Cellular automata for solving mazes. Dr. Dobb’s Journal of Software Tools, 18(2):32, 37–38, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Neeriemer:1999:IPE

[Nee99] Troy Neeriemer. Internet programming: Examining PerL-DAP: Simplifying LDAP access. Dr. Dobb’s Journal of Software Tools, 24(4):72, 74–77, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
04/perldap.txt;   http://
www.ddj.com/ftp/1999/1999_
04/perldap.zip.

Nelson:1990:LIE

[Nel90] Mark R. Nelson. Location is everything! Dr. Dobb’s Journal of Software Tools, 15(1): 124–126, 128, 152–153, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1991:ACS

[Nel91a] Mark R. Nelson. Arithmetic coding and statistical modeling. Dr. Dobb’s Journal of Software Tools, 16(2):16–18, 20, 22, 24, 26, 29, 104, 106–108, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1991:DDC

[Nel91b] Mark R. Nelson. DDJ data compression contest results. Dr. Dobb’s Journal of Software Tools, 16(11):62–??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1992:FVU

[Nel92] Mark R. Nelson. File verification using CRC 32-bit cyclical redundancy check. Dr. Dobb’s Journal of Software Tools, 17(5):64– 67, 92, 94, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1993:FRM

[Nel93a] T. Nelson. Finding run-time memory errors. A sophisticated tool for the thorniest of bugs. Dr. Dobb’s Journal of Software Tools, 18(12):34, 36, 38, 40–41, 92–93, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1993:FRT

[Nel93b] Taed Nelson. Finding run-time memory errors. Dr. Dobb’s Journal of Software Tools, 18 (11):34–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1996:AAF

[Nel96a] Mark R. Nelson. Algorithm alley: Fast string searches with suffix trees. Dr. Dobb’s Journal of Software Tools, 21(8): 115–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1996:DCB

[Nel96b] Mark R. Nelson. Data compression with the Burrows Wheeler transform. Dr. Dobb’s Journal of Software Tools, 21(9):46, 48– 50, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1996:PQS

[Nel96c] Mark R. Nelson. Priority queues and the STL. Dr. Dobb’s Journal of Software Tools, 21(1):18– 20, 22, 24, 26, 96, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1997:ZCL

[Nel97] Mark R. Nelson. The ZLIB compression library — Mark examines zlib, a library of C routines that can be used to compress or expand files using the same deflate algorithm popularized by PKZIP 2.0. Dr. Dobb’s Journal of Software Tools, 22 (1):36–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Nestved:1994:MCC

[Nes94] Timothy D. Nestved. Migrating C code to Unicode. Dr. Dobb’s Journal of Software Tools, 19 (8):28, 30, 32, 91–93, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Neubert:1998:AA

[Neu98] Karl-Dietrich Neubert. Algorithm alley: The Flashsort1 algorithm. Dr. Dobb’s Journal of Software Tools, 23(2):123–125, 131, February 1998.

Newcomer:1993:DRS

[New93a] Joseph M. Newcomer. Debugging real-time systems. Dr. Dobb’s Journal of Software Tools, 18(7):55–56, 58, 60, 62, 98, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Newcomer:1993:FDP

[New93b] Joseph M. Newcomer. Faking DDE with private servers. Dr. Dobb’s Journal of Software Tools, 18(10):37–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Newcomer:1993:MDB

[New93c] Joseph M. Newcomer. Modeless dialog boxes for Windows. Dr. Dobb’s Journal of Software Tools, 18(5):26, 28, 30, 82, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Newcomer:1993:PP

[New93d] Joseph M. Newcomer. Profiling for performance. Dr. Dobb’s Journal of Software Tools, 18 (1):80–81, 85–87, 106, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Newcomer:1994:AWP

[New94a] Joseph M. Newcomer. Avoiding Windows PATH cram. Dr. Dobb’s Journal of Software Tools, 19(??):54–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Newcomer:1994:GAL

[New94b] Joseph M. Newcomer. A generic About … Box handler. Dr. Dobb’s Journal of Software Tools, 19(??):18–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Nickerson:1996:BRJc

[NFP96] Doug Nickerson, David H.

Friedel, Jr., and Anthony Potts. Book review: Java Programming Language Handbook. Dr. Dobb’s Journal of Software Tools, 21(12):134–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Nguyen:1996:EVD

[Ngu96]  Matthew Nguyen.  Extend-

ing the visitor design pattern. Dr. Dobb’s Journal of Software Tools, 21(10):34–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Nicholson:1991:DPG

[Nic91a] Robert T. Nicholson. Designing a portable GUI toolkit (five principles can unravel knotty design problems). Dr. Dobb’s Journal of Software Tools, 16 (1):68, 70–72, 74–75, 117, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Nicolaisen:1991:QAP

[Nic91b] Nancy Nicolaisen. Quick approximations of polygonal areas using BitBlt. Dr. Dobb’s Journal of Software Tools, 16 (12 (supplement)):7–16, December 1991. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Nichol:1992:ETV

[Nic92] Scott Nichol. Extending Turbo vision: Replacing the idle method. Dr. Dobb’s Journal of Software Tools, 17(11):192, 194–196, 198, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Nicholson:1993:PW

[Nic93a] Andy Nicholson. 64-bit programming in a 32-bit world. Dr. Dobb’s Journal of Software Tools, 18(1):34, 39–40, 42, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Nicholson:1993:HHL

[Nic93b] Andy Nicholson. HIPPI and high-performance LANS. Dr. Dobb’s Journal of Software Tools, 18(6):58, 60, 64, 66, 68, 130–131, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Nickerson:1996:BRD

[Nic96a] Doug Nickerson. Book review: David Flanagan’s Java in a Nutshell. Dr. Dobb’s Journal of Software Tools, 21(12):134– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Nickerson:1996:BRP

[Nic96b] Doug Nickerson. Book review: Peter van der Linden, Just Java. Dr. Dobb’s Journal of Software Tools, 21(12):134–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Nilsson:1997:TJ

[Nil97] Stefan Nilsson. Treaps in Java. Dr. Dobb’s Journal of Software Tools, 22(7):40, 42–44, 84, 86, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Nilsen:1998:PSF

[Nil98] Kelvin Nilsen. picoPERC: A small-footprint dialect of Java. Dr. Dobb’s Journal of Software Tools, 23(3):50, 52–54, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
03/javabj.zip.

Nilsson:1998:IPF

[NK98] Stefan Nilsson and Gunnar Karlsson. Internet programming fast IP routing with LCtries: Achieving gbit/sec speed in software. Dr. Dobb’s Journal of Software Tools, 23(8): 70, 72–75, August 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_08/
router.zip.

Nelson:1997:JZF

[NN97] Mark R. Nelson and Mark Nelson. Java and the zip file format: Using Java to compress and extract files. Dr. Dobb’s Journal of Software Tools, 22(12): 50, 52–54, 102, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Noga:1999:DLM

[Nog99] Markus L. Noga. Designing the LegOS multitasking operating system. Dr. Dobb’s Journal of Software Tools, 24(11):66, 68, 70, 73–74, 76, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Nolan:1990:RDA

[Nol90] Tom Nolan. Real-time data acquisition using DMA. Dr. Dobb’s Journal of Software Tools, 15(1):28, 30, 32, 34, 36– 37, 94, 96, January 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Norwood:1991:MWP

[Nor91] John Norwood. Mixed-language Windows programming. Dr. Dobb’s Journal of Software Tools, 16(10):104–108, 130, October 1991. CODEN DDJOEB. ISSN 1044-789X.

North:1993:GSC

[Nor93] Ken North. A generic SQL class library. Dr. Dobb’s Journal of Software Tools, 18(10):69–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

North:1994:DDV

[Nor94] Ken North. Database development and Visual Basic 3.0. Dr. Dobb’s Journal of Software Tools, 19(3):74, 76–80, March 1994. CODEN DDJOEB. ISSN 1044-789X.

North:1996:PTO

[Nor96] Ken North. Performance testing ODBC and native SQL APIs. Dr. Dobb’s Journal of Software Tools, 21(??):17–??, January/February 1996. CODEN DDJOEB. ISSN 1044-789X.

North:1999:EUS

[Nor99] Ken North. Extenders, UDFs, and stored procedures. Dr.

Dobb’s Journal of Software Tools, 24(8):88, 90–93, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
08/sqlexd.txt.

Neal:1999:LSL

[NRG^{+}99] John Neal, Thiadmer Riemersma, Jeff Genender, Torpum Jannak, Richard A. Clarke, William C. Brown, Jeffrey Simmers, and Kemal Gencay. Letters: The Small language; Java 2 graphics; cross-platform independence; version control; median filters; Open Source in Turkey. Dr. Dobb’s Journal of Software Tools, 24(12):10, 12, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Notenboom:1990:BEH

[NV90] Leo Notenboom and Michael Vose. Building an efficient help system (for a hypertext engine). Dr. Dobb’s Journal of Software Tools, 15(6):40–44, 47–48, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Norwood:1994:SMP

[NV94] John Norwood and Shankar Vaidyanathan. Symmetric multiprocessing for PCs. Dr. Dobb’s Journal of Software Tools, 19 (1):80, 82–85, 98–99, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Oblsen:1990:DSA

[Obl90]  Chris Oblsen.  A database

system for automating E-Mail. Dr. Dobb’s Journal of Software Tools, 15(12):72–??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Ochs:1994:PB

[Och94a] Tom Ochs. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20(4):133–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Ochs:1994:RAP

[Och94b] Tom Ochs. Review of Algorithms from P to NP, Volume I: Design and Efficiency, by B. M. E. Moret and H. D. Shapiro. Dr. Dobb’s Journal of Software Tools, 19(4):134–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Ochs:1994:RPC

[Och94c] Tom Ochs. Review of Programming Classics: Implementing the World’s Best Algorithms, by Ian Oliver. Dr. Dobb’s Journal of Software Tools, 19(4):133–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Oestergaard:1997:WIM

[Oes97] Rolf V. Oestergaard. Watchdogs for interrupt monitoring. Dr. Dobb’s Journal of Software Tools, 22(6):60–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Orfali:1995:BSO

[OH95] Robert Orfali and Dan Harkey. Building a SOM OpenDoc part. Dr. Dobb’s Journal of Software Tools, 20(3):131–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Ohlsen:1990:DSA

[Ohl90]  Chris Ohlsen.  A database

system for automating E-mail. Dr. Dobb’s Journal of Software Tools, 15(12):72, 74, 77–80, 104, 106–108, 110, December 1990. CODEN DDJOEB. ISSN 1044-789X.

OMalley:1998:AIA

[OK98] Kevin O’Malley and Terence Kelly. An API for Internet auctions. Dr. Dobb’s Journal of Software Tools, 23(9): 70, 72–74, 96–97, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
09/auction.txt.

Okazaki:1994:WGE

[Oka94] Taku Okazaki. The Windows global EMM import interface. Dr. Dobb’s Journal of Software Tools, 19(10):123–??, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Oldham:1994:DT

[Old94] C. R. Oldham. OS/2 development tools. Dr. Dobb’s Developer Update, 1(5):2–??, May 1994. CODEN ???? ISSN 1079-8595.

Oliver:1991:FRW

[Oli91] Dick Oliver. Fractals in the real world. Dr. Dobb’s Journal of Software Tools, 16(4):68, 71– 72, 74–75, 101–107, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Omohundro:1993:SPL

[Omo93] Stephen M. Omohundro. The Sather programming language. Dr. Dobb’s Journal of Software Tools, 18(11):42–48, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Oney:1992:UDH

[One92] Walter Oney. Using DPMI to hook interrupts in Windows 3. Dr. Dobb’s Journal of Software Tools, 17(2):16–18, 20, 22–23, 78, 80, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Oney:1994:EWS

[One94a] Walter Oney. Examining the Windows Setup Toolkit. Dr. Dobb’s Journal of Software Tools, 19(2):68–72, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Oney:1994:PW

[One94b] Walter Oney. Porting from DOS to Windows. Dr. Dobb’s Journal of Software Tools, 19(3):82, 84– 85, 88–94, 98, 100, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Oney:1994:PB

[One94c] Walter Oney. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(11): 129–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Oney:1994:RWW

[One94d] Walter Oney. Review of Writing Windows Virtual Device Drivers, by David Thielen and Bryan Woodruff. Dr. Dobb’s Journal of Software Tools, 19 (10):129–??, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Orost:1998:BBS

[Oro98] Joseph M. Orost. The Bench++ benchmark suite. Dr. Dobb’s Journal of Software Tools, 23 (10):58, 60–62, 64–65, 112– 113, October 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_10/
benchpp.txt.

Oswald:1993:IU

[Osw93] Rolf Oswald. The Information Utility. Dr. Dobb’s Journal of Software Tools, 18(13):18–20, 24, 26, 28, 30, 104–105, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Oualline:1993:HC

[Oua93] Steve Oualline. Heap checking.

Dr. Dobb’s Journal of Software Tools, 18(12):28, 30, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Ousterhout:1994:TPL

[Ous94] John K. Ousterhout. The Tcl programming language. Dr. Dobb’s Journal of Software Tools, 19(??):40–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Ozick:1991:LLC

[Ozi91] Daniel N. Ozick. A Lisp-style library for C. Dr. Dobb’s Journal of Software Tools, 16(8):36, 38, 40–42, 44, 46, 48, 112–114, 116, 118–126, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Pant:1999:TCP

[Pan99] Lalit Pant. Thread communication in parallel algorithms: Enabling efficient interaction between threads. Dr. Dobb’s Journal of Software Tools, 24 (4):32, 34, 36, 38–39, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        04/parallel.txt.

Papurt:1995:AAI

[Pap95] David M. Papurt. Automating association implementation in C++: Pointer-based association implementation. Dr. Dobb’s Journal of Software Tools, 20 (10):18–20, 22, 23, 24, 108– 110, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Parker:1995:CCE

[Par95a] Tim Parker. C/C++ environments for workstation development. Dr. Dobb’s Developer Update, 2(10):2, 4, October 1995. CODEN ???? ISSN 1079-8595.

Parker:1995:ECC

[Par95b] Tim Parker. Examining C/C++ compilers. Dr. Dobb’s Journal of Software Tools, 20(9):90–91, 93, 95–96, September 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Paterson:1990:ALT

[Pat90a] Tim Paterson. Assembly language tricks of the trade. Dr.

Dobb’s Journal of Software Tools, 15(3):30–36, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Paterson:1990:CDD

[Pat90b] Tim Paterson. Circles and the digital differential analyzer. Dr. Dobb’s Journal of Software Tools, 15(7):30, 32, 34–35, 96, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Patel:1997:MIA

[Pat97]  Mayur Patel.  A memory-

constrained image-processing architecture. Dr. Dobb’s Journal of Software Tools, 22(7):24–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Petit-Bianco:1998:JGC

[PB98a] Alexandre Petit-Bianco. Java garbage collection for realtime systems. Dr. Dobb’s Journal of Software Tools, 23 (10):20–22, 24, 26–29, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
10/kaffe.zip.

Pigeon:1998:AA

[PB98b] Steven Pigeon and Yoshua Bengio. Algorithm alley: Memoryefficient adaptive Huffman coding. Dr. Dobb’s Journal of Software Tools, 23(10):131– 132, 134–135, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
10/Shas/Shas.htm;   http:/
/www.ddj.com/ftp/1998/1998_

                                        10/aa108.txt;   http://www.
                                        ddj.com/ftp/1998/1998_10/
                                        aa108.zip.

Pearson:1991:ABC

[Pea91] Glenn Pearson. Array bounds checking with Turbo C. Dr. Dobb’s Journal of Software Tools, 16(5):72, 74, 78–79, 81– 82, 104–107, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Pearson:1995:BTT

[Pea95] Peter Pearson. Biochemical techniques take on combinatorial problems. Dr. Dobb’s Journal of Software Tools, 20(8): 127–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Pennello:1994:CN

[Pen94] Tom Pennello. C++ namespaces. Dr. Dobb’s Journal of Software Tools, 19(8):44, 46 48– 49, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Perez:1993:TED

[Per93] Christopher Perez. Tools for embedded-systems debugging. Dr. Dobb’s Journal of Software Tools, 18(3):52, 56, 58, 60, 64, 66, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Perez:1999:CDF

[Per99] Cesar A. Gonzalez Perez. The CVS data format. Dr. Dobb’s Journal of Software Tools, 24 (5):50, 52–53, 56–58, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www-gtarpa.usc.es;   http:
                                        //www.ddj.com/9905e/9905e/

9905e.htm;   http://www.ddj.
com/ftp/1999/1999_05/cvs.
txt;   http://www.ddj.com/
ftp/1999/1999_05/cvs.zip.

Pescio:1996:AAM

[Pes96] Carlo Pescio. Algorithm alley: Minimal perfect hashing. Dr. Dobb’s Journal of Software Tools, 21(7):101–102, 104, 116–117, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Paterson:1990:MMDa

[PF90a] Tim Paterson and Steve Flenniken. Managing multiple data segments under Microsoft Windows: Part I. Dr. Dobb’s Journal of Software Tools, 15(2):16– 26, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Paterson:1990:MMDb

[PF90b] Tim Paterson and Steve Flenniken. Managing multiple data segments under Microsoft Windows: Part II. Dr. Dobb’s Journal of Software Tools, 15(3):58– 65, 106–108, March 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Pfefferkorn:1996:PB

[Pfe96] Charles Pfefferkorn. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 21 (1):137–??, January 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Pfleeger:1998:SQ

[Pfl98] Shari Lawrence Pfleeger. Software quality. Dr. Dobb’s Journal of Software Tools, 23(3):22, 24, 28–29, March 1998. CODEN DDJOEB. ISSN 1044-789X.

PerelmanHall:1995:DAG

[PH95a] D. Perelman-Hall. Directed acyclic graph unification. Dr. Dobb’s Journal of Software Tools, 20(4):44, 46, 48, 50–51, 94, 98–99, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Perelman-Hall:1995:DAG

[PH95b]  David Perelman-Hall.  Di-

rected acyclic graph unification. Dr. Dobb’s Journal of Software Tools, 20(4):44–??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Perelman-Hall:1999:JLC

[PH99] David K. Perelman-Hall. Java and lightweight components. Dr. Dobb’s Journal of Software Tools, 24(2):22–24, 26– 28, February 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/lightwei.
                                        txt;   http://www.ddj.com/
                                        ftp/1999/1999_02/lightwei.
                                        zip.

Pickles:1991:SPF

[Pic91] Greg Pickles. Scaling and printing faxes faster. Dr. Dobb’s Journal of Software Tools, 16 (8):94, 97–98, 100–101, 136, 138, 140, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Pickover:1993:RW

[Pic93] Clifford A. Pickover. Recursive worlds. Dr. Dobb’s Journal of Software Tools, 18(9):18–29, 22, 24, 26, September 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Pietrek:1992:IWS

[Pie92a] Matt Pietrek. Inside the Windows scheduler. Dr. Dobb’s Journal of Software Tools, 17 (8):64, 66–68, 70–71, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Pietrek:1992:PD

[Pie92b] Matt Pietrek. Postmortem debugging. Dr. Dobb’s Journal of Software Tools, 17(9):18–20, 22, 24, 26, 31, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Pietrek:1993:EOW

[Pie93a] Matt Pietrek. Examining OPT-LINK for Windows. Dr. Dobb’s Journal of Software Tools, 18 (11):70–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Pietrek:1993:IWM

[Pie93b] Matt Pietrek. Inside the Windows messaging system. Dr. Dobb’s Journal of Software Tools, 18(2):72, 74, 76, 78, 80, 100, 102–103, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Pietrek:1996:DTW

[Pie96] Matt Pietrek. Direct thunking in Windows 95. Dr. Dobb’s Journal of Software Tools, 21 (??):25–??, March/April 1996. CODEN DDJOEB. ISSN 1044-789X.

Pigeon:1999:AAI

[Pig99] Steven Pigeon. Algorithm alley: Image compression with wavelets. Dr. Dobb’s Journal of Software Tools, 24(8): 111–115, August 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL 9908n/9908n.
htm;   http://www.ddj.com/
ftp/1999/1999_08/aa899.txt;
 http://www.ddj.com/ftp/1999/
1999_08/aa899.zip.

Pitt:1998:JQH

[Pit98a] W. David Pitt. Java Q&A: How do I use servlets? Dr. Dobb’s Journal of Software Tools, 23 (12):122, 124–125, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ddj/1998/1998_12/.
./../../ftp/1998/1998_12/
jqa128.txt;   http://www.
ddj.com/ddj/1998/1998_12/.
./../../ftp/1998/1998_12/
jqa128.zip.

Pitt:1998:VPJ

[Pit98b] W. David Pitt. The visitor pattern and a Java grep utility. Dr. Dobb’s Journal of Software Tools, 23(6):30, 32, 93, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Pitt:1999:JQC

[Pit99] W. David Pitt. Java Q&A: Can Java handle exception handling? Dr. Dobb’s Journal of Software Tools, 24(8):103, 106–108, August 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_08/
jqa899.txt;   http://www.
ddj.com/ftp/1999/1999_08/
jqa899.zip.

Prigozhin:1998:LDC

[PJR^{+}98] Alexander Prigozhin, Steve Johnson, Jonathan Rosenne, Sy Wong, Scott Raney, Andreas Eder, and Arthur Nunes. Letters: Design by contract; the standardization process; international Java; to:

<KJPrice@bellatlantic.net>;

the Metacard language; symbolic integration. Dr. Dobb’s Journal of Software Tools, 23 (4):10, 12, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Pamplona:1998:LVC

[PKA^{+}98] August Pamplona, Mike Kurtinitis, Stuart Ambler, Win Carus, Tim McCaffrey, and Peter Sage. Letters: Visual cryptography; EccoFan; help wanted; ternary searches; window sizes and the registry. Dr. Dobb’s Journal of Software Tools, 23(6):10, 12, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Plauger:1992:LCS

[Pla92] P. J. Plauger. Large character sets for C. Dr. Dobb’s Journal of Software Tools, 17(8):16–18, 20, 24, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Plauger:1993:PLG

[Pla93] P. J. Plauger. Programming language guessing games. Dr. Dobb’s Journal of Software Tools, 18(11):16–22, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Plebani:1995:CFA

[Ple95] Louis J. Plebani. Commonfraction approximation of real numbers. Dr. Dobb’s Journal of Software Tools, 20(10):145– ??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Plumb:1994:TRN

[Plu94] Colin Plumb. Truly random numbers. Dr. Dobb’s Journal of Software Tools, 19(13):113– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Pompeii:1996:DOI

[Pom96] John Pompeii. Distributed objects and the Internet. Dr. Dobb’s Journal of Software Tools, 21(3):70, 72, 74–76, 94– 95, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Pomerantz:1998:CNS

[Pom98] Dave Pomerantz. C++ notifiers: Simplifying system development. Dr. Dobb’s Journal of Software Tools, 23(8): 26, 28, 30–31, 89–90, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        08/notifier.txt;   http://
                                        www.ddj.com/ftp/1998/1998_
                                        08/notifier.zip.

Pope:1996:PF

[Pop96] Trevor J. Pope. Password files.

Dr. Dobb’s Journal of Software Tools, 21(1):72, 74, 76, 101, 103– 104, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Potter:1992:DRD

[Pot92] David Potter. Designing a realtime debugger. Dr. Dobb’s Journal of Software Tools, 17(11):70, 72, 74, 76–79, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Pike:1991:DPD

[PPTT91] Rob Pike, Dave Presotto, Ken Thompson, and Howard Trickey. Designing Plan 9 (distributed computing environment). Dr. Dobb’s Journal of Software Tools, 16(1):49–50, 52, 54, 56– 60, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Phillips:1991:CES

[PR91] Stuart G. Phillips and Kevin J.

Rowett. C++ for embedded systems. Dr. Dobb’s Journal of Software Tools, 16(10):76, 78, 80, 83–85, 124, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Prabhakar:1995:IDO

[Pra95]  Ernest N. Prabhakar.  Im-

plementing distributed objects. Dr. Dobb’s Journal of Software Tools, 20(8):80, 82, 84–85, 105– 106, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Preneel:1998:AAP

[PRB98] Bart Preneel, Vincent Rijmen, and Antoon Bosselaers. Algorithm alley: Principles and performance of cryptographic algorithms. Dr. Dobb’s Journal of Software Tools, 23(12): 126–131, December 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/1998/1998_12/shas/
shas.htm.

Price:1994:GA

[Pri94] Kenneth V. Price. Genetic annealing. Dr. Dobb’s Journal of Software Tools, 19(11):127– ??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Prodromou:1992:GIF

[Pro92] Evangelo Prodromou. Graphics import filters for Windows applications. Dr. Dobb’s Journal of Software Tools, 17(7):82, 84, 86–88, 108, 110, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Proffit:1994:MLC

[Pro94]  Brian Proffit.  A multicol-

umn list-box container for OS/2. Dr. Dobb’s Journal of Software Tools, 19(5):90, 92–94, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Proctor:1999:LRT

[Pro99] Frederick M. Proctor. Linux, Real-Time Linux, and IPC. Dr. Dobb’s Journal of Software Tools, 24(11):32, 34, 36, 38–39, November 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_11/
shmem.zip;   http://www.ddj.
com/ftp/1999/1999_11/shmex.
tgz.

Pincus:1997:AAT

[PS97a] Jonathan Pincus and Jerry Schwarz. Algorithm alley: Topological sorting. Dr. Dobb’s Journal of Software Tools, 22 (8):113–116, August 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Price:1997:DES

[PS97b] Kenneth Price and Rainer Storn. Differential evolution: A simple evolution strategy for fast optimization. Dr. Dobb’s Journal of Software Tools, 22 (4):18–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Payne:1998:IAJ

[PSS98a] Jeffery E. Payne, Michael A.

Schatz, and Matthew Schmid. Implementing assertions for Java: Finding bugs early. Dr. Dobb’s Journal of Software Tools, 23(1):40, 42–44, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Plouff:1998:LPQ

[PSS^{+}98b] Wes Plouff, Al Stevens, John Silbert, Ed Guy, Grant Thomsen, Jost Riedel, Ian E. Gorman, and Paul Lu. Letters: PTL queries; finite differences; Y2K challenges; you can take it to the bankers. Dr. Dobb’s Journal of Software Tools, 23 (8):10, 12, August 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Puder:1998:MCC

[Pud98]  Arno Puder.  The MICO

CORBA-compliant system. Dr. Dobb’s Journal of Software Tools, 23(11):44, 46–51, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
11/MICO.TXT;   http://www.
ddj.com/ftp/1998/1998_11/
mico221.zip.

Purdy:1998:JQ

[Pur98] Jason W. Purdy. Java Q&A: How does Java drag-and-drop work? Dr. Dobb’s Journal of Software Tools, 23(11):127– 130, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_11/
JQA118.TXT;   http://www.
ddj.com/ftp/1998/1998_11/
JQA118.ZIP.

Quicksall:1997:SDD

[QG97] Eddy Quicksall and Ken Gibson. Simulation and device-driver development — to ease the process of writing device drivers for new hardware designs, our authors present a simulation environment. to illustrate its use, they use the sample Adaptec AHA-1540 SCSI miniport driver that comes with the Windows NT DDK. Dr. Dobb’s Journal of Software Tools, 22(1):

52–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Quick:1999:LWL

[QTC^{+}99] Ian Quick, Andrew Tucker, Joe Celko, Todd Stephan, Jason Black, and Norman R. Dotti. Letters: And the winner is …; more on Win32 debugging; Quicksort improvements; ah hah; WinCE terminology; real real-time. Dr. Dobb’s Journal of Software Tools, 24(2):10, 12, 16, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm.

Quinn:1997:IMK

[Qui97] Bob Quinn. Internet multicasting: Killer network apps that aren’t network killers. Dr. Dobb’s Journal of Software Tools, 22(10):28, 30–32, 34, 36, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Radtke:1998:WCW

[Rad98] Bruce Radtke. Windows CE Win32 API programming. Dr. Dobb’s Journal of Software Tools, 23(4):62, 64, 66, 68, 70, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1997/1997_
10/udoc/udoc.htm;   http:/
/www.ddj.com/ftp/1998/1998_
04/wince.txt.

Raman:1997:ESE

[Ram97a] T. V. Raman. Emacspeak: A speech-enabling interface: Moving toward auditory user interfaces. Dr. Dobb’s Journal of Software Tools, 22(9):18–20, 22, 23, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Ramm:1997:SIC

[Ram97b] Andy Ramm. Stereoscopic imaging: Creating images in SGI GL and OpenGL. Dr. Dobb’s Journal of Software Tools, 22(9):74, 76, 77, 90, 91, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Ransom:1994:PCC

[Ran94] Jay Frederick Ransom. Pairing C and C++ [letter]. Dr. Dobb’s Journal of Software Tools, 19 (4):10–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Rangaraajan:1999:JQD

[Ran99a] Krishnan Rangaraajan. Java Q&A: Does Java support design by contract? Dr. Dobb’s Journal of Software Tools, 24 (11):113–114, 116, November 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        11/jqa1199.txt.

Rangaraajan:1999:JQH

[Ran99b] Krishnan Rangaraajan. Java Q&A: How can I test Java classes? Dr. Dobb’s Journal of Software Tools, 24(7): 107–109, July 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_07/
                                        jqa799.txt;   http://www.
                                        ddj.com/ftp/1999/1999_07/
                                        jqa799.zip.

Ratcliff:1992:AC

[Rat92] John W. Ratcliff. Audio compression. Dr. Dobb’s Journal of Software Tools, 17(7):32, 37–39, 96, 98–100, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Ratcliff:1993:EPA

[Rat93] John W. Ratcliff. Examining PC audio. Dr. Dobb’s Journal of Software Tools, 18(3):78, 80– 85, 107, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Ratcliff:1994:MAS

[Rat94] John W. Ratcliff. Multimedia audio systems. Dr. Dobb’s Journal of Software Tools, 19(??): 28–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Ratcliff:1995:MSI

[Rat95] John W. Ratcliff. Music and sound for interactive games. Dr. Dobb’s Journal of Software Tools, 20(??):25–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Reichenthal:1997:SSB

[RB97] Steve Reichenthal and Tom Bennett. SBI: The Small Basic Interpreter. Dr. Dobb’s Journal of Software Tools, 22(6):54–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Russinovich:1995:EWL

[RC95] Mark Russinovich and Bryce Cogswell. Examining the Windows 95 Layered File System. Dr. Dobb’s Journal of Software Tools, 20(12):60, 62, 66, 68–70, 108–110, December 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Rissinovich:1996:EVS

[RC96a] M. Rissinovich and B. Cogswell. Examining VxD service hooking. Dr. Dobb’s Journal of Software Tools, 21(5):32, 34, 36–37, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Russinovich:1996:EVS

[RC96b] Mark Russinovich and Bryce Cogswell. Examining VxD service hooking. Dr. Dobb’s Journal of Software Tools, 21(5): 32–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Russinovich:1996:WJP

[RC96c] Mark Russinovich and Bryce Cogswell. Windows 95 journaling and playback. Dr. Dobb’s Journal of Software Tools, 21 (3):46, 48, 50, 93, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Russinovich:1997:EWN

[RC97a] Mark Russinovich and Bryce Cogswell. Examining the Windows NT filesystem — Mark and Bryce open up the inner workings of the NT filesystem by describing how a filesystem request originates in a user’s program and ends up as a disk access. they also present an application called Filemon that monitors and displays all filesystem activity. Dr. Dobb’s Journal of Software Tools, 22(2):42–??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Russinovich:1997:WNS

[RC97b] Mark Russinovich and Bryce Cogswell. Windows NT systemcall hooking — Mark and Bryce present the design and implementation of NTRegmon, a tool that uses hooking to show detailed information about each and every registry access that occurs on a Windows NT system. Dr. Dobb’s Journal of Software Tools, 22(1):42–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Chow:1991:EXW

[RCS91a] James Rhyne, Doris Chow, and Michael Sacks. Enhancing the X-Window system. Dr. Dobb’s Journal of Software Tools, 16 (12):30–??, December 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Rhyne:1991:EXS

[RCS91b] James Rhyne, Doris Chow, and Michael Sacks. Enhancing the X Window System. Dr. Dobb’s Journal of Software Tools, 16(12):30–38, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Russinovich:1996:IS

[RCS96] Mark Russinovich, Bryce Cogswell, and Andrew Schulman. Inside SoftRAM 95. Dr. Dobb’s Journal of Software Tools, 21 (8):52, 54–56, 58, 60–61, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Reagen:1990:RDD

[Rea90] Jeff Reagen. RAM disk driver for Unix. Dr. Dobb’s Journal of Software Tools, 15(10):50, 52, 54–55, 57, October 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Reagen:1991:FFS

[Rea91] Jeff Reagen. FS: A file status utility for UNIX. Dr. Dobb’s Journal of Software Tools, 16 (12):56–62, 96–98, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Recursion:1999:SRW

[Rec99] Earl Augusta Recursion. Simulated recursion: Working around language shortcomings. Dr. Dobb’s Journal of Software Tools, 24(4):40, 42, 44, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        04/simrecur.txt.

Redmond:1995:M

[Red95a] Frank E. Redmond, III. MacHack ’95. Dr. Dobb’s Developer Update, 2(9):2–3, September 1995. CODEN ???? ISSN 1079-8595.

Redmond:1995:VU

[Red95b] Frank E. Redmond, III. A VBX for UDP. Dr. Dobb’s Journal of Software Tools, 20(9):80, 82, 84, 87, 110–111, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Regan:1990:LRS

[Reg90] Shawn M. Regan. LZW revisited (speeding up data compression).

Dr. Dobb’s Journal of Software Tools, 15(6):126–127, 167, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Regelski:1994:BMD

[Reg94] Michael Regelski. Building multimedia databases. Dr. Dobb’s Journal of Software Tools, 19 (13):84, 86–88, 102–103, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Reichenthal:1993:GKT

[Rei93a] Steven Reichenthal. Getting to know TrueType. Dr. Dobb’s Journal of Software Tools, 18 (10):56–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Reilly:1993:ANS

[Rei93b] Douglas Reilly. Accessing Net-Ware SQL files without Net-Ware SQL. Dr. Dobb’s Journal of Software Tools, 18(9):60, 62, 64–66, 92, 94, 96, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Reilly:1993:IBF

[Rei93c] Douglas Reilly. Inside Btrieve files. Dr. Dobb’s Journal of Software Tools, 18(3):68, 70, 72, 77, 106, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Reilly:1993:VAC

[Rei93d]  Douglas Reilly.  A virtual-

array class using C++ templates. Dr. Dobb’s Journal of Software Tools, 18(8):50–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Reilly:1994:CCG

[Rei94a] Douglas Reilly. A C++ class for generating bar codes. Dr. Dobb’s Journal of Software Tools, 19(7):56, 58, 60, 62, 99–100, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Reilly:1994:PB

[Rei94b] Douglas Reilly. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(13):121– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Reimer:1994:DAC

[Rei94c] Jay B. Reimer. DSP and audio compression. Dr. Dobb’s Journal of Software Tools, 19(??): 63–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Remy:1997:DPJ

[Rem97] Martin Remy. Design patterns, Java and Web development. Dr. Dobb’s Journal of Software Tools, 22(6):28, 30, 32, 34, 98–100, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Reznick:1994:CSS

[Rez94a] Assaf Reznick. Character simulation with ScriptX. Dr. Dobb’s Journal of Software Tools, 19 (13):76–??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Reznik:1994:CSS

[Rez94b] Assaf Reznik. Character simulation with ScriptX. Dr. Dobb’s Journal of Software Tools, 19 (13):76, 78, 80, 82, 100–101, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Richman:1991:EHC

[Ric91] Scott Richman. Examining the Hamilton C shell (Unix power for OS/2). Dr. Dobb’s Journal of Software Tools, 16(1): 98, 100, 102, 104–106, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Richter:1998:ESF

[Ric98] H. Thomas Richter. Embedded systems: Fast memory allocation. Dr. Dobb’s Journal of Software Tools, 23(5):78, 80– 82, 84, 86–87, May 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1998/1998_05/fma.
txt;   http://www.ddj.com/
ftp/1998/1998_05/fma.zip.

Ridley:1995:ISP

[Rid95] John Ridley. Identifying serial port IRQs. Dr. Dobb’s Journal of Software Tools, 20(2):94–95, 157, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Riel:1996:NRO

[Rie96] Arthur J. Riel. The newly released Object-Oriented Design Heuristics. Dr. Dobb’s Journal of Software Tools, 21(10): 134–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Riemersma:1999:SSL

[Rie99] Thiadmer Riemersma. The Small scripting language. Dr. Dobb’s Journal of Software Tools, 24(10):68, 70, 72, 74, 76, 78, October 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_10/
small.txt;   http://www.ddj.
com/ftp/1999/1999_10/smallkit.
zip.

Rivest:1995:REA

[Riv95] Ronald L. Rivest. The RC5 encryption algorithm. Dr. Dobb’s Journal of Software Tools, 20 (1):146, 148, January 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Roden:1995:PPR

[RJ95] Thomas A. Roden and Glenn E. Jystad. Plug and play run-time services. Dr. Dobb’s Journal of Software Tools, 20(9):62, 64– 66, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Rosebrugh:1992:MME

[RK92] Christopher Rosebrugh and Eng-Kee Kwang. Multiple microcontrollers in an embedded system: A case study in system architecture and embedded hardware design. Dr. Dobb’s Journal of Software Tools, 17 (1):48, 50, 52, 54, 56–57, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Roach:1991:URC

[Roa91] Kenneth Roach. Using the realtime clock. Dr. Dobb’s Journal of Software Tools, 16(6):26, 28, 30, 32, 88, 90–92, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Roberts:1995:CD

[Rob95a] Dave Roberts. Collision detection. Dr. Dobb’s Journal of Software Tools, 20(??):7–??, May/ June 1995. CODEN DDJOEB. ISSN 1044-789X.

Robinson:1995:ASM

[Rob95b] Mark Robinson. An applicationaccess security model. Dr. Dobb’s Journal of Software Tools, 20(12):82, 84, 86–88, 116, 118, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Robinson:1995:PN

[Rob95c] Mark Robinson. PowerBuilder NVOs. Dr. Dobb’s Journal of Software Tools, 20(10):103– ??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Roberts:1996:DPW

[Rob96] Dale Roberts. Direct port I/O and Windows NT. Dr. Dobb’s Journal of Software Tools, 21 (5):14–16, 18, 20, 22–24, 76–78, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Roberts:1997:EVC

[Rob97] John Roberts. Extending Visual C++. Dr. Dobb’s Journal of Software Tools, 22(5):66–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Roberts:1998:ESI

[Rob98a] Dale Roberts. Embedded systems: Interrupt behavior in Windows NT. Dr. Dobb’s Journal of Software Tools, 23 (4):74, 76, 78, 80, 82, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1998/1998_04/
doirq.txt;   http://www.ddj.
com/ftp/1998/1998_04/doirq.
zip.

Roberts:1998:DTS

[Rob98b] Steven K. Roberts. Designing technomadic systems. Dr. Dobb’s Journal of Software Tools, 23(6):54, 56–58, 60, 62, 64, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Rodriguez:1991:BMP

[Rod91] James Rodriguez. A BRIEF macro package for editing binary files. Dr. Dobb’s Journal of Software Tools, 16(9):82, 84–85, 98– 101, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Roden:1992:HT

[Rod92]  Thomas A. Roden.  High-

resolution timing. Dr. Dobb’s Journal of Software Tools, 17 (9):42, 44, 48, 110, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1994:DGD

[Rod94a] John Rodley. Developing GUIs for database applications. Dr. Dobb’s Journal of Software Tools, 19(6):18–20, 22, 24– 25, 92, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1994:EOE

[Rod94b] John Rodley. Examining OS/2 2.1 executable file formats. Dr. Dobb’s Journal of Software Tools, 19(10):70, 72–75, 97, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1994:OUI

[Rod94c]  John Rodley.  OS/2 and

UnixWare interprocess communication. Dr. Dobb’s Journal of Software Tools, 19(5):78–82, 84, 107–109, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Rodens:1995:ESC

[Rod95a] Ira Rodens. Examining Symantec C++ 7.0. Dr. Dobb’s Journal of Software Tools, 20(8):86– 89, 106–107, August 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Rodens:1995:IOS

[Rod95b] Ira Rodens. Inside the OLE 2 SDK (software development kit). Dr. Dobb’s Journal of Software Tools, 20(5):86–89, 106, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1995:PHA

[Rod95c] John Rodley. Programming HotJava applets. Dr. Dobb’s Journal of Software Tools, 20 (??):7–??, November/December 1995. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1995:TPU

[Rod95d] John Rodley. Thread programming in UnixWare 2.0. Dr. Dobb’s Journal of Software Tools, 20(6):56, 58–61, 102, 104, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Rogers:1995:BOC

[Rog95a] John Rogers. Bit operations with C macros. Dr. Dobb’s Journal of Software Tools, 20 (??):40–??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

Rogers:1995:UMP

[Rog95b] John Rogers. Using the multiple precision library. Dr. Dobb’s Journal of Software Tools, 20 (1):36, 38, 40, 42, 86, 88– 89, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Rogers:1995:YOE

[Rog95c] John Rogers. Your own endian engine. Dr. Dobb’s Journal of Software Tools, 20(11):30, 34– 36, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Rollo:1992:PMW

[Rol92] Carl Rollo. PIE menus for Windows. Dr. Dobb’s Journal of Software Tools, 17(11):30, 32, 34–35, 39, 118–119, 121, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Ross:1993:CCF

[Ros93] John W. Ross. Calling C functions with variably dimensioned arrays. Dr. Dobb’s Journal of Software Tools, 18(8):52, 54, 56, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Ross:1994:PD

[Ros94] Michael Ross. Portability by design. Dr. Dobb’s Journal of Software Tools, 19(3):40, 42–43, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Rosenberg:1995:AAV

[Ros95a]  Robert Rosenberg.  Adding

auxiliary views for Windows apps. Dr. Dobb’s Journal of Software Tools, 20(??):27–??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Ross:1995:OP

[Ros95b] Michael Ross. Optimizing for the PowerPC. Dr. Dobb’s Journal of Software Tools, 20(??):13– ??, September/October 1995. CODEN DDJOEB. ISSN 1044-789X.

Rosenblum:1997:IYP

[Ros97] Bruce D. Rosenblum. Improve your programming with asserts: Ten easy rules for more reliable software. Dr. Dobb’s Journal of Software Tools, 22(12):60, 62– 63, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Rothstein:1990:WDC

[Rot90] Arthur Rothstein. Walking the OS/2 device chain. Dr. Dobb’s Journal of Software Tools, 15 (10):30–??, October 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Roth:1998:PBF

[Rot98]  Peter N. Roth.  Program-

mer’s bookshelf: Fuzzy logic. Dr. Dobb’s Journal of Software Tools, 23(5):141–143, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Roth:1999:PBC

[Rot99]  Peter N. Roth.  Program-

mer’s bookshelf: C++ STL. Dr. Dobb’s Journal of Software Tools, 24(10):127–128, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Rousselle:1998:PTD

[Rou98] Philip Rousselle. Programmer’s toolchest: Dynamic distributed systems in Java. Dr. Dobb’s Journal of Software Tools, 23 (4):88, 90–92, 102–104, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        04/sojournr.txt;   http://
                                        www.ddj.com/ftp/1998/1998_
                                        04/sojournr.zip.

Rowland:1994:CWA

[Row94a] Neil G. Rowland, Jr. Compressing waveform audio files. Dr. Dobb’s Journal of Software Tools, 19(??):20–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Rowland:1994:SSC

[Row94b] Neil G. Rowland, Jr. Speech synthesis in C++. Dr. Dobb’s Journal of Software Tools, 19 (8):84, 86–88, 105–106, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Rowe:1997:BSO

[Row97] Robin Rowe. Building a smart online video application: A software VCR using off-the-shelf parts. Dr. Dobb’s Journal of Software Tools, 22(12):78, 80– 84, 103–105, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Ramanujapuram:1998:DCI

[RR98] Arun Ramanujapuram and Prasad Ram. Digital content and intellectual property rights. Dr. Dobb’s Journal of Software Tools, 23(12):20– 22, 24, 26–27, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
12/../../../ftp/1998/1998_
12/dprl.txt;   http://www.
ddj.com/ddj/1998/1998_12/
schn/schn.htm.

Roden:1997:PPO

[RT97] Tom Roden and Scott Townsend.

Plug-and-play OPROMs and the BIOS boot specification. Dr. Dobb’s Journal of Software Tools, 22(5):38–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Rucker:1996:CC

[Ruc96] Ken Rucker. ColoradOS/2 conference. Dr. Dobb’s Developer Update, 3(1):2, 3, January 1996. CODEN ???? ISSN 1079-8595.

Russell:1993:DCW

[Rus93] John Russell. Digital chart of the world [letter]. Dr. Dobb’s Journal of Software Tools, 18 (3):10–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Rush:1996:COD

[Rus96] Jeff Rush. Component objects and distributed computing. Dr. Dobb’s Journal of Software Tools, 21(10):38–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Russell:1999:BRT

[Rus99] Regan Russell. The BeRays ray tracer. Dr. Dobb’s Journal of Software Tools, 24(11): 62, 64–65, November 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_11/
                                        berays.txt;   http://www.
                                        ddj.com/ftp/1999/1999_11/
                                        rays-x86.zip.

Roberts:1996:UO

[RZ96] Jonathan Roberts and Dan Zigmond. Using OODCE. Dr. Dobb’s Journal of Software Tools, 21(3):98–101, 130–132, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Sadler:1999:FEE

[Sad99] John Sadler. Ficl: An embeddable extension language interpreter. Dr. Dobb’s Journal of Software Tools, 24(1): 70, 72–74, 76, 78, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/ftp/1999/1999_
                                        01/ficl.txt;   http://www.
                                        ddj.com/ddj/ftp/1999/1999_
                                        01/ficl202.zip.

Saelensminde:1995:FCC

[Sae95a]  Kirit Saelensminde.  File-

streaming classes in C++. Dr. Dobb’s Journal of Software Tools, 20(10):58, 60, 118, 120, 122, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Saelensminde:1995:PMM

[Sae95b] Kirit Saelensminde. A pooling memory manager for C++. Dr. Dobb’s Journal of Software Tools, 20(8):34, 36, 38, 40, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Saks:1992:SCS

[Sak92] Dan Saks. Standard C++:

A status report. Dr. Dobb’s Journal of Software Tools, 17 (12 (special supplement)):15s– 20s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Salter:1992:EHW

[Sal92] Brett Salter. An exception handler for Windows 3. Dr. Dobb’s Journal of Software Tools, 17 (9):32, 34, 36, 102, 104–107, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Saltsman:1996:PB

[Sal96] Tom Saltsman. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 21(6):118–??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Sardis:1992:DDB

[Sar92] Robert Sardis. Dynamic dialog boxes and C++. Dr. Dobb’s Journal of Software Tools, 17(11):40, 42, 122, 124– 125, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Sarafin:1994:WT

[Sar94]  Gregory C. Sarafin.  The

WRAPI toolkit. Dr. Dobb’s Journal of Software Tools, 19 (3):26, 28, 30, 32, 34, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Sardis:1995:VBR

[Sar95] Robert Sardis. Visual Basic by remote control. Dr. Dobb’s Journal of Software Tools, 20 (9):48, 50, 52, 104, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Sars:1997:STL

[Sar97] Camillo Sars. The SSH Transport Layer Protocol: Making the Internet secure. Dr. Dobb’s Journal of Software Tools, 22 (10):38, 40, 42–43, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Satchell:1990:P

[Sat90] Stephen Satchell. 68040 programming. Dr. Dobb’s Journal of Software Tools, 15(3):38–44, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Savchenko:1998:AAT

[Sav98] Sergei Savchenko. Algorithm alley: Theorem proving and database querying. Dr. Dobb’s Journal of Software Tools, 23 (8):111–113, 117, August 1998. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.

ddj.com/ftp/1998/1998_08/
aa898.txt;   http://www.ddj.
com/ftp/1998/1998_08/aa898.
zip.

Sawert:1994:ASP

[Saw94] Brian Sawert. The advanced SCSI programming interface. Dr. Dobb’s Journal of Software Tools, 19(3):154, 158–160, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Sax:1992:WCA

[Sax92a] Mike Sax. The Windows communications API: porting your programs from DOS to Windows. Dr. Dobb’s Journal of Software Tools, 17(5):40, 42, 44, 86, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Sax:1992:WT

[Sax92b] Mike Sax. Windows ToolHelp.

Dr. Dobb’s Journal of Software Tools, 17(9):90, 92, 118– 119, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Schroeder:1997:AAP

[SC97] Will Schroeder and Tom Citriniti. Algorithm alley — polygon decimation algorithms reduce the number of polygons in a mesh while maintaining a good approximation to the original data, leading to faster, more realistic 3-D graphics. Dr. Dobb’s Journal of Software Tools, 22 (7):109–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Schatzman:1990:BPO

[Sch90a] Bruce D. Schatzman. Based pointers for optimization. Dr. Dobb’s Journal of Software Tools, 15(8):85–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Schatzman:1990:UOM

[Sch90b] Bruce D. Schatzman. Unraveling optimization in Microsoft C 6.0. Dr. Dobb’s Journal of Software Tools, 15(10):62–??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Schimandle:1990:ECM

[Sch90c]  James Schimandle.  Encap-

sulating C memory allocation. Dr. Dobb’s Journal of Software Tools, 15(8):24–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1990:CSD

[Sch90d] Andrew Schulman. Compiler supported DLLs for DOS. Dr. Dobb’s Journal of Software Tools, 15(5):36–??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1990:EI

[Sch90e] Andrew Schulman. Examining Instant-C. Dr. Dobb’s Journal of Software Tools, 15(6):62– 74, 120, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1990:IWC

[Sch90f]  Andrew Schulman.  Inside

Watcom C 7.0/386 (compiler). Dr. Dobb’s Journal of Software Tools, 15(3):74–83, 115, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1990:OOB

[Sch90g] Andrew Schulman. Opening OS/2’s backdoor. Dr. Dobb’s Journal of Software Tools, 15 (10):28, 30, 32, 34, 36, 38, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1990:SGFa

[Sch90h] Andrew Schulman. Stalking general protection faults: Part I. Dr. Dobb’s Journal of Software Tools, 15(1):74–75, 77–81, 120, 122, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1990:SGFb

[Sch90i] Andrew Schulman. Stalking general protection faults: Part II. Dr. Dobb’s Journal of Software Tools, 15(2):76–86, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Schatzman:1991:EMM

[Sch91a] Bruce D. Schatzman. Examining the Microsoft Mail SDK. Dr. Dobb’s Journal of Software Tools, 16(4):60, 62, 64, 66– 67, 100, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Schmit:1991:MCF

[Sch91b] Mike Schmit. MASM’s changing face. Dr. Dobb’s Journal of Software Tools, 16(7):50, 52, 54–56, 58, 60, 96, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1991:OHF

[Sch91c] Bruce Schneier. One-way hash functions. Dr. Dobb’s Journal of Software Tools, 16(9):148– 151, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Schneider:1992:WPB

[Sch92a] Raymond J. Schneider. Windows programming with Basic. Dr. Dobb’s Journal of Software Tools, 17(4):80, 82, 84–85, 112, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1992:PSG

[Sch92b] Bruce Schneier. Pseudo-random sequence generator for 32-bit CPUs. Dr. Dobb’s Journal of Software Tools, 17(2):34, 37–38, 40, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1992:RT

[Sch92c] Bruce Schneier. Red-black trees. Dr. Dobb’s Journal of Software Tools, 17(4):42, 44–46, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1992:UPC

[Sch92d] Bruce Schneier. Untangling public-key cryptography: the key to secure communications. Dr. Dobb’s Journal of Software Tools, 17(5):16, 17, 20, 22, 24, 26, 28, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1992:PBa

[Sch92e] Andrew Schulman. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 17 (10):157–??, October 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Schulman:1992:PBb

[Sch92f] Andrew Schulman. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 17 (12):147–??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Schneider:1993:DCNb

[Sch93a] Lowell S. Schneider. Distributed computing now: Development environments. Dr. Dobb’s Journal of Software Tools, 18(7):64, 66–68, 71, 100, July 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Schneier:1993:IEA

[Sch93b] Bruce Schneier. The IDEA encryption algorithm. Dr. Dobb’s Journal of Software Tools, 18 (13):50, 52, 54, 56, 106, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:EWA

[Sch93c] Andrew Schulman. Examining the Windows AARD detection code. Dr. Dobb’s Journal of Software Tools, 18(9):42, 44–48, 89, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:IDWb

[Sch93d] Andrew Schulman. Introduction to ‘Documenting Documentation: The Windows .HLP File Format, Part II’ (P. Davis). Dr. Dobb’s Journal of Software Tools, 18(10):143–??, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCb

[Sch93e] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 18(4): 119–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCc

[Sch93f] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 18(5): 125–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCd

[Sch93g] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 18(6): 153–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCe

[Sch93h] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 18(7): 127–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCh

[Sch93i] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 18(11): 143–??, November 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCj

[Sch93j] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 18(13):

123–??, December 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Schulman:1993:IDWa

[Sch93k] Andrew Schulman. Undocumented corner: Introduction to ‘Documenting Documentation: The Windows .HLP File Format, Part I’ (P. Davis). Dr. Dobb’s Journal of Software Tools, 18(9):119–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCf

[Sch93l] Andrew Schulman. Undocumented corner: Introduction to ‘The Windows .RES File Format’ (A. G. Fedorov and D. M. Rogatkin). Dr. Dobb’s Journal of Software Tools, 18(8): 133–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Schmit:1994:OPC

[Sch94a] Mike Schmit. Optimizing Pentium code. Dr. Dobb’s Journal of Software Tools, 19(1):40, 42, 44, 46–49, January 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Schneier:1994:AAa

[Sch94b] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(8):119–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:AAb

[Sch94c] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(11):119–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:AAd

[Sch94d] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(13):113–??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:AAc

[Sch94e] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(13):127–??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:BEA

[Sch94f] Bruce Schneier. The Blowfish encryption algorithm. Dr. Dobb’s Journal of Software Tools, 19(4):38, 40, 98, 99, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:CAW

[Sch94g] Bruce Schneier. The Cambridge algorithms workshop. Dr. Dobb’s Journal of Software Tools, 19(4):18–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:MS

[Sch94h] Bruce Schneier. E-mail security. Dr. Dobb’s Journal of Software Tools, 19(??):18–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:E

[Sch94i] Bruce Schneier. Eurocrypt ’94. Dr. Dobb’s Developer Update, 1 (7):3–??, July 1994. CODEN ???? ISSN 1079-8595.

Schneier:1994:NC

[Sch94j] Bruce Schneier. NP-completeness.

Dr. Dobb’s Journal of Software Tools, 19(10):119–121, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1994:RDS

[Sch94k] Bruce Schneier. RSA data security conference. Dr. Dobb’s Developer Update, 1(4):3–??, April 1994. CODEN ???? ISSN 1079-8595.

Schneier:1994:SL

[Sch94l] Bruce Schneier. Skip lists.

Dr. Dobb’s Journal of Software Tools, 19(1):50, 52, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:IGA

[Sch94m] Andrew Schulman. Introduction to ‘think globally act locally: Inside the Windows instance data manager’ (K. Müller). Dr. Dobb’s Journal of Software Tools, 19(4):125–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:MGS

[Sch94n] Andrew Schulman. Microsoft’s grip on software tightened by antitrust deal. Dr. Dobb’s Journal of Software Tools, 19(11): 143–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCc

[Sch94o] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 19(3):

133–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCd

[Sch94p] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 19(5): 137–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCf

[Sch94q] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 19(7): 123–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCg

[Sch94r] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 19(8): 125–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCh

[Sch94s] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 19(11): 123–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCj

[Sch94t] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 19(13): 143–??, November 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCk

[Sch94u] Andrew Schulman. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 20(4): 125–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCI

[Sch94v] Andrew Schulman. Undocumented corner: Introduction to ‘The Windows 3.1 Virtual Machine Control Block Part 1’ (K. Zytaruk). Dr. Dobb’s Journal of Software Tools, 19(1): 115–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:IWV

[Sch94w] Andrew Schulman. Undocumented corner: Introduction to ‘The Windows 3.1 Virtual Machine Control Block Part 2’ (K. Zytaruk). Dr. Dobb’s Journal of Software Tools, 19(2):107– ??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Schuermann:1994:PAV

[Sch94x] Joachim Schürmann. A program architecture for Visual Basic development. Dr. Dobb’s Journal of Software Tools, 19(??):32–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Scherer:1995:SCG

[Sch95a] Perry W. Scherer. Simplifying C++ GUI development. Dr. Dobb’s Journal of Software Tools, 20(9):40, 42–44, 46, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneider:1995:EPD

[Sch95b] Raymond J. Schneider. Examining the PowerBASIC developer kit. Dr. Dobb’s Journal of Software Tools, 20(3):72, 75–79, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAa

[Sch95c] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 20(3):125–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAc

[Sch95d] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 20(5):119–??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAd

[Sch95e] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 20(6):121–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAe

[Sch95f] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 20(7):121–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAf

[Sch95g] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 20(8):127–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAh

[Sch95h] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 20(10):145–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAi

[Sch95i] Bruce Schneier. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 20(11):133–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAj

[Sch95j] Bruce Schneier. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 20(12):131–??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAB

[Sch95k]  Bruce Schneier.  Algorithm

alley: The Blowfish encryption algorithm: One year later. Dr. Dobb’s Journal of Software Tools, 20(9):137–??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:AAG

[Sch95l] Bruce Schneier. Algorithm alley: The GOST encryption algorithm. Dr. Dobb’s Journal of Software Tools, 20(1):123– ??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1995:CCS

[Sch95m] Bruce Schneier. Computers and communications security. Dr. Dobb’s Developer Update, 2(1): 3–??, January 1995. CODEN ???? ISSN 1079-8595.

Schneier:1995:E

[Sch95n] Bruce Schneier. Eurocrypt ’95.

Dr. Dobb’s Developer Update, 2 (8):2, August 1995. CODEN ???? ISSN 1079-8595.

Schneier:1995:PCC

[Sch95o] Bruce Schneier. A pair of cryptographic conferences. Dr. Dobb’s Developer Update, 2(3): 3, March 1995. CODEN ???? ISSN 1079-8595.

Schulman:1995:UCW

[Sch95p] Andrew Schulman. Undocumented corner: Windows 90+5. Dr. Dobb’s Journal of Software Tools, 20(1):127–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Schauls:1996:WUG

[Sch96a] Troy A. Schauls. Writing userdefinable GUIS. Dr. Dobb’s Journal of Software Tools, 21 (9):36, 38–39, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Scherer:1996:DRD

[Sch96b] Perry Scherer. The DBperl relational database API. Dr. Dobb’s Journal of Software Tools, 21 (11):34–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Scherer:1996:ERT

[Sch96c]  Perry W. Scherer.  Exam-

ining RogueWave’s Tools.h++. Dr. Dobb’s Journal of Software Tools, 21(2):80, 82, 84, 111– 112, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1996:AAa

[Sch96d] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 21(1):123–??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1996:AAb

[Sch96e] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 21(2):135–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1996:AAc

[Sch96f] Bruce Schneier. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 21(3):117–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1996:DLC

[Sch96g] Bruce Schneier. Differential and linear cryptanalysis. Dr. Dobb’s Journal of Software Tools, 21 (1):42, 44, 46, 48, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Schreiber:1996:UWN

[Sch96h] Sven B. Schreiber. Undocumented Windows NT and the NetWare Core Protocol. Dr. Dobb’s Journal of Software Tools, 21(11):70–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Schryver:1996:EWT

[Sch96i] Dot Schryver. Examining the Winbatch Toolkit. Dr. Dobb’s Journal of Software Tools, 21 (12):74–??, December 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Schreiber:1997:IWC

[Sch97a] Sven B. Schreiber. Inside Windows Cabinet files. Dr. Dobb’s Journal of Software Tools, 22 (5):20–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Schreiber:1997:WNC

[Sch97b] Sven B. Schreiber. Win32 Netware Core Protocol, revisited. Dr. Dobb’s Journal of Software Tools, 22(10):60, 62–64, 103– 104, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Schettino:1998:EWC

[Sch98a] John C. Schettino, Jr. Extending Windows CE 2.0 MFC database classes. Dr. Dobb’s Journal of Software Tools, 23 (5):70, 72, 74–75, 110–111, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
05/cequery.txt;   http://
www.ddj.com/ftp/1998/1998_
05/cequery.zip.

Schneier:1998:TEA

[Sch98b] Bruce Schneier. The Twofish encryption algorithm. Dr. Dobb’s Journal of Software Tools, 23 (12):30, 32, 34, 36, 38, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
12/../../../ftp/1998/1998_
12/twofish.zip.

Schreiber:1998:EML

[Sch98c]  Sven B. Schreiber.  Exam-

ining Microsoft’s LDAP API. Dr. Dobb’s Journal of Software Tools, 23(12):108, 110– 111, December 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ddj/1998/1998_12/.
./../../ftp/1998/1998_12/
ntldap.txt;   http://www.
ddj.com/ddj/1998/1998_12/.
./../../ftp/1998/1998_12/
ntldap.zip.

Schaad:1999:PFD

[Sch99a] Rene Schaad. Parallel functional decision trees: Coping with unpredictability. Dr. Dobb’s Journal of Software Tools, 24(4):62, 64–68, 70, April 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9904h/9904h.
htm;   http://www.ddj.com/
ftp/1999/1999_04/insitu05.
zip.

Schaps:1999:PTC

[Sch99b] Gary L. Schaps. Programmer’s toolchest: Compiler construction with Antlr and Java. Dr. Dobb’s Journal of Software Tools, 24(3):84, 86–89, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/antlr.txt;
http://www.ddj.com/ftp/1999/
1999_03/antlr.zip.

Schneier:1999:AT

[Sch99c] Bruce Schneier. Attack trees.

Dr. Dobb’s Journal of Software Tools, 24(12):21–22, 24, 26, 28– 29, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Schreiber:1999:IWN

[Sch99d] Sven B. Schreiber. Inside Windows NT system data. Dr. Dobb’s Journal of Software Tools, 24(11):40, 42, 44–46, 48–49, November 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_11/
ntinfo.txt;   http://www.
ddj.com/ftp/1999/1999_11/
ntinfo.zip.

Schwaderer:1999:DCN

[Sch99e]  Curtis Schwaderer.  Devel-

oping custom network protocols. Dr. Dobb’s Journal of Software Tools, 24(9):62, 64, 66, 68, 70, 72, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/netprot.txt;   http://
www.ddj.com/ftp/1999/1999_
09/netprot.zip.

Silver:1999:LPA

[SDB^{+}99] Stephen Silver, DDJ, Tim Bond, Mike Pidd, Dugald A. Taylor, Chris Clement, Richard Shirey, George Vol, and Andrew Tucker. Letters: Programmer aptitude testing; cross-platform design; A hard day’s night electronic music; so what’s wrong with Cobol?; porting to CE. Dr. Dobb’s Journal of Software Tools, 24(11):10, 12, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Searfoss:1990:BBD

[Sea90] Glenn Searfoss. Bounding box data compression. Dr. Dobb’s Journal of Software Tools, 15 (4):56–64, 108, April 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Selvakumar:1999:ATW

[Sel99] M. Selvakumar. Automated testing for Web applications. Dr. Dobb’s Journal of Software Tools, 24(5):88, 90, 92, 95–96, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/webtest.txt.

Seminatore:1995:REC

[Sem95] Mark Seminatore. A ray-casting engine in C++. Dr. Dobb’s Journal of Software Tools, 20 (7):28, 30, 32, 34, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Seminatore:1996:YOT

[Sem96] Mark Seminatore. Your own two-dimensional gaming engine. Dr. Dobb’s Journal of Software Tools, 21(??):27–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Serrat:1993:NCU

[Ser93] Eduardo M. Serrat. A NetWare chat utility: understanding IPX programming. Dr. Dobb’s Journal of Software Tools, 18(12):60, 64, 67–69, 100, 102, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Seshadri:1998:JQH

[Ses98] Govind Seshadri. Java Q&A: How do I implement callbacks with Java’s RMI? Dr. Dobb’s Journal of Software Tools, 23 (3):123–124, 137–138, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        03/jqa398.txt.

Sain:1997:KRT

[SG97] Biswajit Sain and Timothy A. Gonsalves. Kertesy: A real time event driven microkernel. Dr. Dobb’s Journal of Software Tools, 22(3):48–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Shah:1993:WDR

[Sha93a] Kamal Shah. Windows DDE for real-time applications. Dr. Dobb’s Journal of Software Tools, 18(1):58, 60–61, 64, 94, 96–98, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Sharp:1993:DLU

[Sha93b] Oliver Sharp. Dynamic linking under Berkeley UNIX. Dr. Dobb’s Journal of Software Tools, 18(5):40, 42, 44, 86, 88, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Sharp:1993:NP

[Sha93c] Oliver Sharp. Networking with Perl. Dr. Dobb’s Journal of Software Tools, 18(11):94–102, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1993:DTP

[Sha93d] Dennis Shasha. Database tuning: principles and surprises. Dr. Dobb’s Journal of Software Tools, 18(4 (supplement)): 32S–4, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Sharp:1994:PPL

[Sha94] Oliver Sharp. The Perl programming language. Dr. Dobb’s Journal of Software Tools, 19 (??):12–??, October 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Shankel:1997:IAF

[Sha97] Jason Shankel. Implementing Abstract Factory as an STL container: A practical, reusable structure. Dr. Dobb’s Journal of Software Tools, 22(12):28, 32, 34, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Shankel:1998:SRB

[Sha98a] Jason Shankel. STL’s redblack trees. Dr. Dobb’s Journal of Software Tools, 23(4): 54, 56, 58, 60, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1992/1992_
04/schn/schn.htm;   http:/
/www.ddj.com/ftp/1998/1998_
04/stlrb.txt.

Shasha:1998:DEOg

[Sha98b] Dennis E. Shasha. Dr. Ecco’s omniheurist corner. Dr. Dobb’s Journal of Software Tools, 23 (11):140–141, November 1998. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1998:DEOc

[Sha98c] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Beautiful liars. Dr. Dobb’s Journal of Software Tools, 23(6):121–122, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1998:DEOe

[Sha98d] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Getting even. Dr. Dobb’s Journal of Software Tools, 23(9):123– 124, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1998:DEOf

[Sha98e] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Mapcraft. Dr. Dobb’s Journal of Software Tools, 23(10):137–139, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ddj/1998/1998_
                                        10/index.htm.

Shasha:1998:DEOd

[Sha98f] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Mates. Dr. Dobb’s Journal of Software Tools, 23(8):115–116, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ddj/1998/1998_
                                        08/shas/shas.htm.

Shasha:1998:DEOb

[Sha98g] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Nimmerics. Dr. Dobb’s Journal of Software Tools, 23(5):134–135, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1998:DEOh

[Sha98h] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Subway. Dr. Dobb’s Journal of Software Tools, 23(12):132–133, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1998:DEOa

[Sha98i] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Territory game. Dr. Dobb’s Journal of Software Tools, 23(4):134, 136, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Shankel:1999:LLL

[Sha99a] Jason Shankel. Little languages with Lex, Yacc, and MFC. Dr. Dobb’s Journal of Software Tools, 24(1):28, 30, 32–33, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/ftp/1999/
1999_01/slide.txt;   http:
//www.ddj.com/ddj/ftp/1999/
1999_01/slide1.zip.

Shasha:1999:DEOd

[Sha99b] Dennis E. Shasha. Dr. Ecco’s omniheurist corner. Dr. Dobb’s Journal of Software Tools, 24 (4):114–115, April 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOj

[Sha99c] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Calabaza. Dr. Dobb’s Journal of Software Tools, 24(11):123, 125, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOb

[Sha99d] Dennis E. Shasha. Dr. ecco’s omniheurist corner: Dig. Dr. Dobb’s Journal of Software Tools, 24(2):131–132, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ddj/1999/1999_02/9902o/9902o.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/ecco118.
                                        zip.

Shasha:1999:DEOa

[Sha99e]  Dennis E. Shasha.  Dr.

Ecco’s omniheurist corner: Fair Swedes. Dr. Dobb’s Journal of Software Tools, 24(1):123– 125, January 1999. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOf

[Sha99f] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Flats and steeps. Dr. Dobb’s Journal of Software Tools, 24(7):123–125, July 1999. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOi

[Sha99g] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Inheritance. Dr. Dobb’s Journal of Software Tools, 24(10):123– 124, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOg

[Sha99h] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Laser shuttles. Dr. Dobb’s Journal of Software Tools, 24(8):117, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ercb.com/ddj/1999/ddj.
9908.html.

Shasha:1999:DEOh

[Sha99i] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Monopoles. Dr. Dobb’s Journal of Software Tools, 24(9):115–116, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOk

[Sha99j] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: My Enemy’s enemy. Dr. Dobb’s Journal of Software Tools, 24(12): 123–125, December 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOe

[Sha99k] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Rosetta. Dr. Dobb’s Journal of Software Tools, 24(5):130–131, May 1999. CODEN DDJOEB. ISSN 1044-789X.

Shasha:1999:DEOc

[Sha99l] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Trains for the sultan. Dr. Dobb’s Journal of Software Tools, 24(3): 121–122, March 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm;
 http://www.ddj.com/1999/
9903/9903toc.htm;   http:/
/www.ddj.com/ddj/1999/1999_
03/9903n/9903n.htm.

Sheils:1993:ULM

[She93] Michael Sheils. The undocumented LAN manager and named pipe APIs for DOS and Windows. Dr. Dobb’s Journal of Software Tools, 18(4):119–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Shecter:1999:PTD

[She99] Robb Shecter. Programmer’s toolchest: Design by interface. Dr. Dobb’s Journal of Software Tools, 24(2):96, 98– 101, February 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/design.txt.

Shiffman:1992:TLO

[Shi92] Harris Shiffman. Toward a less object-oriented view of C++. Dr. Dobb’s Journal of Software Tools, 17(12 (special supplement)):35s–38s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Shirah:1994:AWG

[Shi94]  Joe Sam Shirah.  Anima-

tion with the Windows GDI. Dr. Dobb’s Journal of Software Tools, 19(??):48–??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Shinsato:1999:WHP

[Shi99] Harold Shinsato. Writing highperformance graphical Java components. Dr. Dobb’s Journal of Software Tools, 24(9):

50, 52–54, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/javacomp.txt.

Shmidt:1994:RVF

[Shm94] Alex Shmidt. RINGO: VxDs on the fly. Dr. Dobb’s Journal of Software Tools, 19(3):133–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Shoffner:1996:JWO

[Sho96] Michael Shoffner. Java and Web-executable object security. Dr. Dobb’s Journal of Software Tools, 21(11):38, 40, 42, 44, 46, 48, 49, November 1996.

Salo:1999:OPB

[SHR^{+}99] Timo Salo, Justin Hill, Scott Rich, Chuck Bridgham, and Daniel Berg. Object persistence: Beyond serialization. Dr. Dobb’s Journal of Software Tools, 24 (5):19, 22, 24–27, 30–33, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/persist.txt.

Shusman:1996:PM

[Shu96] Dan Shusman. Programming with M. Dr. Dobb’s Journal of Software Tools, 21(??):35–??, January/February 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Siering:1991:OLM

[Sie91] Thomas Siering. OBJ library management. Dr. Dobb’s Journal of Software Tools, 16(9):

58, 60–62, 90–92, 94, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Simone:1995:CA

[Sim95] Cal Simone. CGI and AppleScript. Dr. Dobb’s Journal of Software Tools, 20(??):43–??, November/December 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Sinofsky:1992:DCC

[Sin92]  Steven Sinofsky.  Designing

C++ classes. Dr. Dobb’s Journal of Software Tools, 17(11):52, 54, 56, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Sipe:1996:CDW

[Sip96a] Steve Sipe. Calling 16-bit DLLs from Windows 95. Dr. Dobb’s Journal of Software Tools, 21 (4):??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Sipe:1996:CBD

[Sip96b] Steve Sipe. Calling 16-bit DLLs from Windows 95. Dr. Dobb’s Journal of Software Tools, 21 (4):42, 44, 46, 48–50, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Sipe:1996:CSU

[Sip96c] Steve Sipe. Creating shaped UI objects. Dr. Dobb’s Journal of Software Tools, 21(9):30, 32– 34, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Sippl:1996:SAG

[Sip96d] Roger Sippl. SQL access group’s call-level interface. Dr. Dobb’s Journal of Software Tools, 21 (??):11–??, January/February 1996. CODEN DDJOEB. ISSN 1044-789X.

Sipe:1997:RWN

[Sip97] Steve Sipe. RPC for Windows 95/NT. Dr. Dobb’s Journal of Software Tools, 22(3):42–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Sitarski:1996:AA

[Sit96] Edward Sitarski. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 21(9):107– ??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Singh:1992:DRS

[SJB92] Gurjot Singh, Moses Joseph, and Dave Barnett. Debugging real-time systems. Dr. Dobb’s Journal of Software Tools, 17 (9):70, 72, 74, 76–77, 116– 117, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Subramaniam:1993:PPP

[SKD93] Ramesh Subramaniam, Kiran Kundargi, and J. Dahms. Programming the Pentium processor. Dr. Dobb’s Journal of Software Tools, 18(6):34, 36, 38, 40–42, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Skier:1991:ALM

[Ski91] Ken Skier. Assembly language macros. Dr. Dobb’s Journal of Software Tools, 16(3):28, 30, 32, 34, 74–75, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Schroeder:1996:SC

[SL96a] Will Schroeder and Bill Lorensen. 3-D surface contours. Dr. Dobb’s Journal of Software Tools, 21 (7):??, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Schroeder:1996:DSC

[SL96b] Will Schroeder and Bill Lorensen. 3-D surface contours. Dr. Dobb’s Journal of Software Tools, 21 (7):26, 28, 30, 32, 85–86, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Schneider:1993:DCNa

[SM93] Lowell S. Schneider and Stephen S. Murray. Distributed computing now: Application environments. Dr. Dobb’s Journal of Software Tools, 18(6):82, 84–85, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Smallson:1994:NIP

[Sma94] Fran Smallson. NAFTA’s intellectual property provisions. Dr. Dobb’s Developer Update, 1(13): 2–??, November 1994. CODEN ???? ISSN 1079-8595.

Smedley:1998:VPP

[Sme98] Garth Smedley. Visual programming with Prograph. Dr. Dobb’s Journal of Software Tools, 23(9):76, 78–83, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        09/pgv12b2.exe.

Smith:1993:PCO

[Smi93a] Harry J. Smith. Putting colors in order. Dr. Dobb’s Journal of Software Tools, 18(7):40–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Smith:1993:LPE

[Smi93b] Peter Smith. LUC public-key encryption: A secure alternative to RSA. Dr. Dobb’s Journal of Software Tools, 18(1):44, 46, 48–49, 90–92, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Smith:1994:CE

[Smi94] Peter Smith. Cryptography without exponentiation. Dr. Dobb’s Journal of Software Tools, 19(4):26, 28, 30, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Smith:1997:RTO

[Smi97] Kevin W. Smith. Real-time object-oriented modeling: ObjecTime CASE tool simplifies real-time software development. Dr. Dobb’s Journal of Software Tools, 22(12):64, 66, 68, 70, 72, 74, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Smith:1998:RTW

[Smi98] Richard M. Smith. A real-time weather station. Dr. Dobb’s Journal of Software Tools, 23 (10):40, 42, 44, 46, 109–110, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
10/weather.txt.

Smyth:1997:GGN

[Smy97] Gavin Smyth. GNAT: The GNU New York University Ada Translator: A high-quality, lowcost Ada 95 compiler. Dr. Dobb’s Journal of Software Tools, 22(12):86, 88, 89–90, 105– 107, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Singh:1997:ARL

[SNC97] Satinder Singh, Peter Norvig, and David Cohn. Agents and reinforcement learning. Dr. Dobb’s Journal of Software Tools, 22(3):28–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1993:UCI

[SNH93] Andrew Schulman, Joseph M. Newcomer, and Bruce Horn. Undocumented corner: Inside Windows regions in UC. Dr. Dobb’s Journal of Software Tools, 18(3):129, 130, 132, 134, 136, 137, 150, March 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Soja:1996:IMT

[Soj96] Richard Soja. Inside Motorola’s TPU. Dr. Dobb’s Journal of Software Tools, 21(12):50– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Sood:1998:JQW

[Soo98a]  Mukul Sood.  Java Q&A:

What is Swing? Dr. Dobb’s Journal of Software Tools, 23 (9):111–114, 125, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1998/1998_09/
jqa998.txt;   http://www.

ddj.com/ftp/1998/1998_09/
jqa998.zip.

Sood:1998:PTE

[Soo98b] Mukul Sood. Programmer’s toolchest: Examining JDBC drivers: Current drivers fit into our of four categories. Dr. Dobb’s Journal of Software Tools, 23(1):82, 84–87, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Soukup:1999:DSO

[Sou99] Jiri Soukup. Data structures as objects. Dr. Dobb’s Journal of Software Tools, 24(10): 21–22, 24–26, 28–30, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/datastru.txt.

Spencer:1990:DMA

[Spe90] Lawrence D. Spencer. Debugging memory allocation errors. Dr. Dobb’s Journal of Software Tools, 15(8):80–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Spertus:1997:CGC

[Spe97] Mike Spertus. C++ and garbage collection: Not your father’s garbage collection. Dr. Dobb’s Journal of Software Tools, 22 (12):36, 38, 40–41, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Spillman:1993:GA

[Spi93] Richard Spillman. Genetic algorithms. Dr. Dobb’s Journal of Software Tools, 18(2):26, 28, 30, 90–93, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Spiller:1994:TM

[Spi94] Jeremy Spiller. 3-D texture mapping. Dr. Dobb’s Journal of Software Tools, 19(7):32, 34, 36– 37, 92, 94–95, July 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Schulman:1994:UCW

[SRA^{+}94] Andrew Schulman, Art Rothstein, Roger Alley, et al. Undocumented corner: OS/2 for Windows: IBM’s Patch-O-Rama. Dr. Dobb’s Journal of Software Tools, 19(6):121, 122, 126– 128, 130, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Schneier:1993:CM

[SS93] Bruce Schneier and H. J. Smith. Color models. Dr. Dobb’s Journal of Software Tools, 18(7):38, 40, 42–43, 96, July 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Shapiro:1999:MAE

[SS99] Jim Shapiro and David Shapiro. MMPC: An algorithm for encrypting multiple messages. Dr. Dobb’s Journal of Software Tools, 24(12):32, 34, 36, 38, 40–41, December 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1999/1999_12/mmpc.
                                        txt;   http://www.ddj.com/
                                        ftp/1999/1999_12/mmpc.zip.

Stapfer:1992:TCC

[Sta92] Christian Stapfer. Timed callbacks in C++. Dr. Dobb’s Journal of Software Tools, 17(10): 72–76, 120–126, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1993:RAI

[Sta93a]  William Stallings.  Routing

algorithms for internetworking. Dr. Dobb’s Journal of Software Tools, 18(4):58–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1993:UON

[Sta93b] William Stallings. Understanding OSI network management. Dr. Dobb’s Journal of Software Tools, 18(13):68, 70–72, 74, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1994:PFS

[Sta94a] William Stallings. Packet filtering in the SNMP remote monitor. Dr. Dobb’s Journal of Software Tools, 19(13):70, 72– 74, 99, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1994:PGB

[Sta94b] William Stallings. Password generation by Bloom filters. Dr. Dobb’s Journal of Software Tools, 19(8):119–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1994:SSH

[Sta94c] William Stallings. SHA: the Secure Hash Algorithm. Dr.

Dobb’s Journal of Software Tools, 19(4):32, 34, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1995:CCF

[Sta95a] William Stallings. Congestion control in frame-relay networks. Dr. Dobb’s Journal of Software Tools, 20(3):60, 62, 64, 66, 68, 70, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Stamatakis:1995:UMI

[Sta95b]  William Stamatakis.  Using

MAPI for interapplication communication. Dr. Dobb’s Journal of Software Tools, 20(11):80, 82, 84, 86, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1996:PB

[Sta96] William Stallings. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 21 (7):113–??, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1998:AAE

[Sta98a] William Stallings. Algorithm alley: Exponential smoothing. Dr. Dobb’s Journal of Software Tools, 23(3):127–130, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Stallings:1998:PBF

[Sta98b] William Stallings. Programmer’s bookshelf: In the future with IPv6. Dr. Dobb’s Journal of Software Tools, 23(8): 119–120, August 1998. CODEN DDJOEB. ISSN 1044-789X.

Stanton:1998:TBT

[Sta98c] Scott Stanton. Tclblend: Blending Tcl and Java. Dr. Dobb’s Journal of Software Tools, 23 (2):50, 52–54, 100–101, February 1998.

Stallings:1999:HAK

[Sta99] William Stallings. The HMAC algorithm: Key hashing for message authentication. Dr. Dobb’s Journal of Software Tools, 24 (4):46, 48–49, April 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_04/hmac.
txt.

Stevens:1990:CPC

[Ste90a] A. Stevens. C programming:

CSORT: a sage of a sort. Dr. Dobb’s Journal of Software Tools, 15(4):127–133, 144, 146–147, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPa

[Ste90b] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 15(1):132–??, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPb

[Ste90c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 15(2):123–??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPe

[Ste90d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 15(5):129–??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPf

[Ste90e] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 15(6):135–??, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPg

[Ste90f] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 15(7):131–??, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPh

[Ste90g] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 15(8):149–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPi

[Ste90h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 15(9):127–??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPj

[Ste90i] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 15(10):131–??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPk

[Ste90j] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 15(11):149–??, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPl

[Ste90k] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 15(12):121–??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1990:CPT

[Ste90l]  Al Stevens.  C program-

ming: A thousand CURSES on TEXTSRCH (text retrieval). Dr. Dobb’s Journal of Software Tools, 15(3):127–133, 144–149, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPa

[Ste91a] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 16(1):134–??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPb

[Ste91b] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 16(2):135–??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPc

[Ste91c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 16(3):109–??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPd

[Ste91d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 16(4):117–??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPe

[Ste91e] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 16(5):115–??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPf

[Ste91f] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 16(6):123–??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPg

[Ste91g] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 16(7):113–??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPh

[Ste91h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 16(8):149–??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPi

[Ste91i] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 16(9):??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPj

[Ste91j] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPk

[Ste91k] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 16(11):??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:CPl

[Ste91l] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 16(12):??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCa

[Ste91m] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 16(4):??, April 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCb

[Ste91n] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCc

[Ste91o] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 16(6):??, June 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCd

[Ste91p] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 16(7):??, July 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCe

[Ste91q] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 16(8):??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCf

[Ste91r] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 16(9):??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCg

[Ste91s] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCh

[Ste91t] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 16(11):??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:FCi

[Ste91u] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 16(12):??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:MWS

[Ste91v] Al Stevens. The Mewel Window System. Dr. Dobb’s Journal of Software Tools, 16(3):54, 56– 57, 59–63, 83–84, 86, 88, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1991:VRV

[Ste91w] Al Stevens. Visible results with Visual Basic (Windows tool). Dr. Dobb’s Journal of Software Tools, 16(12):78–87, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Stenger:1992:CDI

[Ste92a]  Allen Stenger.  Converting

dithered images back to gray scale. Dr. Dobb’s Journal of Software Tools, 17(11):64, 66, 68, 133–134, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:ACR

[Ste92b] Al Stevens. Another curmudgeon reflects. Dr. Dobb’s Journal of Software Tools, 17(12): 140–??, December 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPa

[Ste92c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPb

[Ste92d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPc

[Ste92e] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992.

CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPd

[Ste92f] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPe

[Ste92g] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPf

[Ste92h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPg

[Ste92i] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPh

[Ste92j] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPi

[Ste92k] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 17(9):??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPj

[Ste92l] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(9):129–??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPk

[Ste92m] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(10):139–??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPl

[Ste92n] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(11):157–??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CPm

[Ste92o] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 17(12):129–??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CCE

[Ste92p] Al Stevens. Compiler-specific C extensions. Dr. Dobb’s Journal of Software Tools, 17(8):94, 96, 98, 100, 102, 104, 106, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:CBS

[Ste92q] Al Stevens. A conversation with Bjarne Stroustrup. Dr. Dobb’s Journal of Software Tools, 17 (12 (special supplement)):7s– 12s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCRa

[Ste92r] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCRb

[Ste92s] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCRc

[Ste92t] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(11):??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCRd

[Ste92u] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(12):??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCa

[Ste92v] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCb

[Ste92w] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992.

CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCc

[Ste92x] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCd

[Ste92y] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCe

[Ste92z] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCf

[Ste92-27] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCg

[Ste92-28] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCh

[Ste92-29] Al Stevens. ‘D-Flat’ in CP.

Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCi

[Ste92-30] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(9):??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:FCj

[Ste92-31] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 17(10):??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:IBS

[Ste92-32] Al Stevens. Interview with Bjarne Stroustrup. Dr. Dobb’s Journal of Software Tools, 17 (??):7–??, ???? 1992. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1992:LOM

[Ste92-33] Al Stevens. The Lotus Open Message Interface. Dr. Dobb’s Journal of Software Tools, 17 (3):66, 68–71, March 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1992:POC

[Ste92-34] Al Stevens. Persistent objects in C++. Dr. Dobb’s Journal of Software Tools, 17(12):34, 39– 40, 42–44, December 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Steven:1993:CP

[Ste93a] Al Steven. C programming. Dr. Dobb’s Journal of Software Tools, 18(7):115–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:ODM

[Ste93b] A. Stevens. Object-oriented database management systems. Dr. Dobb’s Journal of Software Tools, 18(4):7S–15S, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPa

[Ste93c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 18(1):115–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPb

[Ste93d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 18(2):111–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPc

[Ste93e] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 18(3):113–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPd

[Ste93f] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 18(4):105–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPe

[Ste93g] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 18(5):115–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPf

[Ste93h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 18(6):141–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPg

[Ste93i] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 18(7):??, July 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPh

[Ste93j] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 18(8):123–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPi

[Ste93k] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 18(9):105–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPk

[Ste93l] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 18(11):??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPj

[Ste93m] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 18(11):133–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:CPl

[Ste93n] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 18(13):115–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRa

[Ste93o] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(1):??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRb

[Ste93p] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(2):??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRc

[Ste93q] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(3):??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRd

[Ste93r] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(4):??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRe

[Ste93s] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(6):??, June 1993.

CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRf

[Ste93t] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(11):??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FCRg

[Ste93u] Al Stevens. ‘D-Flat++’: C++ re-implementation of ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(12):??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:FC

[Ste93v] Al Stevens. ‘D-Flat’ in CP. Dr. Dobb’s Journal of Software Tools, 18(2):??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:MAW

[Ste93w] Al Stevens. A multitool approach to Windows development. Dr. Dobb’s Journal of Software Tools, 18(10):7–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Stevens:1993:PBa

[Ste93x] Al Stevens. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18(1):133– ??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1993:PBb

[Ste93y] Al Stevens. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 18(5):137–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:BNL

[Ste94a] Al Stevens. Borland’s new license. Dr. Dobb’s Developer Update, 1(3):2–??, March 1994. CODEN ???? ISSN 1079-8595.

Stevens:1994:CPa

[Ste94b] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(1):105–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPb

[Ste94c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(2):97–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPc

[Ste94d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(3):121–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPd

[Ste94e] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(4):115–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPe

[Ste94f] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(5):117–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPf

[Ste94g] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 19(6):111–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPg

[Ste94h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 19(7):111–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPh

[Ste94i] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 19(8):??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

AlStevens:1994:CP

[Ste94j] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 19(8):113–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPi

[Ste94k] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 19(10):??, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPj

[Ste94l] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 19(11):107–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPk

[Ste94m] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(11):121–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPl

[Ste94n] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(13):109–??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:CPm

[Ste94o] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 19(15):??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:HWH

[Ste94p] Al Stevens. Help for Windows help authors. Dr. Dobb’s Journal of Software Tools, 19(4): 86–91, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:PBa

[Ste94q] Al Stevens. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(5):141–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:PBb

[Ste94r] Al Stevens. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(8):133– ??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QCIa

[Ste94s] Al Stevens. Quincy: A C interpreter project. Dr. Dobb’s Journal of Software Tools, 19(5):??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QCIb

[Ste94t] Al Stevens. Quincy: A C interpreter project. Dr. Dobb’s Journal of Software Tools, 19(6):??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QCIc

[Ste94u] Al Stevens. Quincy: A C interpreter project. Dr. Dobb’s Journal of Software Tools, 19 (8):??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QCId

[Ste94v] Al Stevens. Quincy: A C interpreter project. Dr. Dobb’s Journal of Software Tools, 19 (11):??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QCIe

[Ste94w] Al Stevens. Quincy: A C interpreter project. Dr. Dobb’s Journal of Software Tools, 19(13): ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QCIf

[Ste94x] Al Stevens. Quincy: A C interpreter project. Dr. Dobb’s Journal of Software Tools, 19(15): ??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1994:QAC

[Ste94y] Al Stevens. Quincy: The architecture of A C interpreter. Dr. Dobb’s Journal of Software Tools, 19(??):48–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPa

[Ste95a] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 20(1):115–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPb

[Ste95b] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 20(2):??, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPc

[Ste95c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 20(3):125–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPe

[Ste95d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 20(5):??, May 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPf

[Ste95e] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 20(6):115–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPg

[Ste95f] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(7):115–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPh

[Ste95g] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(8):121–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPi

[Ste95h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(9):123–??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPj

[Ste95i] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(10):137–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPk

[Ste95j] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(11):125–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPl

[Ste95k] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 20(12):125–??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:CPD

[Ste95l] Al Stevens. C programming:

DFWrap: the D-flat C++ wrapper classes. Dr. Dobb’s Journal of Software Tools, 20(5): 115–117, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:PB

[Ste95m] Al Stevens. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20(3):143–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1995:TCG

[Ste95n] Al Stevens. Theatrix: A C++ game class library. Dr. Dobb’s Journal of Software Tools, 20 (??):13–??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Stewart:1995:VWW

[Ste95o] Joe Stewart. VRML and the world wide Web. Dr. Dobb’s Developer Update, 2(6):2–3, June 1995. CODEN ???? ISSN 1079-8595.

Stevens:1996:CPa

[Ste96a] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 21(1):118–??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPb

[Ste96b] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 21(2):121–??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPc

[Ste96c] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 21(3):107–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPd

[Ste96d] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 21(4):105–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPe

[Ste96e] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 21(5):95–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPg

[Ste96f] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 21(6):??, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPi

[Ste96g] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 21(8):105–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPj

[Ste96h] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 21(9):92–??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPk

[Ste96i] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 21(10):105–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPl

[Ste96j] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 21(11):106–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPm

[Ste96k] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 21(12):106–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:CPHb

[Ste96l] Al Stevens. C programming:

Hairy arms for dummies, CD-ROM setups, and Java jive. Dr. Dobb’s Journal of Software Tools, 21(7):91–95, 98–99, 115–116, July 1996. CODEN DDJOEB. ISSN 1044-789X. Compares Java with C++. See comments [Kay96].

Stevens:1996:CPHa

[Ste96m] Al Stevens. C programming:

Herman: document viewer program. Dr. Dobb’s Journal of Software Tools, 21(6):99–??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:QIDa

[Ste96n] Al Stevens. Quincy 96 integrated development environment. Dr. Dobb’s Journal of Software Tools, 21(2):??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:QIDb

[Ste96o] Al Stevens. Quincy 96 integrated development environment. Dr. Dobb’s Journal of Software Tools, 21(3):??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:QIDc

[Ste96p] Al Stevens. Quincy 96 integrated development environment. Dr. Dobb’s Journal of Software Tools, 21(4):??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:QIDd

[Ste96q] Al Stevens. Quincy 96 integrated development environment. Dr. Dobb’s Journal of Software Tools, 21(5):??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1996:QIDe

[Ste96r] Al Stevens. Quincy 96 integrated development environment. Dr. Dobb’s Journal of Software Tools, 21(6):??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Stephens:1997:AAS

[Ste97a] Rod Stephens. Algorithm alley: Searching decision trees. Dr. Dobb’s Journal of Software Tools, 22(12):127–131, 137– 138, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPW

[Ste97b] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 22(3):93–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPb

[Ste97c] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 22(5):91–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPd

[Ste97d] Al Stevens. C programming.

Dr. Dobb’s Journal of Software Tools, 22(6):111–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPR

[Ste97e] Al Stevens. C programming — in rethinking the relationship between C and C++, Al wonders if it is time to put the proverbial cart before the horse. Dr. Dobb’s Journal of Software Tools, 22(7):99–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPU

[Ste97f] Al Stevens. C programming — users expect standard Windows Help for all applications, no matter how small or how intuitive the procedures. to that end, Al updates. Dr. Dobb’s Journal of Software Tools, 22 (1):92–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPAa

[Ste97g] Al Stevens. C programming: Al presents MIDI Xchg, a program that tests the MIDI data stream by sending event messages from one device to another as fast as possible. Dr. Dobb’s Journal of Software Tools, 22(2):97– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPI

[Ste97h] Al Stevens. C programming: Indexing and Microsoft Visual C++ 5.0. Dr. Dobb’s Journal of Software Tools, 22(10):111, 114– 116, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPK

[Ste97i]  Al Stevens.  C program-

ming: Kicking and scripting: JavaScript and CGI. Dr. Dobb’s Journal of Software Tools, 22 (4):92–95, 97, 114, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPAb

[Ste97j] Al Stevens. C programming: McCarthy TeamworX: The path to greatness. Dr. Dobb’s Journal of Software Tools, 22(8): 103–107, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPM

[Ste97k] Al Stevens. C programming: MIDI tech support and Code-Warrior for Win32. Dr. Dobb’s Journal of Software Tools, 22 (11):105–109, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:CPC

[Ste97l] Al Stevens. C programming:

The C mail call. Dr. Dobb’s Journal of Software Tools, 22 (12):113–118, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Stein:1998:FRT

[Ste98a] William M. Stein. Forth and real-time control. Dr. Dobb’s Journal of Software Tools, 23 (10):78, 81–82, 84–88, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
10/rtforth.txt.

Stevens:1998:CPCa

[Ste98b] Al Stevens. C programming:

C^{3}: Conferences, CodeWizard, committees. Dr. Dobb’s Journal of Software Tools, 23(1): 111–113, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:1998:CPCb

[Ste98c] Al Stevens. C programming:

A C++ Standard at last. Dr. Dobb’s Journal of Software Tools, 23(2):115–117, 130–131, February 1998.

Stevens:1998:CPAa

[Ste98d]  Al Stevens.  C program-

ming: AntiPatterns and PTL 2. Dr. Dobb’s Journal of Software Tools, 23(4):114–116, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/ptl2.zip.

Stevens:1998:CPAb

[Ste98e] Al Stevens. C programming: Around the world with the C++ migration. Dr. Dobb’s Journal of Software Tools, 23(10): 121–125, October 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ddj/1998/1998_10/
index.htm.

Stevens:1998:CPD

[Ste98f] Al Stevens. C programming: D-Flat 2000 and Quincy 99: Breathing new life into old projects. Dr. Dobb’s Journal of Software Tools, 23(12):118– 121, December 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://midifitz.com/
alstevens/editor/;   http:
//midifitz.com/alstevens/
quincy99/;   http://www.ddj.
com/ddj/1998/1998_12/../..
/../ftp/1998/1998_12/cpro128.
zip.

Stevens:1998:CPM

[Ste98g] Al Stevens. C programming: MIDIFile: Standard MIDI format file parsing. Dr. Dobb’s Journal of Software Tools, 23 (5):119–121, 137, May 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
05/cprog598.txt;   http://
www.ddj.com/ftp/1998/1998_
05/midifile.zip.

Stevens:1998:CPR

[Ste98h] Al Stevens. C programming: Retro this. Dr. Dobb’s Journal of Software Tools, 23(8): 103–105, August 1998. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1998:CPN

[Ste98i] Al Stevens. C programming:

The next great migration: From C++ to Standard C++. Dr. Dobb’s Journal of Software Tools, 23(9):105–109, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
09/cprog998.txt.

Stevens:1998:CPP

[Ste98j] Al Stevens. C programming:

The persistent template library. Dr. Dobb’s Journal of Software Tools, 23(3):117–120, 137, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
03/cprg398.txt;   http://
www.ddj.com/ftp/1998/1998_
03/ptl1.zip.

Stevens:1998:CPU

[Ste98k] Al Stevens. C programming:

Undo/redo redux. Dr. Dobb’s Journal of Software Tools, 23 (11):119–120, 122, 124–125, November 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/1998/1998_11/CPROG118.
TXT.

Stevens:1998:CPW

[Ste98l] Al Stevens. C programming:

Windows: 98, NT, and CE. Dr. Dobb’s Journal of Software Tools, 23(6):103–105, 123–124, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1999:CPC

[Ste99a] Al Stevens. C programming: C++ for …. Dr. Dobb’s Journal of Software Tools, 24(8): 99–101, August 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL 9908l/9908l.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_08/cprg899.
                                        txt.

Stevens:1999:CPD

[Ste99b]  Al Stevens.  C program-

ming: Doorknob arguments. Dr. Dobb’s Journal of Software Tools, 24(5):115–116, 118, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        05/cprog.txt.

Stevens:1999:CPEa

[Ste99c] Al Stevens. C programming: Editor, an interim project. Dr. Dobb’s Journal of Software Tools, 24(1):107–110, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/ftp/1999/1999_
                                        01/cpro199.txt;   http://
                                        www.ddj.com/ddj/ftp/1999/
                                        1999_01/cpro199.zip.

Stevens:1999:CPH

[Ste99d] Al Stevens. C programming: Herb and me. Dr. Dobb’s Journal of Software Tools, 24 (7):101–102, 104, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/articles/1999/
9907k/9907k.htm.

Stevens:1999:CPJ

[Ste99e] Al Stevens. C programming:

Java jive and scrolling the editor. Dr. Dobb’s Journal of Software Tools, 24(4):96, 98, April 1999. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1999:CPO

[Ste99f] Al Stevens. C programming:

Open Source, open mike. Dr. Dobb’s Journal of Software Tools, 24(9):94, 96–98, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/storch.zip.

Stevens:1999:CPS

[Ste99g] Al Stevens. C programming:

Shareware. Dr. Dobb’s Journal of Software Tools, 24(10):107– 108, 110–111, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1999:CPTb

[Ste99h] Al Stevens. C programming:

Teaching C++ for the sixth time. Dr. Dobb’s Journal of Software Tools, 24(11):107, 108, 110, November 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_11/
cpro1199.zip.

Stevens:1999:CPT

[Ste99i]  Al Stevens.  C program-

ming: Templates and exception handling. Dr. Dobb’s Journal of Software Tools, 24(3): 103–105, March 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm;
 http://www.ddj.com/1999/
9903/9903toc.htm.

Stevens:1999:CPB

[Ste99j] Al Stevens. C programming: The big blow, graphics, and Y2K. Dr. Dobb’s Journal of Software Tools, 24(12):105– 107, December 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_12/
cpro1299.zip.

Stevens:1999:CPEb

[Ste99k] Al Stevens. C programming: The envelope, please. Dr. Dobb’s Journal of Software Tools, 24(2):111–113, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm.

Stinson:1998:VCT

[Sti98] Doug Stinson. Visual cryptography and threshold schemes. Dr. Dobb’s Journal of Software Tools, 23(4):36, 38–43, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Stout:1990:SDE

[Sto90] Robert B. Stout. S-CODER for data encryption. Dr. Dobb’s Journal of Software Tools, 15 (1):52, 54, 56, 58, 110–111, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Stout:1992:CSa

[Sto92a] Robert B. Stout. C snippets. Dr. Dobb’s Journal of Software Tools, 17(9):31, 110, 146, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Stout:1992:CSb

[Sto92b] Robert B. Stout. C snippets.

Dr. Dobb’s Journal of Software Tools, 17(10):124–??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Stout:1992:CSc

[Sto92c] Robert B. Stout. C snippets.

Dr. Dobb’s Journal of Software Tools, 17(11):79, 155, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Stout:1992:CSd

[Sto92d] Robert B. Stout. C snippets.

Dr. Dobb’s Journal of Software Tools, 17(12):61, 119, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Stout:1993:CS

[Sto93] Robert B. Stout. C snippets. Dr. Dobb’s Journal of Software Tools, 18(1):78–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Stockman:1994:OMM

[Sto94] Harlan W. Stockman. Optimizing matrix math on the Pentium. Dr. Dobb’s Journal of Software Tools, 19(5):52, 54, 56, 60, 62, 66, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Storn:1995:CO

[Sto95] Rainer Storn. Constrained optimization. Dr. Dobb’s Journal of Software Tools, 20(5):119–123, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Storn:1997:DTT

[Sto97] Rainer Storn. A debug/trace tool — the debug/trace tool Rainer presents here is useful when you don’t have a powerful debugger at hand, and also when you need to build trace functionality into your code for error-spotting support at the client’s site. Dr. Dobb’s Journal of Software Tools, 22(2):22– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Stroyan:1990:TDG

[Str90]  Michael Stroyan.  Three-

dimensional graphics using the X Window System. Dr. Dobb’s Journal of Software Tools, 15 (2):28–36, February 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Suel:1995:ECD

[Sue95] Patrick Suel. Extending C++ for distributed applications. Dr. Dobb’s Journal of Software Tools, 20(2):50, 52–54, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Sulger:1993:IID

[Sul93] Art Sulger. Indexing image databases. Dr. Dobb’s Journal of Software Tools, 18(8):60, 62– 64, 104–105, August 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Sulger:1994:DMC

[Sul94a] Art Sulger. Database management in C++. Dr. Dobb’s Journal of Software Tools, 19(13):36, 38, 40, 91–93, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Sulger:1994:ERC

[Sul94b] Art Sulger. Extending REXX with C++. Dr. Dobb’s Journal of Software Tools, 19(10):36–37, 68, 86–88, September 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Sulger:1997:DMJ

[Sul97] Art Sulger. Database management and Java. Dr. Dobb’s Journal of Software Tools, 22 (5):28, 30–32, 73, 74, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Summit:1991:CLQa

[Sum91a] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 16(8):78, 93, 111, 140, 172, August 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Summit:1991:CLQb

[Sum91b] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 16(9):75, 101, 141, 147, 151, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Summit:1991:CLQc

[Sum91c] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 16(10):102, 114, 158, 164, 172, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Summit:1991:CLQd

[Sum91d] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 16(11):86– ??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQa

[Sum92a] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(1):130, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQb

[Sum92b] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(2):80, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQc

[Sum92c] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(4):46, 103, 106, 154, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQd

[Sum92d] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(5):147–??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQe

[Sum92e] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(6):95, 164, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQf

[Sum92f] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(7):56, 80, 100, 152, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQg

[Sum92g] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(8):63, 106, 127, 159, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Summit:1992:CLQh

[Sum92h] Steve Summit. C language Q and A. Dr. Dobb’s Journal of Software Tools, 17(9):113, 119, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Samuels:1998:LSA

[SvA^{+}98] Adam D. Samuels, Jerry van Dijk, Dawn Amore, Shlomi Fish, Scott Schwendinger, Arvid R. Hand, Jr., and Howard Mark. Letters: Something in the air; more on Ada; recycling PC’s; server-side scripting; stronger encryption; inner loops; Einstein kudos. Dr. Dobb’s Journal of Software Tools, 23(3):8, 12, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Saade:1995:CSS

[SW95a] Henry Saade and Ann Wallace. Cobol ’97: a status symbol. Dr. Dobb’s Journal of Software Tools, 20(10):52, 54, 115– 116, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Schwaller:1995:DSC

[SW95b] Peter J. Schwaller and John Q. Walker, II. Designing servers with CPI-C. Dr. Dobb’s Journal of Software Tools, 20(2):74, 76– 77, 79–80, 111, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Smith:1995:PA

[SW95c] James E. Smith and Shlomo Weiss. PowerPC 601 and Alpha 21064. Dr. Dobb’s Journal of Software Tools, 20(6):19– ??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1996:UC

[SW96a] George Shepherd and Scot Wingo. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 21(10):123–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1996:UCG

[SW96b] George Shepherd and Scot Wingo. Undocumented corner: George and Scot examine the undocumented areas of CSplitterWnd, one of the most complicated and confusing of all MFC classes. Dr. Dobb’s Journal of Software Tools, 21(12):124– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1997:UCAa

[SW97a] George Shepherd and Scot Wingo. Undocumented corner: Applying CSplitteWnd internals. Dr. Dobb’s Journal of Software Tools, 22(2):111–??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1997:UCAd

[SW97b] George Shepherd and Scot Wingo. Undocumented corner: ATL and dual interfaces. Dr. Dobb’s Journal of Software Tools, 22(12):133–135, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1997:UCAc

[SW97c] George Shepherd and Scot Wingo. Undocumented corner: ATL and the IUknown interface. Dr. Dobb’s Journal of Software Tools, 22(8):119–123, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1997:UCH

[SW97d] George Shepherd and Scot Wingo. Undocumented corner: How MFC does ActiveX connections. Dr. Dobb’s Journal of Software Tools, 22(4):109–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1997:UCM

[SW97e] George Shepherd and Scot Wingo. Undocumented corner: Mini MFC revealed. Dr. Dobb’s Journal of Software Tools, 22(10):125–126, 130, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1997:UCAb

[SW97f] George Shepherd and Scot Wingo. Undocumented corner: The ActiveX template library. Dr. Dobb’s Journal of Software Tools, 22(6):129–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1998:UCAb

[SW98a] George Shepherd and Scot Wingo. Undocumented corner: ATL and connection points. Dr. Dobb’s Journal of Software Tools, 23(6):117–119, 124–125, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Shepherd:1998:UCAa

[SW98b] George Shepherd and Scot Wingo. Undocumented corner: ATL’s ActiveX control architecture. Dr. Dobb’s Journal of Software Tools, 23(2):127–129, 131, February 1998.

Shepherd:1998:UC

[SW98c] George Shepherd and Scot Wingo. Undocumented corner: Inside ATL’s property pages. Dr. Dobb’s Journal of Software Tools, 23(4):130, 132–133, 139, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        04/ud498.txt.

Swaine:1990:PPG

[Swa90a]  M. Swaine.  Programming

paradigms: Getting CLOS (Lisp extension). Dr. Dobb’s Journal of Software Tools, 15(3):

122–125, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPa

[Swa90b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(1): 129–??, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPb

[Swa90c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(2): 118–??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPd

[Swa90d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(4):119–??, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPe

[Swa90e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(5):123–??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPf

[Swa90f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(6):129–??, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPh

[Swa90g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(7):123–??, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPi

[Swa90h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(8): 137–??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPj

[Swa90i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(9):119– ??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPk

[Swa90j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(10): 123–??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPl

[Swa90k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(11):141– ??, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPm

[Swa90l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(12):113– ??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFa

[Swa90m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 15(1):??, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFb

[Swa90n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(2):??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFc

[Swa90o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(3):??, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFd

[Swa90p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(4):??, April 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFe

[Swa90q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(5):??, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFf

[Swa90r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(6):??, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFg

[Swa90s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(7):??, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFh

[Swa90t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(8):??, August 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFi

[Swa90u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(9): ??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFj

[Swa90v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(10):??, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFk

[Swa90w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 15(11): ??, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1990:SFl

[Swa90x]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 15(12): ??, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:FVR

[Swa91a] Michael Swaine. Fire in the valley revisited. Dr. Dobb’s Journal of Software Tools, 16(1): 96J, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPa

[Swa91b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(1): 126–??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPb

[Swa91c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(2): 131–??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPc

[Swa91d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(3):103–??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPd

[Swa91e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(4):111–??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPe

[Swa91f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(5):109–??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPf

[Swa91g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(6):115–??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPg

[Swa91h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(7):105–??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPh

[Swa91i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(8): 143–??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPi

[Swa91j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(9): ??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPj

[Swa91k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPk

[Swa91l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(11): ??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:PPl

[Swa91m] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 16(12): ??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFa

[Swa91n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(1):??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFb

[Swa91o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(2):??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFc

[Swa91p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(3):??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFd

[Swa91q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(4):??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFe

[Swa91r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFf

[Swa91s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(6):??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFg

[Swa91t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(7):??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFh

[Swa91u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(8):??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFi

[Swa91v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(9): ??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFj

[Swa91w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFk

[Swa91x]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(11): ??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1991:SFl

[Swa91y]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 16(12): ??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPa

[Swa92a] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPb

[Swa92b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(2): ??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPc

[Swa92c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPd

[Swa92d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPe

[Swa92e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPf

[Swa92f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPg

[Swa92g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPh

[Swa92h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPi

[Swa92i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(9):123– ??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPj

[Swa92j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(10): 133–??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPk

[Swa92k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(11):151– ??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:PPl

[Swa92l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 17(12):123– ??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFa

[Swa92m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFb

[Swa92n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFc

[Swa92o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFd

[Swa92p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFe

[Swa92q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFf

[Swa92r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFg

[Swa92s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFh

[Swa92t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFi

[Swa92u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(9): ??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFj

[Swa92v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(9):176– ??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFk

[Swa92w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 17(10):??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFl

[Swa92x]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(11):208– ??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1992:SFm

[Swa92y]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 17(12):168– ??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Swain:1993:AA

[Swa93a] Tom Swain. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 18(13):119–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPa

[Swa93b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(1): 109–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPb

[Swa93c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(2): 105–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPc

[Swa93d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(3):109–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPd

[Swa93e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(4):99–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPe

[Swa93f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(5):109–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPf

[Swa93g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(6):137–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPg

[Swa93h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(7):109–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPh

[Swa93i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(8): 117–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPi

[Swa93j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(9):99– ??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPk

[Swa93k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(11): ??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPj

[Swa93l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(11):127– ??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:PPl

[Swa93m] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 18(13 (or 12??)):109–??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFa

[Swa93n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(1):152– ??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFb

[Swa93o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(2):152– ??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFc

[Swa93p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(3):168–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFd

[Swa93q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(4):152–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFe

[Swa93r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(5):152–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFf

[Swa93s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(6):176–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFg

[Swa93t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(7):152–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFh

[Swa93u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(8):160– ??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFi

[Swa93v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(9):144– ??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFj

[Swa93w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 18(11):176– ??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1993:SFk

[Swa93x]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 18(13):144– ??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAa

[Swa93y] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 18(5):119–??, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAb

[Swa93z] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 18(6):147–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAc

[Swa93-27] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 18(7):121–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAd

[Swa93-28] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 18(8):129–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAe

[Swa93-29] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 18(9):115–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAg

[Swa93-30] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 18(11):??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAf

[Swa93-31] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 18(11):139–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Swan:1993:AAh

[Swa93-32] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 18(12):??, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:AIH

[Swa94a] Michael Swaine. Artificial intelligence: Out of hybernation? Dr. Dobb’s Developer Update, 1 (11):1–??, October 1994. CO-DEN ???? ISSN 1079-8595.

Swaine:1994:PPa

[Swa94b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(1): 101–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPb

[Swa94c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(2):91– ??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPc

[Swa94d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(3):115–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPd

[Swa94e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(4):109–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPe

[Swa94f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(5):113–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPf

[Swa94g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(6):105–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPg

[Swa94h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(7):107–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPh

[Swa94i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(8): 109–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPi

[Swa94j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(11): 103–??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPk

[Swa94k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(13):105– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPj

[Swa94l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(13):115– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:PPl

[Swa94m] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 19(15):103– ??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFa

[Swa94n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(1):144– ??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFb

[Swa94o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(2):136– ??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFc

[Swa94p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(3):168–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFd

[Swa94q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(4):152–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFe

[Swa94r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(5):152–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFf

[Swa94s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(6):160–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFg

[Swa94t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(7):152–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFh

[Swa94u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(8):152– ??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFi

[Swa94v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(11):136– ??, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFk

[Swa94w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 19(13):144– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFj

[Swa94x]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(13):160– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1994:SFl

[Swa94y]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 19(15):136– ??, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Swan:1994:AAa

[Swa94z] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(1):111–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Swan:1994:AAb

[Swa94-27] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(2):103–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Swan:1994:AAc

[Swa94-28] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(3):127–??, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Swan:1994:AAe

[Swa94-29] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 19(5):123–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Swan:1994:AAf

[Swa94-30] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 19(6):117–??, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Swan:1994:AAd

[Swa94-31] Tom R. Swan. Algorithm alley.

Dr. Dobb’s Journal of Software Tools, 20(4):121–??, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:ASA

[Swa95a] Michael Swaine. Apple shines at MacWorld expo. Dr. Dobb’s Developer Update, 2(3):1, 6, March 1995. CODEN ???? ISSN 1079-8595.

Swaine:1995:PB

[Swa95b] Michael Swaine. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20(11):145– ??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPa

[Swa95c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(1): 109–??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPb

[Swa95d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(2): ??, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPc

[Swa95e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(3):109–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPe

[Swa95f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(5):??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPd

[Swa95g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(5):109–??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPg

[Swa95h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(7):109–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPh

[Swa95i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(8): 115–??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPi

[Swa95j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(9):119– ??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPj

[Swa95k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(10): 131–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPk

[Swa95l] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(11):121– ??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPl

[Swa95m] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 20(12):121– ??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:PPF

[Swa95n] Michael Swaine. Programming paradigms: Fluid concepts and creative analogies. Dr. Dobb’s Journal of Software Tools, 20 (6):111–113, June 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFa

[Swa95o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(1):152– ??, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFc

[Swa95p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(3):??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFb

[Swa95q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(3):152–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFd

[Swa95r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(4):??, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFe

[Swa95s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(5):152–??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFf

[Swa95t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(6):144–??, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFg

[Swa95u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(7):144–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFh

[Swa95v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(8):152– ??, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFi

[Swa95w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 20(9):152– ??, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFj

[Swa95x]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(10):168– ??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFk

[Swa95y]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(11):152– ??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1995:SFl

[Swa95z]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 20(12):152– ??, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPa

[Swa96a] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(1): 111–??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPc

[Swa96b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(3):103–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPd

[Swa96c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(4):101–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPe

[Swa96d] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(5):9–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPf

[Swa96e] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(6):93–??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPg

[Swa96f] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(7):87–??, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPh

[Swa96g] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(8): 101–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPi

[Swa96h] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(9):88– ??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPj

[Swa96i] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(10): 99–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPk

[Swa96j] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(11):101– ??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPP

[Swa96k] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 21(12): 98–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:PPL

[Swa96l] Michael Swaine. Programming paradigms: Looking for an HTML book. Dr. Dobb’s Journal of Software Tools, 21(2): 117–119, February 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFa

[Swa96m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 21(1):144– ??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFb

[Swa96n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(2):144– ??, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFc

[Swa96o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(3):136–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFd

[Swa96p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(4):136–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFe

[Swa96q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(5):128–??, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFf

[Swa96r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(6):136–??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFg

[Swa96s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(7):128–??, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFh

[Swa96t]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(8):144– ??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFi

[Swa96u]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(9):128– ??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFj

[Swa96v]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 21(10):144– ??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFk

[Swa96w] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 21(12):144– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1996:SFS

[Swa96x]  Michael Swaine.  Swaine’s

flames: A stagnation with big shirts. Dr. Dobb’s Journal of Software Tools, 21(11):136– ??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Swartz:1996:AA

[Swa96y] John Swartz. Algorithm alley: Constructing combinations using LISP: Sometimes, you just need the right tool. Dr. Dobb’s Journal of Software Tools, 21 (11):116, 117, 127, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPa

[Swa97a] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 22(3):89–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPc

[Swa97b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 22(4):87–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPd

[Swa97c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 22(6):107–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPMa

[Swa97d]  Michael Swaine.  Program-

ming paradigms — Michael looks at paradigms past, focusing this month on spreadsheets. Dr. Dobb’s Journal of Software Tools, 22(1):88–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPI

[Swa97e]  Michael Swaine.  Program-

ming paradigms: Isn’t it ironic?

Dr. Dobb’s Journal of Software Tools, 22(8):99–101, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPJb

[Swa97f]  Michael Swaine.  Program-

ming paradigms: Java bashing, books, and beans. Dr. Dobb’s Journal of Software Tools, 22 (7):93–95, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPJa

[Swa97g] Michael Swaine. Programming paradigms: Java jabs, Visi-Corpse, and the blood of Red Smith. Dr. Dobb’s Journal of Software Tools, 22(5):83, 84, 86, 89, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPMb

[Swa97h] Michael Swaine. Programming paradigms: Macs, flaks, and Illiacs. Dr. Dobb’s Journal of Software Tools, 22(10):107–109, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPSa

[Swa97i] Michael Swaine. Programming paradigms: Some observations on Apple and Java. Dr. Dobb’s Journal of Software Tools, 22 (2):91–93, February 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPSc

[Swa97j] Michael Swaine. Programming paradigms: Stocking stuffers. Dr. Dobb’s Journal of Software Tools, 22(12):108–111, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPB

[Swa97k] Michael Swaine. Programming paradigms: The bridges of Santa Clara County. Dr. Dobb’s Journal of Software Tools, 22(9):93– 95, September 1997. CODEN DDJOEB. ISSN 1044-789X. Comments on the role of Java in the Apple’s survival struggle.

Swaine:1997:PPSb

[Swa97l] Michael Swaine. Programming paradigms: The sincerest form of flattery. Dr. Dobb’s Journal of Software Tools, 22(11):101– 103, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFa

[Swa97m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 22(1):128– ??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFb

[Swa97n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 22(2):128– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFd

[Swa97o]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 22(3):128–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFe

[Swa97p]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 22(4):128–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFf

[Swa97q]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 22(5):136–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFg

[Swa97r]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 22(6):144–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFh

[Swa97s]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 22(7):128–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFC

[Swa97t]  Michael Swaine.  Swaine’s

flames: A clock puzzle. Dr. Dobb’s Journal of Software Tools, 22(11):136, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFJ

[Swa97u]  Michael Swaine.  Swaine’s

flames: Just say NOS. Dr. Dobb’s Journal of Software Tools, 22(12):152, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFR

[Swa97v]  Michael Swaine.  Swaine’s

flames: Readers run hot and cold. Dr. Dobb’s Journal of Software Tools, 22(8):136, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:SFW

[Swa97w] Michael Swaine. Swaine’s flames: War and pique. Dr. Dobb’s Journal of Software Tools, 22(10):144, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Swartz:1997:AA

[Swa97x] John Swartz. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 22(6):123–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Swartz:1997:SDA

[Swa97y] John Swartz. Simulating the Denver Airport Automated Baggage System — John uses Xlisp to implement a computer simulation of the essential components of the Denver. Dr. Dobb’s Journal of Software Tools, 22(1):56–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPC

[Swa98a] Michael Swaine. Programming paradigms: A chat with Bob Bemer. Dr. Dobb’s Journal of Software Tools, 23(5):115–117, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPA

[Swa98b]  Michael Swaine.  Program-

ming paradigms: Adventures in benchmarking (my don’t do list). Dr. Dobb’s Journal of Software Tools, 23(3):109–111, 114, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPHa

[Swa98c] Michael Swaine. Programming paradigms: Hill climbing and other paradigms. Dr. Dobb’s Journal of Software Tools, 23 (4):106, 108–109, 111, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPHb

[Swa98d]  Michael Swaine.  Program-

ming paradigms: Hot flashes. Dr. Dobb’s Journal of Software Tools, 23(9):99–100, 102– 103, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPJ

[Swa98e] Michael Swaine. Programming paradigms: Java and other weapons of war. Dr. Dobb’s Journal of Software Tools, 23 (1):107–109, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPL

[Swa98f] Michael Swaine. Programming paradigms: Leo and the lizards. Dr. Dobb’s Journal of Software Tools, 23(7):101–103, July 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPRb

[Swa98g] Michael Swaine. Programming paradigms: Rooting for the little guy. Dr. Dobb’s Journal of Software Tools, 23(10):117–119, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ddj/1998/1998_
                                        10/index.htm.

Swaine:1998:PPSa

[Swa98h] Michael Swaine. Programming paradigms: Show me the pattern. Dr. Dobb’s Journal of Software Tools, 23(6):98–101, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPSc

[Swa98i] Michael Swaine. Programming paradigms: Stocking stuffers. Dr. Dobb’s Journal of Software Tools, 23(12):114, 116– 117, December 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPSb

[Swa98j]  Michael Swaine.  Program-

ming paradigms: Sun dreams of Jini. Dr. Dobb’s Journal of Software Tools, 23(11):113–114, 117, November 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPP

[Swa98k]  Michael Swaine.  Program-

ming paradigms: The philosophical fragments of La Bamba. Dr. Dobb’s Journal of Software Tools, 23(8):99–101, August 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPRa

[Swa98l] Michael Swaine. Programming paradigms: The rest of us. Dr. Dobb’s Journal of Software Tools, 23(2):107–109, February 1998.

Swaine:1998:SFc

[Swa98m] Michael Swaine. Swaine’s flames. Dr. Dobb’s Journal of Software Tools, 23(3):152, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:SFL

[Swa98n]  Michael Swaine.  Swaine’s

flames: 2001: The legacy code. Dr. Dobb’s Journal of Software Tools, 23(5):152, May 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1998:SFb

[Swa98o]  Michael Swaine.  Swaine’s

flames: 2001, Y2K, and other memorable numbers. Dr. Dobb’s Journal of Software Tools, 23 (2):144, February 1998.

Swaine:1998:SFG

[Swa98p]  Michael Swaine.  Swaine’s

flames: A gathering of egos. Dr. Dobb’s Journal of Software Tools, 23(4):152, April 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:SFA

[Swa98q]  Michael Swaine.  Swaine’s

flames: Animal intelligence. Dr. Dobb’s Journal of Software Tools, 23(9):136, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:SFD

[Swa98r]  Michael Swaine.  Swaine’s

flames: Doing facemail at Foo Bar. Dr. Dobb’s Journal of Software Tools, 23(11):152, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/;   http://
www.ddj.com/ddj/1998/1998_
11/#master_top.

Swaine:1998:SFH

[Swa98s]  Michael Swaine.  Swaine’s

flames: Housecleaning at Foo Bar. Dr. Dobb’s Journal of Software Tools, 23(10):152, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/;   http://
www.ddj.com/ddj/1998/1998_
10/#master_top.

Swaine:1998:SFPa

[Swa98t]  Michael Swaine.  Swaine’s

flames: Paranoia. Dr. Dobb’s Journal of Software Tools, 23 (6):130–131, June 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1998:SFPb

[Swa98u]  Michael Swaine.  Swaine’s

flames: Politics. Dr. Dobb’s Journal of Software Tools, 23 (8):128, August 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1998:SFR

[Swa98v]  Michael Swaine.  Swaine’s

flames: The real power behind Linux. Dr. Dobb’s Journal of Software Tools, 23(12):144, December 1998.

CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/;   http://
www.ddj.com/ddj/1998/1998_
12/#master_top.

Swaine:1998:SFS

[Swa98w] Michael Swaine. Swaine’s flames: The second puzzle. Dr. Dobb’s Journal of Software Tools, 23(1):144, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPd

[Swa99a] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 24(12):101– 103, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PP

[Swa99b] Michael Swaine. Programming paradigms: 1984. Dr. Dobb’s Journal of Software Tools, 24 (7):93–97, July 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPB

[Swa99c] Michael Swaine. Programming paradigms: Blarney and Steve Jobs. Dr. Dobb’s Journal of Software Tools, 24(4):91–93, April 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPJa

[Swa99d] Michael Swaine. Programming paradigms: I joined the Rebol alliance. Dr. Dobb’s Journal of Software Tools, 24(2): 107–109, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/paradigm.
                                        txt.

Swaine:1999:PPJb

[Swa99e]  Michael Swaine.  Program-

ming paradigms: Jerry’s kids. Dr. Dobb’s Journal of Software Tools, 24(8):95–97, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPLa

[Swa99f] Michael Swaine. Programming paradigms: Let the fin begin. Dr. Dobb’s Journal of Software Tools, 24(1):101–103, January 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPLb

[Swa99g]  Michael Swaine.  Program-

ming paradigms: Little engines that could. Dr. Dobb’s Journal of Software Tools, 24(3): 99–101, March 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/1999/9902/9902toc.htm.

Swaine:1999:PPM

[Swa99h] Michael Swaine. Programming paradigms: Magic. Dr. Dobb’s Journal of Software Tools, 24 (11):103–105, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPL

[Swa99i] Michael Swaine. Programming paradigms: The long view. Dr. Dobb’s Journal of Software Tools, 24(9):89–90, 92– 93, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:PPS

[Swa99j]  Michael Swaine.  Program-

ming paradigms: The sevenyear itch. Dr. Dobb’s Journal of Software Tools, 24(10): 99–100, 102, 105, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/paradigm.txt.

Swaine:1999:SFA

[Swa99k]  Michael Swaine.  Swaine’s

flames: Another kind of e-book. Dr. Dobb’s Journal of Software Tools, 24(7):136, July 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1999:SFD

[Swa99l]  Michael Swaine.  Swaine’s

flames: Death rattle. Dr. Dobb’s Journal of Software Tools, 24(4):128, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/;   http://
www.ddj.com/ddj/1999/1999_
04/#master_top.

Swaine:1999:SFK

[Swa99m] Michael Swaine. Swaine’s flames: I know what you did last summer. Dr. Dobb’s Journal of Software Tools, 24(10):136, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ddj/privacy.
htm.

Swaine:1999:SFL

[Swa99n]  Michael Swaine.  Swaine’s

flames: Lurking by the Chad Bowl at Trackball 99. Dr. Dobb’s Journal of Software Tools, 24(8):128, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1999:SFM

[Swa99o]  Michael Swaine.  Swaine’s

flames: Moving and chaos. Dr. Dobb’s Journal of Software Tools, 24(5):144, May 1999. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1999:SFT

[Swa99p]  Michael Swaine.  Swaine’s

flames: Out time. Dr. Dobb’s Journal of Software Tools, 24 (11):144, November 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ddj/privacy.htm.

Swaine:1999:SFC

[Swa99q]  Michael Swaine.  Swaine’s

flames: The creative process. Dr. Dobb’s Journal of Software Tools, 24(3):136, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        articles/1999/9903/9903t/
                                        9903t.htm.

Swaine:1999:SFF

[Swa99r]  Michael Swaine.  Swaine’s

flames: The flip side. Dr. Dobb’s Journal of Software Tools, 24 (3):144, March 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
articles/1999/9902/9902u/
9902u.htm.

Swaine:1999:SFG

[Swa99s]  Michael Swaine.  Swaine’s

flames: The ghost in gauze. Dr. Dobb’s Journal of Software Tools, 24(1):136, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/;   http://
www.ddj.com/ddj/1999/1999_
01/#master_top.

Swaine:1999:SFS

[Swa99t]  Michael Swaine.  Swaine’s

flames: The spam festival. Dr. Dobb’s Journal of Software Tools, 24(9):128, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ddj/privacy.
htm.

Swaine:1999:SFW

[Swa99u]  Michael Swaine.  Swaine’s

flames: Whose page is it, anyway? Dr. Dobb’s Journal of Software Tools, 24(12): 136, December 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/privacy.htm.

Sylvester:1999:LMB

[SWF^{+}99] James H. Sylvester, John B.

Willison, Austin Fletcher, Joe Sam Shirah, and John Sadler. Letters: 3D modeling; Bill’s inferno; more on the Crouch-Echlin effect; myopic about file names; Ficl update. Dr. Dobb’s Journal of Software Tools, 24 (4):10, 12, April 1999. CODEN DDJOEB. ISSN 1044-789X.

Switzer:1990:CDB

[Swi90] John Switzer. Closing DOS’s backdoor. Dr. Dobb’s Journal of Software Tools, 15(10):42, 44, 47–48, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Syck:1990:DLL

[Syc90] Gary Syck. Dynamic link libraries for DOS. Dr. Dobb’s Journal of Software Tools, 15 (5):30–39, 104–106, 108–109, May 1990. CODEN DDJOEB. ISSN 1044-789X.

Sylvester:1993:DCA

[Syl93] James H. Sylvester. Differential compression algorithms. Dr. Dobb’s Journal of Software Tools, 18(4):144–146, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Symons:1994:CPT

[Sym94] George J. Symons. Capture/ playback techniques [sidebar]. Dr. Dobb’s Journal of Software Tools, 19(2):42–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Szczerbina:1993:NNC

[Szc93] Pawel Szczerbina. Novell’s Net-Ware core protocol. Dr. Dobb’s Journal of Software Tools, 18 (11):123–??, November 1993.

CODEN DDJOEB. ISSN 1044-789X.

Tarpenning:1991:CMC

[Tar91]  Marc Tarpenning.  Cooper-

ative multitasking in C++. Dr. Dobb’s Journal of Software Tools, 16(4):54, 56, 58–59, 96, 98–99, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Taylor:1998:IPU

[Tay98] Pamela J. Taylor. Internet programming: Using NetRexx. Dr. Dobb’s Journal of Software Tools, 23(4):84–87, 101, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/netrexx.txt.

Taylor:1999:PBL

[Tay99] Jeff Taylor. Programmer’s bookshelf: Linux sysadmin on your own. Dr. Dobb’s Journal of Software Tools, 24(7):127–128, July 1999. CODEN DDJOEB. ISSN 1044-789X.

Teale:1991:PCS

[Tea91]  Steve Teale.  Proposing a

C++ string class standard. Dr. Dobb’s Journal of Software Tools, 16(10):24, 26, 28, 30, 34, 114, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Teig:1998:LJD

[Tei98] Oyvind Teig. Letters: Java deadlock. Dr. Dobb’s Journal of Software Tools, 23(1):10, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Templ:1994:OPL

[Tem94] Josef Templ. The Oberon programming language. Dr. Dobb’s Journal of Software Tools, 19 (??):60–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Tennberg:1998:WND

[Ten98] Patrick Tennberg. Windows NT device driver toolkits. Dr. Dobb’s Journal of Software Tools, 23(3):92, 94–95, 97, 106–107, March 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1998/1998_03/ntdd.
                                        txt;   http://www.ddj.com/
                                        ftp/1998/1998_03/ntdd.zip.

Tessier:1996:SDB

[Tes96a] Tom Tessier. Sharing data between Web page frames using JavaScript: Implementing a hidden-frame technique. Dr. Dobb’s Journal of Software Tools, 21(5):72, 74–75, 86–87, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Tessier:1996:UJC

[Tes96b] Tom Tessier. Using JavaScript to create interactive Web pages. Dr. Dobb’s Journal of Software Tools, 21(3):84, 86–89, 96–97, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Tetewsky:1994:GDR

[Tet94] Avram K. Tetewsky. GUI development for real-time applications. Dr. Dobb’s Journal of Software Tools, 19(6):28, 30, 32, 36, 38, 40–41, June 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Tewell:1997:WPW

[Tew97]  Thomas Tewell.  Writing

portable Win32 SCSI applications: The WinASPI 32-bit DLL emulator module makes it possible. Dr. Dobb’s Journal of Software Tools, 22(12):56, 58– 59, 102–103, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Tucker:1994:FLC

[TFS94] John A. R. Tucker, Phillip E.

Fraley, and Lawrence P. Swanson. Fuzzy logic in C: an update. Dr. Dobb’s Journal of Software Tools, 19(4):56, 58, 60, 62, 101, 102, April 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Thaler:1997:EEM

[Tha97] Les Thaler. Examining extended MAPI 1.0 — the Extended Messaging Application Programming Interface 1.0, not to be confused with the previousgeneration “Simple MAPI,” was created to standardize the interfaces between messaging applications and the underlying messaging systems. Dr. Dobb’s Journal of Software Tools, 22 (1):64–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

LPF:1990:SP

[The90] The League for Programming Freedom. Software patents.

Dr. Dobb’s Journal of Software Tools, 15(11):56–??, November 1990. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVa

[The98a] The DDJ staff. News & views. Dr. Dobb’s Journal of Software Tools, 23(1):18, January 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVO

[The98b] The DDJ staff. News & views: Online database derby results; integrated ICs; Java gets the nod; HP to go; drive capacities go even higher; DSL gets a boost; school days; NC: The network car?; the bug itself; GeezerWare. Dr. Dobb’s Journal of Software Tools, 23(2):18, February 1998.

Thomas:1991:EKD

[Tho91a] Kas Thomas. Entropy: the key to data compression. Dr. Dobb’s Journal of Software Tools, 16 (2):32, 34, 110, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Thorn:1991:FSU

[Tho91b] Dale Thorn. Fast sorting using large string buffers. Dr. Dobb’s Journal of Software Tools, 16 (6):34, 36, 38, 94–97, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Thorpe:1992:VCI

[Tho92] Danny Thorpe. A Video Compatibility Interface for Turbo Debugger. Dr. Dobb’s Journal of Software Tools, 17(9):94, 96, 101, 120–121, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Tichenor:1997:JCB

[Tic97] Charles B. Tichenor. JavaScript cookies; browser-side persistence for Web developers. Dr. Dobb’s Journal of Software Tools, 22(5):42–45, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Taylor:1999:LRT

[TKG^{+}99] Nick Taylor, Robert Konigsberg, Joshua Graham, Peter Roth, Terry J. Reedy, Al Stevens, Joe Ganley, and Doug McIlroy. Letters: Robotics; thread communication and parallel algorithms; analyzing algorithms; command-line arguments; simulated recursion; Y2K bug? Dr. Dobb’s Journal of Software Tools, 24(8): 12, 14, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Litzkow:1995:CDP

[TL95a] Todd Tannenbaum and Michael Litzkow. The Condor distributed processing system. Dr. Dobb’s Journal of Software Tools, 20(2):40–??, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Tannenbaum:1995:CDP

[TL95b] Todd Tannenbaum and Michael Litzkow. The Condor distributed processing system. Dr. Dobb’s Journal of Software Tools, 20(2):40, 42–44, 47–48, 102, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Tryggvesson:1999:JJR

[TMH99] Jörgen Tryggvesson, Torbjörn Mattsson, and Hansruedi Heeb. Jbed: Java for real-time systems. Dr. Dobb’s Journal of Software Tools, 24(11):78, 80, 82–84, 86, November 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1999/1999_11/jbed.
                                        txt;   http://www.ddj.com/
                                        ftp/1999/1999_11/jbed.zip.

Tremblett:1999:LJP

[TMP^{+}99] Paull Tremblett, A. P. Madden, Serguei Patchkovskii, Edd Dumbill, Michael Olson, Tim Kientzle, and Andrew Bowley. Letters: Java provided update; better late than never; CD authoring; full-text searching; taming C++. Dr. Dobb’s Journal of Software Tools, 24(5):10, 12, May 1999. CODEN DDJOEB. ISSN 1044-789X.

Tomson:1996:VHD

[Tom96] Phil Tomson. VHDL for hardware design. Dr. Dobb’s Journal of Software Tools, 21(6):46, 48, 50, 53, 55, 86, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Tonkin:1990:PP

[Ton90] Bruce W. Tonkin. The power in PowerBasic. Dr. Dobb’s Journal of Software Tools, 15(7):72, 74, 76–77, 80, 116–120, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Tonkin:1991:SB

[Ton91] Bruce W. Tonkin. Speedy buffering. Dr. Dobb’s Journal of Software Tools, 16(3):52–53, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Tooke:1990:OD

[Too90] Simon Tooke. Object-oriented debugging. Dr. Dobb’s Journal of Software Tools, 15(11):36, 38, 40, 42, 114–115, November 1990. CODEN DDJOEB. ISSN 1044-789X.

Torelli:1995:MFF

[Tor95] Peter Torelli. The Microsoft flash file system. Dr. Dobb’s Journal of Software Tools, 20 (2):62, 64, 68, 70, 72, 108– 110, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Jean-Fran:1999:HTC

[Tou99] Jean-François Touchette. HTML thin client and transactions. Dr. Dobb’s Journal of Software Tools, 24(10):80, 82, 84–86, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/htmlthin.txt;   http://
www.ddj.com/ftp/1999/1999_
10/htmlthin.zip.

Tracy:1990:ZES

[Tra90] Martin Tracy. ZEN for embedded systems. Dr. Dobb’s Journal of Software Tools, 15(1):38, 40, 42, 44, 46, 98, 100–102, 104– 106, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Tredennick:1993:CSM

[Tre93] Nick Tredennick. Computer science and the microprocessor. Dr. Dobb’s Journal of Software Tools, 18(6):18–20, 22, 24, 28, 30–32, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Tredennick:1994:MHP

[Tre94] Nick Tredennick. Microprocessors hit the performance wall (again) [sidebar]. Dr. Dobb’s Journal of Software Tools, 19 (1):34–??, January 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Tredennick:1995:FPL

[Tre95] Nick Tredennick. The future of programmable logic. Dr. Dobb’s Journal of Software Tools, 20 (7):60, 62, 64, 66, 68, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Tremblett:1998:JRJ

[Tre98] Paul Tremblett. Java Reflection: Not just for tool developers. Dr. Dobb’s Journal of Software Tools, 23(1):36, 38–39, 99–101, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Tremblett:1999:JPA

[Tre99a] Paul Tremblett. The Java provider architecture. Dr. Dobb’s Journal of Software Tools, 24(3):40, 42, 44–47, 49, March 1999. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
1999/9902/9902toc.htm;   http:
//www.ddj.com/ftp/1999/1999_

03/provider.txt;   http://
www.ddj.com/ftp/1999/1999_
03/provider.zip.

Tremblett:1999:JQW

[Tre99b] Paul Tremblett. Java Q&A:

What are JavaServer pages? Dr. Dobb’s Journal of Software Tools, 24(12):109–113, 115, December 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_12/
jqa1299.txt;   http://www.
ddj.com/ftp/1999/1999_12/
jqa1299.zip.

Tremblett:1999:XC

[Tre99c] Paul Tremblett. X.509 certificates. Dr. Dobb’s Journal of Software Tools, 24(7):42, 44, 46, 47–51, July 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_07/x509.
txt.

Troy:1991:RCPa

[Tro91a] Dan Troy. Remote connectivity for portable terminals. part I. Dr. Dobb’s Journal of Software Tools, 16(2):66, 68, 70, 73– 75, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Troy:1991:RCPb

[Tro91b] Dan Troy. Remote connectivity for portable terminals. part II. developing the VT100 terminal emulation application. Dr. Dobb’s Journal of Software Tools, 16(3):70–71, 148– 149, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Troy:1994:ENS

[Tro94] Dan Troy. Emulating Non-DOS systems under MS-DOS. Dr. Dobb’s Journal of Software Tools, 19(3):52, 54, 57–58, 60, 62, March 1994. CODEN DDJOEB. ISSN 1044-789X.

Trout:1997:WNP

[Tro97]  Paul Trout.  Windows NT

printer control. Dr. Dobb’s Journal of Software Tools, 22 (5):53–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Trout:1998:DUT

[Tro98] Paul Trout. Domain usage tracking for Windows NT. Dr. Dobb’s Journal of Software Tools, 23(12):60, 62–65, December 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ddj/1998/1998_
                                        12/../../../ftp/1998/1998_
                                        12/ntdomain.txt;   http://
                                        www.ddj.com/ddj/1998/1998_
                                        12/../../../ftp/1998/1998_
                                        12/ntdomain.zip.

Thomas:1998:IPD

[TSR98] Louis Thomas, Sean Suchter, and Adam Rifkin. Internet programming: Developing peer-to-peer applications for the Internet: Introducing SimulEdit, an distributed text editor. Dr. Dobb’s Journal of Software Tools, 23(1):76, 78–81, 104– 105, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Tucker:1998:WCR

[Tuc98] Andrew Tucker. The Windows CE 2.0 Remote API. Dr. Dobb’s Journal of Software Tools, 23 (9):52, 53–57, 95–96, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
09/ceremote.txt;   http://
www.ddj.com/ftp/1998/1998_
09/ceremote.zip.

Tucker:1999:UIE

[Tuc99] Andrew Tucker. Using Internet Explorer’s HTMLParser. Dr. Dobb’s Journal of Software Tools, 24(8):82, 84–87, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
08/htmlpars.txt;   http://
www.ddj.com/ftp/1999/1999_
08/htmlpars.zip.

Thorson:1993:NAS

[TWH93] Mark Thorson, Forrest Warthman, and Mark Holler. A neural-network audio synthesizer. Dr. Dobb’s Journal of Software Tools, 18(2):50, 52, 54, 58, 60, 62, 64, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Tyma:1996:TJP

[Tym96] Paul Tyma. Tuning Java performance. Dr. Dobb’s Journal of Software Tools, 21(4):52, 55– 56, 58, 90, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Uhler:1999:EBS

[Uhl99] Stephen Uhler. Event-based servers in Tcl. Dr. Dobb’s Journal of Software Tools, 24 (9):74, 76–78, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        09/tclserv.txt;   http://
                                        www.ddj.com/ftp/1999/1999_
                                        09/tclserv.zip.

Vaidyanathan:1993:MFW

[Vai93] Shankar Vaidyanathan. Multitasking Fortran and Windows NT. Dr. Dobb’s Journal of Software Tools, 18(10):21–??, October 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Valdes:1991:FFC

[Val91a] Ray Valdés. Fanning the flames in “The Changing Landscape of Software Development”. Dr. Dobb’s Journal of Software Tools, 16(1):96R, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1991:LLB

[Val91b] Ray Valdés. Little languages, big questions. Dr. Dobb’s Journal of Software Tools, 16(9):16– 18, 20, 22, 25, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1991:TOP

[Val91c]  Ray Valdés.  A technical

overview of PenPoint in PPD. Dr. Dobb’s Journal of Software Tools, 16(11):16–??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1991:WB

[Val91d] Ray Valdés. What is BioComputing? Dr. Dobb’s Journal of Software Tools, 16(4):46, 108–109, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1992:DWT

[Val92a] Ray Valdes. Doing the wrong thing. Dr. Dobb’s Journal of Software Tools, 17(12 (special supplement)):4s–4s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1992:DDH

[Val92b] Ray Valdés. The Dr. Dobb’s handprinting recognition contest. Dr. Dobb’s Journal of Software Tools, 17(7):60–??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1992:E

[Val92c] Ray Valdés. Editorial. Dr.

Dobb’s Journal of Software Tools, 17(??):??, ???? 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Valdes:1992:FSD

[Val92d] Ray Valdés. Finding string distances. Dr. Dobb’s Journal of Software Tools, 17(4):56, 58, 61– 62, 107, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1992:SAF

[Val92e] Ray Valdés. Sizing up application frameworks and class libraries. Dr. Dobb’s Journal of Software Tools, 17(10):18–20, 22–30, 35, 86, 106–115, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1992:SGT

[Val92f] Ray Valdés. Sizing up GUI toolkits. Dr. Dobb’s Journal of Software Tools, 17(11):18–20, 22, 24, 26, 113–114, 116, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1993:DHR

[Val93a] Ray Valdés. DDJ handprinting recognition contest wrap-up. Dr. Dobb’s Journal of Software Tools, 18(1):52, 54, 56, 93, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1993:DWA

[Val93b] Ray Valdés. Debugging Windows applications. Dr. Dobb’s Journal of Software Tools, 18 (11):78–??, November 1993. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1993:LPL

[Val93c] Ray Valdés. The Liana programming language. Dr. Dobb’s Journal of Software Tools, 18 (11):50–52, October 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Valdes:1993:PBa

[Val93d]  Ray Valdes.  Programmer’s

bookshelf. Dr. Dobb’s Journal of Software Tools, 18(3):139–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1993:PBb

[Val93e]  Ray Valdes.  Programmer’s

bookshelf. Dr. Dobb’s Journal of Software Tools, 18(7):133–??, July 1993. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1993:TEA

[Val93f]  Ray Valdés.  Text editors:

Algorithms and architectures. Dr. Dobb’s Journal of Software Tools, 18(4):38, 40, 42–43, 80, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1994:COW

[Val94a] Ray Valdés. Component object wars heat up. Dr. Dobb’s Developer Update, 1(5):1–??, May 1994. CODEN ???? ISSN 1079-8595.

Valdes:1994:IHT

[Val94b] Ray Valdés. The internet: Here today. Dr. Dobb’s Developer Update, 1(8):1–??, August 1994. CODEN ???? ISSN 1079-8595.

Valdes:1994:ISC

[Val94c]  Ray Valdés.  Introducing

ScriptX [sidebar] in “character simulation with ScriptX” (A. Reznik). Dr. Dobb’s Journal of Software Tools, 19(13):78– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1994:LAB

[Val94d] Ray Valdés. Latin America: A boom market waiting to happen? Dr. Dobb’s Developer Update, 1(13):1–??, November 1994. CODEN ???? ISSN 1079-

Valdes:1994:MRC

[Val94e] Ray Valdés. A milestone on the road to Chicago. Dr. Dobb’s Developer Update, 1(7):1–??, July 1994. CODEN ???? ISSN 1079-8595.

Valdes:1994:NGI

[Val94f] Ray Valdés. The next generation of IP: Small step or giant leap. Dr. Dobb’s Developer Update, 1(10):1, 4, December 1994. CODEN ???? ISSN 1079-8595.

Valdes:1994:SPY

[Val94g] Ray Valdés. Software patents: You make the choice. Dr. Dobb’s Developer Update, 1(4): 2–??, April 1994. CODEN ???? ISSN 1079-8595.

Valdes:1994:WAK

[Val94h] Ray Valdés. What’s up at Kaleida. Dr. Dobb’s Developer Update, 1(7):1, 4, September 1994. CODEN ???? ISSN 1079-8595.

Valdes:1994:WL

[Val94i] Ray Valdés. Winners and losers. Dr. Dobb’s Developer Update, 1 (5):2–??, May 1994. CODEN ???? ISSN 1079-8595.

Valdes:1995:AFP

[Val95a] Ray Valdés. App frameworks in a Post-MFC world. Dr. Dobb’s Developer Update, 2(6): 1, 4, June 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:HIW

[Val95b] Ray Valdés. Here’s the Internet where’s the money? Dr.

Dobb’s Developer Update, 2(5): 1, 3, May 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:SBP

[Val95c] Ray Valdés. In search of best practices. Dr. Dobb’s Developer Update, 2(11):1, 3, 4, 6, November 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:MIP

[Val95d] Ray Valdés. More on Internet protocols. Dr. Dobb’s Developer Update, 2(7):4, July 1995. CO-DEN ???? ISSN 1079-8595.

Valdes:1995:NGJ

[Val95e] Ray Valdés. Net gets a Java buzz. Dr. Dobb’s Developer Update, 2(8):1, 3–4, 6, August 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:MM

[Val95f] Ray Valdés. Of milestones and men. Dr. Dobb’s Developer Update, 2(10):1, 3, 6, October 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:PPP

[Val95g] Ray Valdés. P6 processor in the pipeline. Dr. Dobb’s Developer Update, 2(4):1, 4, April 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:PCG

[Val95h] Ray Valdés. Pentium crosses the great divide. Dr. Dobb’s Developer Update, 2(2):1, 4, February 1995. CODEN ???? ISSN 1079-

Valdes:1995:PW

[Val95i] Ray Valdés. The Perl of Wizdom. Dr. Dobb’s Developer Update, 2(12):1, 3, 6, 6, December 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:PB

[Val95j]  Ray Valdes.  Programmer’s

bookshelf. Dr. Dobb’s Journal of Software Tools, 20(10): 149–??, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Valdes:1995:SSV

[Val95k] Ray Valdés. Startup: A silicon valley adventure. Dr. Dobb’s Developer Update, 2(7): 2–3, July 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:SS

[Val95l] Ray Valdés. A surfeit of servers. Dr. Dobb’s Developer Update, 2 (9):1, 3, 4, 6, September 1995. CODEN ???? ISSN 1079-8595.

Valdes:1995:WWW

[Val95m] Ray Valdés. Web wars widen. Dr. Dobb’s Developer Update, 2 (1):1–??, January 1995. CO-DEN ???? ISSN 1079-8595.

VanSickle:1991:CPM

[Van91] Truman T. Van Sickle. C programming for the 68HC05 microcontroller. Dr. Dobb’s Journal of Software Tools, 16(8):66, 68, 70, 74, 128, 132, 134, August 1991. CODEN DDJOEB. ISSN 1044-789X.

VanCamp:1993:WPW

[Van93] David Van Camp. Writing portable Windows applications. Dr. Dobb’s Journal of Software Tools, 18(10):52–??, October 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

VanCamp:1994:USG

[Van94] David Van Camp. Unicode and software globalization. Dr. Dobb’s Journal of Software Tools, 19(3):46, 48–50, March 1994. CODEN DDJOEB. ISSN 1044-789X. URL http://

davidvancamp.com/Articles_
files/SoftwareGlobalization.
DOC.

vanHoff:1995:JIP

[van95] Arthur van Hoff. Java and Internet programming. Dr. Dobb’s Journal of Software Tools, 20 (8):56, 58, 60–61, 101–102, August 1995. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ddj/issues/
j508a.htm.

VanCamp:1997:UUR

[Van97] David Van Camp. An unlimited undo/redo stack pattern for PowerBuilder. Dr. Dobb’s Journal of Software Tools, 22 (6):78–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

vanderWal:1998:PTS

[van98a] Ron van der Wal. Programmer’s toolchest: Source-code profilers for Win32. Dr. Dobb’s Journal of Software Tools, 23 (3):78, 80, 82–88, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        03/calmacst.zip;   http://
                                        www.ddj.com/ftp/1998/1998_
                                        03/mkdep.zip.

vanHoff:1998:JGB

[van98b] Arthur van Hoff. Java: Getting down to business: Is IS ready for real applications? Dr. Dobb’s Journal of Software Tools, 23 (1):20–22, 24, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Vandevoorde:1998:MRP

[Van98c] David Vandevoorde. The maximal rectangle problem. Dr. Dobb’s Journal of Software Tools, 23(4):28, 30–32, 100, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1998/1998_
                                        04/rectang.txt.

Varhol:1991:MCP

[Var91] Peter D. Varhol. ML and colored Petri nets for modeling and simulation. Dr. Dobb’s Journal of Software Tools, 16(9):76–77, 79, 81, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1992:DV

[Var92] P. D. Varhol. Design for visualization. Dr. Dobb’s Journal of Software Tools, 17(6):158, 160, 162, 164, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Varbol:1993:EVL

[Var93a] Peter D. Varbol. Extending A visual language for simulation. Dr. Dobb’s Journal of Software Tools, 18(6):92–??, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Varbol:1993:MSP

[Var93b] Peter D. Varbol. Modeling systems with polynomial networks. Dr. Dobb’s Journal of Software Tools, 18(9):76–??, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1993:EVL

[Var93c] Peter D. Varhol. Extending a visual language for simulation. Dr. Dobb’s Journal of Software Tools, 18(6):92–95, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1993:IDM

[Var93d] Peter D. Varhol. Integrating desktop mapping with geographic data. Dr. Dobb’s Journal of Software Tools, 18(8):84, 86, 88, 90, 114–115, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1993:MSP

[Var93e] Peter D. Varhol. Modeling systems with polynomial networks. Dr. Dobb’s Journal of Software Tools, 18(9):76, 78, 80, 82, 96, September 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Varbol:1994:PB

[Var94a] Peter D. Varbol. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(13):151– ??, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1994:BAV

[Var94b] Peter D. Varhol. Building applications with visual tools. Dr. Dobb’s Developer Update, 1(8): 2–??, August 1994. CODEN ???? ISSN 1079-8595.

Varhol:1994:PBa

[Var94c] Peter D. Varhol. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(1): 131–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1994:PBb

[Var94d] Peter D. Varhol. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 19(2): 117–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1994:TOS

[Var94e] Peter D. Varhol. Trends in operating system design. Dr. Dobb’s Journal of Software Tools, 19 (5):18–20, 22, 26–27, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Varbol:1995:ANS

[Var95a] Peter D. Varbol. An architecture for network simulation. Dr. Dobb’s Journal of Software Tools, 20(7):70–??, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1995:ANS

[Var95b] Peter D. Varhol. An architecture for network simulation. Dr. Dobb’s Journal of Software Tools, 20(7):70, 72, 74, 76, 78, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1995:PQ

[Var95c] Peter D. Varhol. Photon and QNX. Dr. Dobb’s Journal of Software Tools, 20(5):90, 92–94, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Varhol:1995:PB

[Var95d] Peter D. Varhol. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 20(5):127–??, May 1995. CODEN DDJOEB. ISSN 1044-789X.

vanderWal:1996:DDS

[vdW96] Ron van der Wal. Designing a distributed simulation game. Dr. Dobb’s Journal of Software Tools, 21(??):21–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Wal:1997:EWC

[vdW97a] Ron van der Wal. Examining Win32 C++ compilers. Dr. Dobb’s Journal of Software Tools, 22(3):64–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Wal:1997:TTI

[vdW97b] Ron van der Wal. Testing testers — in an ideal world, development progresses smoothly from requirements to completion. in the real world, errors creep in. Ron examines a collection of commercially available automated testing tools that aid in ferreting out those errors. Dr. Dobb’s Journal of Software Tools, 22(2):76–??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Wal:1997:PBR

[vdWS97] Ron van der Wal and William Stallings. Programmer’s bookshelf — Ron examines Stanley Lippman’s Inside the C++ Object Model, while William looks at Peter Wayner’s Disappearing Cryptography. Dr. Dobb’s Journal of Software Tools, 22 (1):116–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Veldhuizen:1997:SCC

[Vel97] Todd Veldhuizen. Scientific computing: C++ versus Fortran: C++ has more than caught up. Dr. Dobb’s Journal of Software Tools, 22(11):34, 36– 38, 91, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Vermeulen:1993:OFS

[Ver93]  Allan Vermeulen.  Object-

oriented finite-element software. Dr. Dobb’s Journal of Software Tools, 18(6):86–91, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Vermeulen:1996:ADP

[Ver96] Allan Vermeulen. An asynchronous design pattern: A pattern for managing concurrency— written in Java! Dr. Dobb’s Journal of Software Tools, 21 (6):42, 44, 84–86, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Vermeulen:1997:JDW

[Ver97] Alain Vermeulen. Java deadlock: The woes of multithreaded design. Dr. Dobb’s Journal of Software Tools, 22(9):52, 54–56, 88, 89, September 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Vigot:1993:FLC

[Vig93] Greg Vigot. Fuzzy logic in C. Dr. Dobb’s Journal of Software Tools, 18(2):40–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Vilot:1995:CSL

[Vil95] Michael J. Vilot. The C++ Standard Library. Dr. Dobb’s Journal of Software Tools, 20 (8):111–113, 144–145, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Vilot:1996:ESL

[Vil96] Michael J. Vilot. Examining the Systems<ToolKit> library. Dr. Dobb’s Journal of Software Tools, 21(11):80–??, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Viot:1993:FLC

[Vio93] Greg Viot. Fuzzy logic in C.

Dr. Dobb’s Journal of Software Tools, 18(2):40, 42, 44–45, 48– 49, 94, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Vlissides:1994:P

[Vli94] John Vlissides. PLoP ’94. Dr. Dobb’s Developer Update, 1(13): 3–??, November 1994. CODEN ???? ISSN 1079-8595.

Vogel:1991:SCG

[Vog91] Karl Vogel. A source code generator for C. Dr. Dobb’s Journal of Software Tools, 16(8):28, 30, 32, 34–35, 102–104, 106– 108, 111, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Voss:1994:PC

[Vos94] Greg Voss. Polymorphic C. Dr. Dobb’s Journal of Software Tools, 19(8):36, 38, 40–42, 94– 95, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Veldhuizen:1996:LAC

[VP96] Todd Veldhuizen and Kumaraswamy Ponnambalam. Linear algebra with C++ template metaprograms. Dr. Dobb’s Journal of Software Tools, 21 (8):38, 40–42, 44, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Walsh:1992:PQ

[Wal92a] Aaron E. Walsh. Programming QuickTime. Dr. Dobb’s Journal of Software Tools, 17(7):76, 78–80, 102, 104–105, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Walsh:1992:SDT

[Wal92b] Aaron E. Walsh. Sound as A data type. Dr. Dobb’s Journal of Software Tools, 17(12):28, 30, 32, 102–103, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Wallace:1995:DSE

[Wal95a] Rob Wallace. Digital sound engineering for game development [sidebar]. Dr. Dobb’s Journal of Software Tools, 20(??):26– ??, May/June 1995. CODEN DDJOEB. ISSN 1044-789X.

Walton:1995:IAS

[Wal95b] Steve Walton. Image authentication for a slippery new age. Dr. Dobb’s Journal of Software Tools, 20(4):18–20, 22, 24, 26, 82, 84–87, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Walworth:1997:JGT

[Wal97] Alan Walworth. Java GUI testing — as Java development takes off, so does the need to test GUI applications written in Java. Alan examines Java GUI testing issues and presents a test harness written in and for Java. Dr. Dobb’s Journal of Software Tools, 22(2):30, 32, 34, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Warren:1991:WPI

[War91] Jim Warren. We the people in the information age. Dr. Dobb’s Journal of Software Tools, 16 (1):96D, January 1991. CODEN DDJOEB. ISSN 1044-789X.

White:1995:PFS

[WB95] Ronald G. White and John Biard. A portable font specification. Dr. Dobb’s Journal of Software Tools, 20(3):28, 30, 32, 34, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Wutka:1997:HJP

[WBB^{+}97] Mark Wutka, David Baker, David Boswell, Ken Cartwright, and Da(??) Liebke. Hacking Java: the professional’s resource kit. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-0935-X. xviii + 852 pp. LCCN QA76.73.J38W88 1997. US$59.99; 119,00 DM. URL

http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm. In-

cludes CD-ROM. See review [You97].

Wilson:1999:PBG

[WC99] Gregory V. Wilson and Steve Chartley. Programmer’s bookshelf: The general, the particular, and the just plain odd. Dr. Dobb’s Journal of Software Tools, 24(4):117–118, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9904q/9904q.
htm.

Weaver:1997:PTM

[Wea97] Mark Weaver. Programmer’s toolchest: Modeling with MAT-LAB: A flexible tool for system solutions. Dr. Dobb’s Journal of Software Tools, 22(11):80–83, 94–95, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Weber:1995:AAC

[Web95] David Weber. Associative arrays in C++. Dr. Dobb’s Journal of Software Tools, 20(??):13– ??, July/August 1995. CODEN DDJOEB. ISSN 1044-789X.

Webber:1996:FIP

[Web96] Victor Webber. Fast interrupt processing in Windows 95. Dr. Dobb’s Journal of Software Tools, 21(??):15–??, March/April 1996. CODEN DDJOEB. ISSN 1044-789X.

Weeks:1990:CFO

[Wee90] Kevin Weeks. C++ file objects.

Dr. Dobb’s Journal of Software Tools, 15(6):50–55, 96–98, 100– 102, 106–112, 115, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Weisfeld:1993:PLE

[Wei93] Matt Weisfeld. A portable library for executing child processes. Dr. Dobb’s Journal of Software Tools, 18(5):46, 48, 50– 52, 90, 92, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Weisfeld:1995:PSH

[Wei95] Matt Weisfeld. Portable screen handling. Dr. Dobb’s Journal of Software Tools, 20(5):142–143, 146, 148–149, May 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Weismann:1996:BES

[Wei96a] Gregg Weismann. Building an expert system. Dr. Dobb’s Journal of Software Tools, 21(8): 84–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Weissman:1996:BES

[Wei96b] Gregg Weissman. Building an expert system. Dr. Dobb’s Journal of Software Tools, 21(8):84, 86–87, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Wells:1994:WSD

[Wel94] Bill Wells. Writing serial drivers for UNIX. Dr. Dobb’s Journal of Software Tools, 19(15):68, 70–71, 97–99, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Welsh:1995:ILK

[Wel95] Matt Welsh. Implementing loadable kernel modules for Linux. Dr. Dobb’s Journal of Software Tools, 20(5):18–20, 22, 24, 96, May 1995. CODEN DDJOEB. ISSN 1044-789X.

WG:1999:LAA

[WEM^{+}99] WG, Toby Everett, Howard Mark, Peter Roth, Ron Verbruggen, Robert Patterson, Aspi Havewala, Kit Adams, and Ed Remmell. Letters: Analyzing algorithms; the version control process; electronic music. Dr. Dobb’s Journal of Software Tools, 24(10):10, 12, October 1999. CODEN DDJOEB. ISSN 1044-789X.

Wendt:1999:MRV

[Wen99]  David Wendt.  Mindstorms

robotics and Visual C++. Dr. Dobb’s Journal of Software Tools, 24(8):52, 54–57, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
08/legocpp.txt;   http://
www.ddj.com/ftp/1999/1999_
08/legocpp.zip.

West:1994:MCA

[Wes94] Alan West. Making a case for animating C++ programs. Dr. Dobb’s Journal of Software Tools, 19(11):54–60, October 1994. CODEN DDJOEB. ISSN 1044-789X.

White:1996:JCA

[Whi96] Greg White. Java command-line arguments. Dr. Dobb’s Journal of Software Tools, 21(2):58, 60, 62, 96, 98–99, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Wicker:1999:SSW

[Wic99] Louis J. Wicker. Simulating severe weather. Dr. Dobb’s Journal of Software Tools, 24(3):18– 20, 22, 24, 26–28, March 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm.

Williams:1990:DG

[Wil90a] Al Williams. DOS + 386 = 4 gigabytes! Dr. Dobb’s Journal of Software Tools, 15(7):62, 64, 66, 68, 70–71, 110–112, 115, July 1990. CODEN DDJOEB. ISSN 1044-789X.

Williams:1990:HDS

[Wil90b] Al Williams. Homegrown debugging — 386 style! Dr. Dobb’s Journal of Software Tools, 15(3):46–57, 96–105, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Williams:1990:RYOa

[Wil90c] Al Williams. Roll your own DOS extender: Part I. Dr. Dobb’s Journal of Software Tools, 15 (10):16–18, 20, 24, October 1990. CODEN DDJOEB. ISSN 1044-789X.

Williams:1990:RYOb

[Wil90d] Al Williams. Roll your own DOS extender: Part II. Dr. Dobb’s Journal of Software Tools, 15 (11):74, 76, 80, 81–83, 85, 122– 128, 130, November 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Williams:1990:SSS

[Wil90e] Walter Williams. Supercharging sequential searches. Dr. Dobb’s Journal of Software Tools, 15(12):54, 56, 58–60, 62, 100, December 1990. CODEN DDJOEB. ISSN 1044-789X.

Williams:1992:PPL

[Wil92a] Al Williams. Programming with Phar Lap’s 286 mod DOSextender. Dr. Dobb’s Journal of Software Tools, 17(2):60, 62, 64, 66, 68, 94, 96, 98–100, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Williams:1992:YOD

[Wil92b] Al Williams. Your own disk duplication program [ER review]. Dr. Dobb’s Journal of Software Tools, 17(1):58, 60, 62, 82, 84, 86–88, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Wilt:1992:ALP

[Wil92c] Nicholas Wilt. Assembly language programming for the 80*87. Dr. Dobb’s Journal of Software Tools, 17(3):36, 38, 40, 42, 88, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Wilt:1992:TC

[Wil92d] Nicholas Wilt. Templates in C++. Dr. Dobb’s Journal of Software Tools, 17(12 (special supplement)):29s–33s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Williams:1993:FM

[Wil93a] Al Williams. 32-bit floatingpoint math. Dr. Dobb’s Journal of Software Tools, 18(6):70, 72, 74, 76, 80, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Williams:1993:QPQ

[Wil93b] Al Williams. A quick port with QuickWin. Dr. Dobb’s Journal of Software Tools, 18(8):92, 94– 95, 98, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Williams:1993:VVW

[Wil93c] Al Williams. VWinL: A virtual window library. Dr. Dobb’s Journal of Software Tools, 18 (10):28–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Williams:1993:CD

[Wil93d] Lowell Williams. Compound documents. Dr. Dobb’s Journal of Software Tools, 18(3):32, 34, 38–39, 101, March 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Wilcox:1994:OD

[Wil94a] Jonathan Wilcox. Object databases. Dr. Dobb’s Journal of Software Tools, 19(13):26, 28– 30, 32, 34, November 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Williams:1994:PW

[Wil94b] Al Williams. DOS pipes for Windows. Dr. Dobb’s Journal of Software Tools, 19(??):26–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Williams:1994:NTM

[Wil94c] Al Williams. NT-Style threads for MS-DOS. Dr. Dobb’s Journal of Software Tools, 19(2):74, 76–77, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Wilmot:1994:IW

[Wil94d] Dick Wilmot. 2PANE illuminates Windows. Dr. Dobb’s Journal of Software Tools, 19 (??):6–??, ???? 1994. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Wilson:1994:ISM

[Wil94e] Les Wilson. Inside OS/2 software motion video. Dr. Dobb’s Journal of Software Tools, 19 (??):34–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Williams:1995:IWP

[Wil95a] Al Williams. Installing Windows 95 programs. Dr. Dobb’s Journal of Software Tools, 20(9):28, 30, 32, 34–36, 100–104, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Williams:1995:SWD

[Wil95b] Al Williams. Simplifying Windows development. Dr. Dobb’s Journal of Software Tools, 20 (??):15–??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Williams:1995:VCD

[Wil95c] Al Williams. Visually constructing Delphi components. Dr. Dobb’s Journal of Software Tools, 20(12):34, 36, 39–41, 102– 105, December 1995. CODEN DDJOEB. ISSN 1044-789X.

Wild:1996:BCT

[Wil96a] Fred Wild. Beyond C++ templates. Dr. Dobb’s Journal of Software Tools, 21(8):16–18, 22–23, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Wild:1996:ICP

[Wil96b] Fred Wild. Instantiating code patterns. Dr. Dobb’s Journal of Software Tools, 21(6):72, 74– 76, 88–91, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Williams:1996:CDA

[Wil96c] Al Williams. Customizing Delphi applications. Dr. Dobb’s Journal of Software Tools, 21 (??):7–??, March/April 1996. CODEN DDJOEB. ISSN 1044-789X.

Williams:1996:CEO

[Wil96d] Al Williams. Customizing the Explorer open dialog. Dr. Dobb’s Journal of Software Tools, 21(9):24, 26–29, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Williams:1996:LL

[Wil96e] Al Williams. Lessons to learn. Dr. Dobb’s Journal of Software Tools, 21(??):51–??, March/ April 1996. CODEN DDJOEB. ISSN 1044-789X.

Williams:1996:PM

[Wil96f] Al Williams. Peace maker. Dr. Dobb’s Journal of Software Tools, 21(??):55–??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Williams:1996:RWC

[Wil96g] Al Williams. Recycling Windows controls for Delphi. Dr. Dobb’s Journal of Software Tools, 21(??):51–??, January/ February 1996. CODEN DDJOEB. ISSN 1044-789X.

Williams:1996:ELR

[Wil96h] Jonathan Williams. Examining the LabObjX real-time chart control. Dr. Dobb’s Journal of Software Tools, 21(10):88– ??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Willston:1996:LEJ

[Wil96i] John B. Willston. Letter to the editor: Java dive. Dr. Dobb’s Journal of Software Tools, 21 (12):8, December 1996.

Wild:1997:AAB

[Wil97a] Fred Wild. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 22(3):105–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Wild:1997:DT

[Wil97b] Fred Wild. Designing for testability. Dr. Dobb’s Journal of Software Tools, 22(6):42–??, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Williams:1997:IPW

[Wil97c] Al Williams. Internet programming — writing ActiveX ISAPI extensions. Dr. Dobb’s Journal of Software Tools, 22(8):72, 74– 77, 95, 98, August 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Williams:1997:VBA

[Wil97d] Al Williams. Visual Basic 5 and ActiveX controls. Dr. Dobb’s Journal of Software Tools, 22 (3):74–??, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1997:JNM

[Wil97e] Andrew Wilson. The Java Native Method Interface and Windows. Dr. Dobb’s Journal of Software Tools, 22(8):46, 48– 50, 94, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1997:SSD

[Wil97f] Andrew Wilson. Software security and the DirectPlay API. Dr. Dobb’s Journal of Software Tools, 22(4):66–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1997:PBMb

[Wil97g] Gregory Wilson. Programmer’s bookshelf — this month, Greg examines Jonathan Rosenberg’s How Debuggers Work, Mark Woodman’s Programming Language Choice, and The Media. Dr. Dobb’s Journal of Software Tools, 22(7):117–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1997:PBMa

[Wil97h] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 22 (3):117–??, March 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Wilson:1997:PBA

[Wil97i] Gregory V. Wilson. Programmer’s bookshelf: From ActiveX to cargo-cult science. Dr. Dobb’s Journal of Software Tools, 22 (9):117–119, September 1997. CODEN DDJOEB. ISSN 1044-789X. Contains reviews of Java books [App97, CM97].

Wilson:1997:PBH

[Wil97j] Gregory V. Wilson. Programmer’s bookshelf: HAL’s legacy and large-scale software design. Dr. Dobb’s Journal of Software Tools, 22(5):124–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1997:PBP

[Wil97k] Gregory V. Wilson. Programmer’s bookshelf: Perusing the bookshelf. Dr. Dobb’s Journal of Software Tools, 22(11):125, 127, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Wild:1998:CIK

[Wil98a] Fred Wild. C++ interfaces:

Keeping interfaces and implementations separate. Dr. Dobb’s Journal of Software Tools, 23 (8):18–20, 22, 89, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
08/cppinter.txt;   http://
www.ddj.com/ftp/1998/1998_
08/cppinter.zip.

Wilson:1998:EMJ

[Wil98b] Andrew Wilson. Examining Microsoft’s J/Direct: Making Java native code easy on Windows 95/NT. Dr. Dobb’s Journal of Software Tools, 23(1):90, 92–96, 104–105, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Wilson:1998:PBA

[Wil98c] Gregory V. Wilson. Programmer’s bookshelf: Acronyms by the bushel: VB, UML, CRC, and GUI. Dr. Dobb’s Journal of Software Tools, 23(1):133– 134, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1998:PB

[Wil98d] Gregory V. Wilson. Programmer’s bookshelf: No engineers wer harmed in the production of this book. Dr. Dobb’s Journal of Software Tools, 23(11):143, 145, November 1998. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1998:WCDa

[Wil98e] James Y. Wilson. Windows CE device driver development, part I. Dr. Dobb’s Journal of Software Tools, 23(10): 66, 68, 70, 72, 74, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
10/index.htm.

Wilson:1998:WCDb

[Wil98f] James Y. Wilson. Windows CE device driver development, part II. Dr. Dobb’s Journal of Software Tools, 23(11):68, 70, 72, 74, 76–78, 80, 82, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
11/WINCEDD.TXT.

Wilson:1999:JQH

[Wil99a] Andrew Wilson. Java Q&A: How do I implement Microsoft’s delegate in Pure Java? Dr. Dobb’s Journal of Software Tools, 24(3):107–110, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/jqa399.txt;
 http://www.ddj.com/ftp/1999/
1999_03/jqa399.zip.

Wilson:1999:JQW

[Wil99b] Andy Wilson. Java Q&A: What is the Java VM Profiler interface? Dr. Dobb’s Journal of Software Tools, 24(9): 103–106, September 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_09/
jqa999.txt;   http://www.
ddj.com/ftp/1999/1999_09/
jqa999.zip.

Wilson:1999:PBJ

[Wil99c] Gregory V. Wilson. Programmer’s bookshelf: A joke too far. Dr. Dobb’s Journal of Software Tools, 24(11):127–128, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1999:PBA

[Wil99d] Gregory V. Wilson. Programmer’s bookshelf: Another basketful of books. Dr. Dobb’s Journal of Software Tools, 24(8):119–120, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/articles/1999/
9908/9908p/9908p.htm.

Wilson:1999:PBS

[Wil99e] Gregory V. Wilson. Programmer’s bookshelf: Small is beautiful — kind of. Dr. Dobb’s Journal of Software Tools, 24 (2):135, 137, February 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm;
 http://www.ercb.com/ddj/
1999/ddj.9902.html.

Winton:1999:EP

[Win99] Greg Winton. Extending the PalmOS. Dr. Dobb’s Journal of Software Tools, 24(11):50, 52–55, November 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_11/
                                        palmos.zip.

Wittenburg:1995:ABG

[Wit95a] Tim Wittenburg. Alpha blending graphic images. Dr. Dobb’s Journal of Software Tools, 20 (8):48, 50–52, 54, 100–101, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Wittenburg:1995:DMC

[Wit95b] Tim Wittenburg. Data models, case tools, and client/server development. Dr. Dobb’s Journal of Software Tools, 20(11): 92–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Withers:1997:REC

[Wit97a]  Wm. Douglas Withers.  A

rapid entropy-coding algorithm. Dr. Dobb’s Journal of Software Tools, 22(4):38–??, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Wittenburg:1997:MBE

[Wit97b] Tim Wittenburg. Motion blur effects. Dr. Dobb’s Journal of Software Tools, 22(7):36–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Wiebel:1992:FLR

[WJ92] M. Wiebel and S. Johnson. The five levels of RAID (discrete event simulation). Dr. Dobb’s Journal of Software Tools, 17 (1):24, 26, 28, 30, 78–80, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Widom:1997:LES

[WL97] Allan Widom and Steen Lehmann.

Letters to the editor: Swaine on Java. Dr. Dobb’s Journal of Software Tools, 22(4):10, 12, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1991:FSR

[Woe91a] Jack J. Woehr. Forth: A status report. Dr. Dobb’s Journal of Software Tools, 16(10): 96–97, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1991:MFM

[Woe91b] Jack J. Woehr. A mediumheavyweight Forth multitasker. Dr. Dobb’s Journal of Software Tools, 16(6):66, 68, 71, 73, 98, 100, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1992:SSI

[Woe92a] Jack J. Woehr. Sculpting on silicon: An interview with Chuck Moore. Dr. Dobb’s Journal of Software Tools, 17(6):48–??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1992:WIS

[Woe92b] Jack J. Woehr. Writable instruction set computers. Dr. Dobb’s Journal of Software Tools, 17 (1):32, 89–92, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1993:PM

[Woe93] Jack J. Woehr. Programming the Motorola 68332. Dr. Dobb’s Journal of Software Tools, 18 (8):66, 68, 70, 106, 108, 110, 112, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1994:LVPb

[Woe94a] Jack Woehr. Lotfi visions, part 2. Dr. Dobb’s Journal of Software Tools, 19(8):52–??, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1994:LVPa

[Woe94b] Jack J. Woehr. Lofti visions, part 1. Dr. Dobb’s Journal of Software Tools, 19(7):44–??, July 1994. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1996:CMC

[Woe96a] Jack J. Woehr. A conversation with Michael Cowlishaw. Dr. Dobb’s Journal of Software Tools, 21(3):36–??, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1996:IDK

[Woe96b] Jack J. Woehr. An interview with Donald Knuth. Dr. Dobb’s Journal of Software Tools, 21 (4):16–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1997:CEB

[Woe97a] Jack Woehr. A conversation with Eva Bozoki. Dr. Dobb’s Journal of Software Tools, 22 (1):16–??, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Woehr:1997:CRR

[Woe97b] Jack Woehr. A conversation with Ron Rivest: How important is cryptography and computer security? Dr. Dobb’s Journal of Software Tools, 22 (10):18–20, 22, 24, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Woehr:1997:CWK

[Woe97c] Jack Woehr. A conversation with William Kahan: How important is numerical accuracy? Dr. Dobb’s Journal of Software Tools, 22(11):18–20, 22, 24, 26, 30, 32, November 1997. CO-DEN DDJOEB. ISSN 1044-789X. Kahan, the father of the IEEE 754 floating-point standard, talks about floating-point arithmetic issues, and numerical deficiencies in Java.

Woehr:1999:CGR

[Woe99] Jack J. Woehr. A conversation with Glenn Reeves. Dr. Dobb’s Journal of Software Tools, 24 (11):21–22, 24, 26–28, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Wong:1992:OPC

[Won92] William G. Wong. Object-Oriented program construction. Dr. Dobb’s Journal of Software Tools, 17(10):36–42, 116– 118, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Wong:1999:JNE

[Won99] Hinkmond Wong. Jini and network-enabled devices. Dr. Dobb’s Journal of Software Tools, 24(7):21–22, 24–26, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        07/jini.txt;   http://www.
                                        ddj.com/ftp/1999/1999_07/
                                        jini.zip.

Woodson:1999:POS

[Woo99] Morgan Woodson. The PowerTV operating system. Dr. Dobb’s Journal of Software Tools, 24(11):56, 58–60, November 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        11/ptvos.zip.

Wright:1994:PDO

[Wri94] Jon Wright. Porting D-Flat++ to OS/2. Dr. Dobb’s Journal of Software Tools, 19(5):86, 88– 89, 110–111, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Wright:1996:MCC

[Wri96] George Wright. A MIDI class in C++. Dr. Dobb’s Journal of Software Tools, 21(??):33– ??, May/June 1996. CODEN DDJOEB. ISSN 1044-789X.

Wright:1999:PSL

[Wri99] Ed Wright. Porting the SPICE library. Dr. Dobb’s Journal of Software Tools, 24(12):94, 96–99, December 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_12/
                                        doc2c.zip;   http://www.ddj.
                                        com/ftp/1999/1999_12/spice.
                                        txt.

Wright:1990:MPA

[WS90] Karl Wright and Rick Schell.

Mixed-language programming with ASM. Dr. Dobb’s Journal of Software Tools, 15(3):84–88, 116–121, March 1990. CODEN DDJOEB. ISSN 1044-789X.

Wingo:1996:CM

[WS96a] Scot Wingo and George Shepherd. Customizing MFC. Dr. Dobb’s Journal of Software Tools, 21(6):??, June 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Wingo:1996:UC

[WS96b] Scot Wingo and George Shepherd. Undocumented corner. Dr. Dobb’s Journal of Software Tools, 21(4):115–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Wingo:1996:UCC

[WS96c] Scot Wingo and George Shepherd. Undocumented corner: Customizing MFC. Dr. Dobb’s Journal of Software Tools, 21 (6):111–??, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Wingo:1996:UCI

[WS96d] Scot Wingo and George Shepherd. Undocumented corner: Inside OLE control property pages. Dr. Dobb’s Journal of Software Tools, 21(8):121– ??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1999:PBR

[WS99] Gregory V. Wilson and William Stallings. Programmer’s bookshelf: A revolution oft-delayed. Dr. Dobb’s Journal of Software Tools, 24(5):133–134, May 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/9905o/9905o.htm.

WooSam:1995:OCI

[WSM95a] J. Woo-Sam and T. Murphy. The Oracle Call Interface and C++. Dr. Dobb’s Journal of Software Tools, 20(11):38, 40– 42, 45, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Woo-Sam:1995:OCI

[WSM95b] Jeremy Woo-Sam and Tony Murphy. The Oracle call interface and C++. Dr. Dobb’s Journal of Software Tools, 20(11): 38–??, November 1995. CODEN DDJOEB. ISSN 1044-789X.

Wu:1997:CSI

[Wu97]  Amy Wu.  Chip shots II.

Dr. Dobb’s Journal of Software Tools, 22(1):12, January 1997. CODEN DDJOEB. ISSN 1044-789X. Short note about Sun’s JavaChip silicon project, and the picoJava I chip.

Varhol:1991:GDV

[WV91a] Marian G. Williams and Peter D. Varhol. Graphical data visualization. Dr. Dobb’s Journal of Software Tools, 16(12): 88–??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Williams:1991:GDV

[WV91b] Marian G. Williams and Peter D. Varhol. Graphical data visualization. Dr. Dobb’s Journal of Software Tools, 16(12):88–90, 93, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1996:TIW

[WV96] Andrew Wilson and Peter D.

Varhol. TCP/IP and Windows 95. Dr. Dobb’s Journal of Software Tools, 21(1):78–??, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Webb:1999:MSL

[WW99] Peter Webb and Gregory V.

Wilson. Matlab as a scripting language: A simple way to do powerful things. Dr. Dobb’s Journal of Software Tools, 24 (1):94, 96–99, January 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/ftp/1999/1999_01/
matlab.txt.

Xiao:1995:VBF

[Xia95] Wei Xiao. A Visual Basic form generator. Dr. Dobb’s Journal of Software Tools, 20(??):23– ??, March/April 1995. CODEN DDJOEB. ISSN 1044-789X.

Yam:1993:EM

[Yam93] Michael Yam. Examining MFC 2.0 [review]. Dr. Dobb’s Journal of Software Tools, 18(6):114, 116–119, 132, 134, June 1993. CODEN DDJOEB. ISSN 1044-789X.

Yam:1994:ESC

[Yam94]  Michael Yam.  Examining

Symantec C++. updating the PT periodic table program. Dr. Dobb’s Journal of Software Tools, 19(6):80, 82–83, 100–101, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Yam:1995:CFD

[Yam95]  Michael Yam.  A C++

framework for DCE threads. Dr. Dobb’s Journal of Software Tools, 20(??):27–??, July/ August 1995. CODEN DDJOEB. ISSN 1044-789X.

Yam:1996:DPV

[Yam96] Michael Yam. DCE pthreads versus NT threads. Michael ports PTF, a C++ class library for DCE pthreads, from HP-UX System 9 to Windows NT. in doing so, he examines the differences between pthreads and NT threads, and describes the porting experience. Dr. Dobb’s Journal of Software Tools, 21(12): 16–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

York:1995:SCU

[Yor95] Kyle A. York. Sound compression using quantized deltas. Dr. Dobb’s Journal of Software Tools, 20(6):121–122, June 1995. CODEN DDJOEB. ISSN 1044-789X.

York:1996:BSN

[Yor96] Kyle A. York. Building a DOS serial network. Dr. Dobb’s Journal of Software Tools, 21(5):38, 40–43, 80, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Young:1992:PW

[You92] Michael J. Young. Printing from Windows 3. Dr. Dobb’s Journal of Software Tools, 17(3):140, 142, 144, 146, March 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Young:1996:PB

[You96] Warren Young. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 21(10): 134–??, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Young:1997:PBJ

[You97] Warren Young. Programmer’s bookshelf: Java pro. Dr. Dobb’s Journal of Software Tools, 22 (10):133–134, October 1997. CODEN DDJOEB. ISSN 1044-789X. Contains book reviews of [CH97a, WBB^{+}97].

Yourst:1998:IJC

[You98] Matt T. Yourst. Inside Java class files: The key to Java’s binary portability. Dr. Dobb’s Journal of Software Tools, 23 (1):46, 49, 51–52, 102–103, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Yu:1996:DMC

[Yu96] Tong Lai Yu. Dynamic Markov compression. Dr. Dobb’s Journal of Software Tools, 21(1):30, 32, 96, 98–100, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Yuen:1994:YOT

[Yue94] Andy Yuen. Your own Token-Ring network manager. Dr. Dobb’s Journal of Software Tools, 19(2):58, 60–62, 64– 65, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Yuen:1996:CSC

[Yue96a] Andy Yuen. Concurrent Small C. Dr. Dobb’s Journal of Software Tools, 21(8):62, 64–66, 68, 71, 73, 94–96, August 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Yuen:1996:TPM

[Yue96b] Andy Yuen. A tiny preemptive multitasking Forth. Dr. Dobb’s Journal of Software Tools, 21 (3):58, 60, 62, 64, 66, 68, 93–94, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Yuen:1997:ESR

[Yue97] Andy Yuen. Embedded systems — retargetable Concurrent Small C. Dr. Dobb’s Journal of Software Tools, 22(8):58, 60, 62, 64, 66, 67, 69, 70, 94, 95, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ia

[Zem91a]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(1):??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ib

[Zem91b]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(2):??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ic

[Zem91c]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(3):??, March 1991.

CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Id

[Zem91d]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(4):??, April 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ie

[Zem91e]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1991:If

[Zem91f]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(6):??, June 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ig

[Zem91g]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(7):??, July 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ih

[Zem91h]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(8):??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ii

[Zem91i]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(9):??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ij

[Zem91j]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(10):??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Ik

[Zem91k]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(11):??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1991:Il

[Zem91l]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 16(12):??, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ia

[Zem92a]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(1):??, January 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ib

[Zem92b]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(2):??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ic

[Zem92c]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Id

[Zem92d]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ie

[Zem92e]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1992:If

[Zem92f]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ig

[Zem92g]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ih

[Zem92h]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(8):??, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ii

[Zem92i]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(9):170–??, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ij

[Zem92j]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(10):??, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Ik

[Zem92k]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(11):200–??, November 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1992:Il

[Zem92l]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 17(12):162–??, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1993:Ia

[Zem93a]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 18(1):146–??, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1993:Ib

[Zem93b]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 18(2):146–??, February 1993. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1993:Ic

[Zem93c]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 18(3):160–??, March 1993. CODEN DDJOEB. ISSN 1044-789X.

Zemel:1993:Id

[Zem93d]  Tami Zemel.  Of interest.

Dr. Dobb’s Journal of Software Tools, 18(4):148–??, April 1993. CODEN DDJOEB. ISSN 1044-789X.

Zhang:1997:DWP

[Zha97] Tony Zhang. A disassembler written in Perl — Tony presents the core subroutines of a disassembler written in Perl. although designed for Intel’s x86 instruction set, you can easily modify or customize the disassembler for your own applications. Dr. Dobb’s Journal of Software Tools, 22(2):36– ??, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Zigon:1990:PC

[Zig90] Robert Zigon. Parametric circles. Dr. Dobb’s Journal of Software Tools, 15(1):60, 63, January 1990. CODEN DDJOEB. ISSN 1044-789X.

Zigmond:1995:GPC

[Zig95] Dan Zigmond. Generic programming and the C++ STL. Dr. Dobb’s Journal of Software Tools, 20(8):18–20, 22, 24, 96, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Zigmond:1996:SA

[Zig96a] Dan Zigmond. STL algorithms.

Dr. Dobb’s Journal of Software Tools, 21(8):32, 34–36, 92– 94, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Zigmond:1996:SI

[Zig96b] Dan Zigmond. STL iterators. Dr. Dobb’s Journal of Software Tools, 21(6):24, 26–28, 30– 31, 84, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Zudock:1996:VAT

[Zud96] Tom Zudock. Virtual audio through ray tracing. Dr. Dobb’s Journal of Software Tools, 21 (12):34–??, December 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Zytaruk:1994:WVMa

[Zyt94a]  Kelly Zytaruk.  The Win-

dows 3.1 virtual machine control block. Part 1. Dr. Dobb’s Journal of Software Tools, 19 (1):115–??, January 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Zytaruk:1994:WVMb

[Zyt94b]  Kelly Zytaruk.  The Win-

dows 3.1 virtual machine control block. Part 2. Dr. Dobb’s Journal of Software Tools, 19 (2):107–??, February 1994. CO-DEN DDJOEB. ISSN 1044-789X.